### v1

# Differential Test Performance and Peer Effects

Abstract
We use variation of test scores measuring closely related skills to isolate peer effects. The intuition for our identification strategy is that the difference in closely related scores eliminates factors common to the performance in either test while retaining idiosyncratic test specific variation. Common factors include unobserved teacher and group effects as well as test invariant ability and factors relevant for peer group formation. Peer effects work through idiosyncratic shocks which have the interpretation of individual and test specific ability or effort. We use education production functions as well as restrictions on the information content of unobserved test taking ability to formalize our approach. An important implication of our identifying assumptions is that we do not need to rely on randomized group assignment. We show that our model restrictions are sufficient for the formulation of linear and quadratic moment conditions that identify the peer effects parameter of interest. We use Project STAR data to empirically measure peer effects in kindergarten through third grade classes. We find evidence of highly significant peer effects with magnitudes that are at the lower end of the range of estimates found in the literature.
1 Introduction
We develop a framework to analyze, identify and estimate the effect of peer groups on performance measures for individuals who are allocated to groups. A leading example are test scores of students who are allocated to classrooms. However, the theory could equally be applied to outcome measures of members of sports teams, teams of workers or other groups. Identification is based on the availability of multiple performance measures of comparable quality or information content. We allow team formation to be endogenous but also cover the case of randomly selected groups. For expositional purposes we refer to groups as classrooms and individuals selected into groups as students with the understanding that our procedure applies to a broader set of scenarios. For the statistical implementation we impose linear functional form restrictions. This allows us to identify the marginal effect of observed and unobserved average peer characteristics.
Our analysis centers around the idea of measuring the performance of individuals in closely related tasks and within a short period of time. An example are aptitude tests administered at the end of the school year and in related areas such as reading, writing and word comprehension. By considering the individual quasi-difference in scores we are able to eliminate unobserved ability and unobserved group effects. Our key identifying assumption postulates that different tests measure similar skills and that variation in differential scores is, apart from variation induced by observed covariates, due to cross-sectional variation in ability or effort of individuals that is not related to prior performance and other systematic and possibly unobserved factors correlated with group formation such as teacher quality or parental support. Our theory builds on the concept of latent performance, ability or effort measures defined in the absence of group interaction.
The first result of our paper shows that our key identifying restriction implies orthogonality conditions that are at the core of our identification strategy. These restrictions are obtained without any assumptions about random group selection or independence of individual characteristics in the population of individuals. Combined with the assumption of linear latent outcomes and mean group peer effects this leads to an empirical model that falls within the class of linear peer effects models that have been prominently studied in the literature, see for example Manski (1993); Calvó-Armengol, Patacchini, and Zenou (2009); Blume et al. (2015) and Angrist (2014) for a critique of these models. We expand on this literature by explicitly accounting for endogenous peer group selection and individual heterogeneity in unobserved test taking ability. Most of the empirical literature uses observable characteristics as well as proxies for unobserved ability to measure the quality of peer groups. Our baseline model is formulated for unobserved, at least to the analyst, peer characteristics. We use multiple performance measures for similar skills to difference out common test taking ability. We allow group selection to depend on unobserved as well as observed characteristics. Recent contributions to the econometrics literature accounting for endogenous group and network formation include Goldsmith-Pinkham and Imbens (2013), Hsieh and Lee (2016) and Griffith (2022) who use an explicit network formation model, Qu and Lee (2015), Johnsson and Moon (2019), Auerbach (2022) who use control function approaches and Kuersteiner and Prucha (2020) who use an instrumental variables approach. [^1] The literature on peer effects is part of a larger literature on network effects. It is well recognized that the adjacency matrix frequently used in modeling peer effect is a special case of the weight matrices used in a class of models introduced by Cliff and Ord (1973, 1981), which were originally intended for modeling spatial network effects; see Anselin (2010) for a review of important contribution of this literature on identification and estimation of network effects, and Kuersteiner and Prucha (2020) for a recent contribution that connects those strands of network literature. .
In our setting, reduced form regression analysis that focuses on exogenous or contextual peer effects leads to biased estimates. We propose instrumental variables estimators to overcome these problems. By exploiting moment restrictions that are implied by our identifying assumptions we are proposing generalized method of moments (GMM) estimators that identify both exogenous as well as endogenous peer effects. Formally, our moments based estimator is a panel type estimator based on quasi-differences between test performance measures. An important difference between our approach and typical panel settings is the lack of a temporal dimension, and where the setup is geared towards the utilization of different test performance measures that are obtained in an essentially simultaneous fashion.
We apply our approach to the Project STAR data set of the Tennessee class size experiment. We exploit variation on closely related test score outcomes for kindergarten to third grade students. Our empirical analysis focuses on the identification, estimation and statistical inference for the parameter determining the marginal effect of unobserved peer quality. More specifically, this parameter measures the impact of a unit increase in peer quality, measured in terms of latent peer outcomes, on individual outcomes. Obtaining data with convincing exogenous peer group variation is generally difficult and may require costly experimental designs. Yet, most of the empirical literature estimating peer group composition effects relies on conditionally randomly assigned peers. An advantage of the approach proposed in this paper is that we do not rely on random group assignment. Our framework enables us to identify the marginal parameter without variation in peer group composition. The random assignment assumption has been criticized for Project STAR data, for example because of attrition in higher grades, see Hanushek (2003). Our approach is designed to work despite these data limitations. In addition, and unlike in related studies identifying similar parameters such as Kelejian and Prucha (2002), Lee (2007b), Graham (2008) or Kuersteiner, Prucha, and Zeng (2023) we do not rely on group size variation or variation in group type variances.
The difficulties of identifying peer effects in models with transitive groups where all peers have the same effect, are well known since Manski (1993). Approaches to overcome identification challenges have focused on random group assignment, for example Sacerdote (2001); Duflo and Saez (2003); Carrell, Sacerdote, and West (2013); Sojourner (2013), exogenous variation in group size as in Lee (2007b) or Kuersteiner, Prucha, and Zeng (2023), instrumentation of peer effects using non-transitive group structures as in Bramoullé, Djebbari, and Fortin (2009), explicit group formation models in Goldsmith-Pinkham and Imbens (2013), variation in group specific heteroskedasticity in Graham (2008), and panel methods in Mas and Moretti (2009), Arcidiacono et al. (2012), Cornelissen, Dustmann, and Schönberg (2017), Miraldo, Propper, and Rose (2021) and Braun and Verdier (2023). The approach pursued in this paper builds on the dynamic panel methods for spatial and social networks developed in Kuersteiner and Prucha (2020).
We use a baseline model of peer effects that is similar to Arcidiacono et al. (2012) who consider peer effects working through unobserved individual characteristics. Arcidiacono et al. (2012) allow for individual specific effects as well as for endogenous group formation and use panel data. Technically, our estimator is different from theirs since we use a quasi-differencing approach to eliminate class and individual fixed effects while they use a non-linear fixed effects estimator. The more important difference lies in the assumptions and empirical implementation of the method we have in mind. While our estimator could be applied to more conventional panel data with performance measures observed at different points in time, we emphasize applications where measures of closely related skills are observed essentially simultaneously. For one, this framework alleviates problems with sample attrition which are well documented for Project STAR. More importantly, our core assumption on unobserved skill or effort centers around a lack of predictive power of additional test results for closely related test measures. For example, a student scoring well on a reading test is likely to score well on a word comprehension test given at the same time, and in this sense the word comprehension test does not reveal additional information about underlying skills or effort. An advantage of our framework is that its credibility can be assessed and influenced by a testing protocol that complies with its basic premise. This is in contrast to more detailed restrictions between various unobserved components as well as on the evolution of individual effects measuring the change in skills over time required in the framework of Arcidiacono et al. (2012). An additional advantage of our procedure is that it does not require estimating unobserved skill or effort measures of individuals.
There is an extensive empirical literature studying educational outcomes using Project STAR data. Peer effects specifically were considered by numerous authors. Boozer and Cacciola (2001) use within and between class variation as well as controls for whether students currently are or previously were in a small classroom to estimate the parameter for endogenous peer effects. Variation in the exposure of classmates to being previously in small classrooms provides variation in peer quality that is used to identify the endogenous peer effect. They find large peer effects in second and third grade and negative but insignificant effects in first grade. Whitmore (2005) studies the effects of variation in the ratio of girls on test scores and finds mixed results depending on grades. A decomposition of the effect into endogenous and exogenous peer effects results in an estimated increase in own test score of 0.6 points for every point increase in the average peer test score. Graham (2008) uses a model similar to ours and estimates endogenous peer effects based on differences of the between and within variances in small and regular size classes. He finds stronger peer effects of being randomly assigned to smaller as compared to larger classes. Kuersteiner, Prucha, and Zeng (2023) interpret the variance approach of Graham (2008) as being part of a class of more general random group effects models, see also Rose (2017) for an approach similar to Graham (2008) using variance restrictions. Chetty et al. (2011) link Project STAR data with 1996-2008 tax records to investigate classroom, teacher and peer effects on future earnings. Their empirical approach relies on random assignment to Kindergarten and elementary school grades which they test using additional individual level data obtained from tax records. Using analysis of variance and regression based methods exploiting within classroom variation they find significant effects of class quality on earnings. In work that looks at a related question Bietenbeck (2020) investigates the link between low ability repeaters in kindergarten classes on educational outcomes later in life and finds positive effects of being exposed to repeaters on the probability to graduate from high school and taking a college entrance exam. Pereda-Fernández (2017) relies on a conditional double randomization assumption resulting in conditional independence restrictions between student and teacher effects. These restrictions imply covariance as well as higher order restrictions that are exploited to estimate both the endogenous peer effect as well as the distribution of teacher effects. The paper finds sizable peer effects and investigates optimal teacher and class size allocation rules. Lewbel, Qu, and Tang (2023) also estimate a linear peer effects model for Project STAR data, but assuming that actual interaction between students is unobserved. They use restrictions on reduced form parameters to recover the endogenous peer effect. Empirically, they find large peer effects for second and third grade math scores.
In our empirical work we use Project STAR data and focus on classrooms in kindergarten to the third grade. We consider four individual SAT scores, mathematics, reading, listening and word study skills. We document the high correlation between pairs of these scores, unconditionally as well as conditionally on a full set of controls, with correlation coefficients ranging from 0.5 to 0.9. The highest correlation is observed for the reading and word study skill scores. The correlation patterns are robust across the four grade levels we consider. The peer effects coefficient of our model is estimated precisely and is stable across the different grades with the exception of first grade where we observe a lower degree of peer effects. We find stronger peer effects for test score pairs that are less highly correlated, in line with our model assumptions where more closely related scores offer better control for unobserved confounding factors. Quantitatively, we estimate that an increase of average unobserved peer quality measured in SAT equivalent scores of 10 SAT points leads to an increase of 2 to 4 SAT points for an individual class mate. This estimate is based on the reading-word score pair. For other score pairs the effects range from a 4 to a 7 SAT point increase for the scores of individual classmates due to a change in average scores. These measurements are quite robust to the inclusion or omission of additional controls. When we measure exogenous peer effects, in other words the effects of average age, race and gender, as well as an indicator of economic background, we find mostly insignificant results on differential scores. This lack of significance is further evidence that our differencing strategy eliminates much of the systematic factors explaining test score variation. However, it does not mean that exogenous controls have no effect on the level of individual scores.
The paper is organized as follows. In Section 2 we develop the identification and estimation strategy. Section 2.1 defines performance measures and group selection, Section 2.2 introduces the education production function and peer effects, Section 2.3 discusses the key identifying restriction and Section 2.4 introduces the estimators we propose. Section 3 contains the empirical analysis of Project STAR data and Section 4 develops formal identification and consistency proofs as well as an asymptotic distribution theory. Section 5 contains conclusions. Proofs and tables are contained in the appendix.
2 Model
We assume that for each student $i=1,...,n\text{ }$we observe results for two different tests, or more generally that we observe two measures of academic achievement, say, $y_{it}$ where $t=1,2$ indexes the test. Our empirical work uses Project STAR data where these measures are scores of the Stanford Achievement Tests (SAT) chosen from math, word study skills, listening and reading tests given to kindergarten through third grade students. Our approach uses two outcome measures to control for unobserved and possibly correlated student characteristics as well as group level effects related to classrooms, teachers and schools.
Let $\tilde{y}_{it}^{*}$ be the unobserved ability or effort of student $i$ taking test $t$. We use * to denote variables that are not observed by the analyst, but may be observed by the individuals, and we use a surmount $\tilde{}$  to denote variables in the original sample before group assignment. The notation $\tilde{y}_{it}^{*}$ emphasizes that the variable refers to the $i$-th student in the population of students prior to group allocation, and that the index $i$ is an individual student identifier which is not directly tied to classrooms, e.g., $i$ may correspond to an alphabetical ordering of all students by name. There are two steps that link the vector of latent outcomes $\left(\tilde{y}_{1t}^{*},...,\tilde{y}_{nt}^{*}\right)$ to a vector of observed performance $\left(y_{1t},...,y_{nt}\right)$. The first step consists of allocating student $i$ in the original sample to classroom $c$ where student $i$ now receives the in-class identifier $r.$ For example, $r$ could be the alphabetical rank of student $i$ in class $c.$ Let $n_{c}$ denote the size of classroom $c$ and suppose there are $C$ classrooms so that $n=n_{1}+...+n_{C}$. We then map the $r$-th student in classroom $c$ into a new index, say, $i^{\prime}=n_{1}+...+n_{c-1}+r$. In the following we will, abusing notation slightly, refer, e.g., to the unobserved ability of the $r$-th student in classroom $c\text{ }$ taking test $t$ either as $y_{crt}^{*}$ or $y_{i^{\prime}t}^{*}$, depending on the context. We will furthermore use $i$ rather than $i^{\prime}$ for the new index for convenience of notation. The second step is an education production function or a more general group interaction model $\psi$ that relates latent ability $y_{crt}^{*}$ of classmates to actual test performance $y_{crt}$. Actual performance depends on class level unobserved effects, own latent ability, observed individual, class and school level covariates, as well as potential peer interaction. In what follows, we use the term class level effects for all factors affecting outcomes that are invariant at the classroom level. Our goal is to isolate the portion of observed performance that is due to peer effects.
2.1 Latent Performance and Group Selection
Selection into classrooms can be correlated with a latent baseline performance measure $\tilde{\zeta}_{i}^{*}$ that can also be thought of as prior test taking ability or latent performance, or effort prior to taking the observed tests. The purpose of introducing baseline test taking ability is to account for unobserved student characteristics that are invariant for the two tests. We account for observed personal student characteristics where $\tilde{v}_{i}^{p}$ and $\tilde{w}_{it}^{p}$ denote row vectors of observed test invariant and test varying characteristics. For notational convenience we collect the performance measures and characteristics for all $n$ students in the sample in the following vectors and matrices of row dimension $n$: $\tilde{\zeta}^{\ast}=(\tilde{\zeta}_{1}^{\ast},...,\tilde{\zeta}_{n}^{\ast})^{\prime}$, $\tilde{y}_{t}^{*}=\left(\tilde{y}_{1t}^{*},\ldots,\tilde{y}_{nt}^{*}\right)^{\prime}$ for $t=1,2$, $\tilde{v}^{p}=\left(\tilde{v}_{1}^{p\prime},\ldots,\tilde{v}_{n}^{p\prime}\right)^{\prime}$, $\tilde{w}_{t}^{p}=\left(\tilde{w}_{1t}^{p\prime},\ldots,\tilde{w}_{nt}^{p\prime}\right)^{\prime}$ for $t=1,2$, and $\tilde{X}^{p}=\left(\tilde{v}^{p},\tilde{w}_{1}^{p},\tilde{w}_{2}^{p}\right)$. The $i$-th row of $\tilde{X}^{p}$ then contains the covariates of student $i$ in the original sample. Selection into classrooms can, in addition to depending on $\tilde{\zeta}^{*}$ and $\tilde{X^{p}}$, depend on observed class level characteristics as well as on a vector $\alpha=\left(\alpha_{1},...,\alpha_{C}\right)^{\prime}$ of unobserved class level characteristics. We do not impose any restrictions on the cross-sectional dependence in $\tilde{\zeta}^{\ast}$, $\tilde{X}^{p}$, $\alpha$ or observed class level characteristics. Latent baseline performance of student $i$ may be correlated with latent performance of student $j$ for a variety of reasons including similar educational, socioeconomic, cultural, religious or geographic backgrounds that may or may not be known to the analyst.
We now describe the selection of students to classrooms. This assignment process results in a reordering of the students characterized by a one-to-one mapping of the original index set to the new index set. Let $\tilde{y}_{t}^{*}=\left(\tilde{y}_{1t}^{*},...,\tilde{y}_{nt}^{*}\right)^{\prime}$ denote the vector of unobserved ability of students corresponding to the the original sample or population, let $S_{c}$ be the $n_{c}\times n\text{ }$ selection matrix that allocates students to classroom $c$, and let $S=\left[S_{1}^{\prime},...,S_{C}^{\prime}\right]^{\prime}$ be the selector matrix for all students. The set of all possible classroom allocations is denoted by $\mathscr{S}$. A random assignment mechanism can be thought of as selecting one element of $\mathscr{S}$ at random. Corresponding to $S$ the vector $y_{t}^{*}=\left(y_{1t}^{*},\ldots,y_{nt}^{*}\right)^{\prime}$ representing the sample or population of unobserved ability for test $t$ of students ordered by classrooms is given by
$y_{t}^{*}=S\tilde{y}_{t}^{*}.$
As remarked above, in slight but obvious abuse of notation, let $i=n_{1}+...+n_{c-1}+r$. Then we denote the $i$-th student’s unobserved ability for test $t$ interchangeably by either $y_{it}^{*}$ or $y_{crt}^{*}$. The latter indexing convention, which we also employ analogously for other variables, is convenient when an analysis at the classroom level is required. Now let $y_{ct}^{*}=\left(y_{c1t}^{*},...,y_{cn_{c}t}^{*}\right)^{\prime},$ then
$y_{ct}^{*}=S_{c}\tilde{y}_{t}^{*}.$
Similarly, we arrange unobserved and observed student characteristics by classroom through the transformation $\zeta^{*}=\left(\zeta_{1}^{*},\ldots,\zeta_{n}^{*}\right)^{\prime}=S\tilde{\zeta}^{\ast}$, $v^{p}=\left(v_{1}^{p\prime},\ldots,v_{n}^{p\prime}\right)^{\prime}=S\tilde{v}^{p}$, $w_{t}^{p}=\left(w_{1t}^{p\prime},\ldots,w_{nt}^{p\prime}\right)^{\prime}=S\tilde{w}_{t}^{p}$ for $t=1,2$, and $X^{p}=\left(x_{1}^{p\prime},\ldots,x_{n}^{p\prime}\right)^{\prime}=S\tilde{X}^{p}=\left(v^{p},w_{1}^{p},w_{2}^{p}\right)$. Thus, the row vector $x_{i}^{p}$ now denotes observable personal characteristics for student $i$ in the sample ordered by classrooms.
We allow for $S$, $\tilde{\zeta}^{\ast}$, $\tilde{y}_{t}^{*}$, for $t=1,2$, and $\alpha$ to be mutually correlated and to be correlated with $\tilde{X}^{p}$ as well as with observed class level characteristics. This assumption includes scenarios where $S$ is selected completely at random, scenarios where $S$ is selected based on student and class level characteristics, as well as scenarios where students respond to the allocation $S$. Our assumptions, formally spelled out in Section 2.3, require that selection happens before testing which is a mild restriction that should be satisfied in most scenarios we have in mind. We do not otherwise specify or restrict the mechanism $S$ that selects students into classrooms. We also do not require that $S$ is observed. This is relevant in situations where we do not have information identifying individual students such as through their names, social security numbers or residential address. On the other hand we do require that we know who, among all test takers, is allocated to the same classroom. For project STAR data, this information is available for the majority of students in grades K through three through a class identifier [^2] The identifier is a teacher ID that is unique for each classroom., but not for students in higher grades.
2.2 Peer Effects
We relate educational outputs to observed and unobserved inputs via an educational production function denoted by $\psi$. Educational production functions were considered by Krueger (1999) and Todd and Wolpin (2003) among others. Our specification of the educational production function explicitly allows for peer effects and models observed performance $y_{crt}$ of student $r$ in classroom $c$ as a function of latent ability or effort $y_{c1t}^{*},\ldots,y_{cn_{c}t}^{*}$ of all students in class $c$. The assumptions we impose on $y_{crt}^{*}$ allow it to result from a class assignment process that depends on both observed and unobserved student and class level characteristics.
We interpret $y_{crt}^{*}$ as the unobserved test taking ability in the absence of peer effects, but accounting for possible selection of student $r$ into classroom $c,$ taking test $t.$ The function $\psi_{crt}\left(.\right)$ accounts for the fact that test performance varies by student and thus depends on $r$ as well as additional student, teacher, classroom and school characteristics not captured by $y_{crt}^{*}$. Observed performance is determined by $y_{crt}=\psi_{crt}\left(y_{c1t}^{*},\ldots,y_{cn_{c}t}^{*}\right)$. The formulation of $\psi_{crt}\left(.\right)$ explicitly accounts for the possibility of peer effects by allowing for individual performance to depend on the characteristics of all peers. In the absence of peer effects the function $\psi_{crt}\left(.\right)$ simplifies to $y_{crt}=\psi_{crt}\left(y_{crt}^{*}\right).$ In other words, the performance of the $r$-th student in class $c$ only depends on own characteristics and class characteristics such as teachers and resources but not on the characteristics of other students in the class.
In empirical applications it may be difficult to work with a non-parametric framework accounting for peer effects. For this reason we focus on linear production functions with linear peer effects. Let $v^{c}$ and $w_{t}^{c},t=1,2$, be matrices with row dimension $n$ that capture observed test invariant and test varying class level characteristics not related to student characteristics, such as the class size, the gender, education and experience of the teacher for subject $t$, and let $X=\left(v^{p},w_{1}^{p},w_{2}^{p},v^{c},w_{1}^{c},w_{2}^{c}\right)$ denote the matrix of observations on all covariates. At the class level let $y_{ct}=\left(y_{c1t},\ldots,y_{cn_{c}t}\right)^{\prime}$, $y_{ct}^{*}=\left(y_{c1t}^{*},\ldots,y_{cn_{c}t}^{*}\right)^{\prime}$, $v_{c}^{p}=\left(v_{c1}^{p\prime},...,v_{cn_{c}}^{p\prime}\right)^{\prime}$, $w_{ct}^{p}=\left(w_{c1t}^{p\prime},...,w_{cn_{c}t}^{p\prime}\right)^{\prime}$ with similar definitions for $v_{c}^{c}$ and $w_{ct}^{c}$, and let $\psi_{ct}=\left(\psi_{c1t},\ldots,\psi_{cn_{c}t}\right)^{\prime}$. We impose the following restriction on the education production function.
Assumption 1.
Let $X_{ct}^{c}=\left(v_{c}^{c},w_{ct}^{c}\right)$, $\beta_{t}^{c}=\left(\beta_{vt}^{c\prime},\beta_{wt}^{c\prime}\right)^{\prime}$, $X_{ct}^{p}=\left(v_{c}^{p},w_{ct}^{p}\right)$, $\beta_{t}^{p}=\left(\beta_{vt}^{p\prime},\beta_{wt}^{p\prime}\right)^{\prime}$. Assume that
$y_{ct}=\psi_{ct}\left(y_{ct}^{*}\right)=\alpha_{c}f_{t}\mathbf{1}_{c}+X_{ct}^{c}\beta_{t}^{c}+\left(I_{c}+\rho M_{c}\right)\left(X_{ct}^{p}\beta_{t}^{p}+y_{ct}^{*}\right),$ (1)
where $\mathbf{1}_{c}=\left(1,....,1\right)^{\prime}$ is an $n_{c}\times 1$ vector and $M_{c}$ is some matrix of dimension $n_{c}\times n_{c}$ with zero diagonal elements and $\rho$ is a fixed parameter. The parameter $f_{t}$ captures test specific effects.
The part of the expression involving $M_{c}$ in (1) models peer effects. The vector $M_{c}y_{ct}^{*}$ contains weighted averages of the latent performance of classroom peers, and the parameter $\rho$ specifies the degree to which peer effects influence actual outcomes.  [^3] In the usual terminology of Cliff-Ord (1973,1981) models, $M_{c}y_{ct}^{*}$ is called a spatial lag and $\rho$ the corresponding parameter. The vector $M_{c}X_{ct}^{p}\beta_{t}^{p}$ represents contextual peer effects.
Linear specifications such as the one in Assumption 1 have been motivated as solutions to Nash games in the literature, see Calvó-Armengol, Patacchini, and Zenou (2009), Blume et al. (2015) or Pereda-Fernández (2017). Pereda-Fernández (2017) considers a full information game of student and teacher effort where student and teacher quality is known to the players and that leads to an equation like (1). Thus, Assumption 1 can be justified by assuming that student ability $y_{ct}^{*}$ is unobserved by the analyst but known to classmates. Calvó-Armengol, Patacchini, and Zenou (2009) consider a network game in effort under full information where agents have linear quadratic utility. Therefore the educational production function (1) can be understood as an optimal response function where $y_{ct}^{*}$ now has the interpretation of effort, unobserved by the analyst but known to peers. Finally, Blume et al. (2015) analyze a Bayes-Nash equilibrium of agents choosing effort and facing a linear quadratic utility function that depends on their own as well as their peer’s effort; cp. also Cohen-Cole, Liu, and Zenou (2018) and Drukker, Egger, and Prucha (2023). Agents maximize expected utility conditional on publicly observable characteristics as well as private information that is only known to the individual player but not the other agents. They argue, see Blume et al. (2015) p. 452, that their model is observationally equivalent to Calvó-Armengol, Patacchini, and Zenou (2009) as long as the row sums of $M_{c}$ are all equal, an assumption we impose below. In particular, as noted by Blume et al. (2015) p. 458, $y_{ct}^{*}$ then can be interpreted as ability, with agents expectations absorbed into a fixed effect. This can be matched to our setting by assuming that $y_{ct}^{*}$ is ability that is unobserved by the analyst.
In line with the spatial and social network literature building on the models introduced by Cliff-Ord (1973,1981), the matrix $M_{c}$ in Assumption 1 could be an arbitrary weight matrix describing peer interaction subject to certain measurability assumptions specified below. For exposition and concreteness we focus on the case where average characteristics of peers determine peer effects. Define
$M_{c}=\left(\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}-I_{c}\right)/\left(n_{c}-1\right)$
where $I_{c}$ is the $n_{c}\times n_{c}$ identity matrix. Then, $M_{c}$ is the operator that computes the leave out average of peer characteristics. Combined with the functional form in Assumption 1 this leads, using scalar notation, to an education production function $y_{crt}=\psi_{crt}\left(y_{ct}^{*}\right)$ for observed outcomes in terms of unobservables and observables as given below ($t=1,2$)
$y_{crt}=\alpha_{c}f_{t}+x_{ct}^{c}\beta_{t}^{c}+x_{crt}^{p}\beta_{t}^{p}+y_{crt}^{*}+\rho\left[(n_{c}-1)^{-1}\sum_{l=1,l\neq r}^{n_{c}}\left(x_{clt}^{p}\beta_{t}^{p}+y_{clt}^{*}\right)\right].$ (2)
2.3 Identification
In this paper we propose an identification strategy for the parameter $\rho$ that does not require randomization over $\mathscr{S}$ or variation in $S_{c}.$ Our argument proceeds in two steps. First, we propose a moment restriction on unobserved test taking ability $y_{it}^{*}$ and show that this restriction implies linear and quadratic moment restrictions for quantities we are able to observe in the data. We then show that $\rho$ can be identified by considering a quasi-differencing transformation of (2) and a GMM estimator that exploits the restrictions on the conditional mean and variance implied by our identifying assumption. The asymptotic theory for a general class of GMM estimators of this type was developed by Kuersteiner and Prucha (2020). Here, we rely on their theory to derive sharp identification results and inference procedures that are specific to this application.
We formalize the information structure of our model in the following definition. Recall that $\zeta^{\ast}$ stands for latent test performance, effort or ability prior to taking tests $t=1,2.$ We use the notation $-t$ in the definition below to denote the other test. We define the following:
Definition 2.1.
Let $y_{-it}^{*}=\left(y_{1t}^{*},...,y_{i-1,t}^{*},y_{i+1,t}^{*},...,y_{n,t}^{*}\right)^{\prime}$ and $y_{t}^{*}=\left(y_{1t}^{*},...,y_{nt}^{*}\right)^{\prime}$. Define the sigma fields (information sets)
$\mathcal{F}_{n,i,t}$ $=\sigma\left(S,\alpha,X,z,\zeta^{\ast},y_{-t}^{\ast},y_{-i,t}^{*}\right)\text{\,for $t=1,2,$}$
where $z$ denotes variables excluded from $X$ that may be used as instruments. Let $\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,X,z,\zeta^{\ast}\right)$ be the sigma field of all conditioning variables and let $\mathcal{Z}_{n}=\sigma\left(X,z\right)\subset\mathcal{Z}_{n}^{*}$ be the subset of observable information.
The information set $\mathcal{Z}_{n}^{*}$ consists both of observable characteristics $X,$ classroom allocations $S,$ excluded variables $z$ that may be used as instruments as well as of unobserved variables $\zeta^{\ast}$ and $\alpha.$ The following proportionality restriction on the conditional means of the latent outcomes is at the core of our proposed method of identifying peer effects.
Assumption 2.
Assume that
$E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=\kappa_{i}f_{t}$
where $\kappa_{i}$ is a random variable that is invariant to $t$ and is measurable with respect to $\mathcal{Z}_{n}^{*}$ and $f_{t}$ is a fixed parameter that only varies with $t$. In addition, assume that the random variable $y_{it}^{*}$ is either bounded, or there exists an $\eta>0$ and a random variable $y$ such that $\left|y_{it}^{*}\right|+\left|\kappa_{i}f_{t}\right|\leq y$ with $E\left[\left|y\right|^{4+\eta}|\mathcal{Z}_{n}^{*}\right]\leq K_{y}<\infty$ for all $i$ and $t.$ In addition, $E\left[\left|\alpha_{c}\right|^{2+\eta}\right]\leq K_{\alpha}<\infty$ for all $c.$
Without loss of generality we normalize $f_{2}=1$. The assumption includes the special case where $f_{t}=1$ for $t=1,2$, which is relevant in situations where tests $t$ measure closely related skills. The interpretation of the condition is that conditional on a hypothetical or actual baseline of test results $\zeta^{*}$, as well as information about group formation $S$, unobserved class level characteristics $\alpha$, as well as other observed characteristics $X$ and $z$, additional test results do not change expected performance, except maybe for a common scale factor $f_{t}$ accounting for systematic differences between tests affecting all test takers in the same fashion. [^4] For simplicity we assume that $f_{t}$ in Assumptions 1 and 2 are the same. This restriction can be relaxed if three test outcomes are used. The factors $f_{t}$ can be treated as random rather than fixed at the expense of slightly more complex asymptotic arguments needed for the analysis of statistical inference, see Kuersteiner and Prucha (2020). It is in this sense that tests $t=1$ and $t=2$ measure essentially identical skills. Note that $E\left[y_{it}^{*}|\mathcal{Z}_{n}^{*}\right]=\kappa_{i}f_{t}$ by iterated expectations and $\mathcal{Z}_{n}^{*}\subset\mathcal{F}_{n,i,t}.$
Relative to a baseline represented by $\mathcal{Z}_{n}^{*}$ the latent test results $y_{it}^{*}$ are in essence as good as randomly assigned. Conditional on observed covariates as well as the information in $\mathcal{Z}_{n}^{*}$, actual outcomes $y_{it}$ which are determined by the educational production function in (1), then are only correlated in the cross-section because of peer effects, i.e. when $\rho\neq 0.$
The variable $\kappa_{i}$ is a function of the entire cross section of latent performance measures $\zeta^{*}$, of the allocation $S$ and observed and unobserved characteristics. It is generally cross-sectionally dependent in ways we do not restrict or specify and captures such common effects as the socioeconomic background of students and classmates, their family background including parental education and support, their interaction with peers, the exposure to their teachers and resources available in their classroom. It also depends on a student’s own ability. The restriction we impose is that these factors do not change between tests $t=1$ and $t=2$, except maybe for a change in scale $f_{t}.$
Baseline performance $\zeta^{*}$ does not necessarily have to be observed or realized and $\zeta^{*}$ can alternatively be interpreted as latent ability prior to taking the actual tests $t=1,2$. In this sense, $\kappa_{i}f_{t}$ and $E\left[y_{it}^{*}|\mathcal{Z}_{n}^{*}\right]$ are also unobserved. A scenario where Assumption 2 is realistic arises when the same type of test is taken multiple times without additional training between iterations of the test, or when several tests are given that focus on related skills, as is the case in Project STAR data for grades K-3 and tests for reading, word study and listening skills. [^5] Another example of a test with multiple scores is the Comprehensive Testing Program (CTP) administered by the Educational Records Bureau and given to Grades 1-11. The CTP test consists of several main categories such as “Auditory Comprehension”, “Reading Comprehension” and “Mathematics” with each category consisting of additional subcategories that each receive separate scores. The assumption is less plausible if comparisons are attempted across different subjects in a high school or college setting, for tests given at different times during the school year, or measuring different skills in a professional or team setting. In the latter case, additional factors can be introduced if multiple measurements per skill category are available. [^6] The multi-factor case is discussed in detail in Kuersteiner and Prucha (2020).
The following theorem formally establishes a representation for $y_{it}^{*}$ that decomposes $y_{it}^{*}$ into its conditional mean, $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]$, that is dependent cross-sectionally as well as across tests, and into uncorrelated idiosyncratic noise $u_{it}$. The theorem formalizes the definition of $u_{it}$
Theorem 2.1.
Suppose Assumption 2 holds. Define
$u_{it}=y_{it}^{*}-E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=y_{it}^{*}-\kappa_{i}f_{t},$
then by construction
$E\left[u_{it}|\mathcal{F}_{n,i,t}\right]=0\ \ \text{and}\ \ \operatorname{Cov}\left(u_{it}u_{js}|\mathcal{F}_{n,i,t}\right)=0$
for any $j\neq i$ or $t\neq s$. Furthermore, let $\mathcal{G}_{n,i,t}$ be any sigma field with $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$, then $E\left[u_{it}|\mathcal{G}_{n,i,t}\right]=0$ and $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{G}_{n,i,t}\right)=0$ for any $j\neq i$ or $t\neq s$, and $\operatorname{Cov}\left(u_{it},\kappa_{i}f_{t}\right)=0,\operatorname{Cov}\left(u_{it},X\right)=0$, $\operatorname{Cov}\left(u_{it},z\right)=0$ etc. If $E\left[y_{it}^{*}-\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]\neq 0$ for some $\mathcal{G}_{n,i,t}$, then Assumption 2 cannot hold.
Examples for the information sets $\mathcal{G}_{n,i,t}$ in the above theorem are $\mathcal{G}_{n,i,t}=\sigma\left(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*s}\right)$ where $y_{-i,t}^{*s}$ is a subvector of $y_{-i,t}^{*}$ or $\mathcal{G}_{n,i,t}=\mathcal{Z}_{n}^{*}.$ The above theorem makes clear that the main restriction we impose on $y_{it}^{*}$ is the assumption that latent test performance is not predictable conditional on observable characteristics and unobserved baseline ability. The properties $E\left[u_{it}|\mathcal{Z}_{n}^{*}\right]=0,$ $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{Z}_{n}^{*}\right)=0$ for any $j\neq i$ or $t\neq s$ and $\operatorname{Cov}\left(u_{it},\kappa_{i}f_{t}\right)=0$, $\operatorname{Cov}\left(u_{it},X\right)=0$, $\operatorname{Cov}\left(u_{it},z\right)=0$ follow directly from Assumption 2 and are not additional assumptions imposed on the distribution of $y_{it}^{*}.$ In other words, Assumption 2 is necessary and sufficient for the decomposition of $y_{it}^{*}=\kappa_{i}f_{t}+u_{it}$ with uncorrelated errors $u_{it}$.
We do not attach a specific economic or behavioral interpretation to the decomposition of $y_{it}^{*}$ into $\kappa_{i}f_{t}$ and $u_{it}$. It is merely a statistical representation for the purpose of isolating variation in test scores that is correlated within groups but invariant across tests and variation that is idiosyncratic to the individual student and test. Normalizing $f_{2}=1$, which is without loss of generality, leads to an immediate consequence of the decomposition: The quasi-difference
$y_{i1}^{*}-y_{i2}^{*}f_{1}=u_{i1}-u_{i2}f_{1}$ (3)
is mean zero and uncorrelated across $i$ conditional on $\mathcal{Z}_{n}^{*}$. We exploit these two moment restrictions to formulate our GMM estimator. They are at the heart of our identification result which is formally presented in Section 4.
Our framework is in contrast with much of the econometrics literature on the identification of peer effects where assumptions about the orthogonality of $u_{it}$ and $\kappa_{i}f_{t}$ and covariates are usually imposed directly. These unobservable quantities often are given economic interpretations. Examples of papers with such assumptions include Lee (2007b) who assumes i.i.d. errors, Graham (2008) who assumes independent classroom and idiosyncratic errors conditional on class type, Arcidiacono et al. (2012) who assume uncorrelated idiosyncratic errors and fixed effects that are orthogonal to idiosyncratic errors, Rose (2017) who assumes uncorrelated idiosyncratic errors, Kuersteiner, Prucha, and Zeng (2023) who assume a random effects specification and Lewbel, Qu, and Tang (2023) who assume exogeneity of group formation. Often these assumptions are justified by random group selection. In our setting we neither assume that individuals are randomly assigned to groups, nor that they are randomly selected from a population or that they have otherwise observed or unobserved characteristics that are independent in the cross-section.
Let $\kappa=\left(\kappa_{1},\ldots,\kappa_{n}\right)^{\prime}$ and $\tilde{\kappa}=\left(\tilde{\kappa}_{1},\ldots,\tilde{\kappa}_{n}\right)^{\prime}=S^{-1}\kappa$, then it is of interest to note that our theory allows for the variables $\tilde{\kappa}_{i}$ to be arbitrarily influential on the process that forms groups. An example of such correlation potentially arises in the Project Star data. Despite random allocation of students and teachers to classrooms in kindergarten it is possible that parents who are unhappy with their allocation try and succeed to move their child to a different class or school, or that those parents provide additional training for their child. In such a scenario individual and classroom specific effects $\tilde{\kappa}_{i}$ are not independent of classroom allocations.
2.4 Estimation
Our estimators are using differential, or quasi-differential test scores, $y_{i1}-y_{i2}f_{1}$. These differential measures purge performance measures from common unobserved influences captured by $\alpha_{c}$ and $\kappa_{c}=\left(\kappa_{c1},\ldots.\kappa_{cn_{c}}\right)^{\prime}=S_{c}\tilde{\kappa}$ and shed light on the existence of peer effects. Define the unobserved class and individual effect as
$\mu_{c}^{\ast}=\alpha_{c}\mathbf{1}_{c}+\left(I_{c}+\rho M_{c}\right)\kappa_{c}$ (4)
such that Assumptions 1, 2 and Theorem 2.1 imply the empirical specification
$y_{ct}$ $=\mu_{c}^{*}f_{t}$ $+v_{c}^{c}\beta_{vt}^{c}+w_{ct}^{c}\beta_{wt}^{c}+\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{vt}^{p}+w_{ct}^{p}\beta_{wt}^{p}+u_{ct}).$ (5)
By quasi-differencing (5), which eliminates classroom and individual fixed effects, we obtain
$y_{c1}-y_{c2}f_{1}$ $=$ $v_{c}^{c}(\beta_{v1}^{c}-f_{1}\beta_{v2}^{c})+w_{c1}^{c}\beta_{w1}^{c}-f_{1}w_{c2}^{c}\beta_{w2}^{c}$
$+$ $\left(I_{c}+\rho M_{c}\right)\left(v_{c}^{p}(\beta_{v1}^{p}-f_{1}\beta_{v2}^{p})+w_{c1}^{p}\beta_{w1}^{p}-f_{1}w_{c2}^{p}\beta_{w2}^{p}\right)+\left(I_{c}+\rho M_{c}\right)(u_{c1}-f_{1}u_{c2}).$
We treat $f_{t}$ as an unknown parameter to be estimated that accounts for differences in average test scores between tests $1$ and $2$. Collecting terms with $f_{1}$ leads to an interpretation of (2.4) where $y_{c2}-v_{c}^{c}\beta_{v2}^{c}-w_{c2}^{c}\beta_{w2}^{c}-\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{v2}^{p}+w_{c2}^{p}\beta_{w2}^{p})$ is used to control for unobserved $\mu_{c}^{\ast}$ similar to a control function approach. The difference to a conventional control function approach is that (5) for $t=2$ cannot be consistently estimated by least squares because unobserved components may be correlated with observed covariates, a problem that our GMM estimators address.
We next write (2.4) more compactly by stacking observations across classrooms. Let $y_{t}$ , $v^{p}$ and $w_{t}^{p}$ be as defined earlier. In similar fashion define $v^{c}$, $w_{t}^{c}$ and $u_{t}$, and let $M=\operatorname{diag}_{c=1}^{C}\left(M_{c}\right)$. Let $\underline{X}=\left(v^{c},w_{1}^{c},w_{2}^{c},v^{p},w_{1}^{p},w_{2}^{p},Mv^{p},Mw_{1}^{p},Mw_{2}^{p}\right)$ and define
$\delta(f_{1},\rho,\beta)=\left(\delta_{v}^{c\prime},\beta_{w1}^{c\prime},-f_{1}\beta_{w2}^{c\prime},\delta_{v}^{p\prime},\beta_{w1}^{p\prime},-f_{1}\beta_{w2}^{p\prime},\rho\delta_{v}^{p\prime},\rho\beta_{w1}^{p\prime},-\rho f_{1}\beta_{w2}^{p\prime}\right)^{\prime},$ (7)
where $\delta_{v}^{c}=\beta_{v1}^{c}-f_{1}\beta_{v2}^{c}$, $\delta_{v}^{p}=\beta_{v1}^{p}-f_{1}\beta_{v2}^{p}$ and $\beta$ is the vector of all $\beta$s above, then we can write (2.4) for the entire sample more compactly as
$y_{1}=f_{1}y_{2}+\underline{X}\delta+(I+\rho M)(u_{1}-f_{1}u_{2})$ (8)
where $\delta=\beta_{1}-f_{1}\beta_{2}$ with $\beta_{t}$ collecting the corresponding coefficients for each variable in $\underline{X}$ and set to 0 if a variable is absent in (5) for given $t$. Similarly we can define $\mu^{\ast}=(\mu_{1}^{\ast\prime},....,\mu_{C}^{\ast\prime})^{\prime}$ and write (5) compactly as
$y_{t}=\mu^{\ast}f_{t}+\text{$\underline{X}$$\beta_{t}+(I+\rho M)u_{t}$},$ (9)
With Assumptions 1 and 2, and standard regularity conditions on $\underline{X}$, we can identify $\delta$, which is sufficient for the identification of $(\rho,f_{1})$. The parameters $\beta_{w1}^{c}$ and $\beta_{w1}^{p}$ can be directly recovered from $\delta$, and $\beta_{w2}^{c}$ and $\beta_{w2}^{p}$ can be identified when $f_{1}\neq 0$. As is common in pure fixed effect panel settings, the effects of test invariant covariates $v$, in our case $\beta_{v1}^{c},\beta_{v2}^{c}$ and $\beta_{v1}^{p},\beta_{v2}^{p}$ are not identified when using within or differencing estimators. In our case, except when $f_{1}=1,$ we are quasi-differencing the equation and allowing for heterogeneity in parameters between $t=1$ and $t=2.$ This leads to a formulation of the model where $v^{c}$ and $v^{p}$ enter the quasi, or fully differenced equation. However, without further assumptions the parameters $\beta_{v1}^{c}$, $\beta_{v2}^{c}$, $\beta_{v1}^{p}$ and $\beta_{v2}^{p}$ are not separately identified. An example of restrictions where the parameters $\beta_{v1}^{c}$ and $\beta_{v2}^{c}$ are identified arises when $f_{1}\neq 1$ and $\beta_{v1}^{c}=\beta_{v2}^{c}$.
In empirical applications it is sometimes sufficient to estimate $\beta$ or even just $\delta,$ for example when $\underline{X}$ contains contextual peer effects. Conventional regression methods applied directly to (5) are invalid because unobserved fixed effects collected in $\mu_{c}^{\ast}$ may be correlated with $\underline{X}.$ We show in Section 4 that instrumenting for $y_{2}$ in (8) identifies the parameters $\delta$ and $f_{1}.$ We show in Corollary 4.1 that when test scores are recorded as non-negative values the classroom fixed effect or an overall constant is a valid instrument. More generally, exogenous covariates that vary at the classroom level but are excluded from $\underline{X}$ are valid instruments by Lemma 4.1.
Valid standard errors are obtained by clustering by classrooms. For estimation of the full set of parameters including the peer effect parameter $\rho$, we use the efficient GMM estimator developed in Kuersteiner and Prucha (2020) for a fairly general class of networks covering both social and spatial network effects. Identification of the parameters $\rho,f_{1}$ and $\delta$ follows from Assumptions 1 and 2 as well as classroom level homoskedasticity and some restrictions on the parameter space that are detailed in Section 4.
Feasible efficient inference requires some additional restrictions on conditional variances that we now describe. We allow for some forms of heteroskedasticity for $u_{it}$ across tests and by class type. Let $\tau_{c}$ be a categorical variable for class type with $\tau_{c}=j$ if classroom $c$ is of type $j\in\left\{1,...,J\right\}.$ Examples of types are small and regular classes, or classes located in urban, suburban or rural areas.
Assumption 3.
Let $u_{it}$ be as defined in Theorem 2.1 and let $i=n_{1}+\ldots+n_{c-1}+r$ with $1\leq r\leq n_{c}$, i.e., $i$ refers to student $r$ in class c. Let $u_{-i,t}=\left(u_{1,t},...,u_{i-1,t},u_{i+1,t},...,u_{n,t}\right)$. Define the filtration
$\mathcal{B}_{n,i,t}=\sigma\left(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t}\right)\text{\,for $t=1,2$}$
where $u_{-t}$ represents information about the other test. Then the conditional variance of $u_{it}$ is given by
$E\left[u_{it}^{2}|\mathcal{B}_{n,i,t}\right]=\sigma_{t}^{2}\rho_{\tau_{c}}^{2},$
where $\tau_{c}$ is $\mathcal{Z}_{n}$-measurable, and $c_{u}\leq\sigma_{t}^{2},\rho_{j}^{2}\leq C_{u}$ for some finite positive bounds.
Assumption 3 implies that for $t=1,2$, $\Sigma_{t}\equiv E\left[u_{t}u_{t}^{\prime}|\mathcal{Z}_{n}^{\ast}\right]=\sigma_{t}^{2}\operatorname{diag}_{c=1}^{C}\left(\rho_{\tau_{c}}^{2}I_{c}\right)$. Observing that $f_{t}$ is a fixed constant, conditional on $S,\zeta^{*},\alpha,X$ and $z$, knowledge of $y_{it}^{*}$ is equivalent with knowledge of $u_{it}$. Thus we have the following equivalence of information sets [^7] See the proof of Lemma B.3 for more detail.
=$\mathcal{F}{}_{n,i,t}=\mathcal{B}_{n,i,t}\,\text{for}\,t=1,2.$
Assumption 3 imposes a similar factor structure on conditional variances as does Assumption 2 for conditional means.
Below we illustrate the nature of our efficient quasi-differencing transformation. By Theorem 2.1 the errors $u_{it}$ are mutually uncorrelated over tests as well as cross-sectionally. Consequently the variance covariance matrix of $u_{1}-f_{1}u_{2}$ is given by,
$\Omega=\operatorname{Var}\left(u_{1}-f_{1}u_{2}|\mathcal{Z}_{n}^{*}\right)=\Sigma_{1}+f_{1}^{2}\Sigma_{2}=\left(\sigma_{1}^{2}+f_{1}^{2}\sigma_{2}^{2}\right)\operatorname{diag}_{c=1}^{C}\left(\rho_{\tau_{c}}^{2}I_{c}\right).$
Define $\gamma_{j}^{2}=\left(\sigma_{1}^{2}+f_{1}^{2}\sigma_{2}^{2}\right)\rho_{j}^{2}$, for $j=1,...,J$ and $\gamma=\left(\gamma_{1},...,\gamma_{J}\right)^{\prime}$. Then $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right).$
The disturbance term in (8) is of the form $(I+\rho M)(u_{1}-f_{1}u_{2})$. Denote the parameter $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$. Premultiplying (8) by $\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$ then yields the following model
$y^{+}(\theta,\gamma)=X^{+}(\theta,\gamma)\delta+u^{+}$ (10)
where,
$y^{+}(\theta,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}(y_{1}-f_{1}y_{2}),$ (11)
$X^{+}(\theta,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}\underline{X},$ (12)
$u^{+}=\Omega(\gamma)^{-1/2}(u_{1}-f_{1}u_{2})$ (13)
and $\operatorname{Var}\left(u^{+}|\mathcal{Z}_{n}^{*}\right)=I_{n}$. In light of (10)-(13) we define
$u^{+}(\theta,\gamma)$ $=y^{+}(\theta,\gamma)-X^{+}(\theta,\gamma)\delta$
$=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}(y_{1}-f_{1}y_{2}-\text{$\underline{X}$$\delta$})$ (14)
and observe that $u^{+}(\theta,\gamma)=u^{+}$ at the true parameters.
The transformation leading to $u^{+}$ consists of three components. Quasi-differencing eliminates classroom and individual effects. The operator $(I+\rho M)^{-1}$ removes cross-sectional correlation and can be understood as a form of spatial Cochrane-Orcutt transformation. Finally, the operator $\Omega(\gamma)^{-1/2}$ scales the spatially uncorrelated residuals to unit variance in the case of heteroskedasticity.
We next provide an outline of our GMM estimation methodology. For clarity we explicitly denote in the following the true parameter vectors as $\theta_{0}=(\rho_{0},f_{1,0},\delta_{0}^{\prime})^{\prime}$ and $\gamma_{0}=(\gamma_{1,0},...,\gamma_{J,0})^{\prime}$. Rigorous consistency and asymptotic normality results as well as additional assumptions needed to establish these properties will be given in Section 4 below. Our GMM estimator is based on both linear and quadratic moment conditions. We consider instrument matrices $H$ and $A$ and require that all elements of $H$ and $A$ are functions of the observable variables that are measurable with respect to $\mathcal{Z}_{n}$. Note that by construction, these variables may depend on the unobserved effects $\alpha$ and $\kappa$. In particular, we set $H=[\underline{X},z]$ as an $n\times q$ instrument matrix.
The moment function consists of a set of linear and quadratic moments for the transformed residuals [^8] The specification can be readily extended to include several quadratic moment conditions along the lines of, e.g., Kuersteiner and Prucha (2020).
$m_{n}(\theta,\gamma)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+\prime}(\theta,\gamma)Au^{+}(\theta,\gamma)\end{array}\right).$
Under the maintained assumptions of this paper the results in Kuersteiner and Prucha (2020), Theorem 1, imply that the linear moment function is uncorrelated with the quadratic moments. We have
$E\left[m_{n}(\theta_{0},\gamma_{0})m_{n}(\theta_{0},\gamma_{0})^{\prime}\right]$ $=\frac{1}{n}\left(\begin{array}[]{cc}E\left[H^{\prime}H\right]&0\\ 0&2E\left[\operatorname{tr}\left(A^{2}\right)\right]\end{array}\right)\equiv\Xi_{n}.$ (17)
[…]
$Q_{n}\left(\theta,\gamma\right)$ $=\frac{1}{n}m_{n}(\theta,\gamma)^{\prime}\hat{\Xi}_{n}^{-1}m_{n}(\theta,\gamma)$
$=\frac{1}{n}u^{+}(\theta,\gamma)^{\prime}H\left(H^{\prime}H\right)^{-1}H^{\prime}u^{+}(\theta,\gamma)+\frac{1}{n}\frac{\left(u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\right)^{2}}{2\operatorname{tr}\left(A^{2}\right)}$ (18)
such that the GMM estimator for $\theta_{0}$ is defined as $\hat{\theta}\left(\gamma\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\gamma\right)$ for a fixed value of $\gamma.$ In applications $\gamma$ is replaced with a first step consistent estimator. We discuss consistent estimation of $\gamma$ and feasible versions of $\hat{\theta}\left(\gamma\right)$ in Section 4.
A simple two-step procedure to obtain initial estimates for $\delta$, $f_{1}$, and $\rho$ may sometimes be sufficient and can be used to estimate $\gamma.$ Such a two-step procedure can be readily defined adapting ideas from the literature on spatial Cliff-Ord models; cp. e.g., Kelejian and Prucha (1998, 1999). The first step relies on linear moment conditions for the identification of the parameters $\delta$ and $f_{1}$ which exploit that $H$ and the disturbance vector $(I+\rho M)(u_{1}-f_{1}u_{2})$ of model (8) are orthogonal. The idea is to concentrate out $\delta$ and $f_{1}$ in the first stage using the linear moment function. Note that by Lemma 2.1 the linear moment condition $E\left[H^{\prime}\left(y_{1}-y_{2}f_{1}-\underline{X}\delta\right)\right]=0$ holds at the true parameter values for $\delta$ and $f_{1}.$ For simplicity consider the case where $z$ is of dimension $n\times 1$. Then this equation corresponds to a regular just identified linear instrumental variables problem with included exogenous covariates $\underline{X}$ and where $z$ is instrumenting for the endogenous variable $y_{2}$. Letting $W=\left(y_{2},\text{$\underline{X}$}\right)$ and solving the sample analog of the moment condition yields
$\left(\begin{array}[]{c}\tilde{f}_{1}\\ \tilde{\delta}\end{array}\right)=\left(H^{\prime}W\right)^{-1}H^{\prime}y_{1}.$ (19)
The estimator for the coefficient $f_{1}$ has the familiar form of just identified two stage least squares with included covariates
$\tilde{f}_{1}=\frac{z^{\prime}Q_{X}y_{1}}{z^{\prime}Q_{X}y_{2}}$ (20)
where $Q_{X}=I-\underline{X}(\text{$\underline{X}$}^{\prime}\text{$\underline{X}$})^{-1}\text{$\underline{X}$}^{\prime}$ is the residual operator of a projection onto $\underline{X}$. With $\tilde{f}_{1}$ determined by (20) $\tilde{\delta}$ is given by
$\tilde{\delta}=(\underline{X}^{\prime}\underline{X})^{-1}\underline{X}^{\prime}(y_{1}-\tilde{f}_{1}y_{2}).$ (21)
In the second step we estimate $\rho$ using a form of the quadratic moment function, where we plug into the residual vector $\tilde{\delta}$ and $\tilde{f}_{1}$ for $\delta$ and $f_{1}$. Let
$\epsilon^{+}\left(\theta\right)=\epsilon^{+}\left(\rho,f_{1},\delta\right)=\left(I+\rho M\right)^{-1}\left(y_{1}-f_{1}y_{2}-\underline{X}\delta\right)$
and consider the quadratic moment vector
$m_{\epsilon}^{q}\left(\rho\right)=\frac{1}{n}\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)^{\prime}A\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)$
with $A$ defined as before. The first stage estimator for $\rho$ then is obtained as
$\tilde{\rho}=\operatorname{argmin}_{\rho}\left(m_{\epsilon}^{q}\left(\rho\right)\right)^{2}.$ (22)
Consistency and the asymptotic distribution of $\tilde{\theta}$ is discussed in Section 4 together with a discussion of the efficient estimator $\hat{\theta}$.
3 Empirical Results
In this section, we apply our identification strategy to the Tennessee’s Project STAR (student-teacher achievement ratio) data, exploring peer effects among students in Kindergarten through Grade 3.999We downloaded the data from Harvard Dataverse (https://dataverse.harvard.edu/dataset.xhtml?persistentId=hdl:1902.1/10766), which differs slightly from the version used by Graham(2008) and Rose(2017). For example, they construct class ID using teacher characteristics while our data has class ID readily available. Their kindergarten sample has 6,172 students, and ours has 6,325. We believe our data set is more trustworthy as it matches the description in Word et al. (1990) in terms of sample size and summary statistics. Project STAR is a randomized experiment aiming at studying the impact of class size reduction on children’s development. The data set has been widely used in studies of class size, peer effects, teacher effectiveness and other education-related topics (e.g., Krueger 1999; Dee 2004; Graham 2008; Chetty et al. 2011; Mueller 2013). Here we briefly discuss the details which are relevant for our study. For a complete description of the project and data, see Word et al. (1990); Mosteller (1995); Boyd-Zaharias et al. (2007).
Project STAR was carried out in Tennessee from 1985 to 1989, with additional data on the participants collected after the project ended. The experiments followed the kindergarten cohort of 1985 in participating schools for four years, from Kindergarten to the third grade. At the start of the 1985 academic year, participating schools randomly assigned kindergarten students and teachers into small classes (with an intended size of 13-17 students), regular classes (with an intended size of 22-25 students) and regular classes with a full-time teacher’s aide. In the actual implementation, small classes included 11 to 20 students, while regular classes (with or without aides) included 15 to 30 students. Randomness of the initial class assignment is ensured by the careful implementation of the random assignment protocol and has been examined by a number of studies, e.g., Krueger (1999); Chetty et al. (2011). However, nonrandom attrition, switching and migration might have happened in higher grades (Hanushek, 1999). Our analysis uses the sample of each of the four grades, as our estimator remains consistent even without random assignment. The bottom panel of Table 1 summarizes the number of students and classes in each grade. While there are 11601 students in the experiment in total, the sample size in each grade is between 6325 and 6840 due to students migrating in or out of the participating schools. The total count of schools is 79 at the Kindergarten level, but it decreases to 76 at the Grade 1 level and subsequently to 75 at the Grade 2-3 level. This decline can be attributed to schools withdrawing from the STAR program. There are between 325 and 340 classes in each grade, with 124 to 140 of them being small classes.
We collect raw Stanford Achievement Tests (SAT) scores in mathematics (math), reading (read), listening (list) and word study skills (word) as test $t=1$ and test $t=2$ in our model. SAT is a nationally standardized test with scores comparable across grades. The test was administered on state specified examination days, occurring between late March and early April each year (Word et al., 1990; Krueger, 1999). The narrow time frame for the tests lends credence to the assumption that student’s ability and preparation remain stable over the testing window. SAT scores for these four subjects are available from Grade K to Grade 3. The mean and standard deviations of scores for the four SAT tests are in the third panel of Table 1. The four tests generally share comparable means, except that the mean listening score is much higher than others in Grade K. Notably, reading and word study skills scores have means that are close in all grades. With four test scores, there are 12 pairs of $y_{1}$ and $y_{2}$. We report results for only six unique combinations, as reversing the order of $y_{1}$ and $y_{2}$ does not change the estimate of $\rho$, transforms $f_{1}$ to its reciprocal, and will scale the coefficients of control variables by $-f_{1}$.
Our main model includes a set of student and teacher characteristics, the mean and standard errors of which are summarized in the top panel of Table 1 by grade. Between $56.0\%$ and $59.2\%$ of the students receive a free lunch during the experimental period, an indicator of low family income. To account for the influence of race, we employ an indicator for black students, as more than 98% of the minority students are of black ethnicity. The share of black students ranges from $32.6\%$ to 34.7%. Girls make up $48.0\%$ to $48.6\%$ of the sample. We calculate age as of April 1st in each year, which is roughly the SAT examination date, using students’ dates of birth. Teacher’s characteristics are weighted by class size. Between 16.5% and 20.9% of the students are taught by black teachers. Meanwhile, 34.6% to 44.2% of the students have teachers with a master’s degree or higher. Average years of experience of the teachers for different grades are between 9.3 and 13.9.
In our preferred specification, we include four types of control variables: school fixed effects (excluding one dummy for the first school), class type fixed effects (excluding the dummy for small classes), student characteristics (including indicators for free lunch, black ethnicity, female gender, and age) and teacher characteristics (including indicators for black teachers, having master’s degree or higher, and years of experience) and peer characteristics – specifically, the leave-out-mean of student characteristics. We start from the specification with only school fixed effects and subsequently introduce other types of control variables incrementally. We allow for heteroscedasticity of $u$ between small classes and regular classes (with or without teacher’s aide). Standard errors are clustered at the class level. The estimators for $f_{1}$ and $\rho$ are summarized in Tables 6 and 7 respectively.
Before proceeding to discussing the main estimates of peer effects, we present several summary statistics in support of our model specifications and assumptions. Our estimator relies on the premise that tests $t=1$ and $t=2$ measure similar skills within a brief time frame. The combination of reading and word study skills aligns most closely with this narrative. To document the close connection between test scores, we use four measures to evaluate the correlation between various SAT scores from Grade K to Grade 3 in Table 2. The first two measures on the left panel are Spearman correlations between $y_{1}$ and $y_{2}$, and between $Q_{X}y_{1}$ and $Q_{X}y_{2}$. Here $Q_{X}=I-\underline{X}(\underline{X}^{\prime}\underline{X})^{-1}\text{$\underline{X}$}^{\prime}$ denotes the residual projection matrix of $\underline{X}$, where $\underline{X}$ encompasses the complete set of control variables: school fixed effects, class type fixed effect, student characteristics, teacher characteristics, and peer characteristics. The right panel of Table 2 reports the pseudo $R^{2}$ from 2SLS, i.e., the Spearman correlation between $y_{1}$ and $\hat{y}_{1}$, or between $y_{2}$ and $\hat{y}_{2}$. Here $\hat{y}_{1}=\tilde{f}_{1}y_{2}+\underline{X}\tilde{\delta}$, with $\tilde{f_{1}}$ and $\tilde{\delta}$ defined in (20) and (21). Similarly, $\hat{y}_{2}$ corresponds to the counterpart obtained by reversing the order of $y_{1}$ and $y_{2}$.  [^10] We also adjust for missing covariate values using a similar method as the one for the GMM estimator. That is replacing the peer averages of the observed individuals $\bar{x}_{c(-r)t}^{obs}=(n_{c}\bar{x}_{ct}-x_{crt}^{obs})/(n_{c}-1)$ with $\tilde{x}_{c(-r)t}^{obs}=(n_{c}\bar{x}_{ct}^{obs}-x_{crt}^{obs})/(n_{c}-1)$, where $\bar{x}_{ct}^{obs}$ is the average of all observed $x$.These adjustments make little difference to the results. We drop individuals with missing outcome variables from the sample. We discuss the procedure in more detail in Appendix E. The different correlation measures paint a consistent picture of the correlation between the six reported pairs as well as across grades. The listening score is generally a bit less correlated with other measures. The math scores are somewhat more correlated with reading and word skills than with listening. We find the strongest connection between word study skills and reading scores which exhibit substantial correlations of approximately 0.9 across all four measures and in all grades. This finding is consistent with these scores measuring skills that are both related to language comprehension. Correlations between other pairs range between 0.5 and 0.8. While a key identifying assumption, Assumption 2 is not directly testable because it relates to unobserved ability, finding strong correlation between test scores is consistent with a setting where test scores have common conditional means. If this interpretation of the results in Table 2 is correct then we would expect the word-read pair of scores to provide the best control for unobserved individual and class effects that may be confounding measures of peer effects at the classroom level. Our model implies that conditional on covariates correlation between tests only is due to unobserved interactive effects $\mu_{c}^{\ast}f_{t}.$ While not a test for correct specification, our results are consistent with $\mu_{c}^{\ast}f_{t}$ accounting for a large fraction of the variation in test scores.
Further evidence of the ability of test score differences to control for unobserved classroom and individual effects is the size of the estimated coefficient $f_{1}$. If scores measure closely related skills we expect $f_{1}$ to be close to one. In that scenario Assumption 2 is interpreted to imply that unobserved individual specific and classroom level effects captured by $\mu_{c}^{\ast}$ affect both scores in the same way. We report estimated values of $f_{1}$ for the set of test score pairs in Table 6. Point estimates for $f_{1}$ are obtained by our efficient GMM estimator for all grades and for specifications with just a school effect in Column (1), school and class type effects in Column (2), with a full set of controls excluding contextual peer effects in Column (3) and the full set of controls in Column (4). The coefficient $f_{1}$ is estimated with high precision in Columns (1) through (3) and somewhat less precisely in Column (4).The point estimates are generally close to one with a few exceptions, such as for the listening scores in grade K and specifications (1)-(3) as well as the listening scores in grades two and three for specification (4). The parameter estimates for the read-word score combination are remarkably stable across all specifications and across all grades, are close to one and estimated precisely. This is further evidence that the read-word score pair is best suited to control for unobserved class level effects.
We explore one additional approach to evaluate the plausibility of Assumption 2. If unobserved individual and class effects $\mu_{c}^{\ast}$ are stable over time then scores from previous years should be good predictors for current year scores. Here, it is important to note that our theory applies to pairs of closely related tests, and that $\mu_{c}^{\ast}$ is not assumed to be the same for different pairs. Thus our theory does not assume $\mu_{c}^{\ast}$ to be time invariant between grades. Conceptually, $\zeta^{*}$ measures unobserved ability just prior to taking the tests in the current year. As a result, $\mu_{c}^{\ast}$ similarly reflects only current year unobserved ability and other classroom level effects coming from teachers and peers. Unobserved effects $\mu_{c}^{\ast}$ are expected to vary year by year for various reasons, including the fact that instruction during the year improves skills, that teachers may change and that students may move to different classes. However, it is expected that at least some components of $\mu_{c}^{\ast}$ that are more closely related to individual student ability are time invariant. We examine this hypothesis by measuring the explanatory power of lagged scores on current scores. The regression sample includes stayers in each grade, defined as individuals participating in the STAR project for both the previous and the current grades to ensure availability of both current and lagged scores. The Grade 1 sample includes 66% stayers, while Grade 2 and 3 samples have 73.8% and 79.6% respectively. Among the stayers in Grade 1, 88.7% stay in the same type of classes (small v.s. regular(w/wo) aide) in the same schools. This proportion exceeds 93% for Grade 2 and Grade 3.111111Based on the authors’ calculation using the STAR data. The share is lower if we consider regular and regular/aid as different types, due to a random reallocation of students in regular and regular/aid classes to these two types within school in Grade 1. However, the literature generally finds no significant impact of teacher’s aide (e.g., Krueger, 1999). The stability of school and class types among the large share of stayers fosters a steady learning environment. While newcomers and leavers might marginally influence peer quality, the allocation of newcomers to different classrooms is random within schools. Consequently, when accounting for school fixed effects, we anticipate a consistent level of peer quality over time.
The relatively stable environment that these students are in suggests that $\mu_{c}^{\ast}$ and $f_{t}$ should contain significant time invariant components that in turn imply high correlation between observed scores and lagged scores. As a result we expect to see relatively high $R^{2}$ in predictive regressions with lagged scores as controls in addition to the observed characteristics. Consequently, we regress test outcomes on the full set of control variables (school and class type fixed effects, students, teacher and peer characteristics and a constant term), and then add lagged outcomes of the same or different type and their peer averages to assess changes in $R^{2}$. The results are summarized in Table 3. The results show the predictive power of lagged scores for the same test in bold, as well as when using other tests instead, for grades one through three. The predictive power of lagged scores is documented by comparing the $R^{2}$ with and without adding lags. For example, for the word score, adding its own lag improves the $R^{2}$ from .244 to .435 in grade one, from .25 to .539 in grade two and from .229 to .488 in grade three. Using the lagged reading score, which is the score of the test most closely related to the word score, leads to increases in the $R^{2}$ that are even slightly higher. Overall, the coefficients on lagged scores when using the own lag of word as well as the lagged reading score are large in magnitude and highly statistically significant. These findings hold up for the other test score pairs with similar magnitudes and statistical significance. As may be expected, the predictive effects are slightly weaker for listening scores than for math, read and word. The results are further evidence that the data distributions are consistent with key features of our model.
Our analysis thus far has documented that pairs of tests scores are highly correlated both contemporaneously as well as over the course of one school year. This evidence is consistent with restrictions imposed by Assumption 2 and the hypothesis that closely related scores depend on the same unobserved components. A caveat is that closely related scores may remove too much common variation when quasi-differences of these scores are used and thus render statistical evidence about peer effects imprecise. Table 4 reports pseudo $R^{2}$ measures of model estimates for the six pairs of scores and model specifications (1) through (4) with column (1) only controlling for school fixed effects, and column (4) being the full specification with school and class type effects, student and teacher characteristics as well as peer characteristics.
The observed pseudo $R^{2}$ measures are generally below 0.2 with a few exceptions for Grade $K$. We also see that the $R^{2}$ is a bit higher for models with listening scores. Moreover and as expected the pseudo $R^{2}$ generally increases with larger models, or in other words by moving from Column (1) to (4). The read-word score pair often has the lowest $R^{2}$ which is an indication that the differential score in these specifications already controls for most of the variation explained by the covariates. We take this as further evidence that read-word best fits the model assumptions we impose.
Table 5 presents standard deviations of $Q_{x}y$ for four test scores and $Q_{x}(y_{1}-f_{1}y_{2})$ for six pairs of test scores for Grade K to 3. The left and right panels are for small and regular classes respectively. Control variables for specifications (1)-(4) are the same as those in Table 4, as discussed above. Estimates of $f_{1}$ are obtained from our efficient GMM estimator and are presented in Table 6. We see that in Grade K and Grade 1, the ratio of standard deviations of $Q_{x}(y_{1}-f_{1}y_{2})$ to that of $Q_{x}y$ is around 1/2 for most pairs of tests and 1/5 to 1/4 for the read-word pair. In Grade 2 and 3, this ratio is around 1/3 for most pairs and about 1/5 for the read-word pair. This shows that quasi-differencing removes a significant part of the variation in the test scores, especially for the read-word pair, consistent with the term $\mu^{*}f_{1}$ playing a major role in determining test performance. At the same time, the results in Table 4 indicate that the inclusion of controls does not eliminate much of the remaining variation not captured by fixed effects. The conclusion is that there is between 20 to 30 percent of test score variation remaining for the identification of $\rho$.
Table 7 presents estimates and standard errors (in parentheses) for the peer effects parameter $\rho$. To provide some context and interpretation of $\rho$ consider the education production function in Equation (1). When $\rho=0$ such that there are no peer effects, it follows that the observed scores are $y_{ct}=\alpha_{c}f_{t}\mathbf{1}_{c}+X_{ct}^{c}\beta_{t}^{c}+(X_{ct}^{p}\beta_{t}^{p}+y_{ct}^{*})$ where $y_{ct}^{*}$ has the interpretation of an unobserved term that is measured in the same units as the test score variable. When $\rho$ is different from zero test score performance of student $i$ changes by $\rho$ times the leave-out-average of $y_{ct}^{*},$ through the term $\rho M_{c}y_{ct}^{*}.$ This interpretation implies that $\rho$ measures the effect on student $i$’s test score as a fraction of the increase in peer ability measured in units of test score performance. Similarly, the effect on test scores as a result of changes in peer characteristics are measured as $\rho\beta^{p}$ times the changes in relevant peer characteristics.
Standard errors are clustered at the class level. Test types of $y_{1}$ and $y_{2}$ are indicated in Columns 1 and 2. Control variables for the four specifications are outlined at the bottom of the table. Specification (1) includes only school fixed effects. Specification (2) further adds class type fixed effects. Specification (3) introduces student characteristics ($free\,lunch$, $black$, $girl$, $age$) and teacher characteristics ($black$, $master$, $experience$). Specification (4) further controls for peer characteristics, defined as the leave-out-mean of the four student characteristics variables.
The comparison of the four specifications demonstrates the robustness of our estimators to additional controls. We now focus on Specification (4), which includes the full set of control variables. Across all grade levels and all test pairings, our results consistently reveal positive and statistically significant peer effects, ranging from 0.2 to 0.7. The results we find are relatively stable across the four specifications in Columns (1)-(4) with the magnitude of the coefficient $\rho$ generally decreasing somewhat for the larger models. The results for read-word are more robust in this sense especially in Grades $K$ and one. Estimated peer effects tend to be largest for specifications involving the listening score as well as the math-word score. The results for Grades $K$, two and three are quite similar, while results for Grade one are generally somewhat smaller in magnitude but remain statistically very significant. Importantly, when examining the pairing of reading and word study skills—a context wherein the tests evaluate closely related abilities as documented by our prior analysis—the effects are notably smaller than for the remaining pairings, ranging from 0.2 to 0.4 across different grade levels. The implications of these results are that improvements in peer quality measured in terms of potential SAT scores $y_{t}^{*}$ translate at the rate of 20 to 40 percent to improvements of individual scores or in other words a 10 point increase in the peer potential SAT score results in a 2 to 4 point increase in individual SAT scores. These effects are at the lower end of the spectrum of results reported in the literature.
The comparison with other results in the literature is complicated by the fact that often a model with endogenous peer effects is estimated. To focus ideas, consider the case without covariates and only one outcome measure $y_{1}$. Then, an endogenous peer effect specification is $y_{1}=\lambda My_{1}+u.$ If peer effects are measured with full rather than leave-one-out means this formulation is identical to our model which is $y_{1}=\left(I+\rho M\right)u$ in this simplified stylized setting. The parameters $\rho$ and $\lambda$ satisfy the one-to-one mapping $\lambda=\rho/(1+\rho)$ which is a consequence of the fact that $\left(I+\rho M\right)=\left(I-\lambda M\right)^{-1}$. The inverse captures multiplier effects of peer performance inherent in the endogenous peer effect formulation, while $\rho$ is a summary measure of all these multiplier effects. When peer effects are measured by leave-one-out rather than full means the relationship between the two formulations is no longer exact. However, it can be shown that $\left(I+\rho M_{c}\right)^{-1}=\left(I-\lambda M_{c}\right)$ continues to hold approximately with an error that is $O\left(n_{c}^{-2}\right)$ where $n_{c}$ is the size of class $c$. Based on this approximation, we convert our measure of $\rho$ to an endogenous peer effect $\lambda=\rho/(1+\rho),$ or equivalently, $\rho=\lambda/\left(1-\lambda\right)$ for the purpose of the discussion that follows.
In comparison with existing results for peer effects employing STAR data, our estimates are relatively small in magnitude but estimated with high precision and statistical significance. Using normalized kindergarten SAT scores, Graham (2008) derives endogenous peer effects of 0.46 and 0.56 for math and reading respectively, which translates to $\rho=0.86$ or $\rho=1.30$ in our context. [^12] Graham (2008) does not directly estimate endogenous peer effects. His model is $y_{ic}=\alpha_{c}+(\gamma-1)\bar{\epsilon}_{c}+\epsilon_{ic}$, with $y_{ic}$ being the outcome of student $i$ in class $c$ and $\bar{\epsilon}_{c}$ is the full mean of unobserved ability $\epsilon_{ic}$ in class $c$. Hence $\gamma-1$ is comparable with our estimate $\rho$, both are coefficients of average peer unobserved ability. Due to the full-mean specification, the equation is equivalent to $y_{ic}=\lambda\bar{y}_{c}+(1-\lambda)\alpha_{c}+\epsilon_{ic}$, where $\lambda=1-\frac{1}{\gamma}$ is the endogenous peer effects. Using normalized Kindergarten SAT math score as outcome, Rose (2017) finds an endogenous peer effect of 0.65 or a correlated effect (measured by coefficients of the peer average errors) of 0.71. When both are considered simultaneously, endogenous and correlated effects are 0.90 and -0.03, albeit with standard deviations over 1.7, indicating weak identification. Boozer and Cacciola (2001), using average percentiles of math and reading scores as outcomes and an instrumental variable methods for estimation, identify endogenous peer effects of 0.30, 0.86 and 0.92 for Grades 1 to 3 respectively, with standard deviations of 1.0, 0.12 and 0.04. Sojourner (2013) uses first-grade averages of SAT percentiles of math, reading and listening as outcome measure, and checks the exogenous peer effects of lagged peer outcomes. He finds an estimate of about 0.35 with a standard deviation around 0.14.
At the upper end of the spectrum are estimates by Lewbel, Qu, and Tang (2023) who report endogenous peer effects of 0.85 and 0.92 for small and regular classes in third-grade SAT math scores, each with standard deviations around 0.02. Translated to our parametrization, these estimates correspond to $\rho=5.6$ and $\rho=11.5$ respectively, using the approximation $\rho=\lambda/(1-\lambda)$. The magnitude of their findings is remarkably high and suggests that a mere 10-point rise in ability equivalent peer SAT scores leads to an increase of individual scores of up to over 100 points.
The observed peer effects in our study exhibit smaller magnitudes compared to most of the results discussed above. It is important to note that estimates in Lewbel, Qu, and Tang (2023), Rose (2017) and Boozer and Cacciola (2001) pertain to coefficients of peer’s average contemporaneous scores, resulting in a multiplier effect as explained above. In contrast, our estimated peer effects are coefficients of average peer unobserved ability. This notion is more akin to the “correlated effects” discussed in Rose (2017). Overall, we find a noticeable difference in peer effects’ magnitude between our study and those of Rose (2017) and Graham (2008).
Our proposed estimator leads to highly precise estimates compared to all the aforementioned estimates, except that of Lewbel, Qu, and Tang (2023). The enhancement in precision can be attributed in part to the elimination of individual fixed effects through differencing. Certain individual characteristics that have been recognized for their significant influence on scores are not observed in the data for Project STAR. These include parents’ education, family income and most importantly unobserved IQ. Our estimator, by removing unobserved class effects and individual fixed effects, effectively eliminates residual variance associated with these characteristics which may be one explanation for the precise estimates we obtain. In addition, our GMM estimators use additional quadratic moment restrictions efficiently.
The aforementioned studies have each focused on one or two specific tests and are often limited to a particular grade. Methods that rely on random peer assignment are typically constrained in where they can be applied. In project STAR the random assignment assumption is more defensible in earlier grades (Boozer and Cacciola, 2001; Graham, 2008; Sojourner, 2013). In contrast, our analysis does not require random assignment and facilitates the estimation of peer effects across diverse test types and grade levels. Encouragingly, our findings consistently fall within a reasonably narrow range, affirming the robustness and reliability of our method.
We also provide the coefficients of peer averages of individual characteristics (age, black, free lunch, and girl) in Table 8. These four variables all pertain to test-invariant individual characteristics. The estimated coefficients can be expressed as $\rho(\beta_{v1}^{p}-f_{1}\beta_{v2}^{p})$, where $\rho\beta_{v1}^{p}$ and $\rho\beta_{v2}^{p}$ represent the exogenous peer effect of the variable on tests $t=1$ and $t=2$ respectively. Thus, these coefficients capture the differential exogenous peer effects across the two tests. We also scale the coefficients for dummies $girl$, $black$ and $free\,lunch$ by 1/100 so that the coefficients are the impact on the outcome of 1 percentage point changes in share of peers who are girl/black/free lunch receivers. The first observation is that, except for kindergarten where the presence of girls among peers positively impacts reading more than word study, none of the four variables show significant differential exogenous peer effects in the closely related word-read pair. Secondly, the share of black peers and the average peer age generally display insignificant differential exogenous peer effects. An exception is in Kindergarten, where peer age more favorably affects listening scores than reading scores. Given our general expectation of a negative impact from peers receiving free lunch and a positive impact from a higher share of girls, the results suggest that these two exogenous peer effects are stronger for reading and word study skills compared to listening and mathematical abilities in Grade 1. Similar patterns emerge in other grade levels, albeit with lower significance and consistency. Overall, the results indicate small and mostly insignificant peer effects on differential scores. This is consistent with a scenario where $y_{2}$ is essentially sufficient to control for both observed and unobserved test invariant determinants of test performance. In fact, when $f_{1}=1$ and $\beta_{v1}^{p}=\beta_{v2}^{p}$ these effects would be zero in our model and the overall evidence points to the conclusion that this is approximately the case in the sample as well.
4 Identification and Inference
4.1 Identification Conditions
We next discuss identification of our model. To establish identification we introduce additional assumptions imposed on the parameter space. We call the model identified at the parameter value $\theta_{0}$ if the population moment conditions underlying the formulation of our estimators are only satisfied at $\theta=\theta_{0}$ for all $\theta\in\Theta$. This definition can be seen as an extension of classical definitions of identification such as Rothenberg (1971) for likelihood based approaches. In general, identification is neither necessary nor sufficient for consistency which is established separately, see Newey and McFadden (1994) for a discussion, and White (1980) and Pötscher and Prucha (1997) for a discussion of the related concept of identifiable uniqueness.
Assumption 4.
We use subscript zero to denote true parameters and treat $\delta_{0}$, ignoring the restrictions given in (7), as an unrestricted parameter vector. Let $\theta_{0}=\left(\rho_{0},f_{1,0},\delta_{0}^{\prime}\right)^{\prime}$ with $\delta_{0}=\delta(f_{1,0},\rho_{0},\beta_{0})$. Then assume $\theta_{0}\in\Theta$, where $\Theta$ is a compact subset of $\Theta_{0}=\left(-1,1\right)\times(-K_{f},K_{f})\times(-K_{X},K_{X})^{p_{x}}$, and where $K_{f}$ and $K_{X}$ are finite positive constants and $p_{x}$ is the column dimension of $\underline{X}$. In addition, let $\gamma_{0}=\left(\gamma_{1,0},...,\gamma_{J,0}\right)^{\prime}$ with $\gamma_{j,0}^{2}=(\sigma_{1,0}^{2}+\sigma_{2,0}^{2}f_{1,0}^{2})\rho_{j,0}^{2}$. Then $\gamma_{0}\in\Gamma$ with $\Gamma=[K_{\gamma}^{-1},K_{\gamma}]^{J}$ where $K_{\gamma}>1$ is a finite constant.
Remark 4.1.
The assumption implies that the parameter space for $\rho$ is a subset of $[-K_{\rho},K_{\rho}]$ for some positive $K_{\rho}<1$.
In addition, we impose the the following assumption on $A$.
Assumption 5.
(i) The matrix $A=(a_{ij})=\operatorname{diag}_{c=1}^{C}(A_{c})$ is measurable with respect to $\mathcal{Z}_{n}$ and has zero diagonal elements. Moreover, $A_{c}$ is a function of $M_{c}$ and is symmetric.
(ii) $\mathbf{1}_{c}^{\prime}A_{c}\mathbf{1}_{c}/n_{c}\geq\underline{K}_{a}>0$ for some constant $\underline{K}_{a}$.
(iii) $\sup_{i}\sum_{j=1}^{n}\left|a_{ij}\right|\leq K_{a}<\infty$ for some constant $K_{a}$.
Two valid choices for $A$ are $A=M$ and $A=M^{\prime}M-\operatorname{diag}\left(M^{\prime}M\right).$
Assumption 6.
Group size $n_{c}$ satisfies $2\leq n_{c}\leq\bar{n}<\infty$ for some constant $\bar{n}$.
Before discussing the general case, we start discussing identification of our model in the case where no additional covariates are present. Note that in the case of no covariates, the coefficients $\delta$ of the covariates are omitted, $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ is reduced to $(\rho,f_{1})^{\prime}$ and (14) becomes $u^{+}(\theta,\gamma)=y^{+}(\theta,\gamma)$.
Observe that $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}(\gamma_{\tau_{c}}^{2}I_{c})$ is diagonal for all admissible $\gamma$ and that $\Omega_{0}=\Omega(\gamma_{0})=\operatorname{Var}(u_{1}-f_{1,0}u_{2}|\mathcal{Z}_{n})$ in light of Assumptions 2 and 3 and Theorem 2.1. Also, $u^{+}(\theta_{0},\gamma)=\Omega(\gamma)^{-1/2}\Omega_{0}^{1/2}u^{+}.$ Thus it is readily seen that for all $A$ satisfying Assumption 5(i) we have
$E\left[u^{+}(\theta_{0},\gamma)|\mathcal{Z}_{n}\right]=0,$ (23)
$E\left[u^{+}(\theta_{0},\gamma)^{\prime}Au^{+}(\theta_{0},\gamma)|\mathcal{Z}_{n}\right]$ $=\operatorname{tr}\left(\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\Omega_{0}\right)$ (24)
$=\operatorname{tr}\left(A\Omega(\gamma)^{-1}\Omega_{0}\right)$ (25)
$=0$ (26)
observing that the diagonal elements of $A\Omega(\gamma)^{-1}\Omega_{0}$ are zero.
We next discuss identification for a given matrix $H=z\equiv\left(z_{1}^{\prime},...,z_{c}^{\prime},...,z_{C}^{\prime}\right)^{\prime}$ of $\mathcal{Z}_{n}$ measurable instruments, where the $z_{c}$ are $n_{c}\times q$ matrices of random variables that are invariant at the classroom level, i.e., $z_{c}=\mathbf{1}_{c}\grave{z}_{c}$, where $\mathbf{1}_{c}$ is an $n_{c}\times 1$ vector with all elements equal to one and $\grave{z}_{c}$ is a $1\times q$ vector of class level characteristics. We focus on the just identified case with $q=1$ for ease of exposition. Additional overidentifying restrictions can be added but do not affect arguments related to identification. Consider the moment vector
$m_{n}\left(\theta,\gamma\right)=n^{-1/2}\left[\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\end{array}\right]$ (27)
where $A$ is a matrix that satisfies the properties listed in Assumption 5.
We next show that $E\left[m_{n}\left(\rho,f_{1},\gamma\right)\right]=0$ if and only if $(\rho,f_{1})=(\rho_{0},f_{1,0})$. Identification of $f_{1,0}$ can be established from an inspection of the linear moment condition. Identification of the social interaction parameter $\rho_{0}$ follows from an analysis of the quadratic moment condition.
Lemma 4.1.
As a special case of (9) without covariates consider the data generating process $y_{t}=\mu^{\ast}f_{t,0}+\left(I_{n}+\rho_{0}M\right)u_{t}$ where $\mu^{*}$ is defined in (4). Suppose Assumptions 1, 2, 3, 4, 5, and 6 hold, let $m_{n}\left(\theta,\gamma\right)$ be defined in (27) and let $H=z$. In addition, assume that $q=1$ and
$n^{-1}\left|E\left[y_{2}^{\prime}\Omega(\gamma)^{-1/2}z\left|\mathcal{Z}_{n}\right.\right]\right|\geq K_{y}>0$ (28)
for all $\gamma\in\Gamma$ and all $n$, where $K_{y}$ is a constant. Then for all $\gamma\in\Gamma$ we have $E\left[m_{n}\left(\rho_{0},f_{1,0},\gamma\right)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{n}\left(\rho,f_{1},\gamma\right)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ a.s. for all $\left(\rho,f_{1}\right)\neq\left(\rho_{0},f_{1,0}\right)$ and $\gamma\in\Gamma$.
A proof of Lemma 4.1 establishing the identifiability of $\left(\rho_{0},f_{1,0}\right)$ is given in the appendix. The subsequent corollary considers a special case of the lemma that is of interest in the context of test scores, and arises when $H=\mathbf{1}_{n}$ and when test scores are non-negative. The latter is typically the case for non-normalized test scores.
Corollary 4.1.
Suppose the assumptions of Lemma 4.1 hold, $H=\mathbf{1}_{n}$, $y_{t}$$\geq 0$, and suppose condition (28) is replaced by $n^{-1}\sum_{c=1}^{C}E\left[y_{c2}^{\prime}\mathbf{1}_{c}\left|\mathcal{Z}_{n}\right.\right]>K_{y}>0$ for some constant $K_{y}.$ Then the conclusions of Lemma 4.1 hold.
We emphasize further that no restrictions on $\mu^{*}$ (other than the fact that $\mu^{*}$ does not depend on $t$ and $\mu^{*}\neq 0$ a.s.) are needed to identify $\rho_{0}$. Also, the analysis in the appendix shows that neither knowledge of, nor restrictions on heteroskedasticity are necessary for the identification of $\theta_{0}$. Thus, initial estimates of $\theta_{0}$ can be obtained with $\gamma_{j}^{2}=1$ for all $j=1,...,J$.
Next we turn to a discussion of identification of $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ when the model contains covariates. We use instruments $H=\left[\underline{X},z\right]$ for linear moment restrictions and a matrix $A$ for quadratic moment restrictions, with $A$ defined as before. Both $H$ and $A$ are $\mathcal{Z}_{n}$-measurable and $\operatorname{diag}(A)=0$. For ease of exposition we focus again on the case where $z$ is a vector. Recall that with covariates, $u^{+}(\theta,\gamma)$ is defined in (14).
We consider the following $(q+1)\times 1$ vector of linear and quadratic moment functions:
$m_{n}(\theta,\gamma)=\left(\begin{array}[]{c}m_{n}^{(l)}(\theta,\gamma)\\ m_{n}^{(q)}(\theta,\gamma)\end{array}\right)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\end{array}\right)$ (29)
where the functions $m_{n}^{(l)}(\theta,\gamma)$ and $m_{n}^{(q)}(\theta,\gamma)$ denote the linear and quadratic moment conditions respectively.
Note that our procedure is not designed to identify $\gamma$, nor is that required for the identification of $\theta$. Lemma 4.2 below shows that for all permissible $\gamma$ the moment condition satisfies $E\left[m_{n}(\theta_{0},\gamma)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{n}(\theta,\gamma)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ for $\theta\neq\theta_{0}$. The parameters $\gamma$ can be pinned down by estimating the variance of $u_{1}-f_{1,0}u_{2}$ for different types of groups. For identification, we further impose the following assumption.
Assumption 7.
(i)The absolute values of the elements $h_{ij}$ of $H=\left[\underline{X},z\right]$ are uniformly bounded in $i,j,n$ by some positive constant $K_{H}<\infty$.
(ii) The smallest eigenvalue of $\underline{X}^{\prime}\underline{X}/n$ is bounded from below by some $\xi_{X}>0$ uniformly in $n$.
(iii) Define $V(\rho,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$ and let
$Q_{V^{1/2}X}(\rho,\gamma)=I-V(\rho,\gamma)^{1/2}\underline{X}\left(\underline{X}^{\prime}V(\rho,\gamma)\underline{X}\right){}^{-1}\underline{X}^{\prime}V(\rho,\gamma)^{1/2}$
be the projection matrix onto the orthogonal complement of the column space of $V^{1/2}\underline{X}$, then
[…]
Remark.
Note that the elements of the matrix $H$ are by construction measurable w.r.t. $\mathcal{Z}_{n}=\sigma(X,z)$.
Since $\underline{X}$ contains $X$, part (i) of the assumption implies that also the elements of $X$ are bounded in absolute value by $K_{H}$. Observing that $(\underline{X},z)$ has full column rank by Assumption 7(iii), as setting $\rho=0$ and $\gamma=(1,...,1)^{\prime}$ in $V(\rho,\gamma)$, (30) becomes $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$, where $Q_{X}=I-\underline{X}\left(\underline{X}^{\prime}\underline{X}\right){}^{-1}\underline{X}^{\prime}$. It is readily seen from an inspection of the proof of Lemma 4.2 that a necessary condition for Assumption 7(iii) is that $\left(\underline{X},E\left[\mu^{*}|\mathcal{Z}_{n}\right]\right)$ has full column rank. The latter condition is equivalent to the condition that $\left(\underline{X},E\left[y_{2}|\mathcal{Z}_{n}\right]\right)$ has full column rank, observing that in light of (9) we have $E\left[y_{2}|\mathcal{Z}_{n}\right]=E\left[\mu^{*}+\underline{X}\beta_{2}|\mathcal{Z}_{n}\right]$ for some vector of parameters $\beta_{2}.$
Condition (30) generalizes condition (28) in Lemma 4.1. Under homoskedasticity or when estimators are not taking heteroskedasticity into account, the latter condition takes the more familiar form $n^{-1}\left|E\left[y_{2}^{\prime}z|\mathcal{Z}_{n}\right]\right|>K_{y}>0$. Further simplifications of (30) are obtained if in addition to heteroskedasticity, spatial correlation is ignored by setting $\rho=0.$ Considering this case is useful when discussing first step estimators for $\delta$ and $f_{1}.$ For the case with covariates (30) simplifies to $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$, where $Q_{X}$ is defined as above. Both versions of the simplified identification conditions are conventional rank conditions for linear instrumental variables estimators. A second scenario under which the conditions can be simplified is when $\gamma$ is evaluated at a fixed constant level. For example, when $\gamma=\bar{\gamma}$ then (28) is replaced by $n^{-1}\left|E\left[y_{2}^{\prime}\Omega(\bar{\gamma})^{-1/2}z\left|\mathcal{Z}_{n}\right.\right]\right|>K_{y}>0$ and a similar simplification obtained for (30).
The following lemma is a generalization of Lemma 4.1.
Lemma 4.2.
Let Assumptions 1, 2, 3, 4, 5, 6 and 7 be satisfied. Let $m_{n}\left(\theta,\gamma\right)$ be defined as in (29). Then for all admissible $\gamma\in\Gamma,$ $E\left[m_{n}\left(\theta_{0},\gamma\right)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{n}\left(\theta,\gamma\right)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ for $\theta\neq\theta_{0}$ a.s.
4.2 Consistency of Parameter Estimators
Formulating the moment conditions using transformed residuals as in (29) requires consistent estimates of the variance parameters $\gamma$. These can be obtained by first obtaining consistent but inefficient parameter estimates. To do so, we set all elements of $\gamma=(\gamma_{1},...,\gamma_{J})^{\prime}$ to one. Lemma 4.2 implies that the parameters $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ are still identified by the moment conditions in (29).
First step estimators $(\tilde{\rho},\tilde{f}_{1},\tilde{\delta}^{\prime})^{\prime}$ are defined in (19) and (22). [^13] Lee (2007a) considers first step and efficient GMM estimators using linear and quadratic moments. To formalize the identification result for these non-efficient first step estimators define the moment vector
$m_{\epsilon}\left(\theta\right)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}(I+\rho M)\epsilon^{+}\left(\theta\right)\\ \epsilon^{+}\left(\theta\right)^{\prime}A\epsilon^{+}\left(\theta\right)\end{array}\right).$ (31)
The next lemma shows that the moment conditions in (31) are sufficient to identify the parameters $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$.
Lemma 4.3.
Let Assumptions 1-6, and Assumption 7(i),(ii) be satisfied. Let $m_{\epsilon}(\theta)$ be defined as in (31). Assume further that $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$ a.s . Then $E\left[m_{\epsilon}\left(\theta_{0}\right)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{\epsilon}\left(\theta\right)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ a.s. if $\theta\neq\theta_{0}$ for all $\theta\in\Theta$
We next show that the initial estimators and the efficient estimators are consistent. To establish consistency we impose assumptions related to the convergence of sample averages and moments of sample averages to well defined limits. These assumptions are adaptations of Assumptions 2, 3 and 5 in Kuersteiner and Prucha (2020) to the present setting which differs somewhat from the framework of sequential exogeneity in that paper.
Let $i=n_{1}+\ldots n_{c-1}+r$ with $1\leq r\leq n_{c}$. Then, with some abuse of notation, denote the $i$-th diagonal element of $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)$ as $\gamma_{(i)}^{2}=\gamma_{\tau_{c}}^{2}$, noting that $\tau_{c}$ varies with $i$. Let the linear and quadratic moment vectors $m_{n}^{(l)}(\theta,\gamma)$ and $m_{n}^{(q)}(\theta,\gamma)$ be as defined in (29). Let $h_{i}=\left[h_{i1},...,h_{iq}\right]$ denote the $i$-th row vector of $H$, and let $a_{ij}$ denote the the $(i,j)$-th element of matrix $A$. Then note that with $u^{+}(\theta_{0},\gamma)=\Omega(\gamma)^{-1}(u_{1}-f_{1,0}u_{2})$, the variance of $m_{n}^{(l)}(\theta_{0},\gamma)$ is
$\frac{1}{n}E\left[H^{\prime}\Omega(\gamma)^{-1}\Omega(\gamma_{0})H\right]=\frac{1}{n}\sum_{i=1}^{n}E\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}h_{i}^{\prime}h_{i}\right]$ (32)
and the variance of $m_{n}^{(q)}(\theta_{0},\gamma)$ is
$\frac{1}{n}E\left[\operatorname{tr}\left(\Omega_{0}\Omega^{-1}A\Omega_{0}\Omega^{-1}A\right)\right]=\frac{1}{n}E\sum_{i=1}^{n}\sum_{j=1}^{n}\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}\frac{\gamma_{(j),0}^{2}}{\gamma_{(j)}^{2}}a_{ij}^{2}\right].$ (33)
The assumption below ensures that for any admissible $\gamma$, $\operatorname{Var}\left(m_{n}^{(l)}(\theta_{0},\gamma)\right)$ and $\operatorname{Var}\left(m_{n}^{(q)}(\theta_{0},\gamma)\right)$ and their corresponding sample analogues converge, respectively, to finite positive definite matrices.
Assumption 8.
The following holds :
$n^{-1}\sum_{i=1}^{n}E\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}h_{i}^{\prime}h_{i}\right]\rightarrow V_{\gamma}^{h},\quad n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}E\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}\frac{\gamma_{(j),0}^{2}}{\gamma_{(j)}^{2}}a_{ij}^{2}\right]\rightarrow V_{\gamma}^{a},$
where the elements of $V_{\gamma}^{h}$ and $V_{\gamma}^{a}$ are finite , and
[…]
The matrix $V_{\gamma}=\operatorname{diag}\left(V_{\gamma}^{h},2V_{\gamma}^{a}\right)$ is positive definite.
Assumption 9.
(i) Let $C\left(\theta,\gamma\right)$ be an $n\times n$ matrix of the form $D$, $DP$, $DAD$, $DADP$, or $PDADP$, where $D$ is an $n\times n$ positive diagonal matrix with elements which are uniformly bounded and measurable w.r.t. $\mathcal{Z}_{n}$ and where $P=\operatorname{diag}_{c}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right)$ and where $p_{c}$ and $q_{c}$ are continuously differentiable functions of the parameters $\theta$ and $\gamma$ as well as of variables generating $\mathcal{Z}_{n}$, and $I_{c}^{\ast}$ and $J_{c}^{\ast}$ are defined in Appendix B. Then
[…]
where $\Upsilon_{a},\Upsilon_{b}$ are selected from the set $\left\{H,MH,\mu^{*},u_{1},u_{2}\right\}$.
(ii) $\lim_{n\rightarrow\infty}\sup_{\theta,\gamma}\left\|E\left[n^{-1}\Upsilon_{a}^{\prime}C\left(\theta,\gamma\right)\Upsilon_{b}\right]-\mathscr{U}_{a,b}\left(\theta,\gamma\right)\right\|=0$ where
$\mathscr{U}_{a,b}\left(\theta,\gamma\right)=\lim_{n\rightarrow\infty}E\left[n^{-1}\Upsilon_{a}^{\prime}C\left(\theta,\gamma\right)\Upsilon_{b}\right]$
is bounded and continuously differentiable in $\theta$ and $\gamma$.
In Kuersteiner, Prucha, and Zeng (2023) we show uniform convergence under more primitive conditions in a closely related situation while directly assuming it here in Assumption 9 to save space. Without further assumptions on cross-sectional dependence, high level assumptions about the convergence of sample averages such as $n^{-1}\underline{X}^{\prime}\underline{X}$ are required irrespective. The next theorem establishes consistency of the GMM estimator.
Theorem 4.1.
Let Assumptions 1- 9 be satisfied. Assume that $\bar{\gamma}_{n}\stackrel{{\scriptstyle p}}{{\rightarrow}}\gamma$ for some sequence $\bar{\gamma}_{n}$ and some $\gamma\in\Gamma.$ Then,
(i) for $\hat{\theta}\left(\gamma\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\gamma\right)$ with $Q_{n}\left(\theta,\gamma\right)$ defined in (18) it follows that $\hat{\theta}\left(\bar{\gamma}_{n}\right)\stackrel{{\scriptstyle p}}{{\rightarrow}}\theta_{0}$.
(ii) for $\tilde{f}_{1}$ and $\tilde{\delta}$ defined in (19) and $\tilde{\rho}$ defined in (22) it follows that $\tilde{f}_{1}\stackrel{{\scriptstyle p}}{{\rightarrow}}f_{1,0}$, $\tilde{\delta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\delta_{0}$ and $\tilde{\rho}\stackrel{{\scriptstyle p}}{{\rightarrow}}\rho_{0}.$
4.3 Asymptotic Normality of Parameter Estimators
In the following we develop an asymptotically justified inference theory for our proposed GMM estimator. For that purpose we first discuss consistent estimators for the variance parameters $\gamma=(\gamma_{1},...,\gamma_{J})^{\prime}$, which are treated as nuisance parameters in the objective function of our GMM estimator.
Assumption 10.
Let $\omega_{j}\equiv\lim{}_{n\rightarrow\infty}n^{-1}\sum_{c=1}^{C}n_{c}1(\tau_{c}=j)$ and assume that $\omega_{j}$ exists and $\omega_{j}>0$ for all $j=1,...,J$.
Assumption 10 guarantees that there are enough data to estimate $\gamma_{j}^{2}$ for each $j$ by requiring that the fraction $\omega_{j}$ of all students in classrooms of type $j$ is asymptotically non-negligible for all types of classrooms.
The first step inefficient but consistent estimators $\tilde{f}_{1}$, $\tilde{\delta}$ and $\tilde{\rho}$ defined in (19) and (22) can be used to obtain consistent estimates of the variance parameters $\gamma.$ Form the residuals
$\tilde{\epsilon}=\epsilon^{+}(\tilde{\theta})=\left(I_{n}+\tilde{\rho}M\right)^{-1}\left(y_{1}-y_{2}\tilde{f}_{1}-\underline{X}\tilde{\delta}\right)$ (34)
and organize the residuals by classrooms $c$ as $\tilde{\epsilon}=\left(\tilde{\epsilon}_{1}^{\prime},....,\tilde{\epsilon}_{C}^{\prime}\right)^{\prime}$. Then construct the variance estimators
$\hat{\gamma}_{j}^{2}=\left(N_{j}-p_{x}-1\right)^{-1}\sum_{c=1}^{C}\tilde{\epsilon}_{c}^{\prime}\tilde{\epsilon}_{c}1\left(\tau_{c}=j\right),$ (35)
where $N_{j}=\sum_{c=1}^{C}n_{c}1\left(\tau_{c}=j\right)$ is the number of students in classrooms of type $j$, and where $p_{x}$ denotes the number of columns in $\underline{X}$. Now set
$\hat{\gamma}=\left(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{J}\right)^{\prime}.$ (36)
Efficient GMM estimators can now be formed by plugging $\hat{\gamma}$ into $u^{+}(\theta,\gamma)$ defined in (14). Recall that $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}(\gamma_{\tau_{c}}^{2}I_{c}).$ In the following we denote with $K_{u}$ a generic finite constant (which is taken, w.o.l.o.g., to be greater than one) and which is invariant over $t=1,2$, $i=1,\ldots,n$, $n\geq 1$. The central limit theorem in Proposition 3 of Kuersteiner and Prucha (2020) can be applied to obtain distributional approximations for estimators based on $u^{+}$. The central limit theorem of Kuersteiner and Prucha (2020) requires that $E\left[\left|u_{it}u_{is}\right|^{1+\eta_{u}}|\mathcal{Z}_{n}\right]\leqslant K_{u}$ with $\eta_{u}>0$ for $t,s=1,2\text{ }$ and all $i$. This condition is implied by Assumption 2, as Lemma B.3 in the appendix demonstrates. When setting
$\tilde{\gamma}=\left(1,1,...,1\right)^{\prime}$ (37)
it follows that $\Omega(\tilde{\gamma})=I_{n}$. The feasible efficient GMM estimator for $\theta_{0}=\left(\rho_{0},f_{1,0},\delta_{0}^{\prime}\right)^{\prime}$ is defined as $\hat{\theta}\left(\hat{\gamma}\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\hat{\gamma}\right)$ where $Q_{n}\left(\theta,\gamma\right)$ is given in (18) and the inefficient counterpart $\hat{\theta}\left(\tilde{\gamma}\right)$ is defined in an analogous way. The next theorem establishes the limiting distribution of $\hat{\theta}\left(\tilde{\gamma}\right)$ and $\hat{\theta}\left(\hat{\gamma}\right).$ The limiting distribution of $\hat{\theta}\left(\tilde{\gamma}\right)$ for $\tilde{\gamma}$ defined in (37) provides a distributional result for an inefficient GMM estimator. The limiting distribution of the first step estimators defined in (19) and (22) can be deduced from the theorem by working out the limiting distribution of $m_{\epsilon}\left(\theta\right).$ We omit the details.
Theorem 4.2.
Let Assumptions 1-10 hold. Let $\hat{\gamma}$ be as defined in (36) and $\tilde{\gamma}$ as defined in (37). For $\gamma_{n}\rightarrow_{p}\gamma$ let $\operatorname{plim}_{n\rightarrow\infty}n^{-1/2}\partial m_{n}\left(\hat{\theta}\left(\gamma_{n}\right),\gamma_{n}\right)/\partial\theta=G(\gamma)$ and $\Xi=\operatorname{plim}_{n\rightarrow\infty}\Xi_{n}$, where $\Xi_{n}$ is defined in (17). Then $\hat{\gamma}$$\rightarrow_{p}\gamma_{0}$, and for $\gamma_{n}=\hat{\gamma}$ or $\gamma_{n}=\tilde{\gamma}$,
$\sqrt{n}\left(\hat{\theta}\left(\gamma_{n}\right)-\theta_{0}\right)\rightarrow_{d}N(0,\Psi_{\gamma})$
where the limiting variance covariance matrix $\Psi_{\gamma}$ has the form
$\Psi_{\gamma}=\left(G^{\prime}\Xi^{-1}G\right)^{-1}G^{\prime}\Xi^{-1}V_{\gamma}\Xi^{-1}G\left(G^{\prime}\Xi^{-1}G\right)^{-1}.$
with $V_{\gamma}$ defined in Assumption 8 and $G$ is the shorthand notation for $G(\gamma)$. When $\gamma_{n}=\hat{\gamma}$ then $V_{\gamma}=\Xi$ and $\Psi_{\gamma}=\left(G(\gamma_{0})^{\prime}\Xi^{-1}G(\gamma_{0})\right)^{-1}$.
In Appendix D we provide explicit formulas for the derivatives $G$. Together with the expressions for $\Xi_{n}$ it is then easy to construct sample based estimators $\hat{G}$ and $\hat{\Xi}$. Letting $\hat{\Psi}_{\gamma}=\left(\hat{G}^{\prime}\hat{\Xi}^{-1}\hat{G}\right)^{-1}$ it follows from a standard Slutsky argument that $\hat{\Psi}_{\gamma}^{-1/2}\sqrt{n}\left(\hat{\theta}\left(\hat{\gamma}\right)-\theta_{0}\right)\rightarrow_{d}N\left(0,I\right).$
5 Conclusions
We identify peer effects using differential scores of closely related subject tests. Our analysis and identification strategy is based on a variance decomposition that distinguishes between test invariant components and unobserved idiosyncratic effects that are not predictable using concurrent test performance in related tests. The additional parametric assumption of linear education production functions and peer effects allows to identify a common endogenous peer effects parameter. Our method depends on the availability of closely related, yet separate test results but has the advantage of being robust to non-random group assignment. We prove identification of model parameters under conditions that are typical in the linear instrumental variables literature. These results depend on a careful exploration of the parametric structure of linear peer effects models. Our empirical results for Project STAR data and classrooms in kindergarten to the third grade show highly significant peer effects, but smaller in magnitude than those found in a number of studies that have looked at the same data.
Appendix A Tables
Table 1: Summary Statistics
Grade K Grade 1 Grade 2 Grade 3
mean sd mean sd mean sd mean sd
indiviudal characteristics
free lunch 0.560 0.496 0.588 0.492 0.592 0.491 0.578 0.494
black 0.326 0.469 0.327 0.469 0.347 0.476 0.333 0.471
girl 0.486 0.500 0.480 0.500 0.484 0.500 0.480 0.500
age 6.011 0.349 7.158 0.492 8.242 0.557 9.285 0.586
teacher characteristics
black teacher 0.165 0.371 0.174 0.379 0.204 0.403 0.209 0.406
master degree 0.347 0.476 0.346 0.476 0.373 0.484 0.442 0.497
years of experience 9.258 5.809 11.633 8.937 13.145 8.655 13.933 8.615
SAT test scores
math 485.377 47.698 530.528 43.109 580.613 44.574 617.970 39.841
read 436.725 31.706 520.787 55.192 583.935 46.043 615.422 38.563
list 537.475 33.140 567.487 33.674 595.476 34.908 624.119 32.265
word 434.179 36.759 513.436 53.316 582.986 50.658 610.136 45.041
Sample Size
number of students 6325 6829 6840 6801
number of classes 325 339 340 336
# small 127 124 133 140
# regular 99 115 100 89
# regular/aide 99 100 107 107
number of schools 79 76 75 75
  • 1. Mean and standard deviations of student characteristics, teacher charcteristics, and SAT scores, as well as number of students and classes in each grade.
Table 2: Rank Correlation between Scores
Rank Correlation between Tests Pseudo R squared of 2SLS
$y_{1}$ $y_{2}$ $y_{1},y_{2}$ $Q_{X}y_{1},Q_{X}y_{2}$ $y_{1},\hat{y}_{1}$ $y_{2},\hat{y}_{2}$
Grade K
list math 0.657 0.584 0.728 0.638
list read 0.621 0.554 0.686 0.691
list word 0.558 0.477 0.637 0.610
math read 0.741 0.672 0.697 0.763
math word 0.656 0.579 0.643 0.727
read word 0.908 0.881 0.915 0.909
Grade 1
list math 0.721 0.654 0.766 0.716
list read 0.624 0.534 0.704 0.601
list word 0.570 0.483 0.645 0.562
math read 0.737 0.669 0.782 0.753
math word 0.686 0.619 0.726 0.715
read word 0.932 0.905 0.936 0.940
Grade 2
list math 0.678 0.611 0.713 0.705
list read 0.667 0.601 0.709 0.699
list word 0.597 0.533 0.639 0.639
math read 0.724 0.673 0.756 0.758
math word 0.625 0.564 0.659 0.670
read word 0.884 0.848 0.892 0.894
Grade 3
list math 0.649 0.610 0.689 0.673
list read 0.652 0.623 0.690 0.683
list word 0.524 0.478 0.581 0.552
math read 0.741 0.706 0.759 0.769
math word 0.683 0.643 0.712 0.708
read word 0.890 0.869 0.901 0.895
  • 1. The left panel is Spearman’s rank correlation between $y_{1}$, $y_{2}$ and between $Q_{X}y_{1}$, $Q_{X}y_{2}$, where $Q_{X}=I-\underline{X}(\underline{X}^{\prime}\underline{X})^{-1}\underline{X}^{\prime}$, $y_{1}$ and $y_{2}$ are raw SAT scores specified in Columns 1 and 2, $\underline{X}$ are the complete set of control variables, including school fixed effects, class type fixed effects, student characteristics, teacher characteristics and peer characteristics.
  • 2. The right panel is pseudo $R^{2}$, i.e., Spearman’s rank correlation between $y_{1}$ and $\hat{y}_{1}$ in Column 3, and between $y_{2}$ and $\hat{y}_{2}$ in Column 4. Here $\hat{y}_{1}=\tilde{f}_{1}y_{2}+\text{$\underline{X}$}\tilde{\delta}$, with $\tilde{f_{1}}$ and $\tilde{\delta}$ defined in (20) and (21), $\underline{X}$ is the complete set of controls as described in note 1. That is, $\hat{y}_{1}$ is the predicted value of $y_{1}$ from 2SLS for $y_{1}=f_{1}y_{2}+\underline{X}\delta+\epsilon$ with $y_{2}$ instrumented by the constant term, $\hat{y}_{2}$ is defined in a similar manner by reversing the order of $y_{1},y_{2}$.
Table 3: Explanatory Power of Lagged Scores
Estiamtes for lagged scores $R^{2}$ w/wo lagged scores
lagged score Grade 1 Grade 2 Grade 3 Grade 1 Grade 2 Grade 3
Dependent Variable: Listening Score
N.A. 0.256 0.297 0.236
list 0.620(0.019) 0.707(0.014) 0.687(0.013) 0.515 0.617 0.599
math 0.362(0.012) 0.465(0.013) 0.394(0.011) 0.435 0.488 0.419
read 0.476(0.020) 0.303(0.009) 0.386(0.011) 0.410 0.448 0.424
word 0.355(0.016) 0.272(0.010) 0.296(0.010) 0.381 0.409 0.378
Dependent Variable: Mathematics Score
N.A. 0.284 0.272 0.258
list 0.616(0.021) 0.679(0.019) 0.606(0.017) 0.454 0.464 0.444
math 0.549(0.015) 0.780(0.016) 0.697(0.013) 0.537 0.614 0.630
read 0.655(0.026) 0.455(0.011) 0.554(0.013) 0.464 0.496 0.513
word 0.505(0.019) 0.421(0.014) 0.396(0.012) 0.442 0.455 0.428
Dependent Variable: Reading Score
N.A. 0.290 0.294 0.246
list 0.592(0.026) 0.662(0.020) 0.628(0.016) 0.391 0.451 0.458
math 0.592(0.019) 0.669(0.018) 0.538(0.013) 0.468 0.521 0.485
read 1.026(0.032) 0.622(0.011) 0.671(0.010) 0.544 0.648 0.644
word 0.809(0.024) 0.582(0.013) 0.512(0.010) 0.522 0.588 0.544
Dependent Variable: Word Study Skills Score
N.A. 0.244 0.250 0.229
list 0.509(0.028) 0.645(0.022) 0.579(0.020) 0.333 0.377 0.363
math 0.511(0.020) 0.645(0.020) 0.579(0.016) 0.399 0.426 0.441
read 0.860(0.033) 0.668(0.012) 0.737(0.013) 0.453 0.588 0.594
word 0.686(0.025) 0.631(0.014) 0.547(0.012) 0.435 0.539 0.488
  • 1. Each panel represents regressions of a specific score type against a comprehensive set of control variables (including school fixed effects, class type fixed effects, student, teacher, and peer characteristics in the main specification and a constant term), along with possible own and peer average lagged scores.
  • 2. The lagged score type is indicated in column 1 and highlighted in bold if it matches the dependent variable type. N.A. means no lagged scores are controlled for.
  • 3. The left panel displays coefficients and standard deviations (in parentheses) of lagged scores for grade 1 to 3. The right panel reports the $R^{2}$ of the regression, where the first row in each panel representing the $R^{2}$ when no lagged scores are included.
  • 4. Coefficients of peer average lagged scores, typically small and statistically insignificant, are omitted for conciseness.
Table 4: Pseudo $R^{2}$: $1-var(u_{1}-f_{1}u_{2})/var(y_{1}-f_{1}y_{2})$
Small Classes Regualr Classes(w/wo Aide)
$y_{1}$ $y_{2}$ (1) (2) (3) (4) (1) (2) (3) (4)
Grade K
list math 0.179 0.180 0.206 0.216 0.177 0.179 0.201 0.208
list read 0.143 0.144 0.153 0.150 0.145 0.146 0.150 0.147
list word 0.124 0.125 0.132 0.129 0.132 0.133 0.136 0.134
math read 0.148 0.148 0.177 0.219 0.126 0.127 0.155 0.192
math word 0.138 0.139 0.164 0.200 0.128 0.129 0.154 0.184
read word 0.104 0.104 0.107 0.113 0.140 0.140 0.142 0.146
Grade 1
list math 0.138 0.140 0.143 0.144 0.113 0.115 0.123 0.124
list read 0.131 0.133 0.154 0.163 0.113 0.116 0.148 0.161
list word 0.105 0.106 0.122 0.132 0.068 0.071 0.094 0.108
math read 0.105 0.105 0.130 0.137 0.070 0.071 0.119 0.132
math word 0.092 0.092 0.114 0.118 0.058 0.059 0.101 0.109
read word 0.088 0.088 0.092 0.090 0.093 0.093 0.098 0.093
Grade 2
list math 0.116 0.117 0.123 0.119 0.104 0.106 0.118 0.109
list read 0.087 0.087 0.144 0.141 0.083 0.084 0.125 0.125
list word 0.084 0.084 0.131 0.132 0.070 0.072 0.107 0.109
math read 0.092 0.092 0.134 0.137 0.086 0.087 0.110 0.109
math word 0.089 0.089 0.124 0.126 0.081 0.081 0.106 0.107
read word 0.064 0.064 0.077 0.082 0.065 0.065 0.074 0.078
Grade 3
list math 0.135 0.136 0.147 0.144 0.085 0.087 0.092 0.089
list read 0.073 0.074 0.133 0.133 0.024 0.027 0.077 0.078
list word 0.090 0.090 0.132 0.131 0.030 0.033 0.065 0.066
math read 0.083 0.083 0.107 0.106 0.050 0.050 0.085 0.086
math word 0.085 0.085 0.107 0.106 0.039 0.040 0.071 0.070
read word 0.093 0.093 0.092 0.096 0.033 0.034 0.035 0.038
school FE Y Y Y Y Y Y Y Y
class type FE Y Y Y Y Y Y
stu&tch char Y Y Y Y
peer char Y Y
  • 1. The table reports one minus the ratio of the variance of $u_{1}-f_{1}u_{2}$ to that of $y_{1}-f_{1}y_{2}$ for small classes (in the left panel) and regular (with or without aide) classes (in the right panel). All models allow for heteroscedasticity across small and regular classes.
  • 2. The scores are raw SAT scores in listening (list), mathematics (math), reading (read), word study skills (word).
  • 3. As indicated at the bottom of the table, the models may control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.
Table 5: Standard Deviations of $Q_{x}y$ and $Q_{x}(y_{1}-f_{1}y_{2})$
Small Classes Regular Classes
$y_{1}$ $y_{2}$ (1) (2) (3) (4) (1) (2) (3) (4)
Grade K
$Q_{x}y$ list 75.810 75.122 37.769 28.508 78.091 76.238 38.954 29.681
$Q_{x}y$ math 75.456 74.762 44.444 39.421 80.333 78.669 48.468 43.319
$Q_{x}y$ read 62.748 62.169 33.699 26.827 66.321 64.827 35.534 29.143
$Q_{x}y$ word 65.196 64.629 38.017 31.883 68.723 67.294 39.693 33.916
$Q_{x}(y_{1}-f_{1}y_{2})$ list math 35.462 35.555 38.504 41.091 38.064 38.131 41.515 44.471
$Q_{x}(y_{1}-f_{1}y_{2})$ list read 30.802 30.801 30.488 27.544 31.701 31.697 31.347 28.060
$Q_{x}(y_{1}-f_{1}y_{2})$ list word 36.916 36.919 36.311 33.052 37.006 37.015 36.336 32.868
$Q_{x}(y_{1}-f_{1}y_{2})$ math read 30.874 30.867 30.338 30.583 31.934 31.927 31.556 32.405
$Q_{x}(y_{1}-f_{1}y_{2})$ math word 35.724 35.696 34.235 32.883 35.731 35.707 34.486 33.819
$Q_{x}(y_{1}-f_{1}y_{2})$ read word 14.055 14.059 13.883 14.896 14.625 14.632 14.422 15.448
Grade 1
$Q_{x}y$ list 78.430 77.682 44.281 29.291 82.153 80.072 43.729 32.263
$Q_{x}y$ math 77.259 76.454 47.817 36.254 81.842 79.872 48.874 39.564
$Q_{x}y$ read 83.720 82.984 57.172 46.261 88.372 86.584 58.298 49.448
$Q_{x}y$ word 81.314 80.624 57.574 46.620 85.880 84.192 57.174 47.750
$Q_{x}(y_{1}-f_{1}y_{2})$ list math 29.766 29.822 29.377 29.449 30.662 30.730 30.227 30.285
$Q_{x}(y_{1}-f_{1}y_{2})$ list read 42.790 42.812 38.711 40.458 43.672 43.660 39.169 40.844
$Q_{x}(y_{1}-f_{1}y_{2})$ list word 44.810 44.856 40.579 42.694 45.258 45.272 40.989 42.897
$Q_{x}(y_{1}-f_{1}y_{2})$ math read 35.658 35.595 32.807 34.226 37.625 37.547 34.073 35.289
$Q_{x}(y_{1}-f_{1}y_{2})$ math word 38.230 38.195 35.500 36.668 39.197 39.143 36.076 36.874
$Q_{x}(y_{1}-f_{1}y_{2})$ read word 21.646 21.649 21.648 21.786 22.530 22.533 22.493 22.616
Grade 2
$Q_{x}y$ list 92.075 90.731 48.088 29.859 85.236 82.708 45.789 32.371
$Q_{x}y$ math 90.878 89.540 53.323 38.275 85.553 83.117 52.234 41.389
$Q_{x}y$ read 93.586 92.295 57.595 39.409 87.831 85.406 54.519 41.740
$Q_{x}y$ word 96.864 95.555 62.274 44.613 90.149 87.782 59.070 46.967
$Q_{x}(y_{1}-f_{1}y_{2})$ list math 32.791 32.843 31.327 28.848 34.666 34.691 32.803 30.147
$Q_{x}(y_{1}-f_{1}y_{2})$ list read 34.210 34.242 29.824 28.460 34.229 34.216 29.939 28.515
$Q_{x}(y_{1}-f_{1}y_{2})$ list word 40.464 40.509 34.546 31.816 39.898 39.907 34.124 31.404
$Q_{x}(y_{1}-f_{1}y_{2})$ math read 31.591 31.571 29.607 30.029 32.541 32.523 31.102 31.550
$Q_{x}(y_{1}-f_{1}y_{2})$ math word 39.107 39.083 35.878 35.793 39.412 39.388 36.700 36.676
$Q_{x}(y_{1}-f_{1}y_{2})$ read word 23.891 23.894 23.273 22.593 24.312 24.318 23.726 23.043
Grade 3
$Q_{x}y$ list 100.934 98.926 46.812 28.881 81.116 78.549 45.571 31.663
$Q_{x}y$ math 99.438 97.593 51.961 35.145 82.400 79.955 50.671 37.816
$Q_{x}y$ read 102.223 100.223 52.813 33.948 85.025 82.325 51.772 38.114
$Q_{x}y$ word 102.419 100.495 56.860 39.320 86.959 84.387 56.262 43.608
$Q_{x}(y_{1}-f_{1}y_{2})$ list math 29.971 29.941 28.331 27.089 30.284 30.259 28.579 27.304
$Q_{x}(y_{1}-f_{1}y_{2})$ list read 29.166 29.145 26.466 25.715 29.828 29.770 26.962 26.078
$Q_{x}(y_{1}-f_{1}y_{2})$ list word 38.217 38.212 34.066 33.314 39.272 39.206 34.788 33.930
$Q_{x}(y_{1}-f_{1}y_{2})$ math read 27.010 27.013 26.326 26.477 26.807 26.801 25.922 26.045
$Q_{x}(y_{1}-f_{1}y_{2})$ math word 32.785 32.796 31.350 31.918 32.844 32.832 31.064 31.683
$Q_{x}(y_{1}-f_{1}y_{2})$ read word 20.631 20.634 20.090 20.639 21.111 21.106 20.485 21.043
  • 1. Standard deviations of $Q_{x}y$ and $Q_{x}(y_{1}-f_{1}y_{2})$ for small (on the left) and regular classes (on the right). For $Q_{x}y$, the test score $y$ is specified in $y_{1}$. For $Q_{x}(y_{1}-f_{1}y_{2})$, $y_{1}$ and $y_{2}$ are specified in columns 2 and 3. The control variables for models (1)-(4) are the same as those for the main models in Tables 6 and 7. The estimates of $f_{1}$ are from our efficient GMM estimations for each specifications and presented in Tables 6.
Table 6: Estimates of $f_{1}$
$y_{1}$ $y_{2}$ (1) (2) (3) (4)
Grade K
list math 1.093(0.011) 1.096(0.010) 1.201(0.027) 1.283(0.243)
list read 1.231(0.008) 1.231(0.008) 1.217(0.021) 1.011(0.109)
list word 1.224(0.007) 1.225(0.006) 1.201(0.023) 1.048(0.128)
math read 1.126(0.005) 1.124(0.004) 1.014(0.018) 0.771(0.140)
math word 1.120(0.008) 1.118(0.008) 0.999(0.021) 0.793(0.152)
read word 0.995(0.004) 0.995(0.005) 0.984(0.009) 1.042(0.055)
Grade 1
list math 1.065(0.009) 1.068(0.008) 1.054(0.012) 1.057(0.070)
list read 1.060(0.008) 1.061(0.008) 0.973(0.014) 1.023(0.077)
list word 1.071(0.009) 1.073(0.009) 0.977(0.015) 1.038(0.087)
math read 0.995(0.005) 0.993(0.006) 0.916(0.012) 0.969(0.072)
math word 1.007(0.004) 1.006(0.005) 0.929(0.012) 0.973(0.084)
read word 1.008(0.005) 1.008(0.006) 1.014(0.009) 1.026(0.045)
Grade 2
list math 1.057(0.005) 1.059(0.005) 1.007(0.012) 0.902(0.055)
list read 1.030(0.006) 1.031(0.006) 0.903(0.009) 0.842(0.050)
list word 1.023(0.006) 1.024(0.007) 0.873(0.011) 0.779(0.052)
math read 0.975(0.006) 0.974(0.006) 0.893(0.010) 0.929(0.048)
math word 0.967(0.006) 0.967(0.007) 0.863(0.010) 0.862(0.055)
read word 0.992(0.004) 0.993(0.004) 0.967(0.007) 0.929(0.030)
Grade 3
list math 1.028(0.002) 1.027(0.002) 0.961(0.008) 0.898(0.055)
list read 1.014(0.003) 1.013(0.003) 0.915(0.008) 0.869(0.045)
list word 1.027(0.004) 1.026(0.004) 0.903(0.010) 0.873(0.053)
math read 0.986(0.003) 0.986(0.003) 0.951(0.008) 0.965(0.047)
math word 0.999(0.004) 0.999(0.004) 0.940(0.009) 0.970(0.051)
read word 1.013(0.002) 1.013(0.002) 0.988(0.005) 1.016(0.032)
school FE Y Y Y Y
class type FE Y Y Y
stu&tch char Y Y
peer char Y
  • 1. Estimates and standard errors (in the parentheses) for $\rho$. Standard errors are clustered at the classroom level. All models allow for heteroscedasticity across small and regular classes. Estimates are adjusted for missing observations.
  • 2. The scores are raw SAT scores in listening (list), mathematics (math), reading (read), word study skills(word).
  • 3. As indicated at the bottom of the table, the models may control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.
Table 7: Estimates of $\rho$
$y_{1}$ $y_{2}$ (1) (2) (3) (4)
Grade K
list math 0.721(0.079) 0.691(0.073) 0.694(0.075) 0.688(0.078)
list read 0.434(0.079) 0.415(0.076) 0.407(0.078) 0.377(0.083)
list word 0.398(0.066) 0.378(0.065) 0.373(0.065) 0.359(0.068)
math read 0.492(0.065) 0.485(0.065) 0.519(0.066) 0.575(0.076)
math word 0.517(0.067) 0.513(0.067) 0.538(0.070) 0.578(0.081)
read word 0.452(0.078) 0.449(0.078) 0.443(0.080) 0.424(0.075)
Grade 1
list math 0.627(0.069) 0.587(0.073) 0.579(0.073) 0.578(0.074)
list read 0.322(0.051) 0.277(0.049) 0.256(0.048) 0.218(0.046)
list word 0.385(0.054) 0.340(0.052) 0.308(0.048) 0.257(0.047)
math read 0.441(0.063) 0.434(0.064) 0.434(0.061) 0.360(0.061)
math word 0.487(0.066) 0.480(0.065) 0.469(0.062) 0.388(0.064)
read word 0.178(0.054) 0.178(0.054) 0.190(0.054) 0.193(0.053)
Grade 2
list math 0.637(0.073) 0.621(0.072) 0.554(0.071) 0.525(0.068)
list read 0.568(0.073) 0.553(0.071) 0.571(0.074) 0.561(0.074)
list word 0.537(0.070) 0.521(0.068) 0.512(0.070) 0.497(0.069)
math read 0.481(0.071) 0.479(0.071) 0.475(0.066) 0.444(0.066)
math word 0.493(0.071) 0.492(0.071) 0.478(0.066) 0.459(0.065)
read word 0.399(0.065) 0.398(0.064) 0.363(0.063) 0.348(0.064)
Grade 3
list math 0.747(0.088) 0.728(0.087) 0.714(0.086) 0.703(0.090)
list read 0.606(0.091) 0.572(0.088) 0.613(0.096) 0.623(0.101)
list word 0.512(0.095) 0.482(0.091) 0.518(0.095) 0.522(0.097)
math read 0.462(0.078) 0.460(0.078) 0.452(0.082) 0.440(0.086)
math word 0.371(0.074) 0.366(0.072) 0.370(0.074) 0.357(0.078)
read word 0.391(0.078) 0.387(0.077) 0.377(0.077) 0.379(0.076)
school FE Y Y Y Y
class type FE Y Y Y
stu&tch char Y Y
peer char Y
  • 1. Estimates and standard errors (in the parentheses) for $\rho$. Standard errors are clustered at the classroom level. All models allow for heteroscedasticity across small and regular classes. Estimates are adjusted for missing observations.
  • 2. The scores are raw SAT scores in listening (list), mathematics (math), reading (read), word study skills(word).
  • 3. As indicated at the bottom of the table, the models may control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.
Table 8: Estimates of Exogenous Peer Effects
$y_{1}$ $y_{2}$ age black(pct.) free lunch(pct.) girl(pct.)
Grade K
list math -4.290(15.730) -0.196(0.244) 0.117(0.099) -0.177(0.102)*
list read 13.784(6.951)** 0.093(0.103) 0.024(0.050) -0.031(0.050)
list word 10.097(8.239) 0.066(0.134) 0.014(0.058) 0.020(0.058)
math read 15.222(9.027)* 0.238(0.159) -0.076(0.065) 0.111(0.063)*
math word 12.705(9.930) 0.200(0.156) -0.082(0.066) 0.159(0.064)**
read word -4.298(3.626) -0.029(0.062) -0.005(0.024) 0.070(0.027)**
Grade 1
list math -0.268(5.134) -0.016(0.095) -0.049(0.068) 0.013(0.079)
list read -1.922(6.076) -0.125(0.124) 0.129(0.084) -0.354(0.095)***
list word -2.212(6.848) -0.186(0.147) 0.160(0.095)* -0.454(0.108)***
math read -2.392(5.540) -0.099(0.103) 0.179(0.076)** -0.330(0.096)***
math word -1.156(6.399) -0.161(0.118) 0.196(0.086)** -0.416(0.107)***
read word -0.719(3.423) -0.033(0.065) 0.038(0.043) -0.037(0.050)
Grade 2
list math 7.639(4.183)* 0.166(0.123) -0.045(0.071) 0.120(0.089)
list read 4.909(4.197) 0.109(0.104) 0.127(0.072)* 0.063(0.082)
list word 7.808(4.551)* -0.001(0.102) 0.150(0.079)* 0.149(0.086)*
math read -2.902(3.933) -0.062(0.107) 0.187(0.068)*** -0.064(0.084)
math word -0.032(4.720) -0.183(0.130) 0.213(0.082)*** 0.025(0.096)
read word 2.977(2.691) -0.134(0.088) 0.025(0.050) 0.105(0.054)*
Grade 3
list math 5.041(4.171) 0.187(0.094)** -0.111(0.055)** -0.059(0.066)
list read 4.026(3.618) 0.030(0.084) -0.079(0.048)* -0.055(0.059)
list word 2.712(4.286) 0.061(0.097) -0.045(0.058) -0.063(0.073)
math read -1.133(3.779) -0.162(0.087)* 0.032(0.047) 0.003(0.054)
math word -2.395(4.101) -0.129(0.104) 0.073(0.057) -0.009(0.058)
read word -2.190(2.593) 0.030(0.062) 0.037(0.040) -0.025(0.045)
  • 1. Estimates and standard errors (in the parentheses) for exogenous peer effects of peer’s age, race, free lunch status and gender. Coefficients for dummies black, free lunch and girl are divided by 100 so they can be interpreted as the impact of one percentage point change in these variables. Note that these four variables are all test-invariant personal characteristics. The reported coefficients are $\rho(\beta_{v1}^{p}-f_{1}\beta_{v2}^{p})$.
  • 2. Standard errors are clustered at the classroom level. All models allow for heteroscedasticity across small and regular classes. Estimates are adjusted for missing observations.
  • 3. The scores are raw SAT scores in listening (list), mathematics (math), reading (read), word study skills(word).
  • 4. All specifications control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.
Appendix B Auxiliary Lemmas
Before proving the main results, we establish a number of preliminary lemmas. Note that many matrices in this paper, e.g., $M$, $\Omega(\gamma)$, $\Sigma_{t}$ defined in Assumption 3 can be written in the form of $\operatorname{diag}_{c=1}^{C}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right),$where $I_{c}^{\ast}=I_{c}-\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}/n_{c}$, $J_{c}^{\ast}=\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}/n_{c}$ are the residual projection matrix and the projection matrix onto $\mathbf{1}_{c}$ respectively. To see this note that
$M$ $=\operatorname{diag}_{c=1}^{C}\left(-\frac{1}{n_{c}-1}I_{c}^{\ast}+J_{c}^{\ast}\right),$ (38)
$I+\rho M$ $=\operatorname{diag}_{c=1}^{C}\left(\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\left(1+\rho\right)J_{c}^{\ast}\right),$ (39)
$\Omega(\gamma)$ $=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}^{\ast}+\gamma_{\tau_{c}}^{2}J_{c}^{\ast}\right),$ (40)
$\Sigma_{t}$ $=\sigma_{0,t}^{2}\operatorname{diag}_{c=1}^{C}\left(\rho_{0,\tau_{c}}^{2}I_{c}\right)=\operatorname{diag}_{c=1}^{C}\left(\rho_{0,\tau_{c}}^{2}\sigma_{0,t}^{2}I_{c}^{\ast}+\rho_{0,\tau_{c}}^{2}\sigma_{0,t}^{2}J_{c}^{\ast}\right),\,t=1,2.$ (41)
The following lemma summarizes some well understood key properties of these matrices for the convenience of the reader. Proofs are straight forward and omitted, and can e.g., be found in Kuersteiner, Prucha, and Zeng (2023).
Lemma B.1.
The matrices $I_{c}^{\ast}$, $J_{c}^{\ast}$ and $pI_{c}^{\ast}+qJ_{c}^{\ast}$ have the following properties:
(i) The matrices $I_{c}^{\ast}$, $J_{c}^{\ast}$ are symmetric, idempotent, orthogonal, $I_{c}^{*}+J_{c}^{*}=I$ and $I_{c}^{\ast}\mathbf{1}_{c}=0$, $J_{c}^{\ast}\mathbf{1}_{c}=\mathbf{1}_{c}$.
(ii) $\operatorname{det}\left(pI_{c}^{\ast}+qJ_{c}^{\ast}\right)=p^{(n_{c}-1)}q$.
(iii) The eigenvalues of $pI_{c}^{\ast}+qJ_{c}^{\ast}$ are $p$ ($n_{c}-1$ times) and $q$ (once).
(iv) For $p_{1}$, $p_{2}$, $q_{1}$ and $q_{2}$ arbitrary constants, $(p_{1}I_{c}^{\ast}+q_{1}J_{c}^{\ast})(p_{2}I_{c}^{\ast}+q_{2}J_{c}^{\ast})=(p_{1}p_{2}I_{c}^{\ast}+q_{1}q_{2}J_{c}^{\ast})$, and thus multiplication of matrices of the form $pI_{c}^{\ast}+qJ_{c}^{\ast}$ is commutative and associative.
(v) If $p\neq 0,$ $q\neq 0$, $(pI_{c}^{\ast}+qJ_{c}^{\ast})$ is invertible with $(pI_{c}^{\ast}+qJ_{c}^{\ast})^{-1}=p^{-1}I_{c}^{\ast}+q^{-1}J_{c}^{\ast}$.
(vi) $\operatorname{tr}(pI_{c}^{\ast}+qJ_{c}^{\ast})=p(n_{c}-1)+q$.
(vii) Both $I_{c}^{\ast}$ and $J_{c}^{\ast}$ have row and column sum uniformly bounded in absolute value. When $p$ and $q$ are bounded in absolute value, the row and column sums of $pI_{c}^{\ast}+qJ_{c}^{\ast}$ are uniformly bounded in absolute value.
Corollary B.1.
Suppose Assumptions 3, 4, 5, 6, 7, and 10 hold. Then $I+\rho M$, $\Omega(\gamma)$, and $\Sigma_{t}$ are non-singular. If $S(\theta,\gamma)$ is an element or product matrix of elements from
$\left\{M,(I+\rho M),(I+\rho M)^{-1},\Omega(\gamma),\Omega(\gamma)^{-1},\Sigma_{t}^{-1},\Sigma_{t}\right\},$ (42)
then $S(\theta,\gamma)$ can be written in the form of $\operatorname{diag}_{c=1}\{p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\}$with $p_{c}$, $q_{c}$ continuously differentiable in $\theta$ and $\gamma$. The matrix $S(\theta,\gamma)A$ is measurable w.r.t $\mathcal{Z}_{n}$ The row and column sums of $S(\theta,\gamma)$ and $S(\theta,\gamma)A$ are uniformly bounded in absolute value. .
Proof.
By Assumptions 3, 4 and 6, $\left(n_{c}-1-\rho\right)/\left(n_{c}-1\right)$, $\left(1+\rho\right)$, $\gamma_{j}^{2}$, $\rho_{0,\tau_{c}}^{2}$, and $\sigma_{0,t}^{2}$ in (39), (40), (41) are positive and uniformly bounded away from 0. Hence by Lemma B.1(v), $I+\rho M$, $\Omega(\gamma)$ and $\Sigma_{t}$ are non-singular. An inspection of Equations (38)-(41) shows that all matrices in the set can be written in the form of $\operatorname{diag}_{c=1}^{C}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right),$ with $p_{c}$ and $q_{c}$ continuously differentiable in $\theta$ and $\gamma$ and uniformly bounded. Also note that by Assumption 5, the row and column sums of $A$ are uniformly bounded in absolute value. Utilizing remark A.1 in Kelejian and Prucha (2004) the corollary thus follows from Lemma B.1. ∎
We use these properties frequently in the derivations that follow.
Lemma B.2.
Suppose Assumptions 3, 4, 5, and 6 hold. Consider the block diagonal matrix $\mathcal{M}\left(\rho\right)=\operatorname{diag}{}_{c=1}^{C}\left(\mathcal{M}_{c}\left(\rho\right)\right)$ with $\mathcal{M}_{c}\left(\rho\right)=\left(I_{c}+\rho M_{c}\right)^{-1}\left(I_{c}+\rho_{0}M_{c}\right)$. Then for all $\gamma\in\Gamma$,
$\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0$
and
$\lim_{n\rightarrow\infty}\frac{1}{n}\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0$
each has only one solution for $\rho\in\left(-1,1\right)$. This solution is given by $\rho=\rho_{0}$.
Proof.
Let $D(\rho,\gamma)=\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}$ and observe that $D(\rho,\gamma)=\operatorname{diag}{}_{c=1}^{C}\left(D_{c}(\rho,\gamma)\right)$ with
$D_{c}(\rho,\gamma)$ $=$ $\mathcal{M}_{c}\left(\rho\right)\left(\gamma_{\tau_{c}}^{2}I_{c}\right)^{-1/2}A_{c}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)^{-1/2}\mathcal{M}_{c}\left(\rho\right)\left(\gamma_{0,\tau_{c}}^{2}I_{c}\right)$
$=$ $\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right).$
Noting that $\operatorname{tr}\left(A_{c}\right)=0$ and hence $\operatorname{tr}\left(A_{c}I_{c}^{\ast}\right)=\operatorname{tr}\left(A_{c}(I-J_{c}^{\ast})\right)=-\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$, and in light of Lemma B.1
$\mathcal{M}_{c}\left(\rho\right)$ $=$ $\left(I_{c}+\rho M_{c}\right)^{-1}\left(I_{c}+\rho_{0}M_{c}\right)=\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}I_{c}^{\ast}+\frac{1+\rho_{0}}{1+\rho}J_{c}^{\ast}.$
We have
$\operatorname{tr}\left(\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right)$ $=$ $\operatorname{tr}\left(A_{c}\left[\left(\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}\right)^{2}I_{c}^{\ast}+\left(\frac{1+\rho_{0}}{1+\rho}\right)^{2}J_{c}^{\ast}\right]\right)$
$=$ $\left[\left(\frac{1+\rho_{0}}{1+\rho}\right)^{2}-\left(\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}\right)^{2}\right]\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$
$=$ $(\rho_{0}-\rho)C(\rho,n_{c})\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$
where
$C(\rho,n_{c})=\left[\left(\frac{1+\rho_{0}}{1+\rho}\right)+\left(\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}\right)\right]\left[\frac{1}{1+\rho}+\frac{1}{n_{c}-1-\rho}\right].$
Hence
$\operatorname{tr}\left(D(\rho,\gamma)\right)=\sum_{c=1}^{C}\operatorname{tr}\left(D_{c}(\rho,\gamma)\right)=(\rho_{0}-\rho)\sum_{c=1}^{C}\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}C(\rho,n_{c})\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right).$
Clearly, at $\rho=\rho_{0}$, for all $\gamma\in\Gamma$, we have $\operatorname{tr}\left(D(\rho_{0},\gamma)\right)=0$ for all $n$.
Next observe that $\operatorname{tr}\left(A_{c}J_{c}^{*}\right)=\mathbf{1}_{c}^{{}^{\prime}}A_{c}\mathbf{1}_{c}/n_{c}\geqslant\text{$\underline{K}$}_{a}>0$ and that for any $n_{c}\geqslant 2$ and $-1<\rho<1$, both $n_{c}-1-\rho$ and $1+\rho$ are positive, and therefore $C(\rho,n_{c})$ is positive for all $\rho\in\left(-1,1\right)$. Therefore, $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]<0$ if $\rho>\rho_{0}$, $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]>0$ if $\rho<\rho_{0}$ and $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]$ has the same sign for all $c$. In all, for any value of $\gamma$ the only solution for
$\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma,\theta)^{-1/2}A\Omega(\gamma,\theta)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0$
is $\rho=\rho_{0}$. Furthermore, by Assumptions 4 and 6 there exists some constants $c_{D}$ and $c_{\gamma}$ such that for any $\rho_{0}$ in the interior of the parameter space, $C(\rho,n_{c})>c_{D}>0$ and $\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}>c_{\gamma}>0$. This implies that,
$\left|\operatorname{tr}(D(\rho))\right|\geqslant|\rho_{0}-\rho|\sum_{c=1}^{C}\text{$\underline{K}$}_{a}c_{D}c_{\gamma},$
$\left|\frac{1}{n}\operatorname{tr}(D(\rho))\right|\geqslant|\rho_{0}-\rho|\text{$\underline{K}$}_{a}c_{D}c_{\gamma}\frac{C}{n}.$
By Assumption 6, $C/n\geq 1/\bar{n}_{c}>0$ such that the conclusion follows. ∎
Lemma B.3.
Let Assumption 2 hold. Then, $u_{it}$ defined in Theorem 2.1 is a martingale difference sequence w.r.t to the filtration $\mathcal{A}_{n,\nu}$ with $\mathcal{A}_{n,0}=\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,X,z,\zeta^{*}\right)$ and
$\mathcal{A}_{n,\left(t-1\right)n+i-1}=\sigma\left(S,\alpha,X,z,\zeta^{*},\left\{u_{j,t-1}\right\}_{j=1}^{n},\left\{u_{jt}\right\}_{j=1}^{i-1}\right)\text{\,for $t=1,2;i=1,\ldots,n$}$ (43)
and where $u_{j,0}=0$ to handle the case $t=1.$ In addition, $\sup_{i,t}E\left[\left|u_{it}\right|^{4+\eta}\right]<\infty$ for some $\eta>0$.
Proof.
First note that the filtrations $\mathcal{A}_{n,\nu}$ are increasing in the sense $\mathcal{A}_{n,0}\subseteq\mathcal{A}_{n,1}....\subseteq\mathcal{A}_{n,\nu}\subseteq\mathcal{A}_{n,\nu+1}....$. By construction $u_{it}=y_{it}^{*}-\kappa_{i}f_{t}$ where $\kappa_{i}f_{t}$ is $\mathcal{A}_{n,0}$-measurable. For general $i$ and $t$ note that since $u_{it}=y_{it}^{*}-\kappa_{i}f_{t}$ and $\kappa_{i}f_{t}$ is $\mathcal{Z}_{n}^{*}$-measurable it follows that $\mathcal{B}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$. Since $y_{it}^{*}=\kappa_{i}f_{t}+u_{it}$ and $\kappa_{i}f_{t}$ is $\mathcal{Z}_{n}^{*}$-measurable it follows that $\mathcal{F}_{n,i,t}\subseteq\mathcal{B}_{n,i,t}$. Consequently $\mathcal{F}_{n,i,t}=\mathcal{B}_{n,i,t}.$ Observing that $\mathcal{A}_{n,\left(t-1\right)n+i-1}\subseteq\mathcal{B}_{n,i,t}$ it follows, using iterated expectations, that
$E\left[u_{it}|\mathcal{A}_{n,\left(t-1\right)n+i-1}\right]$ $=$ $E\left[E\left[u_{it}|\mathcal{B}_{n,i,t}\right]|\mathcal{A}_{n,\left(t-1\right)n+i-1}\right]=0$
observing that $E\left[u_{it}|\mathcal{B}_{n,i,t}\right]=E\left[y_{it}^{*}-\kappa_{i}f_{t}|\mathcal{F}_{n,i,t}\right]=0$ in light of Assumption 2. Finally observe that $u_{it}$ is measurable w.r.t. $\mathcal{A}_{n,\left(t-1\right)n+i}$
Recall that by Assumption 2 there exists a random variable $y$ such that $\left|y_{it}^{*}\right|+\left|\kappa_{i}f_{t}\right|\leq y$ with $E\left[\left|y\right|^{4+\eta}|\mathcal{A}_{n,0}\right]\leq K_{y}<\infty.$ Hence
$E\left[\left|u_{it}\right|^{4+\eta}|\mathcal{A}_{n,0}\right]$ $\leq E\left[\left(\left|y_{it}^{*}\right|+\left|\kappa_{i}f_{t}\right|\right)^{4+\eta}|\mathcal{A}_{n,0}\right]\leq E\left[\left|y\right|^{4+\eta}|\mathcal{A}_{n,0}\right]<K_{y}<\infty.$ (44)
Since the bound on the RHS is uniform in $i$ and $t$ by assumption it follows that $\sup_{i,t}E\left[\left|u_{it}\right|^{4+\eta}\right]<\infty.$
Appendix C Proofs
Proof of Theorem 2.1.
By the assumption on the boundedness of the ($4+\eta$)-th moments all considered expectations and conditional expectations exist. Since $u_{it}=y_{it}^{*}-E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]$ we have $E\left[u_{it}|\mathcal{F}_{n,i,t}\right]=0$ by construction. Observing that $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=\kappa_{i}f_{t}$ and that $\kappa_{i}$ is $\mathcal{Z}_{n}^{*}$-measurable it follows that conditional on $\mathcal{Z}_{n}^{*}$ knowledge of $y_{it}^{*}$ is equivalent to knowledge of $u_{it}$. To see this let $u_{-i,t}=\left(u_{1t},..,u_{i-1,t},u_{i+1,t},...,u_{n,t}\right)^{\prime}$ and use the notation $u_{-t}$ for the unobserved components related to the other test. Define $U=\left(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t},\kappa_{i},f_{t}\right)$ assuming for simplicity of notations that vectors are row vectors, and let $V=\left(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*},\kappa_{i},f_{t}\right)$. Then $V=g(U)$ where $g(.)$ is Borel measurable and thus $\sigma(V)\subseteq\sigma(U)$. Analogously $U=h(V)$ where $h(.)$ is Borel measurable and thus $\sigma(U)\subseteq\sigma(V)$. Consequently $\sigma(U)=\sigma(V)$. Since $\kappa_{i},f_{t}$ are measurable w.r.t. $\sigma(S,\alpha,X,\zeta^{*},z)$ it follows that
$\sigma(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t},\kappa_{i},f_{t})=\sigma(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t})$
and
$\sigma(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*},\kappa_{i},f_{t})=\sigma(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*}),$
which shows that
$\mathcal{B}_{n,i,t}=\sigma\left(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t}\right)=\sigma\left(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*}\right)=\mathcal{F}_{n,i,t}.$
Next observe that the $u_{jt}$ are $\mathcal{F}_{n,i,t}$-measurable for $i\neq j$, and hence by iterated expectations $\operatorname{Cov}\left(u_{it}u_{jt}|\mathcal{F}_{n,i,t}\right)=E\left[u_{jt}E\left(u_{it}|\mathcal{F}_{n,i,t}\right)\right]=0$. Observe further that for $s<t$ the $u_{js}$ are $\mathcal{F}_{n,i,t}$-measurable for all $i$, and hence by iterated expectations $\operatorname{Cov}\left(u_{it}u_{js}|\mathcal{F}_{n,i,t}\right)=E\left[u_{js}E\left[u_{it}|\mathcal{F}_{n,i,t}\right]\right]=0$. The corresponding claims regarding the mean and covariances of the $u_{it}$ conditional on $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$ follow immediately by iterated expectations. This proves the first part of the lemma.
To prove the second part, let $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$ be some information set. Then given that Assumption 2 holds, it follows by iterated expectations that
E$\left[y_{ii}^{*}-\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]$ $=$ $E\left[E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]|\mathcal{G}_{n,i,t}\right]-E\left[\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]=E\left[\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]-E\left[\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]=0.$
Thus by the almost sure uniqueness of conditional expectations, Assumption 2 cannot hold if $E\left[y_{it}^{*}-\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]\neq 0$. ∎
Note that Lemma 4.1, Corollary 4.1, and Lemma 4.3 can be viewed as special cases of Lemma 4.2. We therefore prove Lemma 4.2 first.
Proof of Lemma 4.2.
In light of (9) it follows that,
$y_{1}-f_{1}y_{2}=\mu^{\ast}(f_{1,0}-f_{1})+\underline{X}\delta(f_{1},\rho_{0},\beta_{0})+(I+\rho_{0}M)(u_{1}-f_{1}u_{2}),$ (45)
where $\delta(f_{1},\rho_{0},\beta_{0})$ is defined in (7) and $\delta(f_{1,0},\rho_{0},\beta_{0})=\delta_{0}$. Consequently,
$y_{1}-f_{1}y_{2}-\text{$\underline{X}$}\delta=\mu^{\ast}(f_{1,0}-f_{1})+\underline{X}\left(\delta(f_{1},\rho_{0},\beta_{0})-\delta\right)+(I+\rho_{0}M)(u_{1}-f_{1}u_{2}).$ (46)
Substitution of (46) into (14) yields
$u^{+}(\theta,\gamma)$ $=$ $\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$
×$\times\left\{\left(f_{1,0}-f_{1}\right)\mu^{\ast}+\underline{X}\left(\delta(f_{1},\rho_{0},\beta_{0})-\delta\right)+(I+\rho_{0}M)(u_{1}-f_{1}u_{2})\right\}$
$=$ $V(\rho,\gamma)\left(\begin{array}[]{cc}\underline{X},&\mu^{*}\end{array}\right)\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right)+\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1}u_{2}),$ (51)
recalling that $V(\rho,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$ and $\mathcal{M}(\rho)=(I+\rho M)^{-1}(I+\rho_{0}M)$.
We first analyze the linear moment conditions. Note that
$H^{\prime}u^{+}(\theta,\gamma)=H^{\prime}V(\rho,\gamma)\left(\begin{array}[]{cc}\underline{X},&\mu^{*}\end{array}\right)\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right)+H^{\prime}\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1}u_{2}).$ (52)
Let
$K_{n}(\rho,\gamma)$ $=$ $n^{-1}H^{\prime}V(\rho,\gamma)\begin{array}[]{cc}(\underline{X},&\mu^{*})\end{array}.$ (54)
Observe that
$n^{-1}E\left[H^{\prime}u^{+}(\theta,\gamma)\left|\mathcal{Z}_{n}\right.\right]$ $=$ $E\left(K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right)\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right).$ (57)
where
the r.h.s. is seen to hold in light of (52) and (54), and since $H$ is $\mathcal{Z}_{n}$-measurable and thus
$E\left[H^{\prime}\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1}u_{2})\left|\mathcal{Z}_{n}\right.\right]=0.$
The assumptions imply that $K_{n}$ has full rank a.s. To see this, note that in light of (39), (40), Lemma (B.1)(iv) implies that
$V(\rho,\gamma)=\operatorname{diag}_{c=1}^{C}\left(\frac{1}{\gamma_{\tau_{c}}}\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\frac{1}{\gamma_{\tau_{c}}}\left(1+\rho\right)J_{c}^{\ast}\right).$ (58)
Under Assumptions 4 and 6, there exists some constants $c_{e}$ and $C_{e}$ such that $0<c_{e}<\frac{1}{\gamma_{\tau_{c}}}\frac{n_{c}-1-\rho}{n_{c}-1}<C_{e}<\infty$ and $0<c_{e}<\frac{1}{\gamma_{\tau_{c}}}\left(1+\rho\right)<C_{e}<\infty$. In light of Lemma B.1(iii), the eigenvalues of $V(\rho,\gamma)$ are uniformly bounded below by $c_{e}>0$. By Assumption 7 we have $\lambda_{min}\left(n^{-1}\underline{X}^{\prime}\underline{X}\right)\geq\xi_{X}>0$ and thus $\lambda_{min}\left(n^{-1}\underline{X}^{\prime}V(\rho,\gamma)\underline{X}\right)\geq c_{e}\xi_{X}>0$ uniformly for all $\gamma$, $\rho$, and $n$.
Recalling that $H=\left(\underline{X},z\right)$ and recalling the expression for $K_{n}(\theta,\gamma)$ given in (54) the matrix can be written as
$K_{n}(\rho,\gamma)$ $=\frac{1}{n}\left(\begin{array}[]{cc}\underline{X}^{\prime}V(\rho,\gamma)\underline{X}&\underline{X}^{\prime}V(\rho,\gamma)\mu^{*}\\ z^{\prime}V(\rho,\gamma)\underline{X}&z^{\prime}V(\rho,\gamma)\mu^{*}\end{array}\right).$ (61)
Thus by the determinant for partitioned matrices, and using the shorthand notation $V=V(\rho,\gamma)$,
$\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]\right)\right|$ (62)
$=$ $\left|\operatorname{det}\left(\frac{1}{n}E\left[\underline{X}^{\prime}V\underline{X}\left|\mathcal{Z}_{n}\right.\right]\right)\right|\left|\operatorname{det}\left\{E\left[\left(\frac{1}{n}z^{\prime}V\mu^{*}-\frac{1}{n}z^{\prime}V\underline{X}(\frac{1}{n}\underline{X}^{\prime}V\underline{X})^{-1}\frac{1}{n}\underline{X}^{\prime}V\mu^{*}\right)\left|\mathcal{Z}_{n}\right.\right]\right\}\right|$
$\geqslant$ $\left(\xi_{X}c_{e}\right)^{p_{x}}\left|\frac{1}{n}E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}\mu^{*})\left|\mathcal{Z}_{n}\right.\right]\right|.$
Observe furthermore that
$E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}y_{2})\left|\mathcal{Z}_{n}\right.\right]$ $=$ $E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}\mu^{*})\left|\mathcal{Z}_{n}\right.\right]$
in light of (9) and since $Q_{V^{1/2}X}$ is orthogonal to $V^{1/2}\underline{X}$ and $E\left[u_{i2}|\mathcal{Z}_{n}\right]=0$ by Theorem 2.1. Thus
$\inf_{\gamma\in\Gamma,\rho\in[-K_{\rho},K_{\rho}]}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]\right)\right|$$\geq$ $\left(\xi_{X}c_{e}\right)^{p_{x}}\inf_{\gamma\in\Gamma,\rho\in[-K_{\rho},K_{\rho}]}n^{-1}\left|E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}y_{2})\left|\mathcal{Z}_{n}\right.\right]\right|$
$\geq$ $\left(\xi_{X}c_{e}\right)^{p_{x}}K_{y}>0$
in light of Assumption 7. This proves that $E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]$ in (57) has full rank for all admissible values of $\rho$ and $\gamma$, and consequently $E\left[m_{n}^{(l)}(\theta,\gamma)\right]=0$ if and only if $f_{1}=f_{1,0}$ and $\delta=\delta_{0}$. Therefore $f_{1,0}$ and $\delta_{0}$ are identified from the linear moment condition.
Once $f_{1,0}$ and $\delta_{0}$ are identified, $\rho$ can be identified from the quadratic moment condition. To see this, note that when evaluated at $\delta_{0}$ and $f_{1,0},$ $u^{+}(\theta,\gamma)$ in Equation (C) becomes
$u^{+}(\theta,\gamma)|_{f_{1,0},\delta_{0}}$ $=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}(I+\rho_{0}M)(u_{1}-f_{1,0}u_{2})$
$=\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1,0}u_{2}).$
Then,
$\left.E\left[u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)|\mathcal{Z}_{n}\right]\right|$
$=$ $E\left[(u_{1}-f_{1,0}u_{2})^{\prime}\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)(u_{1}-f_{1,0}u_{2})|\mathcal{Z}_{n}\right]$
$=$ $E\left[\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)|\mathcal{Z}_{n}\right].$
[…]
Proof of Lemma 4.1 and Corollary 4.1.
Lemma 4.1 is a special case of Lemma 4.2 without covariates. Note that Assumptions 1, 2, 3, 4, 5, and 6 hold. Without covariates Equation (61) reduces to $K_{n}\left(\rho,\gamma\right)=z^{\prime}V\left(\rho,\gamma\right)\mu^{*}$ where $V\left(\rho,\gamma\right)=\Omega\left(\gamma\right)^{-1/2}\left(I+\rho M\right)^{-1}.$
Recall that $z=\left[z_{1}^{\prime},...,z_{c}^{\prime},...,z_{C}^{\prime}\right]^{\prime}$ with $z_{c}=\grave{z}_{c}\mathbf{1}_{c}$. Utilizing Lemma B.1,
$(I_{c}+\rho M_{c})^{-1}z_{c}=\left(\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\frac{1}{1+\rho}J_{c}^{\ast}\right)\grave{z}_{c}\mathbf{1}_{c}=\frac{1}{1+\rho}\grave{z}_{c}\mathbf{1}_{c}=\frac{z_{c}}{1+\rho}$
leads to
$n^{-1}E\left[y_{2}^{\prime}\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}z|\mathcal{Z}_{n}\right]$ $=\frac{1}{1+\rho}n^{-1}E\left[y_{2}^{\prime}\Omega(\gamma)^{-1/2}z|\mathcal{Z}_{n}\right]\geq\frac{1}{1+\rho}K_{y}>0$
where the inequality follows from the conditions imposed in the lemma, in particular (28). With $\rho\in(-1,1)$, E$\left[K_{n}\left(\rho,\gamma\right)|\mathcal{Z}_{n}\right]\neq 0$ such that (57) only has one solution, $f_{1}=f_{1,0}$. Identification of $\rho_{0}$ from the quadratic moment is not affected by the absence of covariates. Thus Lemma 4.1 follows from Lemma 4.2.
To prove Corollary 4.1 observe that when $z=\mathbf{1}_{n}$, $n^{-1}\sum_{c=1}^{C}E\left[y_{c2}^{\prime}\mathbf{1}_{c}\right]>K_{y}>0$ implies that (28) holds, observing that
$n^{-1}\left|E\left[y_{2}^{\prime}\Omega(\gamma)^{-1/2}\mathbf{1}_{n}\left|\mathcal{Z}_{n}\right.\right]\right|=n^{-1}E\sum_{c=1}^{C}\left[\frac{y_{c2}^{\prime}\mathbf{1}_{c}}{\gamma_{\tau_{c}}}\left|\mathcal{Z}_{n}\right.\right]\geq\frac{1}{K_{\gamma}}\frac{1}{n}\sum_{c=1}^{C}E\left[y_{c2}^{\prime}\mathbf{1}_{c}\left|\mathcal{Z}_{n}\right.\right]\geq\frac{K_{y}}{K_{\gamma}}>0.$
Proof of Lemma 4.3.
Observe that utilizing (45) we have
$n^{-1}E\left[H^{\prime}(I+\rho M)\epsilon^{+}\left(\rho,f_{1},\delta\right)\left|\mathcal{Z}_{n}\right.\right]$ $=$ $n^{-1}E\left[H^{\prime}\left(y_{1}-f_{1}y_{2}-\underline{X}\delta\right)\left|\mathcal{Z}_{n}\right.\right]$
$=$ $n^{-1}E\left[H^{\prime}\left(\mu^{\ast}(f_{1,0}-f_{1})+\underline{X}(\delta(f_{1},\rho_{0},\beta_{0})-\delta)\right)\left|\mathcal{Z}_{n}\right.\right].$
Let $\tilde{K}_{n}=n^{-1}H^{\prime}\left(\underline{X},\mu^{*}\right)$, then the linear moment condition can be expressed as
$n^{-1}E\left[H^{\prime}(I+\rho M)\epsilon^{+}\left(\rho,f_{1},\delta\right)\left|\mathcal{Z}_{n}\right.\right]=E\left[\tilde{K}_{n}\left|\mathcal{Z}_{n}\right.\right]\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right)=0.$ (64)
Observe that $E\left[\tilde{K}_{n}\right]$ has full rank, as is readily seen from the proof of Lemma 4.2 in light of (C) with $V=I$. This implies that the only solution to (64) is $\delta=\delta_{0}$ and $f_{1}=f_{1,0}$. It now follows that
$\epsilon^{+}\left(\rho,f_{1,0},\delta_{0}\right)=\mathcal{M}\left(\rho\right)(u_{1}-f_{1,0}u_{2})$
such that the quadratic moment condition is
$E\left[\epsilon^{+}\left(\rho,f_{1,0},\delta_{0}\right)^{\prime}A\epsilon^{+}\left(\rho,f_{1,0},\delta_{0}\right)\left|\mathcal{Z}_{n}\right.\right]=\operatorname{tr}\left(\mathcal{M}\left(\rho\right)A\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0.$ (65)
By Lemma B.2 it follows that the only solution to (65) is $\rho=\rho_{0}.$
Proof of Theorem 4.1.
To prove part (i) of the theorem we establish the conditions of Lemma 3.1 in Pötscher and Prucha (1997).
For the linear moment conditions use (52). We have
$n^{-1}H^{\prime}u^{+}(\theta,\gamma)$ $=$ $\left(f_{1,0}-f_{1}\right)n^{-1}H^{\prime}V(\rho,\gamma)\mu^{\ast}$
$+$ $\frac{1}{n}H^{\prime}V(\rho,\gamma)\underline{X}\left(\delta(f_{1},\rho_{0},\beta_{0})-\delta\right)$
$+$ $\frac{1}{n}H^{\prime}\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1}u_{2})$
$\equiv$ $I+II+III.$
Consider $I$ and $II$. Using Assumption 9 and noting that $V(\rho,\gamma)$ defined in (58) satisfies the assumption on $C(\theta,\gamma)$, hence $n^{-1}H^{\prime}V(\rho,\gamma)\underline{X}$ converges uniformly to
$\lim_{n\rightarrow\infty}E\left[n^{-1}H^{\prime}V(\rho,\gamma)\underline{X}\right]\equiv\mathscr{U}_{H,x}\left(\theta,\gamma\right)$ (66)
and $n^{-1}H^{\prime}V(\rho,\gamma)\mu^{\ast}$ converges uniformly to
$\lim_{n\rightarrow\infty}E\left[n^{-1}H^{\prime}V\left(\rho,\gamma\right)\mu^{*}\right]\equiv\mathscr{U}_{H,\mu}\left(\theta,\gamma\right).$ (67)
Consider $III$. From Corollary B.1 and Assumption 9, we have $sup_{\theta,\gamma}|III|\rightarrow_{p}0$.
Now consider the quadratic moment conditions. Let $\vartheta=(\delta^{\prime},f_{1})^{\prime}$, $\vartheta_{0}=(\delta^{\prime}(f_{1},\rho_{0},\beta_{0}),f_{1,0})^{\prime}$ and $\Delta_{1}=[\underline{X},\mu^{\ast}]$. We point out that $\vartheta_{0}$ is a function of $f_{1}$. We omit this dependence for notational convenience and because $f_{1}$ is pinned down by linear moment conditions for $f_{1}$ directly, or in other words by equations related to the second component of $\vartheta$. In light of (C) we have
$n^{-1}u^{+}(\theta,\gamma)Au^{+}(\theta,\gamma)$ $=$ $n^{-1}\left(\vartheta-\vartheta_{0}\right)^{\prime}\Delta_{1}^{\prime}V(\rho,\gamma)^{\prime}AV(\rho,\gamma)\Delta_{1}\left(\vartheta-\vartheta_{0}\right)$ (68)
$-n^{-1}2\left(\vartheta-\vartheta_{0}\right)^{\prime}\Delta_{1}^{\prime}V(\rho,\gamma)^{\prime}A\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)\left(u_{1}-f_{1}u_{2}\right)$
$+n^{-1}\left(u_{1}-f_{1}u_{2}\right)^{\prime}\mathcal{M}(\rho)^{\prime}\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)\left(u_{1}-f_{1}u_{2}\right)$
$=$ $I+II+III.$
By Assumption 9, $n^{-1}\Delta_{1}^{\prime}V(\rho,\gamma)^{\prime}AV(\rho,\gamma)\Delta_{1}$ in $I$ converges uniformly to
$\lim_{n\rightarrow\infty}n^{-1}E\left[\Delta_{1}^{\prime}V(\rho,\gamma)^{\prime}AV(\rho,\gamma)\Delta_{1}\right]\equiv\mathscr{U}_{A,\Upsilon}\left(\theta,\gamma\right)$ (69)
By Corollary B.1 and Assumption 9, $II$ converges uniformly to 0. Consider $III$, let $B(\theta,\gamma)=\mathcal{M}(\rho)^{\prime}\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)$, then
$\frac{1}{n}(u_{1}-f_{1}u_{2})^{\prime}B(\theta,\gamma)(u_{1}-f_{1}u_{2})$ $=\frac{1}{n}u_{1}^{\prime}B(\theta,\gamma)u_{1}+\frac{f_{1}^{2}}{n}u_{2}^{\prime}B(\theta,\gamma)u_{2}-\frac{2}{n}f_{1}u_{1}^{\prime}B(\theta,\gamma)u_{2}.$
Each term on the R.H.S. converges to its mean uniformly in $\theta$ and $\gamma$ under Corollary B.1 and Assumption 9, and the mean is finite and continuously differentiable in $\theta,\gamma$ Consequently, $\frac{1}{n}(u_{1}-f_{1}u_{2})^{\prime}B(\theta,\gamma)(u_{1}-f_{1}u_{2})$ converges uniformly to the limit of its mean
$E\left[\frac{1}{n}(u_{1}-f_{1}u_{2})^{\prime}B(\theta,\gamma)(u_{1}-f_{1}u_{2})\right]$ $=\operatorname{tr}\left[\frac{1}{n}B(\theta,\gamma)\left(\Sigma_{1}+f_{1}^{2}\Sigma_{2}\right)\right]$
$=\frac{1}{n}\operatorname{tr}\left(B(\theta,\gamma)\Omega_{0}\right)+\frac{1}{n}\left(f_{1}^{2}-f_{1,0}^{2}\right)\operatorname{tr}\left(B(\theta,\gamma)\Sigma_{2}\right),$
observing that $\Omega_{0}=\Sigma_{1}+f_{1,0}^{2}\Sigma_{2}$. Next observe that $B(\theta,\gamma)=\mathcal{M}(\rho)^{\prime}\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)$ falls into the class of $C(\theta,\gamma)$ matrices defined in Assumption 9(i). Thus by Assumption 9, there exist bounded and continuously differentiable functions in $\theta$ and $\gamma$ denoted by $\mathscr{U}_{A,\Omega}(\theta,\gamma)$ and $\mathscr{U}_{A,\Sigma}(\theta,\gamma)$ such that
$\lim_{n\rightarrow\infty}\sup_{\theta,\gamma}\left\|E\left[\frac{1}{n}\operatorname{tr}\left(B(\theta,\gamma)\Omega_{0}\right)\right]-\mathscr{U}_{A,\Omega}(\theta,\gamma)\right\|=0$ (70)
and
$\lim_{n\rightarrow\infty}\sup_{\theta,\gamma}\left\|E\left[\frac{1}{n}\operatorname{tr}\left(B(\theta,\gamma)\Sigma_{2}\right)\right]-\mathscr{U}_{A,\Sigma}(\theta,\gamma)\right\|=0$ (71)
From uniform convergence and (66), (69), (70) and (71) it follows that
$\lim_{n\rightarrow\infty}n^{-1/2}E\left[m_{n}\left(\theta,\gamma\right)\right]$ $=\left[\begin{array}[]{c}\left(\mathscr{U}_{H,x}\left(\theta,\gamma\right),\mathscr{U}_{H,\mu}\left(\theta,\gamma\right)\right)\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right)\\ \left(\vartheta-\vartheta_{0}\right)^{\prime}\mathscr{U}_{A,\Upsilon}\left(\theta,\gamma\right)\left(\vartheta-\vartheta_{0}\right)+\mathscr{U}_{A,\Omega}\left(\theta,\gamma\right)+\left(f_{1}^{2}-f_{1,0}^{2}\right)\mathscr{U}_{A,\Sigma}(\theta,\gamma)\end{array}\right]$ (76)
$\equiv\left[\begin{array}[]{c}\mathfrak{m}_{l}\left(\theta,\gamma\right)\\ \mathfrak{m}_{q}\left(\theta,\gamma\right)\end{array}\right]\equiv\mathfrak{m}\left(\theta,\gamma\right).$ (79)
By uniform convergence of $n^{-1/2}m_{n}\left(\theta,\gamma\right)$ it also follows that for $\bar{\gamma}_{n}\rightarrow\gamma_{*}$
$\sup_{\theta}\left\|n^{-1/2}m_{n}\left(\theta,\bar{\gamma}_{n}\right)-\mathfrak{m}\left(\theta,\gamma_{*}\right)\right\|\rightarrow 0\,i.p.$ (80)
We proceed to show, building on our results for finite $n$, that $\theta_{0}$ is also the unique solution vector of the limiting moment condition $\mathfrak{m}\left(\theta,\gamma_{*}\right)=0$. Observe that from (54), (66) and (67) we have $\left[\mathscr{U}_{H,x}\left(\theta,\gamma_{*}\right),\mathscr{U}_{H,\mu}\left(\theta,\gamma_{*}\right)\right]=\lim_{n\rightarrow\infty}E\left[K_{n}(\rho,\gamma_{*})\right],$ and by (C) we have $\inf_{\gamma\in\Gamma,\rho}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\right]\right)\right|\geq C_{K}>0$ for some $C_{K}>0$ and all $n.$ Since $\operatorname{det}(.)$ is a continuous function it follows that for any $\theta,\gamma_{*}$ we have
$\left|\operatorname{det}\left(\mathscr{U}_{H,x}\left(\theta,\gamma_{*}\right),\mathscr{U}_{H,\mu}\left(\theta,\gamma_{*}\right)\right)\right|$ $=\left|\operatorname{det}\left(\lim_{n\rightarrow\infty}E\left[K_{n}(\rho,\gamma_{*})\right]\right)\right|$
$=\lim_{n\rightarrow\infty}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma_{*})\right]\right)\right|\geq C_{K}>0.$
In light of this for any $\rho$ and $\gamma_{*}$
$\mathfrak{m}_{l}\left(\theta,\gamma_{*}\right)$ $=$ $\mathfrak{m}_{l}\left(\rho,f_{1},\delta,\gamma_{*}\right)=0$
if and only if $f_{1}=f_{1,0}$ and $\delta=\delta_{0}$. Note that $\delta(f_{1},\rho_{0},\beta_{0})$ equals $\delta_{0}$ when $f_{1}=f_{1,0}$. In all , $\mathfrak{m}_{l}\left(\theta,\gamma_{*}\right)=0$ if and only if $f_{1}=f_{1,0}$ and $\delta=\delta_{0}$.
To show that $\mathfrak{m}\left(\theta,\gamma_{*}\right)=0$ if and only if $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}=\theta_{0}=(\rho_{0},f_{1,0},\delta_{0}^{\prime})^{\prime}$ it thus suffices to show that $\mathfrak{m}_{q}\left(\rho,f_{1,0},\delta_{0},\gamma_{*}\right)=0$ if and only if $\rho=\rho_{0}$. Recall that
$\mathfrak{m}_{q}\left(\rho,f_{1,0},\delta_{0},\gamma_{*}\right)$ $=\mathscr{U}_{A,\Omega}\left(\rho,f_{1,0},\delta_{0},\gamma_{*}\right)$
$=\lim_{n\rightarrow\infty}\frac{1}{n}\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma_{\ast})^{-1/2}A\Omega(\gamma_{\ast})^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right).$
By Lemma B.2, $\mathfrak{m}_{q}\left(\rho,f_{1,0},\delta_{0},\gamma_{*}\right)=0$ if and only if $\rho=\rho_{0}$, and consequently $\mathfrak{m}\left(\theta,\gamma_{*}\right)=0$ if and only if $\theta=\theta_{0}$. Assumption 9 implies that $\mathfrak{m}\left(\theta,\gamma_{*}\right)$ and thus $\mathfrak{m}\left(\theta,\gamma_{*}\right)^{\prime}\Xi^{-1}\mathfrak{m}\left(\theta,\gamma_{*}\right)$ is continuous. Since the parameter space is compact it follows that $\theta_{0}=[$$\rho_{0},f_{1,0},\delta_{0}^{\prime}]^{\prime}$ is the identifiable unique minimizer of $\mathfrak{m}\left(\theta,\gamma_{*}\right)^{\prime}\Xi^{-1}\mathfrak{m}\left(\theta,\gamma_{*}\right)$; cp. the discussion after Definition 3.1 in Pötscher and Prucha (1997). Consistency now follows immediately from Lemma 3.1 in Pötscher and Prucha (1997).
For part (ii) of the theorem we first consider $\tilde{f}_{1}$. Observing that in light of (8) we have $y_{1}=f_{1,0}y_{2}+\underline{X}\delta_{0}+(I+\rho_{0}M)(u_{1}-f_{1,0}u_{2})$ and that $Q_{X}\underline{X}=0$, it follows from (20) that
$\tilde{f}_{1}$ $=$ $\frac{z^{\prime}Q_{X}y_{1}}{z^{\prime}Q_{X}y_{2}}=f_{1,0}+\frac{n^{-1}z^{\prime}Q_{X}(I+\rho_{0}M)(u_{1}-f_{1,0}u_{2})}{n^{-1}z^{\prime}Q_{X}y_{2}}.$ (81)
It follows that $n^{-1}z^{\prime}Q_{X}(I+\rho_{0}M)(u_{1}-f_{1,0}u_{2})\rightarrow_{p}0$ since $n^{-1}z^{\prime}Q_{X}(I+\rho_{0}M)u_{t}\rightarrow_{p}0$ for $t=1,2$ by Assumption 9.
Observe that for $\rho=0$,$\gamma=(1,...,1)^{\prime}$, $V\left(\rho,\gamma\right)=I_{n}$ and $Q_{V^{1/2}X}=Q_{X}.$ Using (9) implies that
$n^{-1}z^{\prime}Q_{X}y_{2}=n^{-1}z^{\prime}Q_{X}\mu^{*}+n^{-1}z^{\prime}Q_{X}(I+\rho_{0}M)u_{2}$
and $E\left[n^{-1}z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]=E\left[n^{-1}z^{\prime}Q_{X}\mu^{*}|\mathcal{Z}_{n}\right].$ As argued above, $n^{-1}z^{\prime}Q_{X}(I+\rho_{0}M)u_{2}\rightarrow_{p}0.$ By Assumption 9 it follows that $n^{-1}\left|z^{\prime}Q_{X}\mu^{*}-E\left[z^{\prime}Q_{X}\mu^{*}\right]\right|\rightarrow 0$ a.s. and $n^{-1}\left|E\left[z^{\prime}Q_{X}\mu^{*}\right]-\mathcal{\mathscr{U}}_{z\mu}\right|\rightarrow 0$ a.s. Assumption 7(iii) implies that $n^{-1}\left|E\left[z^{\prime}Q_{X}\mu^{*}\right]\right|\geq K_{y}>0$ which in turn implies that $\left|\mathcal{\mathscr{U}}_{z\mu}\right|\geq K_{y}>0$ This shows that the denominator on the r.h.s. of (81) converges to a positive constant. Having shown that the numerator converges to zero it follows that $\tilde{f}_{1}\rightarrow_{p}f_{1,0}.$
Then, turning to
$\tilde{\delta}$ $=$ $\left(\underline{X}^{\prime}\underline{X}\right)^{-1}\underline{X}^{\prime}\left(y_{1}-y_{2}\tilde{f}_{1}\right)$
$=$ $\left(\underline{X}^{\prime}\underline{X}\right)^{-1}\underline{X}^{\prime}\left(\underline{X}\delta_{0}+\left(I_{n}+\rho_{0}M\right)\left(u_{1}-u_{2}f_{1,0}\right)\right)$
$+$ $\left(\underline{X}^{\prime}\underline{X}\right)^{-1}\underline{X}^{\prime}y_{2}\left(f_{1,0}-\tilde{f}_{1}\right)$
$\equiv$ $I+II$
[…]
Now consider the quadratic moment conditions. Let $\Upsilon_{1}=\left[X,y_{2}\right]$, $\tilde{\vartheta}=\left(\tilde{\delta}^{\prime},\tilde{f}_{1}\right)^{\prime}$, and observe that utilizing (8) we have
$\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)$ $=$ $\left(I+\rho M\right)^{-1}[y_{1}-y_{2}\tilde{f}_{1}-\underline{X}\tilde{\delta}]$
$=$ $\left(I+\rho M\right)^{-1}[(I+\rho_{0}M)(u_{1}-f_{1,0}u_{2})-\Upsilon_{1}(\tilde{\vartheta}-\vartheta_{0})],$
where, with slight abuse of notation relative to earlier definitions, we now understand $\vartheta_{0}=(\delta_{0}^{\prime},f_{1,0})^{\prime}.$ Consequently
$n^{-1}\epsilon^{+}(\rho,\tilde{f}_{1},\tilde{\delta})^{\prime}A\epsilon^{+}(\rho,\tilde{f}_{1},\tilde{\delta})$ $=$ $n^{-1}\left(\tilde{\vartheta}-\vartheta_{0}\right)^{\prime}\Upsilon_{1}^{\prime}\left(I+\rho M\right)^{-1}A\left(I+\rho M\right)^{-1}\Upsilon_{1}\left(\tilde{\vartheta}-\vartheta_{0}\right)$
$-n^{-1}2\left(\tilde{\vartheta}-\vartheta_{0}\right)^{\prime}\Upsilon_{1}^{\prime}\left(I+\rho M\right)^{-1}A\mathcal{M}\left(\rho\right)\left(u_{1}-f_{1,0}u_{2}\right)$
$+n^{-1}\left(u_{1}-f_{1,0}u_{2}\right)^{\prime}\mathcal{M}\left(\rho\right)A\mathcal{M}\left(\rho\right)\left(u_{1}-f_{1,0}u_{2}\right)$
$=$ $I+II+III.$
By Assumption 9 it follows that $\sup_{\rho}\left|n^{-1}\Upsilon_{1}^{\prime}(I+\rho M)^{-1}A(I+\rho M)^{-1}\Upsilon_{1}-\mathcal{\mathscr{U_{\mathnormal{\Upsilon_{1}\Upsilon_{1}}}\left(\rho\right))}}\right|\rightarrow 0$ a.s. for some uniformly bounded $U_{\mathnormal{\Upsilon_{1}\Upsilon_{1}}}\left(\rho\right)$. Thus $I$ converges to zero uniformly in probability. Uniform convergence in probability to zero for the second term, $II,$ follows by Assumption 9.
Finally, using Lemma B.2 and Assumption 9 it follows that
[…]
Then the consistency of $\tilde{\rho}$ follows from Lemma 3.1 in Pötscher and Prucha (1997), and analogous arguments as those used for part (i) of the theorem. ∎
Proof of Theorem 4.2.
First show consistency of $\hat{\gamma}$. Define $\Upsilon_{1}=\left[\underline{X},y_{2}\right]$ and $\vartheta=\left(\delta^{\prime},f_{1}\right)^{\prime}$. Observing from (8) that $y_{1}-f_{1,0}y_{2}-\underline{X}\delta_{0}=(I+\rho_{0}M)(u_{1}-f_{1,0}u_{2})$ we have
$\tilde{\epsilon}$ $=$ $\left(I_{n}+\tilde{\rho}M\right)^{-1}\left(y_{1}-\Upsilon_{1}\tilde{\vartheta}\right)$
$=$ $\left(u_{1}-u_{2}f_{1,0}\right)+\left(\mathcal{M}\left(\tilde{\rho}\right)-I\right)\left(u_{1}-u_{2}f_{1,0}\right)$
$+$ $(I+\tilde{\rho}M)\Upsilon_{1}\left(\vartheta_{0}-\tilde{\vartheta}\right)$
such that
$\sum_{c=1}^{C}\tilde{\epsilon}_{c}^{\prime}\tilde{\epsilon}_{c}1(\tau_{c}=j)$ $=$ $\sum_{c=1}^{C}\left(u_{c1}-u_{c2}f_{1,0}\right)^{\prime}\left(u_{c1}-u_{c2}f_{1,0}\right)1(\tau_{c}=j)$
$+\sum_{c=1}^{C}\left(u_{c1}-u_{c2}f_{1,0}\right)^{\prime}\left(\mathcal{M}_{c}\left(\tilde{\rho}\right)-I\right)^{2}\left(u_{c1}-u_{c2}f_{1,0}\right)1(\tau_{c}=j)$
$+\sum_{c=1}^{C}\left(\vartheta_{0}-\tilde{\vartheta}\right)^{\prime}\Upsilon_{1,c}^{\prime}\left(I_{c}+\tilde{\rho}M_{c}^{\prime}\right)^{-1}\left(I_{c}+\tilde{\rho}M_{c}\right)^{-1}\Upsilon_{1,c}\left(\vartheta_{0}-\tilde{\vartheta}\right)1(\tau_{c}=j)$
$+2\sum_{c=1}^{C}\left(u_{c1}-u_{c2}f_{1,0}\right)^{\prime}\left(\mathcal{M}_{c}\left(\tilde{\rho}\right)-I_{c}\right)\left(u_{c1}-u_{c2}f_{1,0}\right)1(\tau_{c}=j)$
$+2\sum_{c=1}^{C}\left(u_{c1}-u_{c2}f_{1,0}\right)^{\prime}\left(\mathcal{M}_{c}\left(\tilde{\rho}\right)-I_{c}\right)^{\prime}\left(I_{n}+\tilde{\rho}M_{c}\right)^{-1}\Upsilon_{1,c}\left(\vartheta_{0}-\tilde{\vartheta}\right)1(\tau_{c}=j)$
$+2\sum_{c=1}^{C}\left(u_{c1}-u_{c2}f_{1,0}\right)^{\prime}\left(I_{c}+\tilde{\rho}M_{c}\right)^{-1}\Upsilon_{1,c}\left(\vartheta_{0}-\tilde{\vartheta}\right)1(\tau_{c}=j)$
$=$ $I+II+III+IV+V+VI.$
Now consider the components of $\left(N_{j}-p_{x}-1\right)^{-1}\sum_{c=1}^{C}\tilde{\epsilon}_{c}^{\prime}\tilde{\epsilon}_{c}1(\tau_{c}=j)$. For $I$, note that
$n^{-1}E\left[I\right]$ $=$ $n^{-1}\sum_{c=1}^{C}E\left[E\left[\left(u_{c1}-u_{c2}f_{1,0}\right)^{\prime}\left(u_{c1}-u_{c2}f_{1,0}\right)|\mathcal{Z}_{n}^{*}\right]1(\tau_{c}=j)\right]$
$=$ $\left(\sigma_{1,0}^{2}+f_{1,0}^{2}\sigma_{2,0}^{2}\right)n^{-1}\sum_{c=1}^{C}n_{c}\rho_{\tau_{c},0}^{2}1(\tau_{c}=j)$
$=$ $\left(\sigma_{1,0}^{2}+f_{1,0}^{2}\sigma_{2,0}^{2}\right)\rho_{j,0}^{2}n^{-1}\sum_{c=1}^{C}n_{c}1(\tau_{c}=j)$
$=$ $\gamma_{j,0}^{2}(N_{j}/n)\rightarrow\gamma_{j,0}^{2}\omega_{j}$
recalling that $N_{j}/n\rightarrow\omega_{j}$ such that
$\left(N_{j}-p_{x}-1\right)^{-1}E\left[I\right]=[n/(N_{j}-p_{x}-1)]n^{-1}E\left[I\right]\rightarrow\left(\sigma_{1,0}^{2}+f_{1,0}^{2}\sigma_{2,0}^{2}\right)\rho_{j,0}^{2}=\gamma_{j,0}^{2}.$
By Assumption 9 it follows that $n^{-1}I\rightarrow\gamma_{j,0}^{2}$ a.s. for $j=1,...,J$. For $II,$ $IV$ and $V$ consider
$\left(\mathcal{M}\left(\tilde{\rho}\right)-I\right)$ $=\left(I+\tilde{\rho}M\right)^{-1}M\left(\rho_{0}-\tilde{\rho}\right)$
$=\operatorname{diag}_{c}\left(-\frac{1}{\left(n_{c}-1-\tilde{\rho}\right)}I_{c}^{*}+\frac{1}{1+\tilde{\rho}}J_{c}^{*}\right)\left(\rho_{0}-\tilde{\rho}\right)$
such that for $II$
$\left(\mathcal{M}\left(\tilde{\rho}\right)-I\right)^{2}=\operatorname{diag}_{c}\left(\frac{1}{\left(n_{c}-1-\tilde{\rho}\right)^{2}}I_{c}^{*}+\frac{1}{\left(1+\tilde{\rho}\right)^{2}}J_{c}^{*}\right)\left(\rho_{0}-\tilde{\rho}\right)^{2}.$
Then using the notation
$C\left(\rho\right)=\operatorname{diag}_{c}\left(\frac{1(\tau_{c}=j)}{\left(n_{c}-1-\rho\right)^{2}}I_{c}^{*}+\frac{1(\tau_{c}=j)}{\left(1+\rho\right)^{2}}J_{c}^{*}\right)$
in line with Assumption 9 shows that
$n^{-1}II$ $=$ $\left(\rho_{0}-\tilde{\rho}\right)^{2}n^{-1}\left(u_{1}-u_{2}f_{1,0}\right)^{\prime}C\left(\tilde{\rho}\right)\left(u_{1}-u_{2}f_{1,0}\right)$
$\leq$ $\left(\rho_{0}-\tilde{\rho}\right)^{2}\sup_{\rho}\left|n^{-1}\left(u_{1}-u_{2}f_{1,0}\right)^{\prime}C\left(\rho\right)\left(u_{1}-u_{2}f_{1,0}\right)-E\left[n^{-1}\operatorname{tr}\left(C\left(\rho\right)\Omega_{0}\right)\right]\right|$
$+\left(\rho_{0}-\tilde{\rho}\right)^{2}\sup_{\rho}\left|\left[E\left[n^{-1}\operatorname{tr}\left(C\left(\rho\right)\Omega_{0}\right)\right]-\lim_{n\rightarrow\infty}E\left[n^{-1}\operatorname{tr}\left(C\left(\rho\right)\Omega_{0}\right)\right]\right]\right|$
$+\left(\rho_{0}-\tilde{\rho}\right)^{2}\sup_{\rho}\left|\lim_{n\rightarrow\infty}E\left[n^{-1}\operatorname{tr}\left(C\left(\rho\right)\Omega_{0}\right)\right]\right|$
$=$ $o_{p}\left(1\right),$
observing that $\rho_{0}-\tilde{\rho}=o_{p}\left(1\right)$ by Theorem 4.1(ii), that the first and second supremum converges to zero a.s. by Assumption 9(i) and that $\lim_{n}E\left[n^{-1}\operatorname{tr}\left(C\left(\rho\right)\Omega_{0}\right)\right]$ is uniformly bounded by Assumption 9(ii). Arguments analogous to these show that the remaining terms of $n^{-1}\sum_{c=1}^{C}\tilde{\epsilon}_{c}^{\prime}\tilde{\epsilon}_{c}1(\tau_{c}=j)$ converge to zero in probability. The details are omitted. This establishes that
$\hat{\gamma}_{j}^{2}=\left(N_{j}-p_{x}-1\right)^{-1}\sum_{c=1}^{C}\tilde{\epsilon}_{c}^{\prime}\tilde{\epsilon}_{c}1(\tau_{c}=j)\rightarrow_{p}\left(\sigma_{0,1}^{2}+f_{1,0}^{2}\sigma_{0,2}^{2}\right)\rho_{j,0}^{2}=\gamma_{j,0}^{2}.$
To obtain the limiting distribution note that $Q_{n}\left(\theta,\gamma\right)$ defined in (18) can be written as
$Q_{n}\left(\theta,\gamma_{n}\right)=\frac{1}{n}m_{n}\left(\theta,\gamma_{n}\right)^{\prime}\hat{\Xi}_{n}^{-1}m_{n}\left(\theta,\gamma_{n}\right)$
where $\hat{\Xi}_{n}$ is defined after (17). Using a mean value expansion of $m_{n}(\hat{\theta},\gamma_{n})$ around $\theta_{0}$ we obtain
00 0 $=\frac{\sqrt{n}}{2}\frac{\partial Q_{n}\left(\hat{\theta},\gamma_{n}\right)}{\partial\theta}=\frac{\partial m_{n}(\hat{\theta},\gamma_{n})^{\prime}}{\sqrt{n}\partial\theta}\hat{\Xi}_{n}^{-1}m_{n}(\hat{\theta},\gamma_{n})$
$=\frac{\partial m_{n}(\hat{\theta},\gamma_{n})^{\prime}}{\sqrt{n}\partial\theta}\hat{\Xi}_{n}^{-1}m_{n}(\theta_{0},\gamma_{n})+\frac{\partial m_{n}(\hat{\theta},\gamma_{n})^{\prime}}{\sqrt{n}\partial\theta}\hat{\Xi}_{n}^{-1}\frac{\partial m_{n}(\bar{\theta},\gamma_{n})}{\sqrt{n}\partial\theta}\sqrt{n}\left(\hat{\theta}-\theta_{0}\right)$
where $\left\|\bar{\theta}-\theta_{0}\right\|\leq\left\|\hat{\theta}-\theta_{0}\right\|$ and with some abuse of notation it is understood that the between value $\bar{\theta}$ differs among rows of $m_{n}\left(.\right).$ Using the explicit derivatives given in Appendix D it follows from Assumption 9(i) and (ii) that for any sequence $\gamma_{n}$ with $\gamma_{n}\rightarrow_{p}\gamma_{*}$ we have $n^{-1/2}\partial m_{n}\left(\hat{\theta},\gamma_{n}\right)/\partial\theta\rightarrow_{p}G(\gamma_{\ast})$ and $n^{-1/2}\partial m_{n}\left(\bar{\theta},\gamma_{n}\right)/\partial\theta\rightarrow_{p}G(\gamma_{\ast})$ where $G(\gamma_{\ast})$ is fixed matrix with full column rank and $\hat{\Xi}_{n}\rightarrow_{p}\Xi$ with $\Xi$ a full-rank square matrix. This implies that
$\sqrt{n}\left(\hat{\theta}-\theta_{0}\right)=-\left(\left(G^{\prime}\Xi^{-1}G\right)^{-1}G^{\prime}\Xi^{-1}+o_{p}\left(1\right)\right)m_{n}\left(\theta_{0},\gamma_{n}\right).$
It remains to establish the limiting distribution of $m_{n}\left(\theta_{0},\gamma_{n}\right)$ where $\gamma_{n}$ is a possibly random sequence with $\gamma_{n}\rightarrow_{p}\gamma_{*}.$ To apply the results of Kuersteiner and Prucha (2020) note that the moment vector $m_{n}\left(.\right)$ can be represented as the sum over the linear and quadratic terms by defining $h_{i}^{*}=\left(h_{i1},\ldots,h_{iq},0\right)$ and $a_{ij}^{\ast}=\left(0,....,0,a_{ij}\right)$, where $(h_{i1},\ldots,h_{iq})$ denotes the $i$-th row of $H$ and $a_{ij}$ the $i,j$-th element of $A.$ Then, $m_{n}\left(\theta,\gamma\right)=n^{-1/2}\sum_{i=1}^{n}h_{i}^{*\prime}u_{i}^{+}(\theta,\gamma)+n^{-1/2}\sum_{j,i=1}^{n}a_{ij}^{*\prime}u_{i}^{+}(\theta,\gamma)u_{j}^{+}(\theta,\gamma)$ has the same form as Equation (23) in Kuersteiner and Prucha (2020). It is then sufficient to check that the conditions for Proposition 3 in Kuersteiner and Prucha (2020) hold. Proposition 3 requires that Assumptions 1-3 of Kuersteiner and Prucha (2020) hold. We abbreviate assumptions in Kuersteiner and Prucha (2020) with A-KP to avoid confusion with assumptions in this paper.
The moment bounds of A-KP 1(i) hold by Assumption 2 and Lemma 2.1 which implies that $u_{it}$ and $\mu_{i}^{\ast}$ have bounded moments, and Assumption 7(i) which bounds the moments of $x_{it}$ and $z_{it}.$ A-KP 1(ii), Eq. (15), holds by Lemma B.3, A-KP 1(ii), Eq. (16) holds by Assumption 3 and A-KP 1(ii), Eq. (17) holds by Lemma B.3, and the parametric restrictions on the conditional cross-sectional variances hold by Assumption 3.
Now turning to A-KP 2, the $\sigma$-field that corresponds to $\mathcal{B}_{n,t}$ is $\mathcal{A}_{n,\left(t-1\right)n}$ defined in (43). Since $\mathcal{Z}_{n}^{*}\subset\mathcal{A}_{n,\left(t-1\right)n}$, the fact that $h_{i}$ and $a_{ij}$ are assumed measurable w.r.t to $\mathcal{Z}_{n}\subset\mathcal{Z}_{n}^{*}$ and elements of $h_{i}$ and the row sums of the absolute elements (and powers thereof) of $A$ are uniformly bounded by Assumptions 5(i)(iii) and 7, A-KP 2(i) and (ii) hold. We also normalize $f_{2}=1$ and restrict the parameter space of $f_{1}$ to a compact interval in Assumption 4 such that A-KP 2(iii) holds.
Finally, consider A-KP 3 which is identical to Assumption 9. In summary, all conditions of Proposition 3 in Kuersteiner and Prucha (2020) hold. It follows that $m_{n}\left(\theta_{0},\gamma\right)\rightarrow^{d}V_{\gamma}^{1/2}\xi\text{ }$ where $\xi\sim N\left(0,I_{q+1}\right)$ and $V_{\gamma}$ is the limit of the variance-covariance matrix of $m_{n}(\theta_{0},\gamma)$ defined in (8). For any conformable, matrix $\mathscr{A}$ it follows that $\mathscr{A}m_{n}\left(\theta_{0},\gamma\right)\rightarrow^{d}\left(\mathscr{A}V\mathscr{A}^{\prime}\right)^{1/2}\xi$. By choosing $\mathscr{A}=\left(G^{\prime}\Xi^{-1}G\right)^{-1}G^{\prime}\Xi^{-1}$ the claim of the theorem is proven. ∎
Appendix D Gradient Vector
In this section, we derive the gradient vector for the two-step-GMM estimator, which is necessary for the limiting variance covariance matrix in Theorem 4.2. Recall that $\theta=\left(\rho,f_{1},\delta^{\prime}\right)^{\prime}$,
$u^{+}(\theta,\gamma)$ $=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}\left(y_{1}-f_{1}y_{2}-\underline{X}\delta\right)$
$=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}\left((f_{10}-f_{1})\mu^{\ast}+\underline{X}(\delta(f_{1},\rho_{0},\beta_{0})-\delta)+(I+\rho_{0}M)(u_{1}-f_{1}u_{2})\right)$
and that $y_{t}=\mu^{*}f_{t,0}+\underline{X}\beta_{0,t}+(I+\rho_{0}M)u_{t}$.
$\frac{\partial m_{n}(\theta,\gamma)/\sqrt{n}}{\partial\theta^{\prime}}$ $=\frac{1}{n}\left[\begin{array}[]{c}H^{\prime}\frac{\partial u^{+}}{\partial\theta^{\prime}}\\ 2u^{+\prime}A\frac{\partial u^{+}}{\partial\theta^{\prime}}\end{array}\right]=\frac{1}{n}\left[\begin{array}[]{ccc}H^{\prime}\frac{\partial u^{+}}{\partial\rho}&H^{\prime}\frac{\partial u^{+}}{\partial f_{1}}&H^{\prime}\frac{\partial u^{+}}{\partial\delta^{\prime}}\\ 2u^{+\prime}A\frac{\partial u^{+}}{\partial\rho}&2u^{+\prime}A\frac{\partial u^{+}}{\partial f_{1}}&2u^{+\prime}A\frac{\partial u^{+}}{\partial\delta^{\prime}}\end{array}\right].$
The derivative of $u^{+}(\theta,\gamma)$ w.r.t $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ are
$\partial\frac{u^{+}(\theta,\gamma)}{\partial f_{1}}$ $=-\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}y_{2}$
$=-\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}[\mu^{\ast}+\underline{X}\beta_{0,2}+(I+\rho_{0}M)u_{2}].$
$\partial\frac{u^{+}(\theta,\gamma)}{\partial\rho}$ $=-\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}M(I+\rho M)^{-1}(y_{1}-f_{1}y_{2}-\underline{X}\delta)$
$=-\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}M(I+\rho M)^{-1}$
×$\times\left((f_{10}-f_{1})\mu^{\ast}+\underline{X}(\delta(f_{1},\rho_{0},\beta_{0})-\delta)+(I+\rho_{0}M)(u_{1}-f_{1}u_{2})\right)$
$\partial\frac{u^{+}(\theta,\gamma)}{\partial\delta}$ $=-\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}\underline{X}.$
It thus is obvious that $u^{+}(\theta,\gamma)$ and each element of $\frac{u^{+}(\theta,\gamma)}{\partial\theta}$ can be written as linear combinations of $S(\theta,\gamma)\Upsilon$, where $S(\theta,\gamma)$ is an element or the product of elements of the matrix set in (42) and $\Upsilon\in\left\{H,\mu^{*},u_{1},u_{2}\right\}$. Applying Corollary B.1 and Assumption 9, the gradient vector converges uniformly to its expected value.
Appendix E GMM with Randomly Missing Variables
If class-level characteristics like teacher’s race are missing and the variable is controlled in the model, then the whole class is dropped from the sample. Under the assumption that these characteristics are missing randomly, dropping these classes should not affect the consistency of our estimator. Below we discuss the case when individual level variables (control variable or outcome variable) are missing.
Suppose class $c$ has $n_{c}$ students, and $n_{c}^{obs}\leqslant n_{c}$ of them are observed, which we define as having non-missing values for all variables. Without loss of generality, we order the vectors for the class such that observed students are at the top and unobserved are at the bottom. That is, $y_{c1}=(y_{c1}^{obs\prime},y_{c1}^{unobs\prime})^{\prime}$, where $y_{c1}$ is the $n_{c}\times 1$ vector of test score 1 for all students in class $c$, $y_{c1}^{obs}$ is the $n_{c}^{obs}\times 1$ vector of test score 1 for all observed students, and $y_{c1}^{unobs}$ is the $n_{c}^{unobs}\times 1$ vector of test score 1 for all unobserved students. Note that $n_{c}^{unobs}=n_{c}-n_{c}^{obs}$. Similarly, $y_{c2}=(y_{c2}^{obs\prime},y_{c2}^{unobs\prime})^{\prime}$, $x_{c}=(x_{c}^{obs\prime},x_{c}^{unobs\prime})$, etc.
The model for class $c$ is
$y_{1c}-f_{1}y_{2c}=\text{$\underline{X}$}_{c}\delta+(I_{c}+\rho M_{c})(u_{1c}-f_{1}u_{2c}).$ (82)
[…]
$y_{1c}-f_{1}y_{2c}=x_{c}^{c}\delta_{c}+x_{c}^{p}\delta_{p}+M_{c}x_{c}^{p}\delta_{m}+(I_{c}+\rho M_{c})(u_{1c}-f_{1}u_{2c}).$ (83)
The composite error term satisfies $E(u_{1c}-f_{1}u_{2c}|\mathcal{Z}_{n}^{\ast})=0$ and $\textrm{var}(u_{1c}-f_{1,0}u_{2c}|\mathcal{Z}_{n}^{\ast})=\gamma_{\tau_{c}}^{2}I_{c}$.
Let $D_{c}=[I_{n_{c}^{obs}\times n_{c}^{obs}},0_{n_{c}^{obs}\times n_{c}^{nob}}]$ be the $n_{c}^{obs}\times n_{c}$ matrix that transforms the matrix for the whole class to that for the observed students. That is $D_{c}y_{c1}=y_{c1}^{obs}$, etc. Premultiplying both sides of (82) by $D_{c}$, we have
$D_{c}y_{1c}-f_{1}D_{c}y_{2c}=D_{c}x_{c}^{c}\delta_{c}+D_{c}x_{c}^{p}\delta_{p}+D_{c}M_{c}x_{c}^{p}\delta_{m}+D_{c}(I_{c}+\rho M_{c})(u_{1c}-f_{1}u_{2c}).$
It is easy to see that $D_{c}y_{1c}=y_{1c}^{obs}$, $D_{c}y_{2c}=y_{2c}^{obs}$, $D_{c}x_{c}^{c}=x_{c}^{c,obs}$ and $D_{c}x_{c}^{p}=x_{c}^{p,obs}$, which are all observed.
For $D_{c}M_{c}x_{c}^{p}$, observing that $D_{c}\iota_{c}=\iota_{n^{obs}}$ , $D_{c}I_{c}=D_{c}$, and $M_{c}=\frac{\iota_{c}\iota_{c}^{\prime}-I_{c}}{n_{c}-1}$, we have
$D_{c}M_{c}x_{c}^{p}=D_{c}\frac{\iota_{c}\iota_{c}^{\prime}-I_{c}}{n_{c}-1}x_{c}^{p}=\frac{1}{n_{c}-1}\left(\iota_{n_{c}^{nobs}}\iota_{c}^{\prime}x_{c}^{p}-x_{c}^{p,obs}\right)=\frac{1}{n_{c}-1}(\iota_{n_{c}^{obs}}n_{c}\bar{x}_{c}^{p}-x_{c}^{p,obs}),$ (84)
where $\bar{x}_{c}^{p}=\frac{1}{n_{c}}\sum_{r=1}^{n_{c}}x_{rc}^{p}$ is the average characteristics of all students. Let $\bar{x}_{c}^{p,obs}$ be the averages of all observed $x$. That is, as long as a variable in $x^{p}$ is observed for a student, it will be used to calculate the average even if the student has other variables missing and counted as unobserved. For example, some classes may have variable “$girl$” observed for all students, while race and test information may be missing for some. In this way we increase the efficiency of our approximation as having missing value for a variable is much less common than having missing values for all the variables. Then $\bar{girl}_{c}^{p,obs}=\bar{girl}_{c}^{p}$ as we use all observed values of "$girl"$. Assuming that observations are missing at random it follows that $E\left[\bar{x}_{c}^{p}-\bar{x}_{c}^{p,obs}\right]=0$ and the difference between the two is very small as personal characteristics don’t have that many missing values. See Table 9 for details. We can therefore approximate $D_{c}M_{c}x_{c}^{p}$ with $\tilde{x}_{c}^{m,obs}=\frac{1}{n_{c}-1}(\iota_{n_{c}^{obs}}n_{c}\bar{x}_{c}^{p,obs}-x_{c}^{p,obs})$. In all,
$y_{1c}^{obs}-f_{1}y_{2c}^{obs}$ $=x_{c}^{c,obs}\delta_{c}+x_{c}^{p,obs}\delta_{p}+\tilde{x}_{c}^{m,obs}\delta_{m}+v_{c},$
[…]
$\epsilon_{c}=D_{c}(I_{c}+\rho M_{c})(u_{1c}-f_{1}u_{2c}),$
$\xi_{c}=\left(D_{c}M_{c}x_{c}^{p}-\tilde{x}_{c}^{m,obs}\right)\delta_{m}=\frac{n_{c}}{n_{c}-1}\iota_{n_{c}^{obs}}(\bar{x}_{c}^{p}-\bar{x}_{c}^{p,obs})\delta_{m}.$
If observations are missing randomly, we have
$E\left[\epsilon_{c}|x_{c}^{c,obs},x_{c}^{p,obs},\tilde{x}_{c}^{m,obs}\right]=0,$
$\Omega^{obs}\equiv E\left[\epsilon_{c}\epsilon_{c}^{\prime}\right]=\sigma_{c}^{2}\left((\frac{n_{c}-1-\rho}{n_{c}-1})^{2}I_{n^{obs}}+\frac{1}{n_{c}}((1+\rho)^{2}-(\frac{n_{c}-1-\rho}{n_{c}-1})^{2})\iota_{n^{obs}}\iota_{n^{obs}}^{\prime}\right).$ (85)
Note that $E\left[v_{c}|\text{$\underline{x}$}_{c}^{obs}\right]$ and $\textrm{var}(v_{c}|\text{$\underline{x}$}_{c}^{obs})$ requires additional assumptions on the distribution of $\xi_{c}$, in particular $E\left[\bar{x}_{c}^{p}-\bar{x}_{c}^{p,obs}|\bar{x}_{c}^{p,obs}\right]=0$. In practice, we ignore the term $\xi_{c}$. As long as the number of missing $x^{p}$ is small, as Table 9 shows, the bias is small and negligible.
The formula for the variance in (85) points to the importance for adjusting for original class size $n_{c}$ even if observations are missing randomly. We can then define
$u_{c}^{+obs}=(\Omega^{obs})^{-1/2}\left(y_{1c}^{obs}-f_{1}y_{2c}^{obs}-\text{$\underline{x}$}_{c}^{obs}\delta\right)$
and then continue with our GMM estimation strategy.
Graham(2008) discusses missing values in the online appendix. For individual characteristics, “Twenty three kindergarten student records were missing information on free and reduced price school lunch eligibility, in these cases the missing values were replaced with either eligibility status for the same student in the closest of first, second or third grade (17 cases) or the median value among kindergarten students in their school (6 cases). In three cases missing student race values were replaced with school median values.” For test scores, he just adjust for the group size like what we do above “The analysis below assumes the pattern of missing test score data is indeed random and hence ignorable. Observed test scores are regarded as a random sample from different classrooms of known size. ”
Table 9: Number of Missing Values for Each Variable
grade K grade 1 grade 2 grade 3
student characteristics
poor 8 32 62 83
black 3 29 89 14
girl 0 13 7 0
age 8 37 7 20
teacher characteristics
tchfemale 0 19 60 50
tchblack 43 19 60 50
tchms 21 19 60 65
tchyears 21 19 101 50
test scores
read 536 434 763 802
list 488 273 797 728
math 454 231 775 725
word 474 857 493 453
Any vairable missing for the test pairs and controls
list&read 643 528 929 966
math&read 589 480 907 946
read&word 608 1,093 883 914
list&math 570 332 926 878
list&word 562 965 928 863
math&word 555 929 904 842
sample size 6,325 6,829 6,840 6,801
  • 1. School ID, class ID, class size and class type have no missing values.

### v2

@@
-Differential Test Performance and Peer Effects
+Differential Test Performance and Peer Effectsthanks: Ying Zeng gratefully acknowledges financial support from the Young Scientists Fund of the Fujian Natural Science Foundation (Grant No. 2023J05010)
 Abstract
@@
 We develop a framework to analyze, identify and estimate the effect of peer groups on performance measures for individuals who are allocated to groups.
-A leading example are test scores of students who are allocated to classrooms.
+Our analysis centers around the idea of measuring the performance of individuals in closely related tasks and within a short period of time.
+An example are aptitude tests administered at the end of the school year and in related areas such as reading, writing and word comprehension.
+By considering the individual quasi-difference in scores we are able to eliminate unobserved ability and unobserved group effects.
+Our key identifying assumption postulates that different tests measure similar skills and that variation in differential scores is, apart from variation induced by observed covariates, due to idiosyncratic cross-sectional variation in ability or effort of individuals.
+That variation is not related to prior performance and other systematic and possibly unobserved factors correlated with group formation such as teacher quality or parental support.
+A leading example where our method can be applied are test scores of students who are allocated to classrooms.
 However, the theory could equally be applied to outcome measures of members of sports teams, teams of workers or other groups.
@@
 For expositional purposes we refer to groups as classrooms and individuals selected into groups as students with the understanding that our procedure applies to a broader set of scenarios.
-For the statistical implementation we impose linear functional form restrictions.
-This allows us to identify the marginal effect of observed and unobserved average peer characteristics.
-Our analysis centers around the idea of measuring the performance of individuals in closely related tasks and within a short period of time.
-An example are aptitude tests administered at the end of the school year and in related areas such as reading, writing and word comprehension.
-By considering the individual quasi-difference in scores we are able to eliminate unobserved ability and unobserved group effects.
-Our key identifying assumption postulates that different tests measure similar skills and that variation in differential scores is, apart from variation induced by observed covariates, due to cross-sectional variation in ability or effort of individuals that is not related to prior performance and other systematic and possibly unobserved factors correlated with group formation such as teacher quality or parental support.
-Our theory builds on the concept of latent performance, ability or effort measures defined in the absence of group interaction.
 The first result of our paper shows that our key identifying restriction implies orthogonality conditions that are at the core of our identification strategy.
@@
 [^1] The literature on peer effects is part of a larger literature on network effects.
-It is well recognized that the adjacency matrix frequently used in modeling peer effect is a special case of the weight matrices used in a class of models introduced by Cliff and Ord (1973, 1981), which were originally intended for modeling spatial network effects; see Anselin (2010) for a review of important contribution of this literature on identification and estimation of network effects, and Kuersteiner and Prucha (2020) for a recent contribution that connects those strands of network literature.
-.
-In our setting, reduced form regression analysis that focuses on exogenous or contextual peer effects leads to biased estimates.
-We propose instrumental variables estimators to overcome these problems.
-By exploiting moment restrictions that are implied by our identifying assumptions we are proposing generalized method of moments (GMM) estimators that identify both exogenous as well as endogenous peer effects.
-Formally, our moments based estimator is a panel type estimator based on quasi-differences between test performance measures.
-An important difference between our approach and typical panel settings is the lack of a temporal dimension, and where the setup is geared towards the utilization of different test performance measures that are obtained in an essentially simultaneous fashion.
+It is well recognized that the adjacency matrix frequently used in modeling peer effect is a special case of the weight matrices used in a class of models introduced by Cliff and Ord (1973, 1981), which were originally intended for modeling spatial network effects; see Anselin (2010) and Lee and Yu (2015) for a review of important contribution of this literature on identification and estimation of network effects from cross sectional and panel data, and Kuersteiner and Prucha (2020) for a recent contribution that connects those strands of network literature.
 We apply our approach to the Project STAR data set of the Tennessee class size experiment.
@@
 In addition, and unlike in related studies identifying similar parameters such as Kelejian and Prucha (2002), Lee (2007b), Graham (2008) or Kuersteiner, Prucha, and Zeng (2023) we do not rely on group size variation or variation in group type variances.
-The difficulties of identifying peer effects in models with transitive groups where all peers have the same effect, are well known since Manski (1993).
-Approaches to overcome identification challenges have focused on random group assignment, for example Sacerdote (2001); Duflo and Saez (2003); Carrell, Sacerdote, and West (2013); Sojourner (2013), exogenous variation in group size as in Lee (2007b) or Kuersteiner, Prucha, and Zeng (2023), instrumentation of peer effects using non-transitive group structures as in Bramoullé, Djebbari, and Fortin (2009), explicit group formation models in Goldsmith-Pinkham and Imbens (2013), variation in group specific heteroskedasticity in Graham (2008), and panel methods in Mas and Moretti (2009), Arcidiacono et al.
+Our estimators are related to methods that use panel data to identify peer effects, see Mas and Moretti (2009), Arcidiacono et al.
 (2012), Cornelissen, Dustmann, and Schönberg (2017), Miraldo, Propper, and Rose (2021) and Braun and Verdier (2023).
-The approach pursued in this paper builds on the dynamic panel methods for spatial and social networks developed in Kuersteiner and Prucha (2020).
-We use a baseline model of peer effects that is similar to Arcidiacono et al.
-(2012) who consider peer effects working through unobserved individual characteristics.
-Arcidiacono et al.
-(2012) allow for individual specific effects as well as for endogenous group formation and use panel data.
-Technically, our estimator is different from theirs since we use a quasi-differencing approach to eliminate class and individual fixed effects while they use a non-linear fixed effects estimator.
-The more important difference lies in the assumptions and empirical implementation of the method we have in mind.
-While our estimator could be applied to more conventional panel data with performance measures observed at different points in time, we emphasize applications where measures of closely related skills are observed essentially simultaneously.
+In contrast with panel methods, we emphasize applications with cross-sectional data where measures of closely related skills are observed essentially simultaneously in time.
 For one, this framework alleviates problems with sample attrition which are well documented for Project STAR.
-More importantly, our core assumption on unobserved skill or effort centers around a lack of predictive power of additional test results for closely related test measures.
-For example, a student scoring well on a reading test is likely to score well on a word comprehension test given at the same time, and in this sense the word comprehension test does not reveal additional information about underlying skills or effort.
+More importantly, within this framework the source of identifying variation is due to remaining randomness, after conditioning on ability, in the responses of an individual to test questions rather than due to random selection of students into classrooms.
 An advantage of our framework is that its credibility can be assessed and influenced by a testing protocol that complies with its basic premise.
-This is in contrast to more detailed restrictions between various unobserved components as well as on the evolution of individual effects measuring the change in skills over time required in the framework of Arcidiacono et al.
-(2012).
-An additional advantage of our procedure is that it does not require estimating unobserved skill or effort measures of individuals.
 There is an extensive empirical literature studying educational outcomes using Project STAR data.
@@
 Empirically, they find large peer effects for second and third grade math scores.
-In our empirical work we use Project STAR data and focus on classrooms in kindergarten to the third grade.
-We consider four individual SAT scores, mathematics, reading, listening and word study skills.
-We document the high correlation between pairs of these scores, unconditionally as well as conditionally on a full set of controls, with correlation coefficients ranging from 0.5 to 0.9.
-The highest correlation is observed for the reading and word study skill scores.
-The correlation patterns are robust across the four grade levels we consider.
-The peer effects coefficient of our model is estimated precisely and is stable across the different grades with the exception of first grade where we observe a lower degree of peer effects.
-We find stronger peer effects for test score pairs that are less highly correlated, in line with our model assumptions where more closely related scores offer better control for unobserved confounding factors.
+In our empirical work we use Project STAR data and focus on students in kindergarten to the third grade.
+We use individual SAT scores in reading and word study skills as our outcome measures.
+We document the high correlation between the two scores, unconditionally as well as conditionally on a full set of controls.
+The correlation coefficients are about 0.9 for all four grades The peer effects coefficient of our model is estimated precisely and is stable across the different grades except for the first grade where we observe a lower degree of peer effects.
 Quantitatively, we estimate that an increase of average unobserved peer quality measured in SAT equivalent scores of 10 SAT points leads to an increase of 2 to 4 SAT points for an individual classmate.
-This estimate is based on the reading-word score pair.
-For other score pairs the effects range from a 4 to a 7 SAT point increase for the scores of individual classmates due to a change in average scores.
 These measurements are quite robust to the inclusion or omission of additional controls.
@@
 The paper is organized as follows.
-In Section 2 we develop the identification and estimation strategy.
-Section 2.1 defines performance measures and group selection, Section 2.2 introduces the education production function and peer effects, Section 2.3 discusses the key identifying restriction and Section 2.4 introduces the estimators we propose.
-Section 3 contains the empirical analysis of Project STAR data and Section 4 develops formal identification and consistency proofs as well as an asymptotic distribution theory.
+In Section 2 we develop the identification strategy for a model without covariates.
+Section 2.1 defines performance measures and group selection, Section 2.2 introduces the education production function and peer effects and Section 2.3 discusses the key identifying restriction that defines closely related skills.
+Section 3 extends the specification to include covariates and introduces the estimators we propose.
+Section 4 contains the empirical analysis of peer effects using Project STAR data.
+We report a battery of specification checks in Section 4.2.
+Our main empirical findings are reported in Section 4.3.
+Section 5 develops formal identification proofs.
 Section 6 contains conclusions.
 Proofs and tables are contained in the appendix.
+An online appendix contains additional lemmas used in the proofs.
+In the online appendix we also give results that formally establish the consistency and asymptotic normality of the considered estimators, as well as some additional empirical results.
 2 Model
 We assume that for each student $i=1,...,n\text{ }$we observe results for two different tests, or more generally that we observe two measures of academic achievement, say, $y_{it}$ where $t=1,2$ indexes the test.
-Our empirical work uses Project STAR data where these measures are scores of the Stanford Achievement Tests (SAT) chosen from math, word study skills, listening and reading tests given to kindergarten through third grade students.
+Our empirical work uses Project STAR data where these measures are scores of the Stanford Achievement Tests (SAT) in word study skills and reading tests given to kindergarten through third grade students.
 Our approach uses two outcome measures to control for unobserved and possibly correlated student characteristics as well as group level effects related to classrooms, teachers and schools.
@@
 2.1 Latent Performance and Group Selection
-Selection into classrooms can be correlated with a latent baseline performance measure $\tilde{\zeta}_{i}^{*}$ that can also be thought of as prior test taking ability or latent performance, or effort prior to taking the observed tests.
+Selection into classrooms can be correlated with a latent baseline performance measure $\tilde{\zeta}_{i}^{*}$ that can also be thought of individual test taking ability or latent performance, or effort prior to taking the observed tests.
 The purpose of introducing baseline test taking ability is to account for unobserved student characteristics that are invariant for the two tests.
-We account for observed personal student characteristics where $\tilde{v}_{i}^{p}$ and $\tilde{w}_{it}^{p}$ denote row vectors of observed test invariant and test varying characteristics.
-For notational convenience we collect the performance measures and characteristics for all $n$ students in the sample in the following vectors and matrices of row dimension $n$: $\tilde{\zeta}^{\ast}=(\tilde{\zeta}_{1}^{\ast},...,\tilde{\zeta}_{n}^{\ast})^{\prime}$, $\tilde{y}_{t}^{*}=\left(\tilde{y}_{1t}^{*},\ldots,\tilde{y}_{nt}^{*}\right)^{\prime}$ for $t=1,2$, $\tilde{v}^{p}=\left(\tilde{v}_{1}^{p\prime},\ldots,\tilde{v}_{n}^{p\prime}\right)^{\prime}$, $\tilde{w}_{t}^{p}=\left(\tilde{w}_{1t}^{p\prime},\ldots,\tilde{w}_{nt}^{p\prime}\right)^{\prime}$ for $t=1,2$, and $\tilde{X}^{p}=\left(\tilde{v}^{p},\tilde{w}_{1}^{p},\tilde{w}_{2}^{p}\right)$.
-The $i$-th row of $\tilde{X}^{p}$ then contains the covariates of student $i$ in the original sample.
-Selection into classrooms can, in addition to depending on $\tilde{\zeta}^{*}$ and $\tilde{X^{p}}$, depend on observed class level characteristics as well as on a vector $\alpha=\left(\alpha_{1},...,\alpha_{C}\right)^{\prime}$ of unobserved class level characteristics.
-We do not impose any restrictions on the cross-sectional dependence in $\tilde{\zeta}^{\ast}$, $\tilde{X}^{p}$, $\alpha$ or observed class level characteristics.
+Our key identifying assumption of the two tests measuring essentially identical skills detailed in Section 2.3 implies that the ability $\tilde{y}_{it}^{*}$ of taking test $t$ can be decomposed into test invariant factors that include baseline ability $\tilde{\zeta}_{i}^{*}$ and idiosyncratic errors that are uncorrelated across students and tests.
+Our method also accounts for observable student and classroom specific characteristics.
+We omit the observed covariates for now to simplify the discussion of our identification strategy.
+We discuss a more complete model that includes covariates in Section 3.
+For notational convenience we collect the latent baseline performance measures for all $n$ students in the sample in the following $n\times 1$ vector: $\tilde{\zeta}^{\ast}=\left(\tilde{\zeta}_{1}^{\ast},...,\tilde{\zeta}_{n}^{\ast}\right)^{\prime}$.
+Selection into classrooms can, in addition to depending on $\tilde{\zeta}^{*}$ and observed individual characteristics, depend on observed class level characteristics as well as on a vector $\alpha=\left(\alpha_{1},...,\alpha_{C}\right)^{\prime}$ of unobserved class level characteristics.
+We do not impose any restrictions on the cross-sectional dependence in $\tilde{\zeta}^{\ast}$, $\alpha$ or observed individual and class level characteristics.
 Latent baseline performance of student $i$ may be correlated with latent performance of student $j$ for a variety of reasons including similar educational, socioeconomic, cultural, religious or geographic backgrounds that may or may not be known to the analyst.
-We now describe the selection of students to classrooms.
+We now describe the selection of students into classrooms.
 This assignment process results in a reordering of the students characterized by a one-to-one mapping of the original index set to the new index set.
-Let $\tilde{y}_{t}^{*}=\left(\tilde{y}_{1t}^{*},...,\tilde{y}_{nt}^{*}\right)^{\prime}$ denote the vector of unobserved ability of students corresponding to the the original sample or population, let $S_{c}$ be the $n_{c}\times n\text{ }$ selection matrix that allocates students to classroom $c$, and let $S=\left[S_{1}^{\prime},...,S_{C}^{\prime}\right]^{\prime}$ be the selector matrix for all students.
+Let $\tilde{y}_{t}^{*}=\left(\tilde{y}_{1t}^{*},...,\tilde{y}_{nt}^{*}\right)^{\prime}$ denote the vector of unobserved test-taking ability of students taking test $t$.
+The vector $\tilde{y}_{t}^{*}$ is indexed corresponding to the the original sample or population.
+Let $S_{c}$ be the $n_{c}\times n\text{ }$ selection matrix that allocates students to classroom $c$, and let $S=\left[S_{1}^{\prime},...,S_{C}^{\prime}\right]^{\prime}$ be the selector matrix for all students.
 The set of all possible classroom allocations is denoted by $\mathscr{S}$.
@@
 $y_{ct}^{*}=S_{c}\tilde{y}_{t}^{*}.$
-Similarly, we arrange unobserved and observed student characteristics by classroom through the transformation $\zeta^{*}=\left(\zeta_{1}^{*},\ldots,\zeta_{n}^{*}\right)^{\prime}=S\tilde{\zeta}^{\ast}$, $v^{p}=\left(v_{1}^{p\prime},\ldots,v_{n}^{p\prime}\right)^{\prime}=S\tilde{v}^{p}$, $w_{t}^{p}=\left(w_{1t}^{p\prime},\ldots,w_{nt}^{p\prime}\right)^{\prime}=S\tilde{w}_{t}^{p}$ for $t=1,2$, and $X^{p}=\left(x_{1}^{p\prime},\ldots,x_{n}^{p\prime}\right)^{\prime}=S\tilde{X}^{p}=\left(v^{p},w_{1}^{p},w_{2}^{p}\right)$.
-Thus, the row vector $x_{i}^{p}$ now denotes observable personal characteristics for student $i$ in the sample ordered by classrooms.
-We allow for $S$, $\tilde{\zeta}^{\ast}$, $\tilde{y}_{t}^{*}$, for $t=1,2$, and $\alpha$ to be mutually correlated and to be correlated with $\tilde{X}^{p}$ as well as with observed class level characteristics.
+The latent performance vector $\zeta^{*}$ is defined analogously.
+We allow for $S$, $\tilde{\zeta}^{\ast}$, $\tilde{y}_{t}^{*}$, for $t=1,2$, and $\alpha$ to be mutually correlated and also to be correlated with observed individual and class level characteristics for the generalized model presented in Section 3.
 This assumption includes scenarios where $S$ is selected completely at random, scenarios where $S$ is selected based on student and class level characteristics, as well as scenarios where students respond to the allocation $S$.
@@
 The formulation of $\psi_{crt}\left(.\right)$ explicitly accounts for the possibility of peer effects by allowing for individual performance to depend on the characteristics of all peers.
-In the absence of peer effects the function $\psi_{crt}\left(.\right)$ simplifies to $y_{crt}=\psi_{crt}\left(y_{crt}^{*}\right).$ In other words, the performance of the $r$-th student in class $c$ only depends on own characteristics and class characteristics such as teachers and resources but not on the characteristics of other students in the class.
-In empirical applications it may be difficult to work with a non-parametric framework accounting for peer effects.
-For this reason we focus on linear production functions with linear peer effects.
-Let $v^{c}$ and $w_{t}^{c},t=1,2$, be matrices with row dimension $n$ that capture observed test invariant and test varying class level characteristics not related to student characteristics, such as the class size, the gender, education and experience of the teacher for subject $t$, and let $X=\left(v^{p},w_{1}^{p},w_{2}^{p},v^{c},w_{1}^{c},w_{2}^{c}\right)$ denote the matrix of observations on all covariates.
-At the class level let $y_{ct}=\left(y_{c1t},\ldots,y_{cn_{c}t}\right)^{\prime}$, $y_{ct}^{*}=\left(y_{c1t}^{*},\ldots,y_{cn_{c}t}^{*}\right)^{\prime}$, $v_{c}^{p}=\left(v_{c1}^{p\prime},...,v_{cn_{c}}^{p\prime}\right)^{\prime}$, $w_{ct}^{p}=\left(w_{c1t}^{p\prime},...,w_{cn_{c}t}^{p\prime}\right)^{\prime}$ with similar definitions for $v_{c}^{c}$ and $w_{ct}^{c}$, and let $\psi_{ct}=\left(\psi_{c1t},\ldots,\psi_{cn_{c}t}\right)^{\prime}$.
-We impose the following restriction on the education production function.
+In the absence of peer effects the function $\psi_{crt}\left(.\right)$ simplifies to $y_{crt}=\psi_{crt}\left(y_{crt}^{*}\right).$ In other words, the performance of the $r$-th student in class $c$ only depends on own characteristics and, for the generalized model presented below, on class characteristics such as teachers and resources but not on the characteristics of other students in the class.
+At the class level let $y_{ct}=\left(y_{c1t},\ldots,y_{cn_{c}t}\right)^{\prime}$, $y_{ct}^{*}=\left(y_{c1t}^{*},\ldots,y_{cn_{c}t}^{*}\right)^{\prime}$, and let $\psi_{ct}=\left(\psi_{c1t},\ldots,\psi_{cn_{c}t}\right)^{\prime}$.
+We postulate the following education production function, which accounts for possible peer effects.
 Assumption 1.
-Let $X_{ct}^{c}=\left(v_{c}^{c},w_{ct}^{c}\right)$, $\beta_{t}^{c}=\left(\beta_{vt}^{c\prime},\beta_{wt}^{c\prime}\right)^{\prime}$, $X_{ct}^{p}=\left(v_{c}^{p},w_{ct}^{p}\right)$, $\beta_{t}^{p}=\left(\beta_{vt}^{p\prime},\beta_{wt}^{p\prime}\right)^{\prime}$.
 Assume that
-$y_{ct}=\psi_{ct}\left(y_{ct}^{*}\right)=\alpha_{c}f_{t}\mathbf{1}_{c}+X_{ct}^{c}\beta_{t}^{c}+\left(I_{c}+\rho M_{c}\right)\left(X_{ct}^{p}\beta_{t}^{p}+y_{ct}^{*}\right),$ (1)
+$y_{ct}=\psi_{ct}\left(y_{ct}^{*}\right)=\alpha_{c}\mathbf{1}_{c}+\varphi_{c}+\left(I_{c}+\rho M_{c}\right)y_{ct}^{*},$ (1)
 where $\mathbf{1}_{c}=\left(1,....,1\right)^{\prime}$ is an $n_{c}\times 1$ vector and $M_{c}$ is some matrix of dimension $n_{c}\times n_{c}$ with zero diagonal elements and $\rho$ is a fixed parameter.
-The parameter $f_{t}$ captures test specific effects.
+The vector $\varphi_{c}=\left(\varphi_{c1},...,\varphi_{cn_{c}}\right)^{\prime}$ contains measurement errors.
 The part of the expression involving $M_{c}$ in (1) models peer effects.
@@
 [^3] In the usual terminology of Cliff-Ord (1973,1981) models, $M_{c}y_{ct}^{*}$ is called a spatial lag and $\rho$ the corresponding parameter.
-The vector $M_{c}X_{ct}^{p}\beta_{t}^{p}$ represents contextual peer effects.
-Linear specifications such as the one in Assumption 1 have been motivated as solutions to Nash games in the literature, see Calvó-Armengol, Patacchini, and Zenou (2009), Blume et al.
+Peer effects specifications such as the one in Assumption 1 have been motivated as solutions to Nash games in the literature, see Calvó-Armengol, Patacchini, and Zenou (2009), Blume et al.
 (2015) or Pereda-Fernández (2017).
-Pereda-Fernández (2017) considers a full information game of student and teacher effort where student and teacher quality is known to the players and that leads to an equation like (1).
-Thus, Assumption 1 can be justified by assuming that student ability $y_{ct}^{*}$ is unobserved by the analyst but known to classmates.
-Calvó-Armengol, Patacchini, and Zenou (2009) consider a network game in effort under full information where agents have linear quadratic utility.
-Therefore the educational production function (1) can be understood as an optimal response function where $y_{ct}^{*}$ now has the interpretation of effort, unobserved by the analyst but known to peers.
-Finally, Blume et al.
-(2015) analyze a Bayes-Nash equilibrium of agents choosing effort and facing a linear quadratic utility function that depends on their own as well as their peer’s effort; cp.
-also Cohen-Cole, Liu, and Zenou (2018) and Drukker, Egger, and Prucha (2023).
-Agents maximize expected utility conditional on publicly observable characteristics as well as private information that is only known to the individual player but not the other agents.
-They argue, see Blume et al.
-(2015) p.
-452, that their model is observationally equivalent to Calvó-Armengol, Patacchini, and Zenou (2009) as long as the row sums of $M_{c}$ are all equal, an assumption we impose below.
-In particular, as noted by Blume et al.
-(2015) p.
-458, $y_{ct}^{*}$ then can be interpreted as ability, with agents expectations absorbed into a fixed effect.
-This can be matched to our setting by assuming that $y_{ct}^{*}$ is ability that is unobserved by the analyst.
+Our model allows for measurement errors $\varphi_{c}$ affecting test scores.
+Consistent with our assumption that tests $t=1$ and $t=2$ measure essentially identical skills, measurement errors are assumed to be test invariant.
+We do not impose any restrictions on whether $\varphi_{c}$ is correlated with $y_{ct}^{*}$ or whether $\varphi_{c}$ is cross-sectionally correlated.
+The model in (1) is equivalent to $y_{ct}=\alpha_{c}\mathbf{1}_{c}+\left(I_{c}+\rho M_{c}\right)\left(y_{ct}^{*}+\varphi_{c}\right)$ where the measurement errors are in terms of ability $y_{ct}^{*}$ rather than outcomes $y_{ct}.$
 In line with the spatial and social network literature building on the models introduced by Cliff-Ord (1973,1981), the matrix $M_{c}$ in Assumption 1 could be an arbitrary weight matrix describing peer interaction subject to certain measurability assumptions specified below.
 For exposition and concreteness we focus on the case where average characteristics of peers determine peer effects.
-Define
-$M_{c}=\left(\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}-I_{c}\right)/\left(n_{c}-1\right)$
-where $I_{c}$ is the $n_{c}\times n_{c}$ identity matrix.
+Define $M_{c}=\left(\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}-I_{c}\right)/\left(n_{c}-1\right)$ where $I_{c}$ is the $n_{c}\times n_{c}$ identity matrix.
 Then, $M_{c}$ is the operator that computes the leave out average of peer characteristics.
-Combined with the functional form in Assumption 1 this leads, using scalar notation, to an education production function $y_{crt}=\psi_{crt}\left(y_{ct}^{*}\right)$ for observed outcomes in terms of unobservables and observables as given below ($t=1,2$)
-$y_{crt}=\alpha_{c}f_{t}+x_{ct}^{c}\beta_{t}^{c}+x_{crt}^{p}\beta_{t}^{p}+y_{crt}^{*}+\rho\left[(n_{c}-1)^{-1}\sum_{l=1,l\neq r}^{n_{c}}\left(x_{clt}^{p}\beta_{t}^{p}+y_{clt}^{*}\right)\right].$ (2)
-2.3 Identification
+Combined with the functional form in Assumption 1 this leads, using scalar notation, to an education production function $y_{crt}=\psi_{crt}\left(y_{ct}^{*}\right)$ for observed outcomes in terms of unobservables as given below ($t=1,2$)
+$y_{crt}=\alpha_{c}+\varphi_{cr}+y_{crt}^{*}+\rho\left[(n_{c}-1)^{-1}\sum_{l=1,l\neq r}^{n_{c}}y_{clt}^{*}\right].$ (2)
+2.3 Closely Related Skills
 In this paper we propose an identification strategy for the parameter $\rho$ that does not require randomization over $\mathscr{S}$ or variation in $S_{c}.$ Our argument proceeds in two steps.
-First, we propose a moment restriction on unobserved test taking ability $y_{it}^{*}$ and show that this restriction implies linear and quadratic moment restrictions for quantities we are able to observe in the data.
+First, we propose moment restrictions on unobserved test taking ability $y_{it}^{*}$ and show that these restrictions imply linear and quadratic moment restrictions for quantities we are able to observe in the data.
 We then show that $\rho$ can be identified by considering a quasi-differencing transformation of (2) and a GMM estimator that exploits the restrictions on the conditional mean and variance implied by our identifying assumption.
@@
 We formalize the information structure of our model in the following definition.
-Recall that $\zeta^{\ast}$ stands for latent test performance, effort or ability prior to taking tests $t=1,2.$ We use the notation $-t$ in the definition below to denote the other test.
+Recall that $\zeta^{*}=\left(\zeta_{1}^{*},\ldots,\zeta_{n}^{*}\right)^{\prime}=S\tilde{\zeta}^{\ast}$, which stands for unobserved individual characteristics, latent test performance, effort or ability prior to taking tests $t=1,2.$ We use the notation $-t$ in the definition below to denote the other test.
 We define the following:
 Definition 2.1.
-Let $y_{-it}^{*}=\left(y_{1t}^{*},...,y_{i-1,t}^{*},y_{i+1,t}^{*},...,y_{n,t}^{*}\right)^{\prime}$ and $y_{t}^{*}=\left(y_{1t}^{*},...,y_{nt}^{*}\right)^{\prime}$.
-Define the sigma fields (information sets)
-$\mathcal{F}_{n,i,t}$ $=\sigma\left(S,\alpha,X,z,\zeta^{\ast},y_{-t}^{\ast},y_{-i,t}^{*}\right)\text{\,for $t=1,2,$}$
-where $z$ denotes variables excluded from $X$ that may be used as instruments.
-Let $\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,X,z,\zeta^{\ast}\right)$ be the sigma field of all conditioning variables and let $\mathcal{Z}_{n}=\sigma\left(X,z\right)\subset\mathcal{Z}_{n}^{*}$ be the subset of observable information.
-The information set $\mathcal{Z}_{n}^{*}$ consists both of observable characteristics $X,$ classroom allocations $S,$ excluded variables $z$ that may be used as instruments as well as of unobserved variables $\zeta^{\ast}$ and $\alpha.$ The following proportionality restriction on the conditional means of the latent outcomes is at the core of our proposed method of identifying peer effects.
+Let $\varphi=\left(\varphi_{1}^{\prime},...,\varphi_{C}^{\prime}\right)^{\prime}$, $y_{-it}^{*}=\left(y_{1t}^{*},...,y_{i-1,t}^{*},y_{i+1,t}^{*},...,y_{n,t}^{*}\right)^{\prime}$ and $y_{t}^{*}=\left(y_{1t}^{*},...,y_{nt}^{*}\right)^{\prime}$.
+Define the sigma fields (information sets) $\mathcal{F}_{n,i,t}=\sigma\left(S,\alpha,\varphi,\zeta^{\ast},y_{-t}^{\ast},y_{-i,t}^{*}\right)\text{\,for $t=1,2$}$.
+Let $\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,\varphi,\zeta^{\ast}\right)$ be the sigma field of all conditioning variables.
+The information set $\mathcal{Z}_{n}^{*}$ consists of classroom allocations $S,$ as well as of unobserved variables $\zeta^{\ast}$, $\alpha$ and $\varphi.$ The following restriction on the conditional means of the latent outcomes is at the core of our proposed method of identifying peer effects.
+We note that the assumption maintains implicitly that the two test scores are measured on the same scale.
 Assumption 2.
-Assume that
-$E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=\kappa_{i}f_{t}$
-where $\kappa_{i}$ is a random variable that is invariant to $t$ and is measurable with respect to $\mathcal{Z}_{n}^{*}$ and $f_{t}$ is a fixed parameter that only varies with $t$.
-In addition, assume that the random variable $y_{it}^{*}$ is either bounded, or there exists an $\eta>0$ and a random variable $y$ such that $\left|y_{it}^{*}\right|+\left|\kappa_{i}f_{t}\right|\leq y$ with $E\left[\left|y\right|^{4+\eta}|\mathcal{Z}_{n}^{*}\right]\leq K_{y}<\infty$ for all $i$ and $t.$ In addition, $E\left[\left|\alpha_{c}\right|^{2+\eta}\right]\leq K_{\alpha}<\infty$ for all $c.$
-Without loss of generality we normalize $f_{2}=1$.
-The assumption includes the special case where $f_{t}=1$ for $t=1,2$, which is relevant in situations where tests $t$ measure closely related skills.
-The interpretation of the condition is that conditional on a hypothetical or actual baseline of test results $\zeta^{*}$, as well as information about group formation $S$, unobserved class level characteristics $\alpha$, as well as other observed characteristics $X$ and $z$, additional test results do not change expected performance, except maybe for a common scale factor $f_{t}$ accounting for systematic differences between tests affecting all test takers in the same fashion.
-[^4] For simplicity we assume that $f_{t}$ in Assumptions 1 and 2 are the same.
-This restriction can be relaxed if three test outcomes are used.
-The factors $f_{t}$ can be treated as random rather than fixed at the expense of slightly more complex asymptotic arguments needed for the analysis of statistical inference, see Kuersteiner and Prucha (2020).
-It is in this sense that tests $t=1$ and $t=2$ measure essentially identical skills.
-Note that $E\left[y_{it}^{*}|\mathcal{Z}_{n}^{*}\right]=\kappa_{i}f_{t}$ by iterated expectations and $\mathcal{Z}_{n}^{*}\subset\mathcal{F}_{n,i,t}.$
+Assume that $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=\kappa_{i}$ where $\kappa_{i}$ is a random variable that is invariant to $t$ and is measurable with respect to $\mathcal{Z}_{n}^{*}$ In addition, assume that the random variable $y_{it}^{*}$ is either bounded, or there exists an $\eta>0$ and a random variable $y$ such that $\left|y_{it}^{*}\right|+\left|\kappa_{i}\right|\leq y$ with $E\left[\left|y\right|^{4+\eta}|\mathcal{Z}_{n}^{*}\right]\leq K_{y}<\infty$ for all $i$ and $t.$ In addition, $E\left[\left|\alpha_{c}\right|^{2+\eta}\right]\leq K_{\alpha}<\infty$ and $E\left[\left|\varphi_{cr}\right|^{2+\eta}\right]\leq K_{\varphi}<\infty$ for all $c$ and $r$.
+The interpretation of the condition is that conditional on a hypothetical or actual baseline of test results $\zeta^{*}$, as well as information about group formation $S$, unobserved class level characteristics $\alpha$, as well as measurement errors $\varphi$ , additional test results do not change expected performance.
+A second implication of Assumption 2 is that expected performance does not depend on the specific test $t.$ In other words, $\kappa_{i}$ only varies by individual $i$ but not by test $t.$ It is in this sense that tests $t=1$ and $t=2$ measure essentially identical skills.
+Note that $E\left[y_{it}^{*}|\mathcal{Z}_{n}^{*}\right]=\kappa_{i}$ by iterated expectations and $\mathcal{Z}_{n}^{*}\subset\mathcal{F}_{n,i,t}.$ If $y_{it}^{*}$ is correlated with measurement errors $\varphi_{c}$ then this dependence is absorbed into $\kappa_{i}$.
+The restriction Assumption 2 imposes on the relationship between measurement error and $y_{it}^{*}$ is that any type of mean dependence is common to both tests $t$.
+A special case covered by Assumption 2 is that $y_{it}^{*}$ may be independent of measurement error.
 Relative to a baseline represented by $\mathcal{Z}_{n}^{*}$ the latent test results $y_{it}^{*}$ are in essence as good as randomly assigned.
@@
 when $\rho\neq 0.$
-The variable $\kappa_{i}$ is a function of the entire cross section of latent performance measures $\zeta^{*}$, of the allocation $S$ and observed and unobserved characteristics.
+The variable $\kappa_{i}$ is a function of the entire cross section of latent performance measures $\zeta^{*}$, of the allocation $S$ and unobserved characteristics that include classroom effects and measurement errors $\varphi$.
 It is generally cross-sectionally dependent in ways we do not restrict or specify and captures such common effects as the socioeconomic background of students and classmates, their family background including parental education and support, their interaction with peers, the exposure to their teachers and resources available in their classroom.
 It also depends on a student’s own ability.
-The restriction we impose is that these factors do not change between tests $t=1$ and $t=2$, except maybe for a change in scale $f_{t}.$
+The restriction we impose is that these factors do not change between tests $t=1$ and $t=2$.
 Baseline performance $\zeta^{*}$ does not necessarily have to be observed or realized and $\zeta^{*}$ can alternatively be interpreted as latent ability prior to taking the actual tests $t=1,2$.
-In this sense, $\kappa_{i}f_{t}$ and $E\left[y_{it}^{*}|\mathcal{Z}_{n}^{*}\right]$ are also unobserved.
-A scenario where Assumption 2 is realistic arises when the same type of test is taken multiple times without additional training between iterations of the test, or when several tests are given that focus on related skills, as is the case in Project STAR data for grades K-3 and tests for reading, word study and listening skills.
+In this sense, $\kappa_{i}$ and $E\left[y_{it}^{*}|\mathcal{Z}_{n}^{*}\right]$ are also unobserved.
+A scenario where Assumption 2 is realistic arises when the same type of test is taken multiple times without additional training between iterations of the test, or when several tests are given that focus on related skills, as is the case in Project STAR data for grades K-3 and tests for reading, and word study skills.
 [^4] Another example of a test with multiple scores is the Comprehensive Testing Program (CTP) administered by the Educational Records Bureau and given to Grades 1-11.
@@
 The assumption is less plausible if comparisons are attempted across different subjects in a high school or college setting, for tests given at different times during the school year, or measuring different skills in a professional or team setting.
-In the latter case, additional factors can be introduced if multiple measurements per skill category are available.
-[^6] The multi-factor case is discussed in detail in Kuersteiner and Prucha (2020).
-The following theorem formally establishes a representation for $y_{it}^{*}$ that decomposes $y_{it}^{*}$ into its conditional mean, $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]$, that is dependent cross-sectionally as well as across tests, and into uncorrelated idiosyncratic noise $u_{it}$.
-The theorem formalizes the definition of $u_{it}$
+The following theorem formally establishes a decomposition of $y_{it}^{*}$ into its conditional mean, $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]$, that is dependent cross-sectionally as well as across tests, and into uncorrelated idiosyncratic noise $u_{it}$.
 Theorem 2.1.
-Suppose Assumption 2 holds. Define
-$u_{it}=y_{it}^{*}-E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=y_{it}^{*}-\kappa_{i}f_{t},$
-then by construction
-$E\left[u_{it}|\mathcal{F}_{n,i,t}\right]=0\ \ \text{and}\ \ \operatorname{Cov}\left(u_{it}u_{js}|\mathcal{F}_{n,i,t}\right)=0$
-for any $j\neq i$ or $t\neq s$.
-Furthermore, let $\mathcal{G}_{n,i,t}$ be any sigma field with $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$, then $E\left[u_{it}|\mathcal{G}_{n,i,t}\right]=0$ and $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{G}_{n,i,t}\right)=0$ for any $j\neq i$ or $t\neq s$, and $\operatorname{Cov}\left(u_{it},\kappa_{i}f_{t}\right)=0,\operatorname{Cov}\left(u_{it},X\right)=0$, $\operatorname{Cov}\left(u_{it},z\right)=0$ etc.
-If $E\left[y_{it}^{*}-\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]\neq 0$ for some $\mathcal{G}_{n,i,t}$, then Assumption 2 cannot hold.
-Examples for the information sets $\mathcal{G}_{n,i,t}$ in the above theorem are $\mathcal{G}_{n,i,t}=\sigma\left(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*s}\right)$ where $y_{-i,t}^{*s}$ is a subvector of $y_{-i,t}^{*}$ or $\mathcal{G}_{n,i,t}=\mathcal{Z}_{n}^{*}.$ The above theorem makes clear that the main restriction we impose on $y_{it}^{*}$ is the assumption that latent test performance is not predictable conditional on observable characteristics and unobserved baseline ability.
-The properties $E\left[u_{it}|\mathcal{Z}_{n}^{*}\right]=0,$ $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{Z}_{n}^{*}\right)=0$ for any $j\neq i$ or $t\neq s$ and $\operatorname{Cov}\left(u_{it},\kappa_{i}f_{t}\right)=0$, $\operatorname{Cov}\left(u_{it},X\right)=0$, $\operatorname{Cov}\left(u_{it},z\right)=0$ follow directly from Assumption 2 and are not additional assumptions imposed on the distribution of $y_{it}^{*}.$ In other words, Assumption 2 is necessary and sufficient for the decomposition of $y_{it}^{*}=\kappa_{i}f_{t}+u_{it}$ with uncorrelated errors $u_{it}$.
-We do not attach a specific economic or behavioral interpretation to the decomposition of $y_{it}^{*}$ into $\kappa_{i}f_{t}$ and $u_{it}$.
+Suppose Assumption 2 holds.
+Define $u_{it}=y_{it}^{*}-E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=y_{it}^{*}-\kappa_{i},$ then by construction $E\left[u_{it}|\mathcal{F}_{n,i,t}\right]=0$ and $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{F}_{n,i,t}\right)=0$ for any $j\neq i$ or $t\neq s$.
+Furthermore, let $\mathcal{G}_{n,i,t}$ be any sigma field with $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$, then $E\left[u_{it}|\mathcal{G}_{n,i,t}\right]=0$ and $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{G}_{n,i,t}\right)=0$ for any $j\neq i$ or $t\neq s$, and $\operatorname{Cov}\left(u_{it},\kappa_{i}\right)=0$ etc.
+If $E\left[y_{it}^{*}-\kappa_{i}|\mathcal{G}_{n,i,t}\right]\neq 0$ for some $\mathcal{G}_{n,i,t}$, then Assumption 2 cannot hold.
+Examples for the information sets $\mathcal{G}_{n,i,t}$ in the above theorem are $\mathcal{G}_{n,i,t}=\sigma\left(S,\alpha,\varphi,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*s}\right)$ where $y_{-i,t}^{*s}$ is a subvector of $y_{-i,t}^{*}$ or $\mathcal{G}_{n,i,t}=\mathcal{Z}_{n}^{*}.$ The above theorem makes clear that the main restriction we impose on $y_{it}^{*}$ is the assumption that latent test performance is not predictable conditional on observable characteristics and unobserved baseline ability.
+The properties $E\left[u_{it}|\mathcal{Z}_{n}^{*}\right]=0,$ $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{Z}_{n}^{*}\right)=0$ for any $j\neq i$ or $t\neq s$ and $\operatorname{Cov}\left(u_{it},\kappa_{i}\right)=0$ follow directly from Assumption 2 and are not additional assumptions imposed on the distribution of $y_{it}^{*}.$ In other words, Assumption 2 is necessary and sufficient for the decomposition of $y_{it}^{*}=\kappa_{i}+u_{it}$ with uncorrelated errors $u_{it}$.
+Note that $M_{c}$ is implicitly a function of class assignment $S$ and thus orthogonal to $u_{it}.$
+We do not attach a specific economic or behavioral interpretation to the decomposition of $y_{it}^{*}$ into $\kappa_{i}$ and $u_{it}$.
 It is merely a statistical representation for the purpose of isolating variation in test scores that is correlated within groups but invariant across tests and variation that is idiosyncratic to the individual student and test.
-Normalizing $f_{2}=1$, which is without loss of generality, leads to an immediate consequence of the decomposition: The quasi-difference
-$y_{i1}^{*}-y_{i2}^{*}f_{1}=u_{i1}-u_{i2}f_{1}$ (3)
-is mean zero and uncorrelated across $i$ conditional on $\mathcal{Z}_{n}^{*}$.
-We exploit these two moment restrictions to formulate our GMM estimator.
-They are at the heart of our identification result which is formally presented in Section 4.
-Our framework is in contrast with much of the econometrics literature on the identification of peer effects where assumptions about the orthogonality of $u_{it}$ and $\kappa_{i}f_{t}$ and covariates are usually imposed directly.
+Let $\kappa_{c}=\left(\kappa_{c1},\ldots.\kappa_{cn_{c}}\right)^{\prime}$, $u_{ct}=(u_{c1t},\ldots,u_{cn_{c}t})^{\prime}$ and combine Assumptions 1 and 2 such that
+$y_{ct}=\mu_{c}^{*}+\left(I_{c}+\rho M_{c}\right)u_{ct}$ (3)
+with the unobserved class and individual effect defined as
+$\mu_{c}^{\ast}=\alpha_{c}\mathbf{1}_{c}+\varphi_{c}+\left(I_{c}+\rho M_{c}\right)\kappa_{c}.$ (4)
+By differencing (3), which eliminates $\mu_{c}^{\ast}$, we obtain
+$y_{c1}-y_{c2}$ $=$ $\left(I_{c}+\rho M_{c}\right)(u_{c1}-u_{c2}).$ (5)
+As seen from the above, differential measures of test scores purge performance measures from common unobserved influences captured by $\mu_{c}^{*}$ while still maintaining relevant information that can shed light on the existence of peer effects.
+Let $y_{t}=\left(y_{1t}^{\prime},\ldots,y_{Ct}^{\prime}\right)^{\prime}$ and $u_{t}=\left(u_{1t}^{\prime},\ldots,u_{Ct}^{\prime}\right)^{\prime}$ for $t=1,2$.
+In similar fashion define $M=\operatorname{diag}_{c=1}^{C}\left(M_{c}\right)$.
+Then we can write (5) for the entire sample more compactly as
+$y_{1}=y_{2}+(I+\rho M)(u_{1}-u_{2}).$ (6)
+Our framework is in contrast with much of the econometrics literature on the identification of peer effects where assumptions about the orthogonality of $u_{it}$ and $\kappa_{i}$ and covariates are usually imposed directly.
 These unobservable quantities often are given economic interpretations.
@@
 In such a scenario individual and classroom specific effects $\tilde{\kappa}_{i}$ are not independent of classroom allocations.
-2.4 Estimation
-Our estimators are using differential, or quasi-differential test scores, $y_{i1}-y_{i2}f_{1}$.
-These differential measures purge performance measures from common unobserved influences captured by $\alpha_{c}$ and $\kappa_{c}=\left(\kappa_{c1},\ldots.\kappa_{cn_{c}}\right)^{\prime}=S_{c}\tilde{\kappa}$ and shed light on the existence of peer effects.
-Define the unobserved class and individual effect as
-$\mu_{c}^{\ast}=\alpha_{c}\mathbf{1}_{c}+\left(I_{c}+\rho M_{c}\right)\kappa_{c}$ (4)
-such that Assumptions 1, 2 and Theorem 2.1 imply the empirical specification
-$y_{ct}$ $=\mu_{c}^{*}f_{t}$ $+v_{c}^{c}\beta_{vt}^{c}+w_{ct}^{c}\beta_{wt}^{c}+\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{vt}^{p}+w_{ct}^{p}\beta_{wt}^{p}+u_{ct}).$ (5)
-By quasi-differencing (5), which eliminates classroom and individual fixed effects, we obtain
+3 Estimation with Covariates
+3.1 Model with Covariates
+We extend the model in Section 2 by adding observable characteristics and by specifying the test average scores in a more flexible way.
+We collect test invariant and test varying student level characteristics for students in class $c$ in the matrices $v_{c}^{p}$ and $w_{ct}^{p},t=1,2$ with row dimension $n_{c}$.
+We also collect observed test invariant and test varying class level characteristics for class $c$ which are not related to student characteristics in the matrices $v_{c}^{c}$ and $w_{ct}^{c},t=1,2$, with row dimension $n_{c}$.
+Class level characteristics may include class size, the gender, education and experience of the teacher for subject $t$, etc.
+In the following let $v^{p}=\left(v_{1}^{p\prime},\ldots,v_{C}^{p\prime}\right)^{\prime}$, $w_{t}^{p}=\left(w_{1t}^{p\prime},\ldots,w_{Ct}^{p\prime}\right)^{\prime}$, $v^{c}=\left(v_{1}^{c\prime},\ldots,v_{C}^{c\prime}\right)^{\prime}$, $w_{t}^{c}=\left(w_{1t}^{c\prime},\ldots,w_{Ct}^{c\prime}\right)^{\prime}$, $t=1,2,$ and let $X=\left(v^{p},w_{1}^{p},w_{2}^{p},v^{c},w_{1}^{c},w_{2}^{c}\right)$ denote the matrix of observations on all covariates.
+Below is an extended version of Assumption 1.
+Assumption 3.
+Let $X_{ct}^{c}=\left(v_{c}^{c},w_{ct}^{c}\right)$, $\beta_{t}^{c}=\left(\beta_{vt}^{c\prime},\beta_{wt}^{c\prime}\right)^{\prime}$, $X_{ct}^{p}=\left(v_{c}^{p},w_{ct}^{p}\right)$, $\beta_{t}^{p}=\left(\beta_{vt}^{p\prime},\beta_{wt}^{p\prime}\right)^{\prime}$.
+Assume that
+$y_{ct}=\psi_{ct}\left(y_{ct}^{*}\right)=\alpha_{c}\mathbf{1}_{c}+\varphi_{c}+X_{ct}^{c}\beta_{t}^{c}+\left(I_{c}+\rho M_{c}\right)\left(X_{ct}^{p}\beta_{t}^{p}+y_{ct}^{*}\right),$ (7)
+where $\mathbf{1}_{c}=\left(1,....,1\right)^{\prime}$ is an $n_{c}\times 1$ vector and $M_{c}$ is some matrix of dimension $n_{c}\times n_{c}$ with zero diagonal elements and $\rho$ is a fixed parameter.
+The vector $M_{c}X_{ct}^{p}\beta_{t}^{p}$ represents contextual peer effects. We update the filtrations given in Definition 2.1 to include observable covariates.
+Definition 3.1.
+Define the sigma fields $\mathcal{F}_{n,i,t}=\sigma\left(S,\alpha,\varphi,X,z,\zeta^{\ast},y_{-t}^{\ast},y_{-i,t}^{*}\right)\text{\,for $t=1,2,$}$ where $z$ denotes variables excluded from $X$ that may be used as instruments.
+Let $\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,\varphi,X,z,\zeta^{\ast}\right)$ be the sigma field of all conditioning variables and let $\mathcal{Z}_{n}=\sigma\left(X,z\right)\subset\mathcal{Z}_{n}^{*}$ be the subset of observable information.
+We maintain Assumption 2 without change, but using the expanded versions of $\mathcal{F}_{n,i,t}$, $\mathcal{Z}_{n}^{*}$ and $\mathcal{Z}_{n}$ in Definition 3.1.
+Then, under Assumptions 2 and 3 the conclusions of Theorem 2.1 continue to hold.
+Furthermore, we have $\operatorname{Cov}\left(u_{it},X\right)=0$ and $\operatorname{Cov}(u_{it},z)=0$.
+We note that Assumptions 2, 3 and Theorem 2.1 imply the peer effects model with covariates for test score outcomes
+$y_{ct}$ $=\mu_{c}^{*}$ $+v_{c}^{c}\beta_{vt}^{c}+w_{ct}^{c}\beta_{wt}^{c}+\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{vt}^{p}+w_{ct}^{p}\beta_{wt}^{p}+u_{ct}).$ (8)
+We note that (8) imposes the restriction that average residualized test scores for both tests $t=1$ and $t=2$ are the same.
+This is in line with our core identification strategy of the two tests testing essentially identical skills.
+3.2 Empirical Specification
+[…]
+$y_{ct}=\mu_{c}^{*}f_{t}+v_{c}^{c}\beta_{vt}^{c}+w_{ct}^{c}\beta_{wt}^{c}+\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{vt}^{p}+w_{ct}^{p}\beta_{wt}^{p}+u_{ct}),$ (9)
+and where our assumptions imply that $f_{t}=1$ for $t=1,2.$ Empirically we treat $f_{1}$ as a free parameter to be estimated, while we maintain $f_{2}=1$ as a normalization.
+An estimated value of $f_{1}$ that differs from $f_{1}=1$ is an indication, although not a rigorous specification test, that Assumption 2 is violated in the data.
+By quasi-differencing (9), which eliminates $\mu_{c}^{\ast}$, we obtain
 $y_{c1}-y_{c2}f_{1}$ $=$ $v_{c}^{c}(\beta_{v1}^{c}-f_{1}\beta_{v2}^{c})+w_{c1}^{c}\beta_{w1}^{c}-f_{1}w_{c2}^{c}\beta_{w2}^{c}$
 $+$ $\left(I_{c}+\rho M_{c}\right)\left(v_{c}^{p}(\beta_{v1}^{p}-f_{1}\beta_{v2}^{p})+w_{c1}^{p}\beta_{w1}^{p}-f_{1}w_{c2}^{p}\beta_{w2}^{p}\right)+\left(I_{c}+\rho M_{c}\right)(u_{c1}-f_{1}u_{c2}).$
-We treat $f_{t}$ as an unknown parameter to be estimated that accounts for differences in average test scores between tests $1$ and $2$.
 Collecting terms with $f_{1}$ leads to an interpretation of (3.2) where $y_{c2}-v_{c}^{c}\beta_{v2}^{c}-w_{c2}^{c}\beta_{w2}^{c}-\left(I_{c}+\rho M_{c}\right)(v_{c}^{p}\beta_{v2}^{p}+w_{c2}^{p}\beta_{w2}^{p})$ is used to control for unobserved $\mu_{c}^{\ast}$ similar to a control function approach.
@@
 We next write (3.2) more compactly by stacking observations across classrooms.
-Let $y_{t}$ , $v^{p}$ and $w_{t}^{p}$ be as defined earlier.
-In similar fashion define $v^{c}$, $w_{t}^{c}$ and $u_{t}$, and let $M=\operatorname{diag}_{c=1}^{C}\left(M_{c}\right)$.
-Let $\underline{X}=\left(v^{c},w_{1}^{c},w_{2}^{c},v^{p},w_{1}^{p},w_{2}^{p},Mv^{p},Mw_{1}^{p},Mw_{2}^{p}\right)$ and define
+Let $v^{p},v^{c}$ and $y_{t},u_{t},w_{t}^{p},w_{t}^{c}$, $t=1,2$, as well as $M$ be as defined above.
+Furthermore, let $\underline{X}=\left(v^{c},w_{1}^{c},w_{2}^{c},v^{p},w_{1}^{p},w_{2}^{p},Mv^{p},Mw_{1}^{p},Mw_{2}^{p}\right)$ and define
 $\delta(f_{1},\rho,\beta)=\left(\delta_{v}^{c\prime},\beta_{w1}^{c\prime},-f_{1}\beta_{w2}^{c\prime},\delta_{v}^{p\prime},\beta_{w1}^{p\prime},-f_{1}\beta_{w2}^{p\prime},\rho\delta_{v}^{p\prime},\rho\beta_{w1}^{p\prime},-\rho f_{1}\beta_{w2}^{p\prime}\right)^{\prime},$ (11)
-where $\delta_{v}^{c}=\beta_{v1}^{c}-f_{1}\beta_{v2}^{c}$, $\delta_{v}^{p}=\beta_{v1}^{p}-f_{1}\beta_{v2}^{p}$ and $\beta$ is the vector of all $\beta$s above, then we can write (2.4) for the entire sample more compactly as
-$y_{1}=f_{1}y_{2}+\underline{X}\delta+(I+\rho M)(u_{1}-f_{1}u_{2})$ (8)
-where $\delta=\beta_{1}-f_{1}\beta_{2}$ with $\beta_{t}$ collecting the corresponding coefficients for each variable in $\underline{X}$ and set to 0 if a variable is absent in (5) for given $t$.
-Similarly we can define $\mu^{\ast}=(\mu_{1}^{\ast\prime},....,\mu_{C}^{\ast\prime})^{\prime}$ and write (5) compactly as
-$y_{t}=\mu^{\ast}f_{t}+\text{$\underline{X}$$\beta_{t}+(I+\rho M)u_{t}$},$ (9)
-With Assumptions 1 and 2, and standard regularity conditions on $\underline{X}$, we can identify $\delta$, which is sufficient for the identification of $(\rho,f_{1})$.
+where $\delta_{v}^{c}=\beta_{v1}^{c}-f_{1}\beta_{v2}^{c}$, $\delta_{v}^{p}=\beta_{v1}^{p}-f_{1}\beta_{v2}^{p}$ and $\beta$ is the vector of all $\beta$s above.
+Then we can write (3.2) for the entire sample more compactly as
+$y_{1}=f_{1}y_{2}+\underline{X}\delta+(I+\rho M)(u_{1}-f_{1}u_{2}).$ (12)
+where $\delta=\beta_{1}-f_{1}\beta_{2}$.
+Here $\beta_{t}$ collects the corresponding coefficients for each variable in $\underline{X}$ with elements of $\beta_{t}$ set to 0 if a variable is absent in (9) for given $t$.
+Let $\mu^{\ast}=(\mu_{1}^{\ast\prime},....,\mu_{C}^{\ast\prime})^{\prime}$.
+Then we can write (9) compactly as
+$y_{t}=\mu^{\ast}f_{t}+\text{$\underline{X}$$\beta_{t}+(I+\rho M)u_{t}$}.$ (13)
+With Assumptions 2 and 3, and full column rank conditions on $\underline{X}$, we can identify $\delta$, which is sufficient for the identification of $(\rho,f_{1})$.
 The parameters $\beta_{w1}^{c}$ and $\beta_{w1}^{p}$ can be directly recovered from $\delta$, and $\beta_{w2}^{c}$ and $\beta_{w2}^{p}$ can be identified when $f_{1}\neq 0$.
@@
 An example of restrictions where the parameters $\beta_{v1}^{c}$ and $\beta_{v2}^{c}$ are identified arises when $f_{1}\neq 1$ and $\beta_{v1}^{c}=\beta_{v2}^{c}$.
+3.3 Estimation
 In empirical applications it is sometimes sufficient to estimate $\beta$ or even just $\delta,$ for example when $\underline{X}$ contains contextual peer effects.
@@
 More generally, exogenous covariates that vary at the classroom level but are excluded from $\underline{X}$ are valid instruments by Lemma 5.1.
-Valid standard errors are obtained by clustering by classrooms.
-For estimation of the full set of parameters including the peer effect parameter $\rho$, we use the efficient GMM estimator developed in Kuersteiner and Prucha (2020) for a fairly general class of networks covering both social and spatial network effects.
-Identification of the parameters $\rho,f_{1}$ and $\delta$ follows from Assumptions 1 and 2 as well as classroom level homoskedasticity and some restrictions on the parameter space that are detailed in Section 4.
-Feasible efficient inference requires some additional restrictions on conditional variances that we now describe.
-We allow for some forms of heteroskedasticity for $u_{it}$ across tests and by class type.
-Let $\tau_{c}$ be a categorical variable for class type with $\tau_{c}=j$ if classroom $c$ is of type $j\in\left\{1,...,J\right\}.$ Examples of types are small and regular classes, or classes located in urban, suburban or rural areas.
-Assumption 3.
-Let $u_{it}$ be as defined in Theorem 2.1 and let $i=n_{1}+\ldots+n_{c-1}+r$ with $1\leq r\leq n_{c}$, i.e., $i$ refers to student $r$ in class c.
-Let $u_{-i,t}=\left(u_{1,t},...,u_{i-1,t},u_{i+1,t},...,u_{n,t}\right)$.
-Define the filtration
-$\mathcal{B}_{n,i,t}=\sigma\left(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t}\right)\text{\,for $t=1,2$}$
-where $u_{-t}$ represents information about the other test. Then the conditional variance of $u_{it}$ is given by
-$E\left[u_{it}^{2}|\mathcal{B}_{n,i,t}\right]=\sigma_{t}^{2}\rho_{\tau_{c}}^{2},$
-where $\tau_{c}$ is $\mathcal{Z}_{n}$-measurable, and $c_{u}\leq\sigma_{t}^{2},\rho_{j}^{2}\leq C_{u}$ for some finite positive bounds.
-Assumption 3 implies that for $t=1,2$, $\Sigma_{t}\equiv E\left[u_{t}u_{t}^{\prime}|\mathcal{Z}_{n}^{\ast}\right]=\sigma_{t}^{2}\operatorname{diag}_{c=1}^{C}\left(\rho_{\tau_{c}}^{2}I_{c}\right)$.
-Observing that $f_{t}$ is a fixed constant, conditional on $S,\zeta^{*},\alpha,X$ and $z$, knowledge of $y_{it}^{*}$ is equivalent with knowledge of $u_{it}$.
-Thus we have the following equivalence of information sets [^7] See the proof of Lemma B.3 for more detail.
-=$\mathcal{F}{}_{n,i,t}=\mathcal{B}_{n,i,t}\,\text{for}\,t=1,2.$
-Assumption 3 imposes a similar factor structure on conditional variances as does Assumption 2 for conditional means.
-Below we illustrate the nature of our efficient quasi-differencing transformation.
-By Theorem 2.1 the errors $u_{it}$ are mutually uncorrelated over tests as well as cross-sectionally.
-Consequently the variance covariance matrix of $u_{1}-f_{1}u_{2}$ is given by,
-$\Omega=\operatorname{Var}\left(u_{1}-f_{1}u_{2}|\mathcal{Z}_{n}^{*}\right)=\Sigma_{1}+f_{1}^{2}\Sigma_{2}=\left(\sigma_{1}^{2}+f_{1}^{2}\sigma_{2}^{2}\right)\operatorname{diag}_{c=1}^{C}\left(\rho_{\tau_{c}}^{2}I_{c}\right).$
-Define $\gamma_{j}^{2}=\left(\sigma_{1}^{2}+f_{1}^{2}\sigma_{2}^{2}\right)\rho_{j}^{2}$, for $j=1,...,J$ and $\gamma=\left(\gamma_{1},...,\gamma_{J}\right)^{\prime}$.
-Then $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right).$
-The disturbance term in (8) is of the form $(I+\rho M)(u_{1}-f_{1}u_{2})$.
-Denote the parameter $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$.
-Premultiplying (8) by $\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$ then yields the following model
-$y^{+}(\theta,\gamma)=X^{+}(\theta,\gamma)\delta+u^{+}$ (10)
-where,
-$y^{+}(\theta,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}(y_{1}-f_{1}y_{2}),$ (11)
-$X^{+}(\theta,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}\underline{X},$ (12)
-$u^{+}=\Omega(\gamma)^{-1/2}(u_{1}-f_{1}u_{2})$ (13)
-and $\operatorname{Var}\left(u^{+}|\mathcal{Z}_{n}^{*}\right)=I_{n}$. In light of (10)-(13) we define
-$u^{+}(\theta,\gamma)$ $=y^{+}(\theta,\gamma)-X^{+}(\theta,\gamma)\delta$
-$=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}(y_{1}-f_{1}y_{2}-\text{$\underline{X}$$\delta$})$ (14)
-and observe that $u^{+}(\theta,\gamma)=u^{+}$ at the true parameters.
-The transformation leading to $u^{+}$ consists of three components.
-Quasi-differencing eliminates classroom and individual effects.
-The operator $(I+\rho M)^{-1}$ removes cross-sectional correlation and can be understood as a form of spatial Cochrane-Orcutt transformation.
-Finally, the operator $\Omega(\gamma)^{-1/2}$ scales the spatially uncorrelated residuals to unit variance in the case of heteroskedasticity.
-We next provide an outline of our GMM estimation methodology.
-For clarity we explicitly denote in the following the true parameter vectors as $\theta_{0}=(\rho_{0},f_{1,0},\delta_{0}^{\prime})^{\prime}$ and $\gamma_{0}=(\gamma_{1,0},...,\gamma_{J,0})^{\prime}$.
-Rigorous consistency and asymptotic normality results as well as additional assumptions needed to establish these properties will be given in Section 4 below.
-Our GMM estimator is based on both linear and quadratic moment conditions.
 We consider instrument matrices $H$ and $A$ and require that all elements of $H$ and $A$ are functions of the observable variables that are measurable with respect to $\mathcal{Z}_{n}$.
@@
 In particular, we set $H=[\underline{X},z]$ as an $n\times q$ instrument matrix.
-The moment function consists of a set of linear and quadratic moments for the transformed residuals [^8] The specification can be readily extended to include several quadratic moment conditions along the lines of, e.g., Kuersteiner and Prucha (2020).
-$m_{n}(\theta,\gamma)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+\prime}(\theta,\gamma)Au^{+}(\theta,\gamma)\end{array}\right).$
-Under the maintained assumptions of this paper the results in Kuersteiner and Prucha (2020), Theorem 1, imply that the linear moment function is uncorrelated with the quadratic moments. We have
-$E\left[m_{n}(\theta_{0},\gamma_{0})m_{n}(\theta_{0},\gamma_{0})^{\prime}\right]$ $=\frac{1}{n}\left(\begin{array}[]{cc}E\left[H^{\prime}H\right]&0\\ 0&2E\left[\operatorname{tr}\left(A^{2}\right)\right]\end{array}\right)\equiv\Xi_{n}.$ (17)
-[…]
-$Q_{n}\left(\theta,\gamma\right)$ $=\frac{1}{n}m_{n}(\theta,\gamma)^{\prime}\hat{\Xi}_{n}^{-1}m_{n}(\theta,\gamma)$
-$=\frac{1}{n}u^{+}(\theta,\gamma)^{\prime}H\left(H^{\prime}H\right)^{-1}H^{\prime}u^{+}(\theta,\gamma)+\frac{1}{n}\frac{\left(u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\right)^{2}}{2\operatorname{tr}\left(A^{2}\right)}$ (18)
-such that the GMM estimator for $\theta_{0}$ is defined as $\hat{\theta}\left(\gamma\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\gamma\right)$ for a fixed value of $\gamma.$ In applications $\gamma$ is replaced with a first step consistent estimator.
-We discuss consistent estimation of $\gamma$ and feasible versions of $\hat{\theta}\left(\gamma\right)$ in Section 4.
-A simple two-step procedure to obtain initial estimates for $\delta$, $f_{1}$, and $\rho$ may sometimes be sufficient and can be used to estimate $\gamma.$ Such a two-step procedure can be readily defined adapting ideas from the literature on spatial Cliff-Ord models; cp.
+A simple two-step procedure is sufficient to obtain consistent but inefficient estimates for $\delta$, $f_{1}$, and $\rho$.
+Such a two-step procedure can be readily defined adapting ideas from the literature on spatial Cliff-Ord models; cp.
 e.g., Kelejian and Prucha (1998, 1999).
 The first step relies on linear moment conditions for the identification of the parameters $\delta$ and $f_{1}$ which exploit that $H$ and the disturbance vector $(I+\rho M)(u_{1}-f_{1}u_{2})$ of model (12) are orthogonal.
-The idea is to concentrate out $\delta$ and $f_{1}$ in the first stage using the linear moment function.
-Note that by Lemma 2.1 the linear moment condition $E\left[H^{\prime}\left(y_{1}-y_{2}f_{1}-\underline{X}\delta\right)\right]=0$ holds at the true parameter values for $\delta$ and $f_{1}.$ For simplicity consider the case where $z$ is of dimension $n\times 1$.
+Note that by Theorem 2.1 the linear moment condition $E\left[H^{\prime}\left(y_{1}-y_{2}f_{1}-\underline{X}\delta\right)\right]=0$ holds at the true parameter values for $\delta$ and $f_{1}.$ For simplicity consider the case where $z$ is of dimension $n\times 1$.
 Then this equation corresponds to a regular just identified linear instrumental variables problem with included exogenous covariates $\underline{X}$ and where $z$ is instrumenting for the endogenous variable $y_{2}$.
-Letting $W=\left(y_{2},\text{$\underline{X}$}\right)$ and solving the sample analog of the moment condition yields
-$\left(\begin{array}[]{c}\tilde{f}_{1}\\ \tilde{\delta}\end{array}\right)=\left(H^{\prime}W\right)^{-1}H^{\prime}y_{1}.$ (19)
+Letting $W=\left(y_{2},\text{$\underline{X}$}\right)$ and solving the sample analog of the moment condition yields $\left(\tilde{f}_{1},\tilde{\delta}^{\prime}\right)^{\prime}=\left(H^{\prime}W\right)^{-1}H^{\prime}y_{1}$.
 The estimator for the coefficient $f_{1}$ has the familiar form of just identified two stage least squares with included covariates
@@
 $\tilde{\delta}=(\underline{X}^{\prime}\underline{X})^{-1}\underline{X}^{\prime}(y_{1}-\tilde{f}_{1}y_{2}).$ (15)
-In the second step we estimate $\rho$ using a form of the quadratic moment function, where we plug into the residual vector $\tilde{\delta}$ and $\tilde{f}_{1}$ for $\delta$ and $f_{1}$. Let
-$\epsilon^{+}\left(\theta\right)=\epsilon^{+}\left(\rho,f_{1},\delta\right)=\left(I+\rho M\right)^{-1}\left(y_{1}-f_{1}y_{2}-\underline{X}\delta\right)$
-and consider the quadratic moment vector
-$m_{\epsilon}^{q}\left(\rho\right)=\frac{1}{n}\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)^{\prime}A\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)$
-with $A$ defined as before. The first stage estimator for $\rho$ then is obtained as
+In the second step we estimate $\rho$ using a form of the quadratic moment function with, for example, $A=M$.
+Quadratic moments are motivated by Theorem 2.1 which indicates that the variance-covariance matrix of residual vector $\epsilon^{+}=u_{1}-f_{1}u_{2}$ is diagonal.
+Hence $E(\epsilon^{+\prime}A\epsilon^{+})=0$ if $A$ has all diagonal elements 0.
+This suggests plugging $\tilde{\delta}$ and $\tilde{f}_{1}$ into $\epsilon^{+}\left(\rho,f_{1},\delta\right)=\left(I+\rho M\right)^{-1}\left(y_{1}-f_{1}y_{2}-\underline{X}\delta\right)$ and forming the quadratic moment vector $m_{\epsilon}^{q}\left(\rho\right)=\frac{1}{n}\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)^{\prime}A\epsilon^{+}\left(\rho,\tilde{f}_{1},\tilde{\delta}\right)$.
+The first stage estimator for $\rho$ then is obtained as
 $\tilde{\rho}=\operatorname{argmin}_{\rho}\left(m_{\epsilon}^{q}\left(\rho\right)\right)^{2}.$ (16)
-Consistency and the asymptotic distribution of $\tilde{\theta}$ is discussed in Section 4 together with a discussion of the efficient estimator $\hat{\theta}$.
-3 Empirical Results
-In this section, we apply our identification strategy to the Tennessee’s Project STAR (student-teacher achievement ratio) data, exploring peer effects among students in Kindergarten through Grade 3.999We downloaded the data from Harvard Dataverse (https://dataverse.harvard.edu/dataset.xhtml?persistentId=hdl:1902.1/10766), which differs slightly from the version used by Graham(2008) and Rose(2017).
+We prove identification in Section 5.
+Consistency and the asymptotic distribution of $\tilde{\theta}=\left(\tilde{\rho},\tilde{f}_{1},\tilde{\delta}^{\prime}\right)^{\prime}$are discussed in Sections C and D in the online appendix together with a discussion of the efficient estimator $\hat{\theta}$ defined below.
+The estimator $\tilde{\theta}$ is inefficient as it is constructed from the residuals $u_{1}-f_{1}u_{2}$ in (12), which are not homoscedastic.
+To develop efficient estimators, denote the variance-covariance matrix of $u_{1}-f_{1}u_{2}$ as $\Omega(\gamma)$, where $\gamma$ are parameters to be specified later.
+Theorem 2.1 indicates that $\Omega(\gamma)$ is a diagonal matrix.
+The form of $\Omega(\gamma)$ and $\gamma$ are discussed under Assumption 5 as well as Assumption 9 in the online appendix.
+Premultiplying (12) by $V\left(\rho,\gamma\right)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$ then yields the following equation
+$y^{+}(\theta,\gamma)=X^{+}(\theta,\gamma)\delta+u^{+}$ (17)
+where $y^{+}(\theta,\gamma)=V\left(\rho,\gamma\right)(y_{1}-f_{1}y_{2}),$ $X^{+}(\theta,\gamma)=V\left(\rho,\gamma\right)\underline{X},$ and $u^{+}=\Omega(\gamma)^{-1/2}(u_{1}-f_{1}u_{2})$ and $\operatorname{Var}\left(u^{+}|\mathcal{Z}_{n}^{*}\right)=I_{n}$.
+In light of (17) we define
+[…]
+and observe that $u^{+}(\theta,\gamma)=u^{+}$ at the true parameters.
+The transformation leading to $u^{+}$ consists of three components.
+Quasi-differencing eliminates classroom and individual effects.
+The operator $(I+\rho M)^{-1}$ removes cross-sectional correlation due to peer effects and can be understood as a form of spatial Cochrane-Orcutt transformation.
+Finally, the operator $\Omega(\gamma)^{-1/2}$ scales the spatially uncorrelated residuals to unit variance in the case of heteroskedasticity.
+We next provide an outline of our GMM estimation methodology.
+For clarity we explicitly denote in the following the true parameter vectors as $\theta_{0}=(\rho_{0},f_{1,0},\delta_{0}^{\prime})^{\prime}$ and $\gamma_{0}$.
+Our GMM estimator [^5] Rigorous consistency and asymptotic normality results as well as additional assumptions needed to establish these properties are given in Sections C and D in the online appendix.
+is based on both linear and quadratic moment conditions.
+The moment function consists of a set of linear and quadratic moments for the transformed residuals [^6] The specification can be readily extended to include several quadratic moment conditions along the lines of, e.g., Kuersteiner and Prucha (2020).
+$m_{n}(\theta,\gamma)=\left(\begin{array}[]{c}m_{n}^{l}(\theta,\gamma)\\ m_{n}^{q}(\theta,\gamma)\end{array}\right)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+\prime}(\theta,\gamma)Au^{+}(\theta,\gamma)\end{array}\right),$ (19)
+where the functions $m_{n}^{l}(\theta,\gamma)$ and $m_{n}^{q}(\theta,\gamma)$ denote the linear and quadratic moment conditions respectively.
+Under the maintained assumptions of this paper the results in Kuersteiner and Prucha (2020), Theorem 1, imply that the linear moment function is uncorrelated with the quadratic moments.
+We have
+$E\left[m_{n}(\theta_{0},\gamma_{0})m_{n}(\theta_{0},\gamma_{0})^{\prime}\right]$ $=\frac{1}{n}\left(\begin{array}[]{cc}E\left[H^{\prime}H\right]&0\\ 0&2E\left[\operatorname{tr}\left(A^{2}\right)\right]\end{array}\right)\equiv\Xi_{n}.$ (22)
+[…]
+$Q_{n}\left(\theta,\gamma\right)$ $=\frac{1}{n}u^{+}(\theta,\gamma)^{\prime}H\left(H^{\prime}H\right)^{-1}H^{\prime}u^{+}(\theta,\gamma)+\frac{1}{n}\frac{\left(u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\right)^{2}}{2\operatorname{tr}\left(A^{2}\right)}$ (23)
+such that the GMM estimator for $\theta_{0}$ is defined as $\hat{\theta}\left(\gamma\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\gamma\right)$ for a fixed value of $\gamma.$ In applications $\gamma$ is replaced with a first step consistent estimator.
+4 Peer Effects in Project STAR
+4.1 Data
+In this section, we apply our identification strategy to the Tennessee’s Project STAR (student-teacher achievement ratio) data, exploring peer effects among students in Kindergarten through Grade 3.777We use STAR data obtained from Harvard Dataverse (https://dataverse.harvard.edu/dataset.xhtml?persistentId=hdl:1902.1/10766), which differs slightly from the version used by Graham(2008) and Rose(2017).
 For example, they construct class ID using teacher characteristics while our data has class ID readily available.
 Their kindergarten sample has 6,172 students, and ours has 6,325.
-We believe our data set is more trustworthy as it matches the description in Word et al.
+We prefer our data as it matches the description in Word et al.
 (1990) in terms of sample size and summary statistics.
@@
 The experiments followed the kindergarten cohort of 1985 in participating schools for four years, from Kindergarten to the third grade.
-At the start of the 1985 academic year, participating schools randomly assigned kindergarten students and teachers into small classes (with an intended size of 13-17 students), regular classes (with an intended size of 22-25 students) and regular classes with a full-time teacher’s aide.
+At the start of the 1985 academic year, participating schools randomly assigned kindergarten students and teachers into small classes , regular classes and regular classes with a full-time teacher’s aide.
 In the actual implementation, small classes included 11 to 20 students, while regular classes (with or without aides) included 15 to 30 students.
@@
 Our analysis uses the sample of each of the four grades, as our estimator remains consistent even without random assignment.
-The bottom panel of Table 1 summarizes the number of students and classes in each grade.
+[^8] A Table with summary statistics for each grade can be found in the online appendix.
 While there are 11601 students in the experiment in total, the sample size in each grade is between 6325 and 6840 due to students migrating in or out of the participating schools.
@@
 There are between 325 and 340 classes in each grade, with 124 to 140 of them being small classes.
-We collect raw Stanford Achievement Tests (SAT) scores in mathematics (math), reading (read), listening (list) and word study skills (word) as test $t=1$ and test $t=2$ in our model.
+Project STAR data contains tests scores for mathematics, reading, listening and word study skills.
+We use raw Stanford Achievement Tests (SAT) scores in word study skills (word) and reading (read) as test $t=1$ and test $t=2$ respectively in our model, as they measure closely related skills and hence are more likely to satisfy our assumptions.
+This choice is supported by diagnostic checks we report in Tables 1 2 and Table 6 in the appendix.
+Reversing the order of $y_{1}$ and $y_{2}$ does not change the estimate of $\rho$.
+Results using scores in mathematics (math) and listening (list) are discussed in the online appendix as a comparison.
 SAT is a nationally standardized test with scores comparable across grades.
@@
 The narrow time frame for the tests lends credence to the assumption that student’s ability and preparation remain stable over the testing window.
-SAT scores for these four subjects are available from Grade K to Grade 3.
-The mean and standard deviations of scores for the four SAT tests are in the third panel of Table 1.
-The four tests generally share comparable means, except that the mean listening score is much higher than others in Grade K.
-Notably, reading and word study skills scores have means that are close in all grades.
-With four test scores, there are 12 pairs of $y_{1}$ and $y_{2}$.
-We report results for only six unique combinations, as reversing the order of $y_{1}$ and $y_{2}$ does not change the estimate of $\rho$, transforms $f_{1}$ to its reciprocal, and will scale the coefficients of control variables by $-f_{1}$.
-Our main model includes a set of student and teacher characteristics, the mean and standard errors of which are summarized in the top panel of Table 1 by grade.
+The reading and word study skills scores have means that are close in all grades.
+Our preferred model includes a set of student and teacher characteristics.
 Between $56.0\%$ and $59.2\%$ of the students receive a free lunch during the experimental period, an indicator of low family income.
@@
 We start from the specification with only school fixed effects and subsequently introduce other types of control variables incrementally.
+We find that our results are not very sensitive to the inclusion of additional controls.
+This is reassuring in that our differencing strategy of closely related scores removes most of the systematic variation in scores that is related to observed and unobserved individual and classroom level factors affecting outcomes.
 We allow for heteroscedasticity of $u$ between small classes and regular classes (with or without teacher’s aide).
 Standard errors are clustered at the class level.
-The estimators for $f_{1}$ and $\rho$ are summarized in Tables 6 and 7 respectively.
+4.2 Plausibility of Assumptions
 Before proceeding to discussing the main estimates of peer effects, we present several summary statistics in support of our model specifications and assumptions.
-Our estimator relies on the premise that tests $t=1$ and $t=2$ measure similar skills within a brief time frame.
-The combination of reading and word study skills aligns most closely with this narrative.
-To document the close connection between test scores, we use four measures to evaluate the correlation between various SAT scores from Grade K to Grade 3 in Table 2.
+Our estimator relies on the premise that word score $y_{1}$ and reading score $y_{2}$ measure similar skills within a brief time frame.
+To document the close connection between test scores, we use four measures to evaluate the correlation between $y_{1}$ and $y_{2}$ from Grade K to Grade 3 in Table 1.
 The first two measures on the left panel are Spearman correlations between $y_{1}$ and $y_{2}$, and between $Q_{X}y_{1}$ and $Q_{X}y_{2}$.
-Here $Q_{X}=I-\underline{X}(\underline{X}^{\prime}\underline{X})^{-1}\text{$\underline{X}$}^{\prime}$ denotes the residual projection matrix of $\underline{X}$, where $\underline{X}$ encompasses the complete set of control variables: school fixed effects, class type fixed effect, student characteristics, teacher characteristics, and peer characteristics.
+Here $Q_{X}=I-\underline{X}(\underline{X}^{\prime}\underline{X})^{-1}\text{$\underline{X}$}^{\prime}$ denotes the residual projection matrix of X, where X encompasses the complete set of control variables: school fixed effects, class type fixed effect, student characteristics, teacher characteristics, and peer characteristics.
 The right panel of Table 1 reports the pseudo $R^{2}$ from 2SLS, i.e., the Spearman correlation between $y_{1}$ and $\hat{y}_{1}$, or between $y_{2}$ and $\hat{y}_{2}$.
-Here $\hat{y}_{1}=\tilde{f}_{1}y_{2}+\underline{X}\tilde{\delta}$, with $\tilde{f_{1}}$ and $\tilde{\delta}$ defined in (20) and (21).
+Here $\hat{y}_{1}=\tilde{f}_{1}y_{2}+\underline{X}\tilde{\delta}$, with $\tilde{f_{1}}$ and $\tilde{\delta}$ defined in (14) and (15) being the 2SLS estimator when $y_{2}$ is instrumented by 1.
 Similarly, $\hat{y}_{2}$ corresponds to the counterpart obtained by reversing the order of $y_{1}$ and $y_{2}$.
@@
 We drop individuals with missing outcome variables from the sample.
-We discuss the procedure in more detail in Appendix E.
-The different correlation measures paint a consistent picture of the correlation between the six reported pairs as well as across grades.
-The listening score is generally a bit less correlated with other measures.
-The math scores are somewhat more correlated with reading and word skills than with listening.
-We find the strongest connection between word study skills and reading scores which exhibit substantial correlations of approximately 0.9 across all four measures and in all grades.
+We discuss the procedure in more detail in Section G of the online appendix.
+We find that the word study skills and reading scores exhibit substantial correlations of approximately 0.9 across all four measures and in all grades.
 This finding is consistent with these scores measuring skills that are both related to language comprehension.
-Correlations between other pairs range between 0.5 and 0.8.
-While a key identifying assumption, Assumption 2 is not directly testable because it relates to unobserved ability, finding strong correlation between test scores is consistent with a setting where test scores have common conditional means.
-If this interpretation of the results in Table 2 is correct then we would expect the word-read pair of scores to provide the best control for unobserved individual and class effects that may be confounding measures of peer effects at the classroom level.
-Our model implies that conditional on covariates correlation between tests only is due to unobserved interactive effects $\mu_{c}^{\ast}f_{t}.$ While not a test for correct specification, our results are consistent with $\mu_{c}^{\ast}f_{t}$ accounting for a large fraction of the variation in test scores.
-Further evidence of the ability of test score differences to control for unobserved classroom and individual effects is the size of the estimated coefficient $f_{1}$.
-If scores measure closely related skills we expect $f_{1}$ to be close to one.
-In that scenario Assumption 2 is interpreted to imply that unobserved individual specific and classroom level effects captured by $\mu_{c}^{\ast}$ affect both scores in the same way.
-We report estimated values of $f_{1}$ for the set of test score pairs in Table 6.
-Point estimates for $f_{1}$ are obtained by our efficient GMM estimator for all grades and for specifications with just a school effect in Column (1), school and class type effects in Column (2), with a full set of controls excluding contextual peer effects in Column (3) and the full set of controls in Column (4).
-The coefficient $f_{1}$ is estimated with high precision in Columns (1) through (3) and somewhat less precisely in Column (4).The point estimates are generally close to one with a few exceptions, such as for the listening scores in grade K and specifications (1)-(3) as well as the listening scores in grades two and three for specification (4).
-The parameter estimates for the read-word score combination are remarkably stable across all specifications and across all grades, are close to one and estimated precisely.
-This is further evidence that the read-word score pair is best suited to control for unobserved class level effects.
-We explore one additional approach to evaluate the plausibility of Assumption 2.
+On the other hand, correlations between other pairs are smaller and range between 0.5 and 0.8, as shown in Table 6 in the Appendix.
+While a key identifying assumption, Assumption 2 is not directly testable because it relates to unobserved ability, finding strong correlation between test scores is consistent with a setting where test scores have common unobserved individual and class effects that may be confounding measures of peer effects at the classroom level.
+Our model implies that conditional on covariates correlation between tests only is due to unobserved fixed effects $\mu_{c}^{\ast}$ and peer effects working through idiosyncratic errors $u_{it}.$ While not a test for correct specification, our results are consistent with $\mu_{c}^{\ast}$ accounting for a large fraction of the variation in test scores.
+We next explore an additional approach [^10] Two more specification checks are reported in the online appendix.
+to evaluate the plausibility of Assumption 2.
 If unobserved individual and class effects $\mu_{c}^{\ast}$ are stable over time then scores from previous years should be good predictors for current year scores.
@@
 Consequently, when accounting for school fixed effects, we anticipate a consistent level of peer quality over time.
-The relatively stable environment that these students are in suggests that $\mu_{c}^{\ast}$ and $f_{t}$ should contain significant time invariant components that in turn imply high correlation between observed scores and lagged scores.
+The relatively stable environment that these students are in suggests that $\mu_{c}^{\ast}$ should contain significant time invariant components that in turn imply high correlation between observed scores and lagged scores.
 As a result we expect to see relatively high $R^{2}$ in predictive regressions with lagged scores as controls in addition to the observed characteristics.
@@
 The predictive power of lagged scores is documented by comparing the $R^{2}$ with and without adding lags.
-For example, for the word score, adding its own lag improves the $R^{2}$ from .244 to .435 in grade one, from .25 to .539 in grade two and from .229 to .488 in grade three.
-Using the lagged reading score, which is the score of the test most closely related to the word score, leads to increases in the $R^{2}$ that are even slightly higher.
+For example, for the word score, adding its own lag improves the $R^{2}$ from 0.244 to 0.435 in grade one, from 0.25 to 0.539 in grade two and from 0.229 to 0.488 in grade three.
+Using the lagged reading score leads to increases in the $R^{2}$ that are even slightly higher.
 Overall, the coefficients on lagged scores when using the own lag of word as well as the lagged reading score are large in magnitude and highly statistically significant.
-These findings hold up for the other test score pairs with similar magnitudes and statistical significance.
-As may be expected, the predictive effects are slightly weaker for listening scores than for math, read and word.
 The results are further evidence that the data distributions are consistent with key features of our model.
-Our analysis thus far has documented that pairs of tests scores are highly correlated both contemporaneously as well as over the course of one school year.
-This evidence is consistent with restrictions imposed by Assumption 2 and the hypothesis that closely related scores depend on the same unobserved components.
-A caveat is that closely related scores may remove too much common variation when quasi-differences of these scores are used and thus render statistical evidence about peer effects imprecise.
-Table 4 reports pseudo $R^{2}$ measures of model estimates for the six pairs of scores and model specifications (1) through (4) with column (1) only controlling for school fixed effects, and column (4) being the full specification with school and class type effects, student and teacher characteristics as well as peer characteristics.
-The observed pseudo $R^{2}$ measures are generally below 0.2 with a few exceptions for Grade $K$.
-We also see that the $R^{2}$ is a bit higher for models with listening scores.
-Moreover and as expected the pseudo $R^{2}$ generally increases with larger models, or in other words by moving from Column (1) to (4).
-The read-word score pair often has the lowest $R^{2}$ which is an indication that the differential score in these specifications already controls for most of the variation explained by the covariates.
-We take this as further evidence that read-word best fits the model assumptions we impose.
-Table 5 presents standard deviations of $Q_{x}y$ for four test scores and $Q_{x}(y_{1}-f_{1}y_{2})$ for six pairs of test scores for Grade K to 3.
-The left and right panels are for small and regular classes respectively.
-Control variables for specifications (1)-(4) are the same as those in Table 4, as discussed above.
-Estimates of $f_{1}$ are obtained from our efficient GMM estimator and are presented in Table 6.
-We see that in Grade K and Grade 1, the ratio of standard deviations of $Q_{x}(y_{1}-f_{1}y_{2})$ to that of $Q_{x}y$ is around 1/2 for most pairs of tests and 1/5 to 1/4 for the read-word pair.
-In Grade 2 and 3, this ratio is around 1/3 for most pairs and about 1/5 for the read-word pair.
-This shows that quasi-differencing removes a significant part of the variation in the test scores, especially for the read-word pair, consistent with the term $\mu^{*}f_{1}$ playing a major role in determining test performance.
-At the same time, the results in Table 4 indicate that the inclusion of controls does not eliminate much of the remaining variation not captured by fixed effects.
-The conclusion is that there is between 20 to 30 percent of test score variation remaining for the identification of $\rho$.
-Table 7 presents estimates and standard errors (in parentheses) for the peer effects parameter $\rho$.
+Table 3 presents estimates and standard errors (in parentheses) for $f_{1}$ (left panel) and the peer effects parameter $\rho$ (right panel).
+[^12] Results for contextual peer effects are reported in Table 10 in the online appendix.
+Estimates for other pairs of scores are in Table 7 in the online appendix.
+These estimates are obtained by our efficient GMM estimator for all grades and for specifications with just a school effect in Column (1), school and class type effects in Column (2), with a full set of controls excluding contextual peer effects in Column (3) and the full set of controls in Column (4).
+The point estimates of $f_{1}$ are remarkably stable across all specifications and across all grades.
+The size of the estimated coefficient $f_{1}$ is close to one and estimated precisely.
+This is further evidence that word and reading scores measure closely related skills and that the test score differences control unobserved classroom and individual effects.
+Assumption 2 postulates that unobserved individual specific and classroom level effects captured by $\mu_{c}^{\ast}$ affect both scores in the same way.
+This is credible when both scores measure essentially the same skills.
+Our empirical finding of $f_{1}$ close to one for the word-read pair is consistent with this interpretation of the data.
+Other score pairings reported in the appendix have estimated $f_{1}$ parameters that are less reliably close to one, lending further support for our choice of focusing on word-read as our preferred specification.
+4.3 Peer Effects
 To provide some context and interpretation of $\rho$ consider the education production function in Equation (7).
@@
 Similarly, the effect on test scores as a result of changes in peer characteristics are measured as $\rho\beta^{p}$ times the changes in relevant peer characteristics.
-Standard errors are clustered at the class level.
-Test types of $y_{1}$ and $y_{2}$ are indicated in Columns 1 and 2.
-Control variables for the four specifications are outlined at the bottom of the table.
-Specification (1) includes only school fixed effects.
-Specification (2) further adds class type fixed effects.
-Specification (3) introduces student characteristics ($free\,lunch$, $black$, $girl$, $age$) and teacher characteristics ($black$, $master$, $experience$).
-Specification (4) further controls for peer characteristics, defined as the leave-out-mean of the four student characteristics variables.
 The comparison of the four specifications demonstrates the robustness of our estimators to additional controls.
 We now focus on Specification (4), which includes the full set of control variables.
-Across all grade levels and all test pairings, our results consistently reveal positive and statistically significant peer effects, ranging from 0.2 to 0.7.
-The results we find are relatively stable across the four specifications in Columns (1)-(4) with the magnitude of the coefficient $\rho$ generally decreasing somewhat for the larger models.
-The results for read-word are more robust in this sense especially in Grades $K$ and one.
-Estimated peer effects tend to be largest for specifications involving the listening score as well as the math-word score.
-The results for Grades $K$, two and three are quite similar, while results for Grade one are generally somewhat smaller in magnitude but remain statistically very significant.
-Importantly, when examining the pairing of reading and word study skills—a context wherein the tests evaluate closely related abilities as documented by our prior analysis—the effects are notably smaller than for the remaining pairings, ranging from 0.2 to 0.4 across different grade levels.
+Across all grade levels, our results consistently reveal positive and statistically significant peer effects, ranging from 0.2 to 0.4.
+The results we find are relatively stable across the four specifications in Columns (1)-(4).
+The results for Grades $K$, two and three are quite similar, while results for Grade one are smaller in magnitude but remain statistically very significant.
 The implications of these results are that improvements in peer quality measured in terms of potential SAT scores $y_{t}^{*}$ translate at the rate of 20 to 40 percent to improvements of individual scores or in other words a 10 point increase in the peer potential SAT score results in a 2 to 4 point increase in individual SAT scores.
@@
 To focus ideas, consider the case without covariates and only one outcome measure $y_{1}$.
-Then, an endogenous peer effect specification is $y_{1}=\lambda My_{1}+u.$ If peer effects are measured with full rather than leave-one-out means this formulation is identical to our model which is $y_{1}=\left(I+\rho M\right)u$ in this simplified stylized setting.
-The parameters $\rho$ and $\lambda$ satisfy the one-to-one mapping $\lambda=\rho/(1+\rho)$ which is a consequence of the fact that $\left(I+\rho M\right)=\left(I-\lambda M\right)^{-1}$.
+Then, an endogenous peer effect specification for class room $c$ is $y_{ct}=\lambda M_{c}y_{ct}+u_{ct}.$ If peer effects are measured with full rather than leave-one-out means, i.e., if $M_{c}=n^{-1}\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}$, this formulation is identical to our model, which is $y_{ct}=\left(I+\rho M_{c}\right)u_{ct}$ in this simplified stylized setting.
+The parameters $\rho$ and $\lambda$ satisfy the one-to-one mapping $\lambda=\rho/(1+\rho)$ which is a consequence of the fact that $\left(I+\rho M_{c}\right)=\left(I-\lambda M_{c}\right)^{-1}$.
 The inverse captures multiplier effects of peer performance inherent in the endogenous peer effect formulation, while $\rho$ is a summary measure of all these multiplier effects.
 When peer effects are measured by leave-one-out rather than full means the relationship between the two formulations is no longer exact.
-However, it can be shown that $\left(I+\rho M_{c}\right)^{-1}=\left(I-\lambda M_{c}\right)$ continues to hold approximately with an error that is $O\left(n_{c}^{-2}\right)$ where $n_{c}$ is the size of class $c$.
+However, it can be shown that $\left(I+\rho M_{c}\right)=\left(I-\lambda M_{c}\right)^{-1}$ continues to hold approximately with an error that is $O\left(n_{c}^{-2}\right)$ where $n_{c}$ is the size of class $c$.
 Based on this approximation, we convert our measure of $\rho$ to an endogenous peer effect $\lambda=\rho/(1+\rho),$ or equivalently, $\rho=\lambda/\left(1-\lambda\right)$ for the purpose of the discussion that follows.
+Our estimates for $\rho$ in the range of $0.2$ to $0.4$ then translate to values for $\lambda$ that range from $0.16$ to $0.28$.
 In comparison with existing results for peer effects employing STAR data, our estimates are relatively small in magnitude but estimated with high precision and statistical significance.
@@
 Using normalized Kindergarten SAT math score as outcome, Rose (2017) finds an endogenous peer effect of 0.65 or a correlated effect (measured by coefficients of the peer average errors) of 0.71.
-When both are considered simultaneously, endogenous and correlated effects are 0.90 and -0.03, albeit with standard deviations over 1.7, indicating weak identification.
+When both are considered simultaneously, endogenous and correlated effects are 0.90 and -0.03, albeit with standard errors over 1.7, indicating weak identification.
 Boozer and Cacciola (2001), using average percentiles of math and reading scores as outcomes and an instrumental variable methods for estimation, identify endogenous peer effects of 0.30, 0.86 and 0.92 for Grades 1 to 3 respectively, with standard deviations of 1.0, 0.12 and 0.04.
@@
 He finds an estimate of about 0.35 with a standard deviation around 0.14.
-At the upper end of the spectrum are estimates by Lewbel, Qu, and Tang (2023) who report endogenous peer effects of 0.85 and 0.92 for small and regular classes in third-grade SAT math scores, each with standard deviations around 0.02.
-Translated to our parametrization, these estimates correspond to $\rho=5.6$ and $\rho=11.5$ respectively, using the approximation $\rho=\lambda/(1-\lambda)$.
-The magnitude of their findings is remarkably high and suggests that a mere 10-point rise in ability equivalent peer SAT scores leads to an increase of individual scores of up to over 100 points.
 The observed peer effects in our study exhibit smaller magnitudes compared to most of the results discussed above.
@@
 Overall, we find a noticeable difference in peer effects’ magnitude between our study and those of Rose (2017) and Graham (2008).
-Our proposed estimator leads to highly precise estimates compared to all the aforementioned estimates, except that of Lewbel, Qu, and Tang (2023).
+Our proposed estimator leads to highly precise estimates compared to all the aforementioned estimates, except that of Lewbel, Qu, and Tang (2023) who also report precise estimates.
 The enhancement in precision can be attributed in part to the elimination of individual fixed effects through differencing.
 Certain individual characteristics that have been recognized for their significant influence on scores are not observed in the data for Project STAR.
-These include parents’ education, family income and most importantly unobserved IQ.
+These include parent’s education, family income and most importantly unobserved IQ.
 Our estimator, by removing unobserved class effects and individual fixed effects, effectively eliminates residual variance associated with these characteristics which may be one explanation for the precise estimates we obtain.
 In addition, our GMM estimators use additional quadratic moment restrictions efficiently.
-The aforementioned studies have each focused on one or two specific tests and are often limited to a particular grade.
+The aforementioned studies are often limited to a particular grade.
 Methods that rely on random peer assignment are typically constrained in where they can be applied.
 In project STAR the random assignment assumption is more defensible in earlier grades (Boozer and Cacciola, 2001; Graham, 2008; Sojourner, 2013).
-In contrast, our analysis does not require random assignment and facilitates the estimation of peer effects across diverse test types and grade levels.
+In contrast, our analysis does not require random assignment and facilitates the estimation of peer effects across grade levels.
 Encouragingly, our findings consistently fall within a reasonably narrow range, affirming the robustness and reliability of our method.
-We also provide the coefficients of peer averages of individual characteristics (age, black, free lunch, and girl) in Table 8.
-These four variables all pertain to test-invariant individual characteristics.
-The estimated coefficients can be expressed as $\rho(\beta_{v1}^{p}-f_{1}\beta_{v2}^{p})$, where $\rho\beta_{v1}^{p}$ and $\rho\beta_{v2}^{p}$ represent the exogenous peer effect of the variable on tests $t=1$ and $t=2$ respectively.
-Thus, these coefficients capture the differential exogenous peer effects across the two tests.
-We also scale the coefficients for dummies $girl$, $black$ and $free\,lunch$ by 1/100 so that the coefficients are the impact on the outcome of 1 percentage point changes in share of peers who are girl/black/free lunch receivers.
-The first observation is that, except for kindergarten where the presence of girls among peers positively impacts reading more than word study, none of the four variables show significant differential exogenous peer effects in the closely related word-read pair.
-Secondly, the share of black peers and the average peer age generally display insignificant differential exogenous peer effects.
-An exception is in Kindergarten, where peer age more favorably affects listening scores than reading scores.
-Given our general expectation of a negative impact from peers receiving free lunch and a positive impact from a higher share of girls, the results suggest that these two exogenous peer effects are stronger for reading and word study skills compared to listening and mathematical abilities in Grade 1.
-Similar patterns emerge in other grade levels, albeit with lower significance and consistency.
-Overall, the results indicate small and mostly insignificant peer effects on differential scores.
-This is consistent with a scenario where $y_{2}$ is essentially sufficient to control for both observed and unobserved test invariant determinants of test performance.
-In fact, when $f_{1}=1$ and $\beta_{v1}^{p}=\beta_{v2}^{p}$ these effects would be zero in our model and the overall evidence points to the conclusion that this is approximately the case in the sample as well.
-4 Identification and Inference
-4.1 Identification Conditions
+5 Identification
 We next discuss identification of our model.
@@
 We call the model identified at the parameter value $\theta_{0}$ if the population moment conditions underlying the formulation of our estimators are only satisfied at $\theta=\theta_{0}$ for all $\theta\in\Theta$.
-This definition can be seen as an extension of classical definitions of identification such as Rothenberg (1971) for likelihood based approaches.
-In general, identification is neither necessary nor sufficient for consistency which is established separately, see Newey and McFadden (1994) for a discussion, and White (1980) and Pötscher and Prucha (1997) for a discussion of the related concept of identifiable uniqueness.
+Throughout this section we treat $f_{1}$ as an unknown parameter to be estimated.
+To accommodate the empirical specification in Section 3.2 and to justify the specification checks reported in Section 4.2 we establish that the parameter $f_{1}$ in (9) can be consistently estimated irrespective of whether $f_{1}=1$ holds or not.
+For this purpose we modify Assumption 2 to enlarge the class of models we consider to cases where $f_{1}\neq 1.$
+Assumption 4.
+Assume that Assumption 2 holds except that $E\left[y_{it}^{*}|F_{n,i,t}\right]=\kappa_{i}f_{t}$ for $t=1,2$ where $\kappa_{i}$ is measurable with respect to $\mathcal{Z}_{n}^{*}$ and $f_{t}$ is a constant.
+Assumption 4 implies that the correctly specified model is (9) where we impose the normalization $f_{2}=1$ without loss of generality.
+When in addition $f_{1}=1$ then the correctly specified model reduces to (8).
+It is easy to see that the conclusions of Theorem 2.1 continue to hold under Assumption 4 when $u_{it}=y_{it}^{*}-\kappa_{i}f_{t}.$ The next assumption defines the parameter space.
 Assumption 5.
@@
 Then assume $\theta_{0}\in\Theta$, where $\Theta$ is a compact subset of $\Theta_{0}=\left(-1,1\right)\times(-K_{f},K_{f})\times(-K_{X},K_{X})^{p_{x}}$, and where $K_{f}$ and $K_{X}$ are finite positive constants and $p_{x}$ is the column dimension of $\underline{X}$.
-In addition, let $\gamma_{0}=\left(\gamma_{1,0},...,\gamma_{J,0}\right)^{\prime}$ with $\gamma_{j,0}^{2}=(\sigma_{1,0}^{2}+\sigma_{2,0}^{2}f_{1,0}^{2})\rho_{j,0}^{2}$.
-Then $\gamma_{0}\in\Gamma$ with $\Gamma=[K_{\gamma}^{-1},K_{\gamma}]^{J}$ where $K_{\gamma}>1$ is a finite constant.
-Remark 4.1.
-The assumption implies that the parameter space for $\rho$ is a subset of $[-K_{\rho},K_{\rho}]$ for some positive $K_{\rho}<1$.
+In addition, suppose there are $J$ class types denoted by $\tau_{c}\in\{1,\ldots,J\},$ and $\tau_{c}$ is assumed to be measurable w.r.t.
+$\mathcal{Z}_{n}$.
+Furthermore, let $\Omega(\gamma)=\operatorname{diag}_{c}^{C}\left\{\gamma_{\tau_{c}}^{2}I_{c}\right\}$ with $\gamma=(\gamma_{1},\ldots,\gamma_{J})^{\prime},$ then we assume that $\operatorname{Var}(u_{1}-f_{1,0}u_{2}|\mathcal{Z}_{n}^{*})=\Omega(\gamma_{0})$ for some $\gamma_{0}=(\gamma_{1,0},\ldots,\gamma_{J,0})^{\prime}\in\Gamma$ with $\Gamma=[K_{\gamma}^{-1},K_{\gamma}]^{J}$ where $K_{\gamma}>1$ is a finite constant.
+The assumption implies that the parameter space for $\rho$ is a subset of $[-K_{\rho},K_{\rho}]$ for some positive constant $K_{\rho}<1$.
+The conditions postulated for $\Omega(\gamma)$ are satisfied under Assumption 9 in Appendix B regarding $u_{t}$.
 In addition, we impose the the following assumption on $A$.
@@
 Note that in the case of no covariates, the coefficients $\delta$ of the covariates are omitted, $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ is reduced to $(\rho,f_{1})^{\prime}$ and (18) becomes $u^{+}(\theta,\gamma)=y^{+}(\theta,\gamma)$.
-Observe that $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}(\gamma_{\tau_{c}}^{2}I_{c})$ is diagonal for all admissible $\gamma$ and that $\Omega_{0}=\Omega(\gamma_{0})=\operatorname{Var}(u_{1}-f_{1,0}u_{2}|\mathcal{Z}_{n})$ in light of Assumptions 2 and 3 and Theorem 2.1.
-Also, $u^{+}(\theta_{0},\gamma)=\Omega(\gamma)^{-1/2}\Omega_{0}^{1/2}u^{+}.$ Thus it is readily seen that for all $A$ satisfying Assumption 5(i) we have
-$E\left[u^{+}(\theta_{0},\gamma)|\mathcal{Z}_{n}\right]=0,$ (23)
-$E\left[u^{+}(\theta_{0},\gamma)^{\prime}Au^{+}(\theta_{0},\gamma)|\mathcal{Z}_{n}\right]$ $=\operatorname{tr}\left(\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\Omega_{0}\right)$ (24)
-$=\operatorname{tr}\left(A\Omega(\gamma)^{-1}\Omega_{0}\right)$ (25)
-$=0$ (26)
-observing that the diagonal elements of $A\Omega(\gamma)^{-1}\Omega_{0}$ are zero.
 We next discuss identification for a given matrix $H=z\equiv\left(z_{1}^{\prime},...,z_{c}^{\prime},...,z_{C}^{\prime}\right)^{\prime}$ of $\mathcal{Z}_{n}$ measurable instruments, where the $z_{c}$ are $n_{c}\times q$ matrices of random variables that are invariant at the classroom level, i.e., $z_{c}=\mathbf{1}_{c}\grave{z}_{c}$, where $\mathbf{1}_{c}$ is an $n_{c}\times 1$ vector with all elements equal to one and $\grave{z}_{c}$ is a $1\times q$ vector of class level characteristics.
@@
 Additional overidentifying restrictions can be added but do not affect arguments related to identification.
-Consider the moment vector
-$m_{n}\left(\theta,\gamma\right)=n^{-1/2}\left[\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\end{array}\right]$ (27)
-where $A$ is a matrix that satisfies the properties listed in Assumption 5.
+Consider the moment vector $m_{n}\left(\theta,\gamma\right)$ defined in (19) where $A$ is a matrix that satisfies the properties listed in Assumption 6.
+Observe that $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}(\gamma_{\tau_{c}}^{2}I_{c})$ is diagonal for all admissible $\gamma$ and that $\Omega_{0}=\Omega(\gamma_{0})=\operatorname{Var}(u_{1}-f_{1,0}u_{2}|\mathcal{Z}_{n})$ in light of Assumptions 4 and 5 and Theorem 2.1.
+Also, $u^{+}(\theta_{0},\gamma)=\Omega(\gamma)^{-1/2}\Omega_{0}^{1/2}u^{+}.$ Thus it is readily seen that for all $A$ satisfying Assumption 6(i) we have $E\left[m_{n}\left(\theta_{0},\gamma\right)\right]=0$, where the expectation of the quadratic moments is $\operatorname{tr}\left(A\Omega(\gamma)^{-1}\Omega_{0}\right)$ which is zero because diagonal elements of $A\Omega(\gamma)^{-1}\Omega_{0}$ are zero.
 We next show that $E\left[m_{n}\left(\rho,f_{1},\gamma\right)\right]=0$ if and only if $(\rho,f_{1})=(\rho_{0},f_{1,0})$.
@@
 As a special case of (13) without covariates consider the data generating process $y_{t}=\mu^{\ast}f_{t,0}+\left(I_{n}+\rho_{0}M\right)u_{t}$ where $\mu^{*}$ is defined in (4).
-Suppose Assumptions 1, 2, 3, 4, 5, and 6 hold, let $m_{n}\left(\theta,\gamma\right)$ be defined in (27) and let $H=z$.
+Suppose Assumptions 3, 4, 5, 6, and 7 hold, let $m_{n}\left(\theta,\gamma\right)$ be defined in (19) and let $H=z$ where $z$ is invariant at the classroom level.
 In addition, assume that $q=1$ and
@@
 for all $\left(\rho,f_{1}\right)\neq\left(\rho_{0},f_{1,0}\right)$ and $\gamma\in\Gamma$.
-A proof of Lemma 4.1 establishing the identifiability of $\left(\rho_{0},f_{1,0}\right)$ is given in the appendix.
+A proof of Lemma 5.1 establishing the identifiability of $\left(\rho_{0},f_{1,0}\right)$ is given in the online appendix.
 The subsequent corollary considers a special case of the lemma that is of interest in the context of test scores, and arises when $H=\mathbf{1}_{n}$ and when test scores are non-negative.
@@
 Recall that with covariates, $u^{+}(\theta,\gamma)$ is defined in (18).
-We consider the following $(q+1)\times 1$ vector of linear and quadratic moment functions:
-$m_{n}(\theta,\gamma)=\left(\begin{array}[]{c}m_{n}^{(l)}(\theta,\gamma)\\ m_{n}^{(q)}(\theta,\gamma)\end{array}\right)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\end{array}\right)$ (29)
-where the functions $m_{n}^{(l)}(\theta,\gamma)$ and $m_{n}^{(q)}(\theta,\gamma)$ denote the linear and quadratic moment conditions respectively.
+We consider the $(q+1)\times 1$ vector of linear and quadratic moment functions $m_{n}(\theta,\gamma)$ defined in (19).
 Note that our procedure is not designed to identify $\gamma$, nor is that required for the identification of $\theta$.
-Lemma 4.2 below shows that for all permissible $\gamma$ the moment condition satisfies $E\left[m_{n}(\theta_{0},\gamma)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{n}(\theta,\gamma)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ for $\theta\neq\theta_{0}$.
+Theorem 5.1 below shows that for all permissible $\gamma$ the moment condition satisfies $E\left[m_{n}(\theta_{0},\gamma)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{n}(\theta,\gamma)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ for $\theta\neq\theta_{0}$.
 The parameters $\gamma$ can be pinned down by estimating the variance of $u_{1}-f_{1,0}u_{2}$ for different types of groups.
@@
 […]
-Remark.
-Note that the elements of the matrix $H$ are by construction measurable w.r.t. $\mathcal{Z}_{n}=\sigma(X,z)$.
+Note that the elements of the matrix $H$ are by construction measurable w.r.t.
+$\mathcal{Z}_{n}=\sigma(X,z)$.
 Since $\underline{X}$ contains $X$, part (i) of the assumption implies that also the elements of $X$ are bounded in absolute value by $K_{H}$.
 Observing that $(\underline{X},z)$ has full column rank by Assumption 8(iii), as setting $\rho=0$ and $\gamma=(1,...,1)^{\prime}$ in $V(\rho,\gamma)$, (25) becomes $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$, where $Q_{X}=I-\underline{X}\left(\underline{X}^{\prime}\underline{X}\right){}^{-1}\underline{X}^{\prime}$.
-It is readily seen from an inspection of the proof of Lemma 4.2 that a necessary condition for Assumption 7(iii) is that $\left(\underline{X},E\left[\mu^{*}|\mathcal{Z}_{n}\right]\right)$ has full column rank.
+It is readily seen from an inspection of the proof of Theorem 5.1 below that a necessary condition for Assumption 8(iii) is that $\left(\underline{X},E\left[\mu^{*}|\mathcal{Z}_{n}\right]\right)$ has full column rank.
 The latter condition is equivalent to the condition that $\left(\underline{X},E\left[y_{2}|\mathcal{Z}_{n}\right]\right)$ has full column rank, observing that in light of (13) we have $E\left[y_{2}|\mathcal{Z}_{n}\right]=E\left[\mu^{*}+\underline{X}\beta_{2}|\mathcal{Z}_{n}\right]$ for some vector of parameters $\beta_{2}.$
@@
 For example, when $\gamma=\bar{\gamma}$ then (24) is replaced by $n^{-1}\left|E\left[y_{2}^{\prime}\Omega(\bar{\gamma})^{-1/2}z\left|\mathcal{Z}_{n}\right.\right]\right|>K_{y}>0$ and a similar simplification obtained for (25).
-The following lemma is a generalization of Lemma 4.1.
-Lemma 4.2.
-Let Assumptions 1, 2, 3, 4, 5, 6 and 7 be satisfied.
+The following theorem is a generalization of Lemma 5.1.
+Theorem 5.1.
+Let Assumptions 3, 4, 5, 6, 7 and 8 be satisfied.
 Let $m_{n}\left(\theta,\gamma\right)$ be defined as in (19).
 Then for all admissible $\gamma\in\Gamma,$ $E\left[m_{n}\left(\theta_{0},\gamma\right)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{n}\left(\theta,\gamma\right)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ for $\theta\neq\theta_{0}$ a.s.
-4.2 Consistency of Parameter Estimators
-Formulating the moment conditions using transformed residuals as in (29) requires consistent estimates of the variance parameters $\gamma$.
-These can be obtained by first obtaining consistent but inefficient parameter estimates.
-To do so, we set all elements of $\gamma=(\gamma_{1},...,\gamma_{J})^{\prime}$ to one.
-Lemma 4.2 implies that the parameters $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ are still identified by the moment conditions in (29).
-First step estimators $(\tilde{\rho},\tilde{f}_{1},\tilde{\delta}^{\prime})^{\prime}$ are defined in (19) and (22).
-[^13] Lee (2007a) considers first step and efficient GMM estimators using linear and quadratic moments.
-To formalize the identification result for these non-efficient first step estimators define the moment vector
-$m_{\epsilon}\left(\theta\right)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}(I+\rho M)\epsilon^{+}\left(\theta\right)\\ \epsilon^{+}\left(\theta\right)^{\prime}A\epsilon^{+}\left(\theta\right)\end{array}\right).$ (31)
-The next lemma shows that the moment conditions in (31) are sufficient to identify the parameters $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$.
-Lemma 4.3.
-Let Assumptions 1-6, and Assumption 7(i),(ii) be satisfied.
-Let $m_{\epsilon}(\theta)$ be defined as in (31).
-Assume further that $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$ a.s .
-Then $E\left[m_{\epsilon}\left(\theta_{0}\right)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{\epsilon}\left(\theta\right)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ a.s.
-if $\theta\neq\theta_{0}$ for all $\theta\in\Theta$
-We next show that the initial estimators and the efficient estimators are consistent.
-To establish consistency we impose assumptions related to the convergence of sample averages and moments of sample averages to well defined limits.
-These assumptions are adaptations of Assumptions 2, 3 and 5 in Kuersteiner and Prucha (2020) to the present setting which differs somewhat from the framework of sequential exogeneity in that paper.
-Let $i=n_{1}+\ldots n_{c-1}+r$ with $1\leq r\leq n_{c}$.
-Then, with some abuse of notation, denote the $i$-th diagonal element of $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)$ as $\gamma_{(i)}^{2}=\gamma_{\tau_{c}}^{2}$, noting that $\tau_{c}$ varies with $i$.
-Let the linear and quadratic moment vectors $m_{n}^{(l)}(\theta,\gamma)$ and $m_{n}^{(q)}(\theta,\gamma)$ be as defined in (29).
-Let $h_{i}=\left[h_{i1},...,h_{iq}\right]$ denote the $i$-th row vector of $H$, and let $a_{ij}$ denote the the $(i,j)$-th element of matrix $A$.
-Then note that with $u^{+}(\theta_{0},\gamma)=\Omega(\gamma)^{-1}(u_{1}-f_{1,0}u_{2})$, the variance of $m_{n}^{(l)}(\theta_{0},\gamma)$ is
-$\frac{1}{n}E\left[H^{\prime}\Omega(\gamma)^{-1}\Omega(\gamma_{0})H\right]=\frac{1}{n}\sum_{i=1}^{n}E\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}h_{i}^{\prime}h_{i}\right]$ (32)
-and the variance of $m_{n}^{(q)}(\theta_{0},\gamma)$ is
-$\frac{1}{n}E\left[\operatorname{tr}\left(\Omega_{0}\Omega^{-1}A\Omega_{0}\Omega^{-1}A\right)\right]=\frac{1}{n}E\sum_{i=1}^{n}\sum_{j=1}^{n}\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}\frac{\gamma_{(j),0}^{2}}{\gamma_{(j)}^{2}}a_{ij}^{2}\right].$ (33)
-The assumption below ensures that for any admissible $\gamma$, $\operatorname{Var}\left(m_{n}^{(l)}(\theta_{0},\gamma)\right)$ and $\operatorname{Var}\left(m_{n}^{(q)}(\theta_{0},\gamma)\right)$ and their corresponding sample analogues converge, respectively, to finite positive definite matrices.
-Assumption 8.
-The following holds :
-$n^{-1}\sum_{i=1}^{n}E\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}h_{i}^{\prime}h_{i}\right]\rightarrow V_{\gamma}^{h},\quad n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}E\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}\frac{\gamma_{(j),0}^{2}}{\gamma_{(j)}^{2}}a_{ij}^{2}\right]\rightarrow V_{\gamma}^{a},$
-where the elements of $V_{\gamma}^{h}$ and $V_{\gamma}^{a}$ are finite , and
-[…]
-The matrix $V_{\gamma}=\operatorname{diag}\left(V_{\gamma}^{h},2V_{\gamma}^{a}\right)$ is positive definite.
-Assumption 9.
-(i) Let $C\left(\theta,\gamma\right)$ be an $n\times n$ matrix of the form $D$, $DP$, $DAD$, $DADP$, or $PDADP$, where $D$ is an $n\times n$ positive diagonal matrix with elements which are uniformly bounded and measurable w.r.t.
-$\mathcal{Z}_{n}$ and where $P=\operatorname{diag}_{c}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right)$ and where $p_{c}$ and $q_{c}$ are continuously differentiable functions of the parameters $\theta$ and $\gamma$ as well as of variables generating $\mathcal{Z}_{n}$, and $I_{c}^{\ast}$ and $J_{c}^{\ast}$ are defined in Appendix B.
-Then
-[…]
-where $\Upsilon_{a},\Upsilon_{b}$ are selected from the set $\left\{H,MH,\mu^{*},u_{1},u_{2}\right\}$.
-(ii) $\lim_{n\rightarrow\infty}\sup_{\theta,\gamma}\left\|E\left[n^{-1}\Upsilon_{a}^{\prime}C\left(\theta,\gamma\right)\Upsilon_{b}\right]-\mathscr{U}_{a,b}\left(\theta,\gamma\right)\right\|=0$ where
-$\mathscr{U}_{a,b}\left(\theta,\gamma\right)=\lim_{n\rightarrow\infty}E\left[n^{-1}\Upsilon_{a}^{\prime}C\left(\theta,\gamma\right)\Upsilon_{b}\right]$
-is bounded and continuously differentiable in $\theta$ and $\gamma$.
-In Kuersteiner, Prucha, and Zeng (2023) we show uniform convergence under more primitive conditions in a closely related situation while directly assuming it here in Assumption 9 to save space.
-Without further assumptions on cross-sectional dependence, high level assumptions about the convergence of sample averages such as $n^{-1}\underline{X}^{\prime}\underline{X}$ are required irrespective.
-The next theorem establishes consistency of the GMM estimator.
-Theorem 4.1.
-Let Assumptions 1- 9 be satisfied. Assume that $\bar{\gamma}_{n}\stackrel{{\scriptstyle p}}{{\rightarrow}}\gamma$ for some sequence $\bar{\gamma}_{n}$ and some $\gamma\in\Gamma.$ Then,
-(i) for $\hat{\theta}\left(\gamma\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\gamma\right)$ with $Q_{n}\left(\theta,\gamma\right)$ defined in (18) it follows that $\hat{\theta}\left(\bar{\gamma}_{n}\right)\stackrel{{\scriptstyle p}}{{\rightarrow}}\theta_{0}$.
-(ii) for $\tilde{f}_{1}$ and $\tilde{\delta}$ defined in (19) and $\tilde{\rho}$ defined in (22) it follows that $\tilde{f}_{1}\stackrel{{\scriptstyle p}}{{\rightarrow}}f_{1,0}$, $\tilde{\delta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\delta_{0}$ and $\tilde{\rho}\stackrel{{\scriptstyle p}}{{\rightarrow}}\rho_{0}.$
-4.3 Asymptotic Normality of Parameter Estimators
-In the following we develop an asymptotically justified inference theory for our proposed GMM estimator.
-For that purpose we first discuss consistent estimators for the variance parameters $\gamma=(\gamma_{1},...,\gamma_{J})^{\prime}$, which are treated as nuisance parameters in the objective function of our GMM estimator.
-Assumption 10.
-Let $\omega_{j}\equiv\lim{}_{n\rightarrow\infty}n^{-1}\sum_{c=1}^{C}n_{c}1(\tau_{c}=j)$ and assume that $\omega_{j}$ exists and $\omega_{j}>0$ for all $j=1,...,J$.
-Assumption 10 guarantees that there are enough data to estimate $\gamma_{j}^{2}$ for each $j$ by requiring that the fraction $\omega_{j}$ of all students in classrooms of type $j$ is asymptotically non-negligible for all types of classrooms.
-The first step inefficient but consistent estimators $\tilde{f}_{1}$, $\tilde{\delta}$ and $\tilde{\rho}$ defined in (19) and (22) can be used to obtain consistent estimates of the variance parameters $\gamma.$ Form the residuals
-$\tilde{\epsilon}=\epsilon^{+}(\tilde{\theta})=\left(I_{n}+\tilde{\rho}M\right)^{-1}\left(y_{1}-y_{2}\tilde{f}_{1}-\underline{X}\tilde{\delta}\right)$ (34)
-and organize the residuals by classrooms $c$ as $\tilde{\epsilon}=\left(\tilde{\epsilon}_{1}^{\prime},....,\tilde{\epsilon}_{C}^{\prime}\right)^{\prime}$. Then construct the variance estimators
-$\hat{\gamma}_{j}^{2}=\left(N_{j}-p_{x}-1\right)^{-1}\sum_{c=1}^{C}\tilde{\epsilon}_{c}^{\prime}\tilde{\epsilon}_{c}1\left(\tau_{c}=j\right),$ (35)
-where $N_{j}=\sum_{c=1}^{C}n_{c}1\left(\tau_{c}=j\right)$ is the number of students in classrooms of type $j$, and where $p_{x}$ denotes the number of columns in $\underline{X}$. Now set
-$\hat{\gamma}=\left(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{J}\right)^{\prime}.$ (36)
-Efficient GMM estimators can now be formed by plugging $\hat{\gamma}$ into $u^{+}(\theta,\gamma)$ defined in (14).
-Recall that $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}(\gamma_{\tau_{c}}^{2}I_{c}).$ In the following we denote with $K_{u}$ a generic finite constant (which is taken, w.o.l.o.g., to be greater than one) and which is invariant over $t=1,2$, $i=1,\ldots,n$, $n\geq 1$.
-The central limit theorem in Proposition 3 of Kuersteiner and Prucha (2020) can be applied to obtain distributional approximations for estimators based on $u^{+}$.
-The central limit theorem of Kuersteiner and Prucha (2020) requires that $E\left[\left|u_{it}u_{is}\right|^{1+\eta_{u}}|\mathcal{Z}_{n}\right]\leqslant K_{u}$ with $\eta_{u}>0$ for $t,s=1,2\text{ }$ and all $i$.
-This condition is implied by Assumption 2, as Lemma B.3 in the appendix demonstrates.
-When setting
-$\tilde{\gamma}=\left(1,1,...,1\right)^{\prime}$ (37)
-it follows that $\Omega(\tilde{\gamma})=I_{n}$.
-The feasible efficient GMM estimator for $\theta_{0}=\left(\rho_{0},f_{1,0},\delta_{0}^{\prime}\right)^{\prime}$ is defined as $\hat{\theta}\left(\hat{\gamma}\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\hat{\gamma}\right)$ where $Q_{n}\left(\theta,\gamma\right)$ is given in (18) and the inefficient counterpart $\hat{\theta}\left(\tilde{\gamma}\right)$ is defined in an analogous way.
-The next theorem establishes the limiting distribution of $\hat{\theta}\left(\tilde{\gamma}\right)$ and $\hat{\theta}\left(\hat{\gamma}\right).$ The limiting distribution of $\hat{\theta}\left(\tilde{\gamma}\right)$ for $\tilde{\gamma}$ defined in (37) provides a distributional result for an inefficient GMM estimator.
-The limiting distribution of the first step estimators defined in (19) and (22) can be deduced from the theorem by working out the limiting distribution of $m_{\epsilon}\left(\theta\right).$ We omit the details.
-Theorem 4.2.
-Let Assumptions 1-10 hold.
-Let $\hat{\gamma}$ be as defined in (36) and $\tilde{\gamma}$ as defined in (37).
-For $\gamma_{n}\rightarrow_{p}\gamma$ let $\operatorname{plim}_{n\rightarrow\infty}n^{-1/2}\partial m_{n}\left(\hat{\theta}\left(\gamma_{n}\right),\gamma_{n}\right)/\partial\theta=G(\gamma)$ and $\Xi=\operatorname{plim}_{n\rightarrow\infty}\Xi_{n}$, where $\Xi_{n}$ is defined in (17).
-Then $\hat{\gamma}$$\rightarrow_{p}\gamma_{0}$, and for $\gamma_{n}=\hat{\gamma}$ or $\gamma_{n}=\tilde{\gamma}$,
-$\sqrt{n}\left(\hat{\theta}\left(\gamma_{n}\right)-\theta_{0}\right)\rightarrow_{d}N(0,\Psi_{\gamma})$
-where the limiting variance covariance matrix $\Psi_{\gamma}$ has the form
-$\Psi_{\gamma}=\left(G^{\prime}\Xi^{-1}G\right)^{-1}G^{\prime}\Xi^{-1}V_{\gamma}\Xi^{-1}G\left(G^{\prime}\Xi^{-1}G\right)^{-1}.$
-with $V_{\gamma}$ defined in Assumption 8 and $G$ is the shorthand notation for $G(\gamma)$.
-When $\gamma_{n}=\hat{\gamma}$ then $V_{\gamma}=\Xi$ and $\Psi_{\gamma}=\left(G(\gamma_{0})^{\prime}\Xi^{-1}G(\gamma_{0})\right)^{-1}$.
-In Appendix D we provide explicit formulas for the derivatives $G$.
-Together with the expressions for $\Xi_{n}$ it is then easy to construct sample based estimators $\hat{G}$ and $\hat{\Xi}$.
-Letting $\hat{\Psi}_{\gamma}=\left(\hat{G}^{\prime}\hat{\Xi}^{-1}\hat{G}\right)^{-1}$ it follows from a standard Slutsky argument that $\hat{\Psi}_{\gamma}^{-1/2}\sqrt{n}\left(\hat{\theta}\left(\hat{\gamma}\right)-\theta_{0}\right)\rightarrow_{d}N\left(0,I\right).$
+We report results that establish the consistency and asymptotic normality of our estimators in an online appendix where we also discuss consistent estimation of $\gamma$ and standard errors.
 6 Conclusions
@@
 Our empirical results for Project STAR data and classrooms in kindergarten to the third grade show highly significant peer effects, but smaller in magnitude than those found in a number of studies that have looked at the same data.
-Appendix A Tables
-Table 1: Summary Statistics
-Grade K Grade 1 Grade 2 Grade 3
-mean sd mean sd mean sd mean sd
-indiviudal characteristics
-free lunch 0.560 0.496 0.588 0.492 0.592 0.491 0.578 0.494
-black 0.326 0.469 0.327 0.469 0.347 0.476 0.333 0.471
-girl 0.486 0.500 0.480 0.500 0.484 0.500 0.480 0.500
-age 6.011 0.349 7.158 0.492 8.242 0.557 9.285 0.586
-teacher characteristics
-black teacher 0.165 0.371 0.174 0.379 0.204 0.403 0.209 0.406
-master degree 0.347 0.476 0.346 0.476 0.373 0.484 0.442 0.497
-years of experience 9.258 5.809 11.633 8.937 13.145 8.655 13.933 8.615
-SAT test scores
-math 485.377 47.698 530.528 43.109 580.613 44.574 617.970 39.841
-read 436.725 31.706 520.787 55.192 583.935 46.043 615.422 38.563
-list 537.475 33.140 567.487 33.674 595.476 34.908 624.119 32.265
-word 434.179 36.759 513.436 53.316 582.986 50.658 610.136 45.041
-Sample Size
-number of students 6325 6829 6840 6801
-number of classes 325 339 340 336
-# small 127 124 133 140
-# regular 99 115 100 89
-# regular/aide 99 100 107 107
-number of schools 79 76 75 75
-  • 1. Mean and standard deviations of student characteristics, teacher charcteristics, and SAT scores, as well as number of students and classes in each grade.
-Table 2: Rank Correlation between Scores
-Rank Correlation between Tests Pseudo R squared of 2SLS
-$y_{1}$ $y_{2}$ $y_{1},y_{2}$ $Q_{X}y_{1},Q_{X}y_{2}$ $y_{1},\hat{y}_{1}$ $y_{2},\hat{y}_{2}$
-Grade K
-list math 0.657 0.584 0.728 0.638
-list read 0.621 0.554 0.686 0.691
-list word 0.558 0.477 0.637 0.610
-math read 0.741 0.672 0.697 0.763
-math word 0.656 0.579 0.643 0.727
-read word 0.908 0.881 0.915 0.909
-Grade 1
-list math 0.721 0.654 0.766 0.716
-list read 0.624 0.534 0.704 0.601
-list word 0.570 0.483 0.645 0.562
-math read 0.737 0.669 0.782 0.753
-math word 0.686 0.619 0.726 0.715
-read word 0.932 0.905 0.936 0.940
-Grade 2
-list math 0.678 0.611 0.713 0.705
-list read 0.667 0.601 0.709 0.699
-list word 0.597 0.533 0.639 0.639
-math read 0.724 0.673 0.756 0.758
-math word 0.625 0.564 0.659 0.670
-read word 0.884 0.848 0.892 0.894
-Grade 3
-list math 0.649 0.610 0.689 0.673
-list read 0.652 0.623 0.690 0.683
-list word 0.524 0.478 0.581 0.552
-math read 0.741 0.706 0.759 0.769
-math word 0.683 0.643 0.712 0.708
-read word 0.890 0.869 0.901 0.895
-  • 1.
-The left panel is Spearman’s rank correlation between $y_{1}$, $y_{2}$ and between $Q_{X}y_{1}$, $Q_{X}y_{2}$, where $Q_{X}=I-\underline{X}(\underline{X}^{\prime}\underline{X})^{-1}\underline{X}^{\prime}$, $y_{1}$ and $y_{2}$ are raw SAT scores specified in Columns 1 and 2, $\underline{X}$ are the complete set of control variables, including school fixed effects, class type fixed effects, student characteristics, teacher characteristics and peer characteristics.
-  • 2.
-The right panel is pseudo $R^{2}$, i.e., Spearman’s rank correlation between $y_{1}$ and $\hat{y}_{1}$ in Column 3, and between $y_{2}$ and $\hat{y}_{2}$ in Column 4.
-Here $\hat{y}_{1}=\tilde{f}_{1}y_{2}+\text{$\underline{X}$}\tilde{\delta}$, with $\tilde{f_{1}}$ and $\tilde{\delta}$ defined in (20) and (21), $\underline{X}$ is the complete set of controls as described in note 1.
-That is, $\hat{y}_{1}$ is the predicted value of $y_{1}$ from 2SLS for $y_{1}=f_{1}y_{2}+\underline{X}\delta+\epsilon$ with $y_{2}$ instrumented by the constant term, $\hat{y}_{2}$ is defined in a similar manner by reversing the order of $y_{1},y_{2}$.
-Table 3: Explanatory Power of Lagged Scores
-Estiamtes for lagged scores $R^{2}$ w/wo lagged scores
-lagged score Grade 1 Grade 2 Grade 3 Grade 1 Grade 2 Grade 3
-Dependent Variable: Listening Score
-N.A. 0.256 0.297 0.236
-list 0.620(0.019) 0.707(0.014) 0.687(0.013) 0.515 0.617 0.599
-math 0.362(0.012) 0.465(0.013) 0.394(0.011) 0.435 0.488 0.419
-read 0.476(0.020) 0.303(0.009) 0.386(0.011) 0.410 0.448 0.424
-word 0.355(0.016) 0.272(0.010) 0.296(0.010) 0.381 0.409 0.378
-Dependent Variable: Mathematics Score
-N.A. 0.284 0.272 0.258
-list 0.616(0.021) 0.679(0.019) 0.606(0.017) 0.454 0.464 0.444
-math 0.549(0.015) 0.780(0.016) 0.697(0.013) 0.537 0.614 0.630
-read 0.655(0.026) 0.455(0.011) 0.554(0.013) 0.464 0.496 0.513
-word 0.505(0.019) 0.421(0.014) 0.396(0.012) 0.442 0.455 0.428
-Dependent Variable: Reading Score
-N.A. 0.290 0.294 0.246
-list 0.592(0.026) 0.662(0.020) 0.628(0.016) 0.391 0.451 0.458
-math 0.592(0.019) 0.669(0.018) 0.538(0.013) 0.468 0.521 0.485
-read 1.026(0.032) 0.622(0.011) 0.671(0.010) 0.544 0.648 0.644
-word 0.809(0.024) 0.582(0.013) 0.512(0.010) 0.522 0.588 0.544
-Dependent Variable: Word Study Skills Score
-N.A. 0.244 0.250 0.229
-list 0.509(0.028) 0.645(0.022) 0.579(0.020) 0.333 0.377 0.363
-math 0.511(0.020) 0.645(0.020) 0.579(0.016) 0.399 0.426 0.441
-read 0.860(0.033) 0.668(0.012) 0.737(0.013) 0.453 0.588 0.594
-word 0.686(0.025) 0.631(0.014) 0.547(0.012) 0.435 0.539 0.488
-  • 1.
-Each panel represents regressions of a specific score type against a comprehensive set of control variables (including school fixed effects, class type fixed effects, student, teacher, and peer characteristics in the main specification and a constant term), along with possible own and peer average lagged scores.
-  • 2. The lagged score type is indicated in column 1 and highlighted in bold if it matches the dependent variable type. N.A. means no lagged scores are controlled for.
-  • 3.
-The left panel displays coefficients and standard deviations (in parentheses) of lagged scores for grade 1 to 3.
-The right panel reports the $R^{2}$ of the regression, where the first row in each panel representing the $R^{2}$ when no lagged scores are included.
-  • 4. Coefficients of peer average lagged scores, typically small and statistically insignificant, are omitted for conciseness.
-Table 4: Pseudo $R^{2}$: $1-var(u_{1}-f_{1}u_{2})/var(y_{1}-f_{1}y_{2})$
-Small Classes Regualr Classes(w/wo Aide)
-$y_{1}$ $y_{2}$ (1) (2) (3) (4) (1) (2) (3) (4)
-Grade K
-list math 0.179 0.180 0.206 0.216 0.177 0.179 0.201 0.208
-list read 0.143 0.144 0.153 0.150 0.145 0.146 0.150 0.147
-list word 0.124 0.125 0.132 0.129 0.132 0.133 0.136 0.134
-math read 0.148 0.148 0.177 0.219 0.126 0.127 0.155 0.192
-math word 0.138 0.139 0.164 0.200 0.128 0.129 0.154 0.184
-read word 0.104 0.104 0.107 0.113 0.140 0.140 0.142 0.146
-Grade 1
-list math 0.138 0.140 0.143 0.144 0.113 0.115 0.123 0.124
-list read 0.131 0.133 0.154 0.163 0.113 0.116 0.148 0.161
-list word 0.105 0.106 0.122 0.132 0.068 0.071 0.094 0.108
-math read 0.105 0.105 0.130 0.137 0.070 0.071 0.119 0.132
-math word 0.092 0.092 0.114 0.118 0.058 0.059 0.101 0.109
-read word 0.088 0.088 0.092 0.090 0.093 0.093 0.098 0.093
-Grade 2
-list math 0.116 0.117 0.123 0.119 0.104 0.106 0.118 0.109
-list read 0.087 0.087 0.144 0.141 0.083 0.084 0.125 0.125
-list word 0.084 0.084 0.131 0.132 0.070 0.072 0.107 0.109
-math read 0.092 0.092 0.134 0.137 0.086 0.087 0.110 0.109
-math word 0.089 0.089 0.124 0.126 0.081 0.081 0.106 0.107
-read word 0.064 0.064 0.077 0.082 0.065 0.065 0.074 0.078
-Grade 3
-list math 0.135 0.136 0.147 0.144 0.085 0.087 0.092 0.089
-list read 0.073 0.074 0.133 0.133 0.024 0.027 0.077 0.078
-list word 0.090 0.090 0.132 0.131 0.030 0.033 0.065 0.066
-math read 0.083 0.083 0.107 0.106 0.050 0.050 0.085 0.086
-math word 0.085 0.085 0.107 0.106 0.039 0.040 0.071 0.070
-read word 0.093 0.093 0.092 0.096 0.033 0.034 0.035 0.038
-school FE Y Y Y Y Y Y Y Y
-class type FE Y Y Y Y Y Y
-stu&tch char Y Y Y Y
-peer char Y Y
-  • 1.
-The table reports one minus the ratio of the variance of $u_{1}-f_{1}u_{2}$ to that of $y_{1}-f_{1}y_{2}$ for small classes (in the left panel) and regular (with or without aide) classes (in the right panel).
-All models allow for heteroscedasticity across small and regular classes.
-  • 2. The scores are raw SAT scores in listening (list), mathematics (math), reading (read), word study skills (word).
-  • 3.
-As indicated at the bottom of the table, the models may control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.
-Table 5: Standard Deviations of $Q_{x}y$ and $Q_{x}(y_{1}-f_{1}y_{2})$
-Small Classes Regular Classes
-$y_{1}$ $y_{2}$ (1) (2) (3) (4) (1) (2) (3) (4)
-Grade K
-$Q_{x}y$ list 75.810 75.122 37.769 28.508 78.091 76.238 38.954 29.681
-$Q_{x}y$ math 75.456 74.762 44.444 39.421 80.333 78.669 48.468 43.319
-$Q_{x}y$ read 62.748 62.169 33.699 26.827 66.321 64.827 35.534 29.143
-$Q_{x}y$ word 65.196 64.629 38.017 31.883 68.723 67.294 39.693 33.916
-$Q_{x}(y_{1}-f_{1}y_{2})$ list math 35.462 35.555 38.504 41.091 38.064 38.131 41.515 44.471
-$Q_{x}(y_{1}-f_{1}y_{2})$ list read 30.802 30.801 30.488 27.544 31.701 31.697 31.347 28.060
-$Q_{x}(y_{1}-f_{1}y_{2})$ list word 36.916 36.919 36.311 33.052 37.006 37.015 36.336 32.868
-$Q_{x}(y_{1}-f_{1}y_{2})$ math read 30.874 30.867 30.338 30.583 31.934 31.927 31.556 32.405
-$Q_{x}(y_{1}-f_{1}y_{2})$ math word 35.724 35.696 34.235 32.883 35.731 35.707 34.486 33.819
-$Q_{x}(y_{1}-f_{1}y_{2})$ read word 14.055 14.059 13.883 14.896 14.625 14.632 14.422 15.448
-Grade 1
-$Q_{x}y$ list 78.430 77.682 44.281 29.291 82.153 80.072 43.729 32.263
-$Q_{x}y$ math 77.259 76.454 47.817 36.254 81.842 79.872 48.874 39.564
-$Q_{x}y$ read 83.720 82.984 57.172 46.261 88.372 86.584 58.298 49.448
-$Q_{x}y$ word 81.314 80.624 57.574 46.620 85.880 84.192 57.174 47.750
-$Q_{x}(y_{1}-f_{1}y_{2})$ list math 29.766 29.822 29.377 29.449 30.662 30.730 30.227 30.285
-$Q_{x}(y_{1}-f_{1}y_{2})$ list read 42.790 42.812 38.711 40.458 43.672 43.660 39.169 40.844
-$Q_{x}(y_{1}-f_{1}y_{2})$ list word 44.810 44.856 40.579 42.694 45.258 45.272 40.989 42.897
-$Q_{x}(y_{1}-f_{1}y_{2})$ math read 35.658 35.595 32.807 34.226 37.625 37.547 34.073 35.289
-$Q_{x}(y_{1}-f_{1}y_{2})$ math word 38.230 38.195 35.500 36.668 39.197 39.143 36.076 36.874
-$Q_{x}(y_{1}-f_{1}y_{2})$ read word 21.646 21.649 21.648 21.786 22.530 22.533 22.493 22.616
-Grade 2
-$Q_{x}y$ list 92.075 90.731 48.088 29.859 85.236 82.708 45.789 32.371
-$Q_{x}y$ math 90.878 89.540 53.323 38.275 85.553 83.117 52.234 41.389
-$Q_{x}y$ read 93.586 92.295 57.595 39.409 87.831 85.406 54.519 41.740
-$Q_{x}y$ word 96.864 95.555 62.274 44.613 90.149 87.782 59.070 46.967
-$Q_{x}(y_{1}-f_{1}y_{2})$ list math 32.791 32.843 31.327 28.848 34.666 34.691 32.803 30.147
-$Q_{x}(y_{1}-f_{1}y_{2})$ list read 34.210 34.242 29.824 28.460 34.229 34.216 29.939 28.515
-$Q_{x}(y_{1}-f_{1}y_{2})$ list word 40.464 40.509 34.546 31.816 39.898 39.907 34.124 31.404
-$Q_{x}(y_{1}-f_{1}y_{2})$ math read 31.591 31.571 29.607 30.029 32.541 32.523 31.102 31.550
-$Q_{x}(y_{1}-f_{1}y_{2})$ math word 39.107 39.083 35.878 35.793 39.412 39.388 36.700 36.676
-$Q_{x}(y_{1}-f_{1}y_{2})$ read word 23.891 23.894 23.273 22.593 24.312 24.318 23.726 23.043
-Grade 3
-$Q_{x}y$ list 100.934 98.926 46.812 28.881 81.116 78.549 45.571 31.663
-$Q_{x}y$ math 99.438 97.593 51.961 35.145 82.400 79.955 50.671 37.816
-$Q_{x}y$ read 102.223 100.223 52.813 33.948 85.025 82.325 51.772 38.114
-$Q_{x}y$ word 102.419 100.495 56.860 39.320 86.959 84.387 56.262 43.608
-$Q_{x}(y_{1}-f_{1}y_{2})$ list math 29.971 29.941 28.331 27.089 30.284 30.259 28.579 27.304
-$Q_{x}(y_{1}-f_{1}y_{2})$ list read 29.166 29.145 26.466 25.715 29.828 29.770 26.962 26.078
-$Q_{x}(y_{1}-f_{1}y_{2})$ list word 38.217 38.212 34.066 33.314 39.272 39.206 34.788 33.930
-$Q_{x}(y_{1}-f_{1}y_{2})$ math read 27.010 27.013 26.326 26.477 26.807 26.801 25.922 26.045
-$Q_{x}(y_{1}-f_{1}y_{2})$ math word 32.785 32.796 31.350 31.918 32.844 32.832 31.064 31.683
-$Q_{x}(y_{1}-f_{1}y_{2})$ read word 20.631 20.634 20.090 20.639 21.111 21.106 20.485 21.043
-  • 1.
-Standard deviations of $Q_{x}y$ and $Q_{x}(y_{1}-f_{1}y_{2})$ for small (on the left) and regular classes (on the right).
-For $Q_{x}y$, the test score $y$ is specified in $y_{1}$.
-For $Q_{x}(y_{1}-f_{1}y_{2})$, $y_{1}$ and $y_{2}$ are specified in columns 2 and 3.
-The control variables for models (1)-(4) are the same as those for the main models in Tables 6 and 7.
-The estimates of $f_{1}$ are from our efficient GMM estimations for each specifications and presented in Tables 6.
-Table 6: Estimates of $f_{1}$
-$y_{1}$ $y_{2}$ (1) (2) (3) (4)
-Grade K
-list math 1.093(0.011) 1.096(0.010) 1.201(0.027) 1.283(0.243)
-list read 1.231(0.008) 1.231(0.008) 1.217(0.021) 1.011(0.109)
-list word 1.224(0.007) 1.225(0.006) 1.201(0.023) 1.048(0.128)
-math read 1.126(0.005) 1.124(0.004) 1.014(0.018) 0.771(0.140)
-math word 1.120(0.008) 1.118(0.008) 0.999(0.021) 0.793(0.152)
-read word 0.995(0.004) 0.995(0.005) 0.984(0.009) 1.042(0.055)
-Grade 1
-list math 1.065(0.009) 1.068(0.008) 1.054(0.012) 1.057(0.070)
-list read 1.060(0.008) 1.061(0.008) 0.973(0.014) 1.023(0.077)
-list word 1.071(0.009) 1.073(0.009) 0.977(0.015) 1.038(0.087)
-math read 0.995(0.005) 0.993(0.006) 0.916(0.012) 0.969(0.072)
-math word 1.007(0.004) 1.006(0.005) 0.929(0.012) 0.973(0.084)
-read word 1.008(0.005) 1.008(0.006) 1.014(0.009) 1.026(0.045)
-Grade 2
-list math 1.057(0.005) 1.059(0.005) 1.007(0.012) 0.902(0.055)
-list read 1.030(0.006) 1.031(0.006) 0.903(0.009) 0.842(0.050)
-list word 1.023(0.006) 1.024(0.007) 0.873(0.011) 0.779(0.052)
-math read 0.975(0.006) 0.974(0.006) 0.893(0.010) 0.929(0.048)
-math word 0.967(0.006) 0.967(0.007) 0.863(0.010) 0.862(0.055)
-read word 0.992(0.004) 0.993(0.004) 0.967(0.007) 0.929(0.030)
-Grade 3
-list math 1.028(0.002) 1.027(0.002) 0.961(0.008) 0.898(0.055)
-list read 1.014(0.003) 1.013(0.003) 0.915(0.008) 0.869(0.045)
-list word 1.027(0.004) 1.026(0.004) 0.903(0.010) 0.873(0.053)
-math read 0.986(0.003) 0.986(0.003) 0.951(0.008) 0.965(0.047)
-math word 0.999(0.004) 0.999(0.004) 0.940(0.009) 0.970(0.051)
-read word 1.013(0.002) 1.013(0.002) 0.988(0.005) 1.016(0.032)
-school FE Y Y Y Y
-class type FE Y Y Y
-stu&tch char Y Y
-peer char Y
-  • 1.
-Estimates and standard errors (in the parentheses) for $\rho$.
-Standard errors are clustered at the classroom level.
-All models allow for heteroscedasticity across small and regular classes.
-Estimates are adjusted for missing observations.
-  • 2. The scores are raw SAT scores in listening (list), mathematics (math), reading (read), word study skills(word).
-  • 3.
-As indicated at the bottom of the table, the models may control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.
-Table 7: Estimates of $\rho$
-$y_{1}$ $y_{2}$ (1) (2) (3) (4)
-Grade K
-list math 0.721(0.079) 0.691(0.073) 0.694(0.075) 0.688(0.078)
-list read 0.434(0.079) 0.415(0.076) 0.407(0.078) 0.377(0.083)
-list word 0.398(0.066) 0.378(0.065) 0.373(0.065) 0.359(0.068)
-math read 0.492(0.065) 0.485(0.065) 0.519(0.066) 0.575(0.076)
-math word 0.517(0.067) 0.513(0.067) 0.538(0.070) 0.578(0.081)
-read word 0.452(0.078) 0.449(0.078) 0.443(0.080) 0.424(0.075)
-Grade 1
-list math 0.627(0.069) 0.587(0.073) 0.579(0.073) 0.578(0.074)
-list read 0.322(0.051) 0.277(0.049) 0.256(0.048) 0.218(0.046)
-list word 0.385(0.054) 0.340(0.052) 0.308(0.048) 0.257(0.047)
-math read 0.441(0.063) 0.434(0.064) 0.434(0.061) 0.360(0.061)
-math word 0.487(0.066) 0.480(0.065) 0.469(0.062) 0.388(0.064)
-read word 0.178(0.054) 0.178(0.054) 0.190(0.054) 0.193(0.053)
-Grade 2
-list math 0.637(0.073) 0.621(0.072) 0.554(0.071) 0.525(0.068)
-list read 0.568(0.073) 0.553(0.071) 0.571(0.074) 0.561(0.074)
-list word 0.537(0.070) 0.521(0.068) 0.512(0.070) 0.497(0.069)
-math read 0.481(0.071) 0.479(0.071) 0.475(0.066) 0.444(0.066)
-math word 0.493(0.071) 0.492(0.071) 0.478(0.066) 0.459(0.065)
-read word 0.399(0.065) 0.398(0.064) 0.363(0.063) 0.348(0.064)
-Grade 3
-list math 0.747(0.088) 0.728(0.087) 0.714(0.086) 0.703(0.090)
-list read 0.606(0.091) 0.572(0.088) 0.613(0.096) 0.623(0.101)
-list word 0.512(0.095) 0.482(0.091) 0.518(0.095) 0.522(0.097)
-math read 0.462(0.078) 0.460(0.078) 0.452(0.082) 0.440(0.086)
-math word 0.371(0.074) 0.366(0.072) 0.370(0.074) 0.357(0.078)
-read word 0.391(0.078) 0.387(0.077) 0.377(0.077) 0.379(0.076)
-school FE Y Y Y Y
-class type FE Y Y Y
-stu&tch char Y Y
-peer char Y
-  • 1.
-Estimates and standard errors (in the parentheses) for $\rho$.
-Standard errors are clustered at the classroom level.
-All models allow for heteroscedasticity across small and regular classes.
-Estimates are adjusted for missing observations.
-  • 2. The scores are raw SAT scores in listening (list), mathematics (math), reading (read), word study skills(word).
-  • 3.
-As indicated at the bottom of the table, the models may control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.
-Table 8: Estimates of Exogenous Peer Effects
-$y_{1}$ $y_{2}$ age black(pct.) free lunch(pct.) girl(pct.)
-Grade K
-list math -4.290(15.730) -0.196(0.244) 0.117(0.099) -0.177(0.102)*
-list read 13.784(6.951)** 0.093(0.103) 0.024(0.050) -0.031(0.050)
-list word 10.097(8.239) 0.066(0.134) 0.014(0.058) 0.020(0.058)
-math read 15.222(9.027)* 0.238(0.159) -0.076(0.065) 0.111(0.063)*
-math word 12.705(9.930) 0.200(0.156) -0.082(0.066) 0.159(0.064)**
-read word -4.298(3.626) -0.029(0.062) -0.005(0.024) 0.070(0.027)**
-Grade 1
-list math -0.268(5.134) -0.016(0.095) -0.049(0.068) 0.013(0.079)
-list read -1.922(6.076) -0.125(0.124) 0.129(0.084) -0.354(0.095)***
-list word -2.212(6.848) -0.186(0.147) 0.160(0.095)* -0.454(0.108)***
-math read -2.392(5.540) -0.099(0.103) 0.179(0.076)** -0.330(0.096)***
-math word -1.156(6.399) -0.161(0.118) 0.196(0.086)** -0.416(0.107)***
-read word -0.719(3.423) -0.033(0.065) 0.038(0.043) -0.037(0.050)
-Grade 2
-list math 7.639(4.183)* 0.166(0.123) -0.045(0.071) 0.120(0.089)
-list read 4.909(4.197) 0.109(0.104) 0.127(0.072)* 0.063(0.082)
-list word 7.808(4.551)* -0.001(0.102) 0.150(0.079)* 0.149(0.086)*
-math read -2.902(3.933) -0.062(0.107) 0.187(0.068)*** -0.064(0.084)
-math word -0.032(4.720) -0.183(0.130) 0.213(0.082)*** 0.025(0.096)
-read word 2.977(2.691) -0.134(0.088) 0.025(0.050) 0.105(0.054)*
-Grade 3
-list math 5.041(4.171) 0.187(0.094)** -0.111(0.055)** -0.059(0.066)
-list read 4.026(3.618) 0.030(0.084) -0.079(0.048)* -0.055(0.059)
-list word 2.712(4.286) 0.061(0.097) -0.045(0.058) -0.063(0.073)
-math read -1.133(3.779) -0.162(0.087)* 0.032(0.047) 0.003(0.054)
-math word -2.395(4.101) -0.129(0.104) 0.073(0.057) -0.009(0.058)
-read word -2.190(2.593) 0.030(0.062) 0.037(0.040) -0.025(0.045)
-  • 1.
-Estimates and standard errors (in the parentheses) for exogenous peer effects of peer’s age, race, free lunch status and gender.
-Coefficients for dummies black, free lunch and girl are divided by 100 so they can be interpreted as the impact of one percentage point change in these variables.
-Note that these four variables are all test-invariant personal characteristics.
-The reported coefficients are $\rho(\beta_{v1}^{p}-f_{1}\beta_{v2}^{p})$.
-  • 2. Standard errors are clustered at the classroom level. All models allow for heteroscedasticity across small and regular classes. Estimates are adjusted for missing observations.
-  • 3. The scores are raw SAT scores in listening (list), mathematics (math), reading (read), word study skills(word).
-  • 4.
-All specifications control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.
-Appendix B Auxiliary Lemmas
-Before proving the main results, we establish a number of preliminary lemmas.
-Note that many matrices in this paper, e.g., $M$, $\Omega(\gamma)$, $\Sigma_{t}$ defined in Assumption 3 can be written in the form of $\operatorname{diag}_{c=1}^{C}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right),$where $I_{c}^{\ast}=I_{c}-\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}/n_{c}$, $J_{c}^{\ast}=\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}/n_{c}$ are the residual projection matrix and the projection matrix onto $\mathbf{1}_{c}$ respectively.
-To see this note that
-$M$ $=\operatorname{diag}_{c=1}^{C}\left(-\frac{1}{n_{c}-1}I_{c}^{\ast}+J_{c}^{\ast}\right),$ (38)
-$I+\rho M$ $=\operatorname{diag}_{c=1}^{C}\left(\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\left(1+\rho\right)J_{c}^{\ast}\right),$ (39)
-$\Omega(\gamma)$ $=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}^{\ast}+\gamma_{\tau_{c}}^{2}J_{c}^{\ast}\right),$ (40)
-$\Sigma_{t}$ $=\sigma_{0,t}^{2}\operatorname{diag}_{c=1}^{C}\left(\rho_{0,\tau_{c}}^{2}I_{c}\right)=\operatorname{diag}_{c=1}^{C}\left(\rho_{0,\tau_{c}}^{2}\sigma_{0,t}^{2}I_{c}^{\ast}+\rho_{0,\tau_{c}}^{2}\sigma_{0,t}^{2}J_{c}^{\ast}\right),\,t=1,2.$ (41)
-The following lemma summarizes some well understood key properties of these matrices for the convenience of the reader.
-Proofs are straight forward and omitted, and can e.g., be found in Kuersteiner, Prucha, and Zeng (2023).
-Lemma B.1.
-The matrices $I_{c}^{\ast}$, $J_{c}^{\ast}$ and $pI_{c}^{\ast}+qJ_{c}^{\ast}$ have the following properties:
-(i) The matrices $I_{c}^{\ast}$, $J_{c}^{\ast}$ are symmetric, idempotent, orthogonal, $I_{c}^{*}+J_{c}^{*}=I$ and $I_{c}^{\ast}\mathbf{1}_{c}=0$, $J_{c}^{\ast}\mathbf{1}_{c}=\mathbf{1}_{c}$.
-(ii) $\operatorname{det}\left(pI_{c}^{\ast}+qJ_{c}^{\ast}\right)=p^{(n_{c}-1)}q$.
-(iii) The eigenvalues of $pI_{c}^{\ast}+qJ_{c}^{\ast}$ are $p$ ($n_{c}-1$ times) and $q$ (once).
-(iv) For $p_{1}$, $p_{2}$, $q_{1}$ and $q_{2}$ arbitrary constants, $(p_{1}I_{c}^{\ast}+q_{1}J_{c}^{\ast})(p_{2}I_{c}^{\ast}+q_{2}J_{c}^{\ast})=(p_{1}p_{2}I_{c}^{\ast}+q_{1}q_{2}J_{c}^{\ast})$, and thus multiplication of matrices of the form $pI_{c}^{\ast}+qJ_{c}^{\ast}$ is commutative and associative.
-(v) If $p\neq 0,$ $q\neq 0$, $(pI_{c}^{\ast}+qJ_{c}^{\ast})$ is invertible with $(pI_{c}^{\ast}+qJ_{c}^{\ast})^{-1}=p^{-1}I_{c}^{\ast}+q^{-1}J_{c}^{\ast}$.
-(vi) $\operatorname{tr}(pI_{c}^{\ast}+qJ_{c}^{\ast})=p(n_{c}-1)+q$.
-(vii) Both $I_{c}^{\ast}$ and $J_{c}^{\ast}$ have row and column sum uniformly bounded in absolute value.
-When $p$ and $q$ are bounded in absolute value, the row and column sums of $pI_{c}^{\ast}+qJ_{c}^{\ast}$ are uniformly bounded in absolute value.
-Corollary B.1.
-Suppose Assumptions 3, 4, 5, 6, 7, and 10 hold. Then $I+\rho M$, $\Omega(\gamma)$, and $\Sigma_{t}$ are non-singular. If $S(\theta,\gamma)$ is an element or product matrix of elements from
-$\left\{M,(I+\rho M),(I+\rho M)^{-1},\Omega(\gamma),\Omega(\gamma)^{-1},\Sigma_{t}^{-1},\Sigma_{t}\right\},$ (42)
-then $S(\theta,\gamma)$ can be written in the form of $\operatorname{diag}_{c=1}\{p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\}$with $p_{c}$, $q_{c}$ continuously differentiable in $\theta$ and $\gamma$.
-The matrix $S(\theta,\gamma)A$ is measurable w.r.t $\mathcal{Z}_{n}$ The row and column sums of $S(\theta,\gamma)$ and $S(\theta,\gamma)A$ are uniformly bounded in absolute value.
-.
-Proof.
-By Assumptions 3, 4 and 6, $\left(n_{c}-1-\rho\right)/\left(n_{c}-1\right)$, $\left(1+\rho\right)$, $\gamma_{j}^{2}$, $\rho_{0,\tau_{c}}^{2}$, and $\sigma_{0,t}^{2}$ in (39), (40), (41) are positive and uniformly bounded away from 0.
-Hence by Lemma B.1(v), $I+\rho M$, $\Omega(\gamma)$ and $\Sigma_{t}$ are non-singular.
-An inspection of Equations (38)-(41) shows that all matrices in the set can be written in the form of $\operatorname{diag}_{c=1}^{C}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right),$ with $p_{c}$ and $q_{c}$ continuously differentiable in $\theta$ and $\gamma$ and uniformly bounded.
-Also note that by Assumption 5, the row and column sums of $A$ are uniformly bounded in absolute value.
-Utilizing remark A.1 in Kelejian and Prucha (2004) the corollary thus follows from Lemma B.1.
-∎
-We use these properties frequently in the derivations that follow.
-Lemma B.2.
-Suppose Assumptions 3, 4, 5, and 6 hold.
-Consider the block diagonal matrix $\mathcal{M}\left(\rho\right)=\operatorname{diag}{}_{c=1}^{C}\left(\mathcal{M}_{c}\left(\rho\right)\right)$ with $\mathcal{M}_{c}\left(\rho\right)=\left(I_{c}+\rho M_{c}\right)^{-1}\left(I_{c}+\rho_{0}M_{c}\right)$.
-Then for all $\gamma\in\Gamma$,
-$\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0$
-and
-$\lim_{n\rightarrow\infty}\frac{1}{n}\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0$
-each has only one solution for $\rho\in\left(-1,1\right)$. This solution is given by $\rho=\rho_{0}$.
-Proof.
-Let $D(\rho,\gamma)=\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}$ and observe that $D(\rho,\gamma)=\operatorname{diag}{}_{c=1}^{C}\left(D_{c}(\rho,\gamma)\right)$ with
-$D_{c}(\rho,\gamma)$ $=$ $\mathcal{M}_{c}\left(\rho\right)\left(\gamma_{\tau_{c}}^{2}I_{c}\right)^{-1/2}A_{c}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)^{-1/2}\mathcal{M}_{c}\left(\rho\right)\left(\gamma_{0,\tau_{c}}^{2}I_{c}\right)$
-$=$ $\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right).$
-Noting that $\operatorname{tr}\left(A_{c}\right)=0$ and hence $\operatorname{tr}\left(A_{c}I_{c}^{\ast}\right)=\operatorname{tr}\left(A_{c}(I-J_{c}^{\ast})\right)=-\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$, and in light of Lemma B.1
-$\mathcal{M}_{c}\left(\rho\right)$ $=$ $\left(I_{c}+\rho M_{c}\right)^{-1}\left(I_{c}+\rho_{0}M_{c}\right)=\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}I_{c}^{\ast}+\frac{1+\rho_{0}}{1+\rho}J_{c}^{\ast}.$
-We have
-$\operatorname{tr}\left(\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right)$ $=$ $\operatorname{tr}\left(A_{c}\left[\left(\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}\right)^{2}I_{c}^{\ast}+\left(\frac{1+\rho_{0}}{1+\rho}\right)^{2}J_{c}^{\ast}\right]\right)$
-$=$ $\left[\left(\frac{1+\rho_{0}}{1+\rho}\right)^{2}-\left(\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}\right)^{2}\right]\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$
-$=$ $(\rho_{0}-\rho)C(\rho,n_{c})\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$
-where
-$C(\rho,n_{c})=\left[\left(\frac{1+\rho_{0}}{1+\rho}\right)+\left(\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}\right)\right]\left[\frac{1}{1+\rho}+\frac{1}{n_{c}-1-\rho}\right].$
-Hence
-$\operatorname{tr}\left(D(\rho,\gamma)\right)=\sum_{c=1}^{C}\operatorname{tr}\left(D_{c}(\rho,\gamma)\right)=(\rho_{0}-\rho)\sum_{c=1}^{C}\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}C(\rho,n_{c})\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right).$
-Clearly, at $\rho=\rho_{0}$, for all $\gamma\in\Gamma$, we have $\operatorname{tr}\left(D(\rho_{0},\gamma)\right)=0$ for all $n$.
-Next observe that $\operatorname{tr}\left(A_{c}J_{c}^{*}\right)=\mathbf{1}_{c}^{{}^{\prime}}A_{c}\mathbf{1}_{c}/n_{c}\geqslant\text{$\underline{K}$}_{a}>0$ and that for any $n_{c}\geqslant 2$ and $-1<\rho<1$, both $n_{c}-1-\rho$ and $1+\rho$ are positive, and therefore $C(\rho,n_{c})$ is positive for all $\rho\in\left(-1,1\right)$.
-Therefore, $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]<0$ if $\rho>\rho_{0}$, $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]>0$ if $\rho<\rho_{0}$ and $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]$ has the same sign for all $c$.
-In all, for any value of $\gamma$ the only solution for
-$\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma,\theta)^{-1/2}A\Omega(\gamma,\theta)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0$
-is $\rho=\rho_{0}$.
-Furthermore, by Assumptions 4 and 6 there exists some constants $c_{D}$ and $c_{\gamma}$ such that for any $\rho_{0}$ in the interior of the parameter space, $C(\rho,n_{c})>c_{D}>0$ and $\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}>c_{\gamma}>0$.
-This implies that,
-$\left|\operatorname{tr}(D(\rho))\right|\geqslant|\rho_{0}-\rho|\sum_{c=1}^{C}\text{$\underline{K}$}_{a}c_{D}c_{\gamma},$
-$\left|\frac{1}{n}\operatorname{tr}(D(\rho))\right|\geqslant|\rho_{0}-\rho|\text{$\underline{K}$}_{a}c_{D}c_{\gamma}\frac{C}{n}.$
-By Assumption 6, $C/n\geq 1/\bar{n}_{c}>0$ such that the conclusion follows. ∎
+Appendix
+Appendix A Proofs
+Proof of Theorem 2.1.
+By the assumption on the boundedness of the $\left(4+\eta\right)$-th moments all considered expectations and conditional expectations exist.
+Since $u_{it}=y_{it}^{*}-E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]$ we have $E\left[u_{it}|\mathcal{F}_{n,i,t}\right]=0$ by construction.
+Observing that $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=\kappa_{i}$ and that $\kappa_{i}$ is $\mathcal{Z}_{n}^{*}$-measurable it follows that conditional on $\mathcal{Z}_{n}^{*}$ knowledge of $y_{it}^{*}$ is equivalent to knowledge of $u_{it}$.
+To see this let $u_{-i,t}=\left(u_{1t},..,u_{i-1,t},u_{i+1,t},...,u_{n,t}\right)^{\prime}$ and use the notation $u_{-t}$ for the unobserved components related to the other test.
+Define $U=\left(S,\alpha,\varphi,\zeta^{*},u_{-t},u_{-i,t},\kappa_{i}\right)$ assuming for simplicity of notations that vectors are row vectors, and let $V=\left(S,\alpha,\varphi,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*},\kappa_{i}\right)$.
+Then $V=g(U)$ where $g(.)$ is Borel measurable and thus $\sigma(V)\subseteq\sigma(U)$.
+Analogously $U=h(V)$ where $h(.)$ is Borel measurable and thus $\sigma(U)\subseteq\sigma(V)$.
+Consequently $\sigma(U)=\sigma(V)$.
+Since $\kappa_{i}$ is measurable w.r.t.
+$\sigma(S,\alpha,\varphi,\zeta^{*})$ it follows that
+$\sigma(S,\alpha,\varphi,\zeta^{*},u_{-t},u_{-i,t},\kappa_{i})=\sigma(S,\alpha,\varphi,\zeta^{*},u_{-t},u_{-i,t})$
+and $\sigma(S,\alpha,\varphi,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*},\kappa_{i})=\sigma(S,\alpha,\varphi,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*}),$ which shows that
+$\mathcal{B}_{n,i,t}=\sigma\left(S,\alpha,\varphi,\zeta^{*},u_{-t},u_{-i,t}\right)=\sigma\left(S,\alpha,\varphi,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*}\right)=\mathcal{F}_{n,i,t}.$
+Next observe that the $u_{jt}$ are $\mathcal{F}_{n,i,t}$-measurable for $i\neq j$, and hence by iterated expectations $\operatorname{Cov}\left(u_{it},u_{jt}|\mathcal{F}_{n,i,t}\right)=E\left[u_{jt}E\left(u_{it}|\mathcal{F}_{n,i,t}\right)\right]=0$.
+Observe further that for $s\neq t$ the $u_{js}$ are $\mathcal{F}_{n,i,t}$-measurable for all $i$, and hence by iterated expectations $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{F}_{n,i,t}\right)=E\left[u_{js}E\left[u_{it}|\mathcal{F}_{n,i,t}\right]\right]=0$.
+The corresponding claims regarding the mean and covariances of the $u_{it}$ conditional on $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$ follow immediately by iterated expectations.
+This proves the first part of the theorem.
+To prove the second part, let $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$ be some information set. Then given that Assumption 2 holds, it follows by iterated expectations that
+E$\left[y_{it}^{*}-\kappa_{i}|\mathcal{G}_{n,i,t}\right]$ $=$ $E\left[E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]|\mathcal{G}_{n,i,t}\right]-E\left[\kappa_{i}|\mathcal{G}_{n,i,t}\right]=E\left[\kappa_{i}|\mathcal{G}_{n,i,t}\right]-E\left[\kappa_{i}|\mathcal{G}_{n,i,t}\right]=0.$
+Thus by the almost sure uniqueness of conditional expectations, Assumption 2 cannot hold if $E\left[y_{it}^{*}-\kappa_{i}|\mathcal{G}_{n,i,t}\right]\neq 0$. ∎
+Note that Lemma 5.1 and Corollary 5.1 can be viewed as special cases of Theorem 5.1.
+We therefore only give an explicit proof of Theorem 5.1 here.
+Proofs of the additional results are given in the online appendix.
+[^14] Equations labeled 32 and higher, as well as Appendices B and higher can be found in the online appendix.
+Proof of Theorem 5.1.
+In light of (13) it follows that,
+$y_{1}-f_{1}y_{2}=\mu^{\ast}(f_{1,0}-f_{1})+\underline{X}\delta(f_{1},\rho_{0},\beta_{0})+(I+\rho_{0}M)(u_{1}-f_{1}u_{2}),$ (26)
+where $\delta(f_{1},\rho_{0},\beta_{0})$ is defined in (11) and $\delta(f_{1,0},\rho_{0},\beta_{0})=\delta_{0}$. Consequently,
+$y_{1}-f_{1}y_{2}-\text{\text@underline{X}}\delta=\mu^{\ast}(f_{1,0}-f_{1})+\underline{X}\left(\delta(f_{1},\rho_{0},\beta_{0})-\delta\right)+(I+\rho_{0}M)(u_{1}-f_{1}u_{2}).$ (27)
+Let $\mathcal{M}(\rho)=(I+\rho M)^{-1}(I+\rho_{0}M)$. Substitution of (27) into (18) yields
+$u^{+}(\theta,\gamma)$ $=$ $\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$
+×$\times\left\{\left(f_{1,0}-f_{1}\right)\mu^{\ast}+\underline{X}\left(\delta(f_{1},\rho_{0},\beta_{0})-\delta\right)+(I+\rho_{0}M)(u_{1}-f_{1}u_{2})\right\}$
+$=$ $V(\rho,\gamma)\left(\begin{array}[]{cc}\underline{X},&\mu^{*}\end{array}\right)\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right)+\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1}u_{2}),$ (32)
+recalling that $V(\rho,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$
+We first analyze the linear moment conditions. Note that
+$H^{\prime}u^{+}(\theta,\gamma)=H^{\prime}V(\rho,\gamma)\left(\begin{array}[]{cc}\underline{X},&\mu^{*}\end{array}\right)\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right)+H^{\prime}\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1}u_{2}).$ (33)
+Let
+$K_{n}(\rho,\gamma)$ $=$ $n^{-1}H^{\prime}V(\rho,\gamma)\begin{array}[]{cc}(\underline{X},&\mu^{*})\end{array}.$ (35)
+Observe that
+$n^{-1}E\left[H^{\prime}u^{+}(\theta,\gamma)\left|\mathcal{Z}_{n}\right.\right]$ $=$ $E\left(K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right)\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right).$ (38)
+where the r.h.s. is seen to hold in light of (33) and (35), and since $H$ is $\mathcal{Z}_{n}$-measurable and thus
+$E\left[H^{\prime}\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1}u_{2})\left|\mathcal{Z}_{n}\right.\right]=0.$
+The assumptions imply that $K_{n}$ has full rank a.s. To see this, note that in light of (73), (74), Lemma E.1(iv) in online Appendix E implies that
+$V(\rho,\gamma)=\operatorname{diag}_{c=1}^{C}\left(\frac{1}{\gamma_{\tau_{c}}}\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\frac{1}{\gamma_{\tau_{c}}}\left(1+\rho\right)J_{c}^{\ast}\right),$ (39)
+Under Assumptions 5 and 7, there exists some constants $c_{e}$ and $C_{e}$ such that $0<c_{e}<\frac{1}{\gamma_{\tau_{c}}}\frac{n_{c}-1-\rho}{n_{c}-1}<C_{e}<\infty$ and $0<c_{e}<\frac{1}{\gamma_{\tau_{c}}}\left(1+\rho\right)<C_{e}<\infty$.
+In light of Lemma E.1(iii), the eigenvalues of $V(\rho,\gamma)$ are uniformly bounded below by $c_{e}>0$.
+By Assumption 8 we have $\lambda_{min}\left(n^{-1}\underline{X}^{\prime}\underline{X}\right)\geq\xi_{X}>0$ and thus $\lambda_{min}\left(n^{-1}\underline{X}^{\prime}V(\rho,\gamma)\underline{X}\right)\geq c_{e}\xi_{X}>0$ uniformly for all $\gamma$, $\rho$, and $n$.
+Recalling that $H=\left(\underline{X},z\right)$ and recalling the expression for $K_{n}(\theta,\gamma)$ given in (35) the matrix can be written as
+$K_{n}(\rho,\gamma)$ $=\frac{1}{n}\left(\begin{array}[]{cc}\underline{X}^{\prime}V(\rho,\gamma)\underline{X}&\underline{X}^{\prime}V(\rho,\gamma)\mu^{*}\\ z^{\prime}V(\rho,\gamma)\underline{X}&z^{\prime}V(\rho,\gamma)\mu^{*}\end{array}\right).$ (42)
+Thus by the determinant for partitioned matrices, and using the shorthand notation $V=V(\rho,\gamma)$,
+$\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]\right)\right|$ (43)
+$=$ $\left|\operatorname{det}\left(\frac{1}{n}E\left[\underline{X}^{\prime}V\underline{X}\left|\mathcal{Z}_{n}\right.\right]\right)\right|\left|\operatorname{det}\left\{E\left[\left(\frac{1}{n}z^{\prime}V\mu^{*}-\frac{1}{n}z^{\prime}V\underline{X}(\frac{1}{n}\underline{X}^{\prime}V\underline{X})^{-1}\frac{1}{n}\underline{X}^{\prime}V\mu^{*}\right)\left|\mathcal{Z}_{n}\right.\right]\right\}\right|$
+$\geqslant$ $\left(\xi_{X}c_{e}\right)^{p_{x}}\left|\frac{1}{n}E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}\mu^{*})\left|\mathcal{Z}_{n}\right.\right]\right|.$
+Observe furthermore that $E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}y_{2})\left|\mathcal{Z}_{n}\right.\right]=E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}\mu^{*})\left|\mathcal{Z}_{n}\right.\right]$ in light of (13) and since $Q_{V^{1/2}X}$ is orthogonal to $V^{1/2}\underline{X}$ and $E\left[u_{i2}|\mathcal{Z}_{n}\right]=0$ by Theorem 2.1.
+Thus
+$\inf_{\gamma\in\Gamma,\rho\in[-K_{\rho},K_{\rho}]}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]\right)\right|\geq$ (44)
+$\left(\xi_{X}c_{e}\right)\left(\inf_{\gamma\in\Gamma,\rho\in[-K_{\rho},K_{\rho}]}n^{-1}\left|E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}y_{2})\left|\mathcal{Z}_{n}\right.\right]\right|\right)$
+$\geq\left(\xi_{X}c_{e}\right)^{p_{x}}K_{y}>0$
+in light of Assumption 8.
+This proves that $E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]$ in (38) has full rank for all admissible values of $\rho$ and $\gamma$, and consequently $E\left[m_{n}^{(l)}(\theta,\gamma)\right]=0$ if and only if $f_{1}=f_{1,0}$ and $\delta=\delta_{0}$.
+Therefore $f_{1,0}$ and $\delta_{0}$ are identified from the linear moment condition.
+Once $f_{1,0}$ and $\delta_{0}$ are identified, $\rho$ can be identified from the quadratic moment condition.
+To see this, note that when evaluated at $\delta_{0}$ and $f_{1,0},$ $u^{+}(\theta,\gamma)$ in Equation (A) becomes
+$u^{+}(\theta,\gamma)|_{f_{1,0},\delta_{0}}$ $=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}(I+\rho_{0}M)(u_{1}-f_{1,0}u_{2})$
+$=\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1,0}u_{2}).$
+Then,
+$\left.E\left[u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)|\mathcal{Z}_{n}\right]\right|$
+$=$ $E\left[(u_{1}-f_{1,0}u_{2})^{\prime}\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)(u_{1}-f_{1,0}u_{2})|\mathcal{Z}_{n}\right]$
+$=$ $E\left[\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)|\mathcal{Z}_{n}\right].$
+[…]
+Online Appendix for “Differential Test Performance and Peer Effects” by Guido M. Kuersteiner, Ingmar R. Prucha and Ying Zeng
+Appendix B Assumptions and Results for $u_{it}$
+We impose the following formal assumption about the variance of the error structure.
+Assumption 9.
+Let $u_{it}=y_{it}^{*}-\kappa_{i}f_{t}$ and let $i=n_{1}+\ldots+n_{c-1}+r$ with $1\leq r\leq n_{c}$, i.e., $i$ refers to student $r$ in class c.
+Let $u_{-i,t}=\left(u_{1,t},...,u_{i-1,t},u_{i+1,t},...,u_{n,t}\right)$.
+Define the filtration $\mathcal{B}_{n,i,t}=\sigma\left(S,\alpha,\varphi,X,z,\zeta^{*},u_{-t},u_{-i,t}\right)\text{\,for $t=1,2$}$ where $u_{-t}$ represents information about the other test.
+Then the conditional variance of $u_{it}$ is given by $E\left[u_{it}^{2}|\mathcal{B}_{n,i,t}\right]=\sigma_{t}^{2}\rho_{\tau_{c}}^{2},$where $\tau_{c}$ is $\mathcal{Z}_{n}$-measurable and $\tau_{c}\in\{1,\ldots,J\}$, and $c_{u}\leq\sigma_{t}^{2},\rho_{j}^{2}\leq C_{u}$ for some finite positive bounds.
+Assumption 9 implies that for $t=1,2$, $\Sigma_{t}\equiv E\left[u_{t}u_{t}^{\prime}|\mathcal{Z}_{n}^{\ast}\right]=\sigma_{t}^{2}\operatorname{diag}_{c=1}^{C}\left(\rho_{\tau_{c}}^{2}I_{c}\right)$.
+Observing that $f_{t}$ is a fixed constant, conditional on $S,\zeta^{*},\alpha,X$ and $z$, knowledge of $y_{it}^{*}$ is equivalent to knowledge of $u_{it}$.
+Thus we have the following equivalence of information sets [^15] See the proof of Lemma B.1 for more detail.
+=$\mathcal{F}{}_{n,i,t}=\mathcal{B}_{n,i,t}\,\text{for}\,t=1,2.$
+Assumption 9 imposes a similar factor structure on conditional variances as does Assumption 2 for conditional means.
+By Theorem 2.1 the errors $u_{it}$ are mutually uncorrelated over tests as well as cross-sectionally.
+Consequently the variance covariance matrix of $u_{1}-f_{1}u_{2}$ is given by,
+$\Omega=\operatorname{Var}\left(u_{1}-f_{1}u_{2}|\mathcal{Z}_{n}^{*}\right)=\Sigma_{1}+f_{1}^{2}\Sigma_{2}=\left(\sigma_{1}^{2}+f_{1}^{2}\sigma_{2}^{2}\right)\operatorname{diag}_{c=1}^{C}\left(\rho_{\tau_{c}}^{2}I_{c}\right).$
+Define $\gamma_{j}^{2}=\left(\sigma_{1}^{2}+f_{1}^{2}\sigma_{2}^{2}\right)\rho_{j}^{2}$, for $j=1,...,J$ and $\gamma=\left(\gamma_{1},...,\gamma_{J}\right)^{\prime}$.
+Then $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)$ and the conditions about $\Omega(\gamma)$ and $\gamma$ in Assumption 5 hold.
 Lemma B.1.
 Let Assumption 4 hold.
-Then, $u_{it}$ defined in Theorem 2.1 is a martingale difference sequence w.r.t to the filtration $\mathcal{A}_{n,\nu}$ with $\mathcal{A}_{n,0}=\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,X,z,\zeta^{*}\right)$ and
-$\mathcal{A}_{n,\left(t-1\right)n+i-1}=\sigma\left(S,\alpha,X,z,\zeta^{*},\left\{u_{j,t-1}\right\}_{j=1}^{n},\left\{u_{jt}\right\}_{j=1}^{i-1}\right)\text{\,for $t=1,2;i=1,\ldots,n$}$ (43)
+Then, $u_{it}$ defined in Theorem 2.1 is a martingale difference sequence w.r.t to the filtration $\mathcal{A}_{n,\nu}$ with $\mathcal{A}_{n,0}=\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,\varphi,X,z,\zeta^{*}\right)$ and
+$\mathcal{A}_{n,\left(t-1\right)n+i-1}=\sigma\left(S,\alpha,\varphi,X,z,\zeta^{*},\left\{u_{j,t-1}\right\}_{j=1}^{n},\left\{u_{jt}\right\}_{j=1}^{i-1}\right)\text{\,for $t=1,2;i=1,\ldots,n$}$ (45)
 and where $u_{j,0}=0$ to handle the case $t=1.$ In addition, $\sup_{i,t}E\left[\left|u_{it}\right|^{4+\eta}\right]<\infty$ for some $\eta>0$.
@@
 Since the bound on the RHS is uniform in $i$ and $t$ by assumption it follows that $\sup_{i,t}E\left[\left|u_{it}\right|^{4+\eta}\right]<\infty.$
-Appendix C Proofs
-Proof of Theorem 2.1.
-By the assumption on the boundedness of the ($4+\eta$)-th moments all considered expectations and conditional expectations exist.
-Since $u_{it}=y_{it}^{*}-E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]$ we have $E\left[u_{it}|\mathcal{F}_{n,i,t}\right]=0$ by construction.
-Observing that $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=\kappa_{i}f_{t}$ and that $\kappa_{i}$ is $\mathcal{Z}_{n}^{*}$-measurable it follows that conditional on $\mathcal{Z}_{n}^{*}$ knowledge of $y_{it}^{*}$ is equivalent to knowledge of $u_{it}$.
-To see this let $u_{-i,t}=\left(u_{1t},..,u_{i-1,t},u_{i+1,t},...,u_{n,t}\right)^{\prime}$ and use the notation $u_{-t}$ for the unobserved components related to the other test.
-Define $U=\left(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t},\kappa_{i},f_{t}\right)$ assuming for simplicity of notations that vectors are row vectors, and let $V=\left(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*},\kappa_{i},f_{t}\right)$.
-Then $V=g(U)$ where $g(.)$ is Borel measurable and thus $\sigma(V)\subseteq\sigma(U)$.
-Analogously $U=h(V)$ where $h(.)$ is Borel measurable and thus $\sigma(U)\subseteq\sigma(V)$.
-Consequently $\sigma(U)=\sigma(V)$.
-Since $\kappa_{i},f_{t}$ are measurable w.r.t.
-$\sigma(S,\alpha,X,\zeta^{*},z)$ it follows that
-$\sigma(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t},\kappa_{i},f_{t})=\sigma(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t})$
-and
-$\sigma(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*},\kappa_{i},f_{t})=\sigma(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*}),$
-which shows that
-$\mathcal{B}_{n,i,t}=\sigma\left(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t}\right)=\sigma\left(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*}\right)=\mathcal{F}_{n,i,t}.$
-Next observe that the $u_{jt}$ are $\mathcal{F}_{n,i,t}$-measurable for $i\neq j$, and hence by iterated expectations $\operatorname{Cov}\left(u_{it}u_{jt}|\mathcal{F}_{n,i,t}\right)=E\left[u_{jt}E\left(u_{it}|\mathcal{F}_{n,i,t}\right)\right]=0$.
-Observe further that for $s<t$ the $u_{js}$ are $\mathcal{F}_{n,i,t}$-measurable for all $i$, and hence by iterated expectations $\operatorname{Cov}\left(u_{it}u_{js}|\mathcal{F}_{n,i,t}\right)=E\left[u_{js}E\left[u_{it}|\mathcal{F}_{n,i,t}\right]\right]=0$.
-The corresponding claims regarding the mean and covariances of the $u_{it}$ conditional on $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$ follow immediately by iterated expectations.
-This proves the first part of the lemma.
-To prove the second part, let $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$ be some information set. Then given that Assumption 2 holds, it follows by iterated expectations that
-E$\left[y_{ii}^{*}-\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]$ $=$ $E\left[E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]|\mathcal{G}_{n,i,t}\right]-E\left[\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]=E\left[\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]-E\left[\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]=0.$
-Thus by the almost sure uniqueness of conditional expectations, Assumption 2 cannot hold if $E\left[y_{it}^{*}-\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]\neq 0$. ∎
-Note that Lemma 4.1, Corollary 4.1, and Lemma 4.3 can be viewed as special cases of Lemma 4.2. We therefore prove Lemma 4.2 first.
-Proof of Lemma 4.2.
-In light of (9) it follows that,
-$y_{1}-f_{1}y_{2}=\mu^{\ast}(f_{1,0}-f_{1})+\underline{X}\delta(f_{1},\rho_{0},\beta_{0})+(I+\rho_{0}M)(u_{1}-f_{1}u_{2}),$ (45)
-where $\delta(f_{1},\rho_{0},\beta_{0})$ is defined in (7) and $\delta(f_{1,0},\rho_{0},\beta_{0})=\delta_{0}$. Consequently,
-$y_{1}-f_{1}y_{2}-\text{$\underline{X}$}\delta=\mu^{\ast}(f_{1,0}-f_{1})+\underline{X}\left(\delta(f_{1},\rho_{0},\beta_{0})-\delta\right)+(I+\rho_{0}M)(u_{1}-f_{1}u_{2}).$ (46)
-Substitution of (46) into (14) yields
-$u^{+}(\theta,\gamma)$ $=$ $\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$
-×$\times\left\{\left(f_{1,0}-f_{1}\right)\mu^{\ast}+\underline{X}\left(\delta(f_{1},\rho_{0},\beta_{0})-\delta\right)+(I+\rho_{0}M)(u_{1}-f_{1}u_{2})\right\}$
-$=$ $V(\rho,\gamma)\left(\begin{array}[]{cc}\underline{X},&\mu^{*}\end{array}\right)\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right)+\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1}u_{2}),$ (51)
-recalling that $V(\rho,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$ and $\mathcal{M}(\rho)=(I+\rho M)^{-1}(I+\rho_{0}M)$.
-We first analyze the linear moment conditions. Note that
-$H^{\prime}u^{+}(\theta,\gamma)=H^{\prime}V(\rho,\gamma)\left(\begin{array}[]{cc}\underline{X},&\mu^{*}\end{array}\right)\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right)+H^{\prime}\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1}u_{2}).$ (52)
-Let
-$K_{n}(\rho,\gamma)$ $=$ $n^{-1}H^{\prime}V(\rho,\gamma)\begin{array}[]{cc}(\underline{X},&\mu^{*})\end{array}.$ (54)
-Observe that
-$n^{-1}E\left[H^{\prime}u^{+}(\theta,\gamma)\left|\mathcal{Z}_{n}\right.\right]$ $=$ $E\left(K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right)\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right).$ (57)
-where
-the r.h.s. is seen to hold in light of (52) and (54), and since $H$ is $\mathcal{Z}_{n}$-measurable and thus
-$E\left[H^{\prime}\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1}u_{2})\left|\mathcal{Z}_{n}\right.\right]=0.$
-The assumptions imply that $K_{n}$ has full rank a.s. To see this, note that in light of (39), (40), Lemma (B.1)(iv) implies that
-$V(\rho,\gamma)=\operatorname{diag}_{c=1}^{C}\left(\frac{1}{\gamma_{\tau_{c}}}\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\frac{1}{\gamma_{\tau_{c}}}\left(1+\rho\right)J_{c}^{\ast}\right).$ (58)
-Under Assumptions 4 and 6, there exists some constants $c_{e}$ and $C_{e}$ such that $0<c_{e}<\frac{1}{\gamma_{\tau_{c}}}\frac{n_{c}-1-\rho}{n_{c}-1}<C_{e}<\infty$ and $0<c_{e}<\frac{1}{\gamma_{\tau_{c}}}\left(1+\rho\right)<C_{e}<\infty$.
-In light of Lemma B.1(iii), the eigenvalues of $V(\rho,\gamma)$ are uniformly bounded below by $c_{e}>0$.
-By Assumption 7 we have $\lambda_{min}\left(n^{-1}\underline{X}^{\prime}\underline{X}\right)\geq\xi_{X}>0$ and thus $\lambda_{min}\left(n^{-1}\underline{X}^{\prime}V(\rho,\gamma)\underline{X}\right)\geq c_{e}\xi_{X}>0$ uniformly for all $\gamma$, $\rho$, and $n$.
-Recalling that $H=\left(\underline{X},z\right)$ and recalling the expression for $K_{n}(\theta,\gamma)$ given in (54) the matrix can be written as
-$K_{n}(\rho,\gamma)$ $=\frac{1}{n}\left(\begin{array}[]{cc}\underline{X}^{\prime}V(\rho,\gamma)\underline{X}&\underline{X}^{\prime}V(\rho,\gamma)\mu^{*}\\ z^{\prime}V(\rho,\gamma)\underline{X}&z^{\prime}V(\rho,\gamma)\mu^{*}\end{array}\right).$ (61)
-Thus by the determinant for partitioned matrices, and using the shorthand notation $V=V(\rho,\gamma)$,
-$\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]\right)\right|$ (62)
-$=$ $\left|\operatorname{det}\left(\frac{1}{n}E\left[\underline{X}^{\prime}V\underline{X}\left|\mathcal{Z}_{n}\right.\right]\right)\right|\left|\operatorname{det}\left\{E\left[\left(\frac{1}{n}z^{\prime}V\mu^{*}-\frac{1}{n}z^{\prime}V\underline{X}(\frac{1}{n}\underline{X}^{\prime}V\underline{X})^{-1}\frac{1}{n}\underline{X}^{\prime}V\mu^{*}\right)\left|\mathcal{Z}_{n}\right.\right]\right\}\right|$
-$\geqslant$ $\left(\xi_{X}c_{e}\right)^{p_{x}}\left|\frac{1}{n}E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}\mu^{*})\left|\mathcal{Z}_{n}\right.\right]\right|.$
-Observe furthermore that
-$E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}y_{2})\left|\mathcal{Z}_{n}\right.\right]$ $=$ $E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}\mu^{*})\left|\mathcal{Z}_{n}\right.\right]$
-in light of (9) and since $Q_{V^{1/2}X}$ is orthogonal to $V^{1/2}\underline{X}$ and $E\left[u_{i2}|\mathcal{Z}_{n}\right]=0$ by Theorem 2.1. Thus
-$\inf_{\gamma\in\Gamma,\rho\in[-K_{\rho},K_{\rho}]}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]\right)\right|$$\geq$ $\left(\xi_{X}c_{e}\right)^{p_{x}}\inf_{\gamma\in\Gamma,\rho\in[-K_{\rho},K_{\rho}]}n^{-1}\left|E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}y_{2})\left|\mathcal{Z}_{n}\right.\right]\right|$
-$\geq$ $\left(\xi_{X}c_{e}\right)^{p_{x}}K_{y}>0$
-in light of Assumption 7.
-This proves that $E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]$ in (57) has full rank for all admissible values of $\rho$ and $\gamma$, and consequently $E\left[m_{n}^{(l)}(\theta,\gamma)\right]=0$ if and only if $f_{1}=f_{1,0}$ and $\delta=\delta_{0}$.
-Therefore $f_{1,0}$ and $\delta_{0}$ are identified from the linear moment condition.
-Once $f_{1,0}$ and $\delta_{0}$ are identified, $\rho$ can be identified from the quadratic moment condition.
-To see this, note that when evaluated at $\delta_{0}$ and $f_{1,0},$ $u^{+}(\theta,\gamma)$ in Equation (C) becomes
-$u^{+}(\theta,\gamma)|_{f_{1,0},\delta_{0}}$ $=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}(I+\rho_{0}M)(u_{1}-f_{1,0}u_{2})$
-$=\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1,0}u_{2}).$
-Then,
-$\left.E\left[u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)|\mathcal{Z}_{n}\right]\right|$
-$=$ $E\left[(u_{1}-f_{1,0}u_{2})^{\prime}\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)(u_{1}-f_{1,0}u_{2})|\mathcal{Z}_{n}\right]$
-$=$ $E\left[\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)|\mathcal{Z}_{n}\right].$
-[…]
-Proof of Lemma 4.1 and Corollary 4.1.
-Lemma 4.1 is a special case of Lemma 4.2 without covariates.
-Note that Assumptions 1, 2, 3, 4, 5, and 6 hold.
-Without covariates Equation (61) reduces to $K_{n}\left(\rho,\gamma\right)=z^{\prime}V\left(\rho,\gamma\right)\mu^{*}$ where $V\left(\rho,\gamma\right)=\Omega\left(\gamma\right)^{-1/2}\left(I+\rho M\right)^{-1}.$
-Recall that $z=\left[z_{1}^{\prime},...,z_{c}^{\prime},...,z_{C}^{\prime}\right]^{\prime}$ with $z_{c}=\grave{z}_{c}\mathbf{1}_{c}$. Utilizing Lemma B.1,
-$(I_{c}+\rho M_{c})^{-1}z_{c}=\left(\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\frac{1}{1+\rho}J_{c}^{\ast}\right)\grave{z}_{c}\mathbf{1}_{c}=\frac{1}{1+\rho}\grave{z}_{c}\mathbf{1}_{c}=\frac{z_{c}}{1+\rho}$
-leads to
-$n^{-1}E\left[y_{2}^{\prime}\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}z|\mathcal{Z}_{n}\right]$ $=\frac{1}{1+\rho}n^{-1}E\left[y_{2}^{\prime}\Omega(\gamma)^{-1/2}z|\mathcal{Z}_{n}\right]\geq\frac{1}{1+\rho}K_{y}>0$
-where the inequality follows from the conditions imposed in the lemma, in particular (28).
-With $\rho\in(-1,1)$, E$\left[K_{n}\left(\rho,\gamma\right)|\mathcal{Z}_{n}\right]\neq 0$ such that (57) only has one solution, $f_{1}=f_{1,0}$.
-Identification of $\rho_{0}$ from the quadratic moment is not affected by the absence of covariates.
-Thus Lemma 4.1 follows from Lemma 4.2.
-To prove Corollary 4.1 observe that when $z=\mathbf{1}_{n}$, $n^{-1}\sum_{c=1}^{C}E\left[y_{c2}^{\prime}\mathbf{1}_{c}\right]>K_{y}>0$ implies that (28) holds, observing that
-$n^{-1}\left|E\left[y_{2}^{\prime}\Omega(\gamma)^{-1/2}\mathbf{1}_{n}\left|\mathcal{Z}_{n}\right.\right]\right|=n^{-1}E\sum_{c=1}^{C}\left[\frac{y_{c2}^{\prime}\mathbf{1}_{c}}{\gamma_{\tau_{c}}}\left|\mathcal{Z}_{n}\right.\right]\geq\frac{1}{K_{\gamma}}\frac{1}{n}\sum_{c=1}^{C}E\left[y_{c2}^{\prime}\mathbf{1}_{c}\left|\mathcal{Z}_{n}\right.\right]\geq\frac{K_{y}}{K_{\gamma}}>0.$
-Proof of Lemma 4.3.
+Appendix C Consistency of Parameter Estimators
+Formulating the moment conditions using transformed residuals as in (19) requires consistent estimates of the variance parameters $\gamma$.
+These can be obtained by first obtaining consistent but inefficient parameter estimates.
+To do so, we set all elements of $\gamma=(\gamma_{1},...,\gamma_{J})^{\prime}$ to one.
+Theorem 5.1 implies that the parameters $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ are still identified by the moment conditions in (19).
+First step estimators $(\tilde{\rho},\tilde{f}_{1},\tilde{\delta}^{\prime})^{\prime}$ are defined in (14), (15) and (16).
+[^16] Lee (2007a) considers first step and efficient GMM estimators using linear and quadratic moments.
+To formalize the identification result for these non-efficient first step estimators define the moment vector
+$m_{\epsilon}\left(\theta\right)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}(I+\rho M)\epsilon^{+}\left(\theta\right)\\ \epsilon^{+}\left(\theta\right)^{\prime}A\epsilon^{+}\left(\theta\right)\end{array}\right).$ (47)
+The next lemma shows that the moment conditions in (47) are sufficient to identify the parameters $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$.
+Lemma C.1.
+Let Assumptions 3-7, and Assumption 8(i),(ii) be satisfied.
+Let $m_{\epsilon}(\theta)$ be defined as in (47).
+Assume further that $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$ a.s .
+Then $E\left[m_{\epsilon}\left(\theta_{0}\right)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{\epsilon}\left(\theta\right)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ a.s.
+if $\theta\neq\theta_{0}$ for all $\theta\in\Theta$.
+Proof of Lemma C.1.
 Observe that utilizing (26) we have
@@
 $n^{-1}E\left[H^{\prime}(I+\rho M)\epsilon^{+}\left(\rho,f_{1},\delta\right)\left|\mathcal{Z}_{n}\right.\right]=E\left[\tilde{K}_{n}\left|\mathcal{Z}_{n}\right.\right]\left(\begin{array}[]{c}\delta(f_{1},\rho_{0},\beta_{0})-\delta\\ f_{1,0}-f_{1}\end{array}\right)=0.$ (48)
-Observe that $E\left[\tilde{K}_{n}\right]$ has full rank, as is readily seen from the proof of Lemma 4.2 in light of (C) with $V=I$.
+Observe that $E\left[\tilde{K}_{n}\right]$ has full rank, as is readily seen from the proof of Theorem 5.1 in light of (44) with $V=I$.
 This implies that the only solution to (48) is $\delta=\delta_{0}$ and $f_{1}=f_{1,0}$.
-It now follows that
+With $\mathcal{M}(\rho)=(I+\rho M)^{-1}(I+\rho_{0}M)$, it now follows that
 $\epsilon^{+}\left(\rho,f_{1,0},\delta_{0}\right)=\mathcal{M}\left(\rho\right)(u_{1}-f_{1,0}u_{2})$
@@
 $E\left[\epsilon^{+}\left(\rho,f_{1,0},\delta_{0}\right)^{\prime}A\epsilon^{+}\left(\rho,f_{1,0},\delta_{0}\right)\left|\mathcal{Z}_{n}\right.\right]=\operatorname{tr}\left(\mathcal{M}\left(\rho\right)A\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0.$ (49)
-By Lemma B.2 it follows that the only solution to (65) is $\rho=\rho_{0}.$
-Proof of Theorem 4.1.
-To prove part (i) of the theorem we establish the conditions of Lemma 3.1 in Pötscher and Prucha (1997).
-For the linear moment conditions use (52). We have
+By Lemma E.2 it follows that the only solution to (49) is $\rho=\rho_{0}.$
+We next show that the initial estimators and the efficient estimators are consistent.
+To establish consistency we impose assumptions related to the convergence of sample averages and moments of sample averages to well defined limits.
+These assumptions are adaptations of Assumptions 2, 3 and 5 in Kuersteiner and Prucha (2020) to the present setting which differs somewhat from the framework of sequential exogeneity in that paper.
+Let $i=n_{1}+\ldots n_{c-1}+r$ with $1\leq r\leq n_{c}$.
+Then, with some abuse of notation, denote the $i$-th diagonal element of $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)$ as $\gamma_{(i)}^{2}=\gamma_{\tau_{c}}^{2}$, noting that $\tau_{c}$ varies with $i$.
+Let the linear and quadratic moment vectors $m_{n}^{(l)}(\theta,\gamma)$ and $m_{n}^{(q)}(\theta,\gamma)$ be as defined in (19).
+Let $h_{i}=\left[h_{i1},...,h_{iq}\right]$ denote the $i$-th row vector of $H$, and let $a_{ij}$ denote the the $(i,j)$-th element of matrix $A$.
+Then note that with $u^{+}(\theta_{0},\gamma)=\Omega(\gamma)^{-1/2}(u_{1}-f_{1,0}u_{2})$, the variance of $m_{n}^{(l)}(\theta_{0},\gamma)$ is
+$\frac{1}{n}E\left[H^{\prime}\Omega(\gamma)^{-1}\Omega(\gamma_{0})H\right]=\frac{1}{n}\sum_{i=1}^{n}E\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}h_{i}^{\prime}h_{i}\right]$ (50)
+and the variance of $m_{n}^{(q)}(\theta_{0},\gamma)$ is
+$\frac{2}{n}E\left[\operatorname{tr}\left(\Omega_{0}\Omega^{-1}A\Omega_{0}\Omega^{-1}A\right)\right]=\frac{2}{n}E\sum_{i=1}^{n}\sum_{j=1}^{n}\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}\frac{\gamma_{(j),0}^{2}}{\gamma_{(j)}^{2}}a_{ij}^{2}\right].$ (51)
+The assumption below ensures that for any admissible $\gamma$, $\operatorname{Var}\left(m_{n}^{(l)}(\theta_{0},\gamma)\right)$ and $\operatorname{Var}\left(m_{n}^{(q)}(\theta_{0},\gamma)\right)$ and their corresponding sample analogues converge, respectively, to finite positive definite matrices.
+Assumption 10.
+The following holds :
+$n^{-1}\sum_{i=1}^{n}E\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}h_{i}^{\prime}h_{i}\right]\rightarrow V_{\gamma}^{h},\quad n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}E\left[\frac{\gamma_{(i),0}^{2}}{\gamma_{(i)}^{2}}\frac{\gamma_{(j),0}^{2}}{\gamma_{(j)}^{2}}a_{ij}^{2}\right]\rightarrow V_{\gamma}^{a},$
+where the elements of $V_{\gamma}^{h}$ and $V_{\gamma}^{a}$ are finite , and
+[…]
+The matrix $V_{\gamma}=\operatorname{diag}\left(V_{\gamma}^{h},2V_{\gamma}^{a}\right)$ is positive definite.
+Assumption 11.
+(i) Let $C\left(\theta,\gamma\right)$ be an $n\times n$ matrix of the form $D$, $DP$, $DAD$, $DADP$, or $PDADP$, where $D$ is an $n\times n$ positive diagonal matrix with elements which are uniformly bounded and measurable w.r.t.
+$\mathcal{Z}_{n}$ and where $P=\operatorname{diag}_{c}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right)$ and where $p_{c}$ and $q_{c}$ are continuously differentiable functions of the parameters $\theta$ and $\gamma$ as well as of variables generating $\mathcal{Z}_{n}$, and $I_{c}^{\ast}$ and $J_{c}^{\ast}$ are defined in Appendix E.
+Then
+[…]
+where $\Upsilon_{a},\Upsilon_{b}$ are selected from the set $\left\{H,MH,\mu^{*},u_{1},u_{2}\right\}$.
+(ii) $\lim_{n\rightarrow\infty}\sup_{\theta,\gamma}\left\|E\left[n^{-1}\Upsilon_{a}^{\prime}C\left(\theta,\gamma\right)\Upsilon_{b}\right]-\mathscr{U}_{a,b}\left(\theta,\gamma\right)\right\|=0$ where
+$\mathscr{U}_{a,b}\left(\theta,\gamma\right)=\lim_{n\rightarrow\infty}E\left[n^{-1}\Upsilon_{a}^{\prime}C\left(\theta,\gamma\right)\Upsilon_{b}\right]$
+is bounded and continuously differentiable in $\theta$ and $\gamma$.
+In Kuersteiner, Prucha, and Zeng (2023) we show uniform convergence under more primitive conditions in a closely related situation while directly assuming it here in Assumption 11 to save space.
+Without further assumptions on cross-sectional dependence, high level assumptions about the convergence of sample averages such as $n^{-1}\underline{X}^{\prime}\underline{X}$ are required irrespective.
+The next theorem establishes consistency of the GMM estimator.
+Theorem C.1.
+Let Assumptions 3- 11 be satisfied. Assume that $\bar{\gamma}_{n}\stackrel{{\scriptstyle p}}{{\rightarrow}}\gamma$ for some sequence $\bar{\gamma}_{n}$ and some $\gamma\in\Gamma.$ Then,
+(i) for $\hat{\theta}\left(\gamma\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\gamma\right)$ with $Q_{n}\left(\theta,\gamma\right)$ defined in (23) it follows that $\hat{\theta}\left(\bar{\gamma}_{n}\right)\stackrel{{\scriptstyle p}}{{\rightarrow}}\theta_{0}$.
+(ii) for $\tilde{f}_{1}$ and $\tilde{\delta}$ defined in (14) and (15) and $\tilde{\rho}$ defined in (16) it follows that $\tilde{f}_{1}\stackrel{{\scriptstyle p}}{{\rightarrow}}f_{1,0}$, $\tilde{\delta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\delta_{0}$ and $\tilde{\rho}\stackrel{{\scriptstyle p}}{{\rightarrow}}\rho_{0}.$
+Proof.
+To prove part (i) of the theorem we establish the conditions of Lemma 3.1 in Pötscher and Prucha (1997). For the linear moment conditions use (33). We have
 $n^{-1}H^{\prime}u^{+}(\theta,\gamma)$ $=$ $\left(f_{1,0}-f_{1}\right)n^{-1}H^{\prime}V(\rho,\gamma)\mu^{\ast}$
@@
 $\lim_{n\rightarrow\infty}E\left[n^{-1}H^{\prime}V\left(\rho,\gamma\right)\mu^{*}\right]\equiv\mathscr{U}_{H,\mu}\left(\theta,\gamma\right).$ (53)
-Consider $III$. From Corollary B.1 and Assumption 9, we have $sup_{\theta,\gamma}|III|\rightarrow_{p}0$.
+Consider $III$. From Corollary E.1 and Assumption 11, we have $sup_{\theta,\gamma}|III|\rightarrow_{p}0$.
 Now consider the quadratic moment conditions.
@@
 We omit this dependence for notational convenience and because $f_{1}$ is pinned down by linear moment conditions for $f_{1}$ directly, or in other words by equations related to the second component of $\vartheta$.
-In light of (C) we have
+In light of (A) we have
 $n^{-1}u^{+}(\theta,\gamma)Au^{+}(\theta,\gamma)$ $=$ $n^{-1}\left(\vartheta-\vartheta_{0}\right)^{\prime}\Delta_{1}^{\prime}V(\rho,\gamma)^{\prime}AV(\rho,\gamma)\Delta_{1}\left(\vartheta-\vartheta_{0}\right)$ (54)
@@
 $\lim_{n\rightarrow\infty}n^{-1}E\left[\Delta_{1}^{\prime}V(\rho,\gamma)^{\prime}AV(\rho,\gamma)\Delta_{1}\right]\equiv\mathscr{U}_{A,\Upsilon}\left(\theta,\gamma\right)$ (55)
-By Corollary B.1 and Assumption 9, $II$ converges uniformly to 0. Consider $III$, let $B(\theta,\gamma)=\mathcal{M}(\rho)^{\prime}\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)$, then
+By Corollary E.1 and Assumption 11, $II$ converges uniformly to 0. Consider $III$, let $B(\theta,\gamma)=\mathcal{M}(\rho)^{\prime}\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)$, then
 $\frac{1}{n}(u_{1}-f_{1}u_{2})^{\prime}B(\theta,\gamma)(u_{1}-f_{1}u_{2})$ $=\frac{1}{n}u_{1}^{\prime}B(\theta,\gamma)u_{1}+\frac{f_{1}^{2}}{n}u_{2}^{\prime}B(\theta,\gamma)u_{2}-\frac{2}{n}f_{1}u_{1}^{\prime}B(\theta,\gamma)u_{2}.$
 Each term on the R.H.S.
-converges to its mean uniformly in $\theta$ and $\gamma$ under Corollary B.1 and Assumption 9, and the mean is finite and continuously differentiable in $\theta,\gamma$ Consequently, $\frac{1}{n}(u_{1}-f_{1}u_{2})^{\prime}B(\theta,\gamma)(u_{1}-f_{1}u_{2})$ converges uniformly to the limit of its mean
+converges to its mean uniformly in $\theta$ and $\gamma$ under Corollary E.1 and Assumption 11, and the mean is finite and continuously differentiable in $\theta,\gamma$ Consequently, $\frac{1}{n}(u_{1}-f_{1}u_{2})^{\prime}B(\theta,\gamma)(u_{1}-f_{1}u_{2})$ converges uniformly to the limit of its mean
 $E\left[\frac{1}{n}(u_{1}-f_{1}u_{2})^{\prime}B(\theta,\gamma)(u_{1}-f_{1}u_{2})\right]$ $=\operatorname{tr}\left[\frac{1}{n}B(\theta,\gamma)\left(\Sigma_{1}+f_{1}^{2}\Sigma_{2}\right)\right]$
@@
 We proceed to show, building on our results for finite $n$, that $\theta_{0}$ is also the unique solution vector of the limiting moment condition $\mathfrak{m}\left(\theta,\gamma_{*}\right)=0$.
-Observe that from (54), (66) and (67) we have $\left[\mathscr{U}_{H,x}\left(\theta,\gamma_{*}\right),\mathscr{U}_{H,\mu}\left(\theta,\gamma_{*}\right)\right]=\lim_{n\rightarrow\infty}E\left[K_{n}(\rho,\gamma_{*})\right],$ and by (C) we have $\inf_{\gamma\in\Gamma,\rho}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\right]\right)\right|\geq C_{K}>0$ for some $C_{K}>0$ and all $n.$ Since $\operatorname{det}(.)$ is a continuous function it follows that for any $\theta,\gamma_{*}$ we have
+Observe that from (35), (52) and (53) we have $\left[\mathscr{U}_{H,x}\left(\theta,\gamma_{*}\right),\mathscr{U}_{H,\mu}\left(\theta,\gamma_{*}\right)\right]=\lim_{n\rightarrow\infty}E\left[K_{n}(\rho,\gamma_{*})\right],$ and by (44) we have $\inf_{\gamma\in\Gamma,\rho}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\right]\right)\right|\geq C_{K}>0$ for some $C_{K}>0$ and all $n.$ Since $\operatorname{det}(.)$ is a continuous function it follows that for any $\theta,\gamma_{*}$ we have
 $\left|\operatorname{det}\left(\mathscr{U}_{H,x}\left(\theta,\gamma_{*}\right),\mathscr{U}_{H,\mu}\left(\theta,\gamma_{*}\right)\right)\right|$ $=\left|\operatorname{det}\left(\lim_{n\rightarrow\infty}E\left[K_{n}(\rho,\gamma_{*})\right]\right)\right|$
@@
 $=\lim_{n\rightarrow\infty}\frac{1}{n}\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma_{\ast})^{-1/2}A\Omega(\gamma_{\ast})^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right).$
-By Lemma B.2, $\mathfrak{m}_{q}\left(\rho,f_{1,0},\delta_{0},\gamma_{*}\right)=0$ if and only if $\rho=\rho_{0}$, and consequently $\mathfrak{m}\left(\theta,\gamma_{*}\right)=0$ if and only if $\theta=\theta_{0}$.
+By Lemma E.2, $\mathfrak{m}_{q}\left(\rho,f_{1,0},\delta_{0},\gamma_{*}\right)=0$ if and only if $\rho=\rho_{0}$, and consequently $\mathfrak{m}\left(\theta,\gamma_{*}\right)=0$ if and only if $\theta=\theta_{0}$.
 Assumption 11 implies that $\mathfrak{m}\left(\theta,\gamma_{*}\right)$ and thus $\mathfrak{m}\left(\theta,\gamma_{*}\right)^{\prime}\Xi^{-1}\mathfrak{m}\left(\theta,\gamma_{*}\right)$ is continuous.
@@
 $\equiv$ $I+II$
-[…]
+where we used (12) for $y_{1}-f_{1,0}y_{2}$.
+By Assumptions 8 and 11 as well as (13) it follows that $n^{-1}\underline{X}^{\prime}\underline{X}\rightarrow_{p}\mathscr{U}_{XX}$ where $\mathscr{U}_{XX}$ is positive definite and symmetric and $n^{-1}\underline{X}^{\prime}y_{2}\rightarrow_{p}\mathscr{U}_{Xy}$ where $\mathscr{U}_{Xy}$ is bounded.
+This implies that $II\rightarrow_{p}0.$ For $I$ recall from above that $n^{-1}\underline{X}^{\prime}\left(I_{n}+\rho_{0}M\right)\left(u_{1}-u_{2}f_{1,0}\right)\rightarrow_{p}0.$ This shows that $\tilde{\delta}\rightarrow_{p}\delta_{0}.$
 Now consider the quadratic moment conditions.
@@
 Uniform convergence in probability to zero for the second term, $II,$ follows by Assumption 11.
-Finally, using Lemma B.2 and Assumption 9 it follows that
+Finally, using Lemma E.2 and Assumption 11 it follows that
 […]
 Then the consistency of $\tilde{\rho}$ follows from Lemma 3.1 in Pötscher and Prucha (1997), and analogous arguments as those used for part (i) of the theorem. ∎
-Proof of Theorem 4.2.
+Appendix D Asymptotic Normality of Parameter Estimators
+In the following we develop an asymptotically justified inference theory for our proposed GMM estimator.
+For that purpose we first discuss consistent estimators for the variance parameters $\gamma=(\gamma_{1},...,\gamma_{J})^{\prime}$, which are treated as nuisance parameters in the objective function of our GMM estimator.
+Assumption 12.
+Let $\omega_{j}\equiv\lim{}_{n\rightarrow\infty}n^{-1}\sum_{c=1}^{C}n_{c}1(\tau_{c}=j)$ and assume that $\omega_{j}$ exists and $\omega_{j}>0$ for all $j=1,...,J$.
+Assumption 12 guarantees that there are enough data to estimate $\gamma_{j}^{2}$ for each $j$ by requiring that the fraction $\omega_{j}$ of all students in classrooms of type $j$ is asymptotically non-negligible for all types of classrooms.
+The first step inefficient but consistent estimators $\tilde{f}_{1}$, $\tilde{\delta}$ and $\tilde{\rho}$ defined in (14), (15) and (16) can be used to obtain consistent estimates of the variance parameters $\gamma.$ Form the residuals
+$\tilde{\epsilon}=\epsilon^{+}(\tilde{\theta})=\left(I_{n}+\tilde{\rho}M\right)^{-1}\left(y_{1}-y_{2}\tilde{f}_{1}-\underline{X}\tilde{\delta}\right)$ (68)
+and organize the residuals by classrooms $c$ as $\tilde{\epsilon}=\left(\tilde{\epsilon}_{1}^{\prime},....,\tilde{\epsilon}_{C}^{\prime}\right)^{\prime}$. Then construct the variance estimators
+$\hat{\gamma}_{j}^{2}=\left(N_{j}-p_{x}-1\right)^{-1}\sum_{c=1}^{C}\tilde{\epsilon}_{c}^{\prime}\tilde{\epsilon}_{c}1\left(\tau_{c}=j\right),$ (69)
+where $N_{j}=\sum_{c=1}^{C}n_{c}1\left(\tau_{c}=j\right)$ is the number of students in classrooms of type $j$, and where $p_{x}$ denotes the number of columns in $\underline{X}$. Now set
+$\hat{\gamma}=\left(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{J}\right)^{\prime}.$ (70)
+Efficient GMM estimators can now be formed by plugging $\hat{\gamma}$ into $u^{+}(\theta,\gamma)$ defined in (18).
+Recall that $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}(\gamma_{\tau_{c}}^{2}I_{c}).$ In the following we denote with $K_{u}$ a generic finite constant (which is taken, w.o.l.o.g., to be greater than one) and which is invariant over $t=1,2$, $i=1,\ldots,n$, $n\geq 1$.
+The central limit theorem in Proposition 3 of Kuersteiner and Prucha (2020) can be applied to obtain distributional approximations for estimators based on $u^{+}$.
+The central limit theorem of Kuersteiner and Prucha (2020) requires that $E\left[\left|u_{it}u_{is}\right|^{1+\eta_{u}}|\mathcal{Z}_{n}\right]\leqslant K_{u}$ with $\eta_{u}>0$ for $t,s=1,2\text{ }$ and all $i$.
+This condition is implied by Assumption 2, as Lemma B.1 demonstrates.
+When setting
+$\tilde{\gamma}=\left(1,1,...,1\right)^{\prime}$ (71)
+it follows that $\Omega(\tilde{\gamma})=I_{n}$.
+The feasible efficient GMM estimator for $\theta_{0}=\left(\rho_{0},f_{1,0},\delta_{0}^{\prime}\right)^{\prime}$ is defined as $\hat{\theta}\left(\hat{\gamma}\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\hat{\gamma}\right)$ where $Q_{n}\left(\theta,\gamma\right)$ is given in (23) and the inefficient counterpart $\hat{\theta}\left(\tilde{\gamma}\right)$ is defined in an analogous way.
+The next theorem establishes the limiting distribution of $\hat{\theta}\left(\tilde{\gamma}\right)$ and $\hat{\theta}\left(\hat{\gamma}\right).$ The limiting distribution of $\hat{\theta}\left(\tilde{\gamma}\right)$ for $\tilde{\gamma}$ defined in (71) provides a distributional result for an inefficient GMM estimator.
+The limiting distribution of the first step estimators defined in (14), (15) and (16) can be deduced from the theorem by working out the limiting distribution of $m_{\epsilon}\left(\theta\right).$ We omit the details.
+Theorem D.1.
+Let Assumptions 3-12 hold.
+Let $\hat{\gamma}$ be as defined in (70) and $\tilde{\gamma}$ as defined in (71).
+For $\gamma_{n}\rightarrow_{p}\gamma$ let $\operatorname{plim}_{n\rightarrow\infty}n^{-1/2}\partial m_{n}\left(\hat{\theta}\left(\gamma_{n}\right),\gamma_{n}\right)/\partial\theta=G(\gamma)$ and $\Xi=\operatorname{plim}_{n\rightarrow\infty}\Xi_{n}$, where $\Xi_{n}$ is defined in (22).
+Then $\hat{\gamma}$$\rightarrow_{p}\gamma_{0}$, and for $\gamma_{n}=\hat{\gamma}$ or $\gamma_{n}=\tilde{\gamma}$,
+$\sqrt{n}\left(\hat{\theta}\left(\gamma_{n}\right)-\theta_{0}\right)\rightarrow_{d}N(0,\Psi_{\gamma})$
+where the limiting variance covariance matrix $\Psi_{\gamma}$ has the form
+$\Psi_{\gamma}=\left(G^{\prime}\Xi^{-1}G\right)^{-1}G^{\prime}\Xi^{-1}V_{\gamma}\Xi^{-1}G\left(G^{\prime}\Xi^{-1}G\right)^{-1}.$
+with $V_{\gamma}$ defined in Assumption 10 and $G$ is the shorthand notation for $G(\gamma)$.
+When $\gamma_{n}=\hat{\gamma}$ then $V_{\gamma}=\Xi$ and $\Psi_{\gamma}=\left(G(\gamma_{0})^{\prime}\Xi^{-1}G(\gamma_{0})\right)^{-1}$.
+In Appendix F we provide explicit formulas for the derivatives $G$.
+Together with the expressions for $\Xi_{n}$ it is then easy to construct sample based estimators $\hat{G}$ and $\hat{\Xi}$.
+Letting $\hat{\Psi}_{\gamma}=\left(\hat{G}^{\prime}\hat{\Xi}^{-1}\hat{G}\right)^{-1}$ it follows from a standard Slutsky argument that $\hat{\Psi}_{\gamma}^{-1/2}\sqrt{n}\left(\hat{\theta}\left(\hat{\gamma}\right)-\theta_{0}\right)\rightarrow_{d}N\left(0,I\right).$
+Proof of Theorem D.1.
 First show consistency of $\hat{\gamma}$.
@@
 $=$ $o_{p}\left(1\right),$
-observing that $\rho_{0}-\tilde{\rho}=o_{p}\left(1\right)$ by Theorem 4.1(ii), that the first and second supremum converges to zero a.s.
+observing that $\rho_{0}-\tilde{\rho}=o_{p}\left(1\right)$ by Theorem C.1(ii), that the first and second supremum converges to zero a.s.
 by Assumption 11(i) and that $\lim_{n}E\left[n^{-1}\operatorname{tr}\left(C\left(\rho\right)\Omega_{0}\right)\right]$ is uniformly bounded by Assumption 11(ii).
@@
 $=\frac{\partial m_{n}(\hat{\theta},\gamma_{n})^{\prime}}{\sqrt{n}\partial\theta}\hat{\Xi}_{n}^{-1}m_{n}(\theta_{0},\gamma_{n})+\frac{\partial m_{n}(\hat{\theta},\gamma_{n})^{\prime}}{\sqrt{n}\partial\theta}\hat{\Xi}_{n}^{-1}\frac{\partial m_{n}(\bar{\theta},\gamma_{n})}{\sqrt{n}\partial\theta}\sqrt{n}\left(\hat{\theta}-\theta_{0}\right)$
-where $\left\|\bar{\theta}-\theta_{0}\right\|\leq\left\|\hat{\theta}-\theta_{0}\right\|$ and with some abuse of notation it is understood that the between value $\bar{\theta}$ differs among rows of $m_{n}\left(.\right).$ Using the explicit derivatives given in Appendix D it follows from Assumption 9(i) and (ii) that for any sequence $\gamma_{n}$ with $\gamma_{n}\rightarrow_{p}\gamma_{*}$ we have $n^{-1/2}\partial m_{n}\left(\hat{\theta},\gamma_{n}\right)/\partial\theta\rightarrow_{p}G(\gamma_{\ast})$ and $n^{-1/2}\partial m_{n}\left(\bar{\theta},\gamma_{n}\right)/\partial\theta\rightarrow_{p}G(\gamma_{\ast})$ where $G(\gamma_{\ast})$ is fixed matrix with full column rank and $\hat{\Xi}_{n}\rightarrow_{p}\Xi$ with $\Xi$ a full-rank square matrix.
+where $\left\|\bar{\theta}-\theta_{0}\right\|\leq\left\|\hat{\theta}-\theta_{0}\right\|$ and with some abuse of notation it is understood that the between value $\bar{\theta}$ differs among rows of $m_{n}\left(.\right).$ Using the explicit derivatives given in Appendix F it follows from Assumption 11(i) and (ii) that for any sequence $\gamma_{n}$ with $\gamma_{n}\rightarrow_{p}\gamma_{*}$ we have $n^{-1/2}\partial m_{n}\left(\hat{\theta},\gamma_{n}\right)/\partial\theta\rightarrow_{p}G(\gamma_{\ast})$ and $n^{-1/2}\partial m_{n}\left(\bar{\theta},\gamma_{n}\right)/\partial\theta\rightarrow_{p}G(\gamma_{\ast})$ where $G(\gamma_{\ast})$ is fixed matrix with full column rank and $\hat{\Xi}_{n}\rightarrow_{p}\Xi$ with $\Xi$ a full-rank square matrix.
 This implies that
@@
 We abbreviate assumptions in Kuersteiner and Prucha (2020) with A-KP to avoid confusion with assumptions in this paper.
-The moment bounds of A-KP 1(i) hold by Assumption 2 and Lemma 2.1 which implies that $u_{it}$ and $\mu_{i}^{\ast}$ have bounded moments, and Assumption 7(i) which bounds the moments of $x_{it}$ and $z_{it}.$ A-KP 1(ii), Eq.
+The moment bounds of A-KP 1(i) hold by Assumption 2 and Theorem 2.1 which implies that $u_{it}$ and $\mu_{i}^{\ast}$ have bounded moments, and Assumption 8(i) which bounds the moments of $x_{it}$ and $z_{it}.$ A-KP 1(ii), Eq.
 (15), holds by Lemma B.1, A-KP 1(ii), Eq.
@@
 Now turning to A-KP 2, the $\sigma$-field that corresponds to $\mathcal{B}_{n,t}$ is $\mathcal{A}_{n,\left(t-1\right)n}$ defined in (45).
-Since $\mathcal{Z}_{n}^{*}\subset\mathcal{A}_{n,\left(t-1\right)n}$, the fact that $h_{i}$ and $a_{ij}$ are assumed measurable w.r.t to $\mathcal{Z}_{n}\subset\mathcal{Z}_{n}^{*}$ and elements of $h_{i}$ and the row sums of the absolute elements (and powers thereof) of $A$ are uniformly bounded by Assumptions 5(i)(iii) and 7, A-KP 2(i) and (ii) hold.
+Since $\mathcal{Z}_{n}^{*}\subset\mathcal{A}_{n,\left(t-1\right)n}$, the fact that $h_{i}$ and $a_{ij}$ are assumed measurable w.r.t to $\mathcal{Z}_{n}\subset\mathcal{Z}_{n}^{*}$ and elements of $h_{i}$ and the row sums of the absolute elements (and powers thereof) of $A$ are uniformly bounded by Assumptions 6(i)(iii) and 8, A-KP 2(i) and (ii) hold.
 We also normalize $f_{2}=1$ and restrict the parameter space of $f_{1}$ to a compact interval in Assumption 5 such that A-KP 2(iii) holds.
@@
 ∎
-Appendix D Gradient Vector
-In this section, we derive the gradient vector for the two-step-GMM estimator, which is necessary for the limiting variance covariance matrix in Theorem 4.2.
+Appendix E Auxiliary Lemmas
+To prove Theorems C.1 and D.1, we establish two preliminary lemmas.
+We also provide proofs for Lemma 5.1 and Corollary 5.1 in this section.
+Note that many matrices in this paper, e.g., $M$, $\Omega(\gamma)$, $\Sigma_{t}$ defined in Assumption 9 can be written in the form of $\operatorname{diag}_{c=1}^{C}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right),$where $I_{c}^{\ast}=I_{c}-\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}/n_{c}$, $J_{c}^{\ast}=\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}/n_{c}$ are the residual projection matrix and the projection matrix onto $\mathbf{1}_{c}$ respectively.
+To see this note that
+$M$ $=\operatorname{diag}_{c=1}^{C}\left(-\frac{1}{n_{c}-1}I_{c}^{\ast}+J_{c}^{\ast}\right),$ (72)
+$I+\rho M$ $=\operatorname{diag}_{c=1}^{C}\left(\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\left(1+\rho\right)J_{c}^{\ast}\right),$ (73)
+$\Omega(\gamma)$ $=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)=\operatorname{diag}_{c=1}^{C}\left(\gamma_{\tau_{c}}^{2}I_{c}^{\ast}+\gamma_{\tau_{c}}^{2}J_{c}^{\ast}\right),$ (74)
+$\Sigma_{t}$ $=\sigma_{0,t}^{2}\operatorname{diag}_{c=1}^{C}\left(\rho_{0,\tau_{c}}^{2}I_{c}\right)=\operatorname{diag}_{c=1}^{C}\left(\rho_{0,\tau_{c}}^{2}\sigma_{0,t}^{2}I_{c}^{\ast}+\rho_{0,\tau_{c}}^{2}\sigma_{0,t}^{2}J_{c}^{\ast}\right),\,t=1,2.$ (75)
+The following lemma summarizes some well understood key properties of these matrices for the convenience of the reader.
+Proofs are straight forward and omitted, and can e.g., be found in Kuersteiner, Prucha, and Zeng (2023).
+Lemma E.1.
+The matrices $I_{c}^{\ast}$, $J_{c}^{\ast}$ and $pI_{c}^{\ast}+qJ_{c}^{\ast}$ have the following properties:
+(i) The matrices $I_{c}^{\ast}$, $J_{c}^{\ast}$ are symmetric, idempotent, orthogonal, $I_{c}^{*}+J_{c}^{*}=I$ and $I_{c}^{\ast}\mathbf{1}_{c}=0$, $J_{c}^{\ast}\mathbf{1}_{c}=\mathbf{1}_{c}$.
+(ii) $\operatorname{det}\left(pI_{c}^{\ast}+qJ_{c}^{\ast}\right)=p^{(n_{c}-1)}q$.
+(iii) The eigenvalues of $pI_{c}^{\ast}+qJ_{c}^{\ast}$ are $p$ ($n_{c}-1$ times) and $q$ (once).
+(iv) For $p_{1}$, $p_{2}$, $q_{1}$ and $q_{2}$ arbitrary constants, $(p_{1}I_{c}^{\ast}+q_{1}J_{c}^{\ast})(p_{2}I_{c}^{\ast}+q_{2}J_{c}^{\ast})=(p_{1}p_{2}I_{c}^{\ast}+q_{1}q_{2}J_{c}^{\ast})$, and thus multiplication of matrices of the form $pI_{c}^{\ast}+qJ_{c}^{\ast}$ is commutative and associative.
+(v) If $p\neq 0,$ $q\neq 0$, $(pI_{c}^{\ast}+qJ_{c}^{\ast})$ is invertible with $(pI_{c}^{\ast}+qJ_{c}^{\ast})^{-1}=p^{-1}I_{c}^{\ast}+q^{-1}J_{c}^{\ast}$.
+(vi) $\operatorname{tr}(pI_{c}^{\ast}+qJ_{c}^{\ast})=p(n_{c}-1)+q$.
+(vii) Both $I_{c}^{\ast}$ and $J_{c}^{\ast}$ have row and column sum uniformly bounded in absolute value.
+When $p$ and $q$ are bounded in absolute value, the row and column sums of $pI_{c}^{\ast}+qJ_{c}^{\ast}$ are uniformly bounded in absolute value.
+Corollary E.1.
+Suppose Assumptions 5, 6, 7, 8, 9, and 12 hold. Then $I+\rho M$, $\Omega(\gamma)$, and $\Sigma_{t}$ are non-singular. If $S(\theta,\gamma)$ is an element or product matrix of elements from
+$\left\{M,(I+\rho M),(I+\rho M)^{-1},\Omega(\gamma),\Omega(\gamma)^{-1},\Sigma_{t}^{-1},\Sigma_{t}\right\},$ (76)
+then $S(\theta,\gamma)$ can be written in the form of $\operatorname{diag}_{c=1}\{p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\}$with $p_{c}$, $q_{c}$ continuously differentiable in $\theta$ and $\gamma$.
+The matrix $S(\theta,\gamma)A$ is measurable w.r.t $\mathcal{Z}_{n}$ The row and column sums of $S(\theta,\gamma)$ and $S(\theta,\gamma)A$ are uniformly bounded in absolute value.
+.
+Proof.
+By Assumptions 5, 7 and 9, $\left(n_{c}-1-\rho\right)/\left(n_{c}-1\right)$, $\left(1+\rho\right)$, $\gamma_{j}^{2}$, $\rho_{0,\tau_{c}}^{2}$, and $\sigma_{0,t}^{2}$ in (73), (74), (75) are positive and uniformly bounded away from 0.
+Hence by Lemma E.1(v), $I+\rho M$, $\Omega(\gamma)$ and $\Sigma_{t}$ are non-singular.
+An inspection of Equations (72)-(75) shows that all matrices in the set can be written in the form of $\operatorname{diag}_{c=1}^{C}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right),$ with $p_{c}$ and $q_{c}$ continuously differentiable in $\theta$ and $\gamma$ and uniformly bounded.
+Also note that by Assumption 6, the row and column sums of $A$ are uniformly bounded in absolute value.
+Utilizing remark A.1 in Kelejian and Prucha (2004) the corollary thus follows from Lemma E.1.
+∎
+Lemma E.2.
+Suppose Assumptions 5, 6, 7 and 9 hold.
+Consider the block diagonal matrix $\mathcal{M}\left(\rho\right)=\operatorname{diag}{}_{c=1}^{C}\left(\mathcal{M}_{c}\left(\rho\right)\right)$ with $\mathcal{M}_{c}\left(\rho\right)=\left(I_{c}+\rho M_{c}\right)^{-1}\left(I_{c}+\rho_{0}M_{c}\right)$.
+Then for all $\gamma\in\Gamma$,
+$\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0$
+and
+$\lim_{n\rightarrow\infty}\frac{1}{n}\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0$
+each has only one solution for $\rho\in\left(-1,1\right)$. This solution is given by $\rho=\rho_{0}$.
+Proof.
+Let $D(\rho,\gamma)=\mathcal{M}\left(\rho\right)\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}$ and observe that $D(\rho,\gamma)=\operatorname{diag}{}_{c=1}^{C}\left(D_{c}(\rho,\gamma)\right)$ with
+$D_{c}(\rho,\gamma)$ $=$ $\mathcal{M}_{c}\left(\rho\right)\left(\gamma_{\tau_{c}}^{2}I_{c}\right)^{-1/2}A_{c}\left(\gamma_{\tau_{c}}^{2}I_{c}\right)^{-1/2}\mathcal{M}_{c}\left(\rho\right)\left(\gamma_{0,\tau_{c}}^{2}I_{c}\right)$
+$=$ $\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right).$
+Noting that $\operatorname{tr}\left(A_{c}\right)=0$ and hence $\operatorname{tr}\left(A_{c}I_{c}^{\ast}\right)=\operatorname{tr}\left(A_{c}(I-J_{c}^{\ast})\right)=-\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$, and in light of Lemma E.1
+$\mathcal{M}_{c}\left(\rho\right)$ $=$ $\left(I_{c}+\rho M_{c}\right)^{-1}\left(I_{c}+\rho_{0}M_{c}\right)=\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}I_{c}^{\ast}+\frac{1+\rho_{0}}{1+\rho}J_{c}^{\ast}.$
+We have
+$\operatorname{tr}\left(\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right)$ $=$ $\operatorname{tr}\left(A_{c}\left[\left(\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}\right)^{2}I_{c}^{\ast}+\left(\frac{1+\rho_{0}}{1+\rho}\right)^{2}J_{c}^{\ast}\right]\right)$
+$=$ $\left[\left(\frac{1+\rho_{0}}{1+\rho}\right)^{2}-\left(\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}\right)^{2}\right]\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$
+$=$ $(\rho_{0}-\rho)C(\rho,n_{c})\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$
+where
+$C(\rho,n_{c})=\left[\left(\frac{1+\rho_{0}}{1+\rho}\right)+\left(\frac{n_{c}-1-\rho_{0}}{n_{c}-1-\rho}\right)\right]\left[\frac{1}{1+\rho}+\frac{1}{n_{c}-1-\rho}\right].$
+Hence
+$\operatorname{tr}\left(D(\rho,\gamma)\right)=\sum_{c=1}^{C}\operatorname{tr}\left(D_{c}(\rho,\gamma)\right)=(\rho_{0}-\rho)\sum_{c=1}^{C}\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}C(\rho,n_{c})\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right).$
+Clearly, at $\rho=\rho_{0}$, for all $\gamma\in\Gamma$, we have $\operatorname{tr}\left(D(\rho_{0},\gamma)\right)=0$ for all $n$.
+Next observe that $\operatorname{tr}\left(A_{c}J_{c}^{*}\right)=\mathbf{1}_{c}^{{}^{\prime}}A_{c}\mathbf{1}_{c}/n_{c}\geqslant\text{\text@underline{K}}_{a}>0$ and that for any $n_{c}\geqslant 2$ and $-1<\rho<1$, both $n_{c}-1-\rho$ and $1+\rho$ are positive, and therefore $C(\rho,n_{c})$ is positive for all $\rho\in\left(-1,1\right)$.
+Therefore, $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]<0$ if $\rho>\rho_{0}$, $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]>0$ if $\rho<\rho_{0}$ and $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]$ has the same sign for all $c$.
+In all, for any value of $\gamma$ the only solution for
+$\operatorname{tr}\left(\mathcal{M}\left(\rho\right)\Omega(\gamma,\theta)^{-1/2}A\Omega(\gamma,\theta)^{-1/2}\mathcal{M}\left(\rho\right)\Omega_{0}\right)=0$
+is $\rho=\rho_{0}$.
+Furthermore, by Assumptions 5 and 7 there exists some constants $c_{D}$ and $c_{\gamma}$ such that for any $\rho_{0}$ in the interior of the parameter space, $C(\rho,n_{c})>c_{D}>0$ and $\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}>c_{\gamma}>0$.
+This implies that,
+$\left|\operatorname{tr}(D(\rho))\right|\geqslant|\rho_{0}-\rho|\sum_{c=1}^{C}\text{\text@underline{K}}_{a}c_{D}c_{\gamma},$
+$\left|\frac{1}{n}\operatorname{tr}(D(\rho))\right|\geqslant|\rho_{0}-\rho|\text{\text@underline{K}}_{a}c_{D}c_{\gamma}\frac{C}{n}.$
+By Assumption 7, $C/n\geq 1/\bar{n}_{c}>0$ such that the conclusion follows. ∎
+Proofs of Additional Identification Results.
+Proof of Lemma 5.1 and Corollary 5.1.
+Lemma 5.1 is a special case of Theorem 5.1 without covariates.
+Note that Assumptions 3, 4, 5, 6, 7 and 9 hold.
+Without covariates Equation (42) reduces to $K_{n}\left(\rho,\gamma\right)=z^{\prime}V\left(\rho,\gamma\right)\mu^{*}$ where $V\left(\rho,\gamma\right)=\Omega\left(\gamma\right)^{-1/2}\left(I+\rho M\right)^{-1}.$
+Recall that $z=\left[z_{1}^{\prime},...,z_{c}^{\prime},...,z_{C}^{\prime}\right]^{\prime}$ with $z_{c}=\grave{z}_{c}\mathbf{1}_{c}$. Utilizing Lemma E.1,
+$(I_{c}+\rho M_{c})^{-1}z_{c}=\left(\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\frac{1}{1+\rho}J_{c}^{\ast}\right)\grave{z}_{c}\mathbf{1}_{c}=\frac{1}{1+\rho}\grave{z}_{c}\mathbf{1}_{c}=\frac{z_{c}}{1+\rho}$
+leads to
+$n^{-1}E\left[y_{2}^{\prime}\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}z|\mathcal{Z}_{n}\right]$ $=\frac{1}{1+\rho}n^{-1}E\left[y_{2}^{\prime}\Omega(\gamma)^{-1/2}z|\mathcal{Z}_{n}\right]\geq\frac{1}{1+\rho}K_{y}>0$
+where the inequality follows from the conditions imposed in the lemma, in particular (24).
+With $\rho\in(-1,1)$, $E\left[K_{n}\left(\rho,\gamma\right)|\mathcal{Z}_{n}\right]\neq 0$ such that (38) only has one solution, $f_{1}=f_{1,0}$.
+Identification of $\rho_{0}$ from the quadratic moment is not affected by the absence of covariates.
+Thus Lemma 5.1 follows from Theorem 5.1.
+To prove Corollary 5.1 observe that when $z=\mathbf{1}_{n}$, $n^{-1}\sum_{c=1}^{C}E\left[y_{c2}^{\prime}\mathbf{1}_{c}\right]>K_{y}>0$ implies that (24) holds, observing that
+$n^{-1}\left|E\left[y_{2}^{\prime}\Omega(\gamma)^{-1/2}\mathbf{1}_{n}\left|\mathcal{Z}_{n}\right.\right]\right|=n^{-1}E\sum_{c=1}^{C}\left[\frac{y_{c2}^{\prime}\mathbf{1}_{c}}{\gamma_{\tau_{c}}}\left|\mathcal{Z}_{n}\right.\right]\geq\frac{1}{K_{\gamma}}\frac{1}{n}\sum_{c=1}^{C}E\left[y_{c2}^{\prime}\mathbf{1}_{c}\left|\mathcal{Z}_{n}\right.\right]\geq\frac{K_{y}}{K_{\gamma}}>0.$
+Appendix F Gradient Vector
+In this section, we derive the gradient vector for the two-step-GMM estimator, which is a component of the limiting variance covariance matrix in Theorem D.1.
 Recall that $\theta=\left(\rho,f_{1},\delta^{\prime}\right)^{\prime}$,
@@
 It thus is obvious that $u^{+}(\theta,\gamma)$ and each element of $\frac{u^{+}(\theta,\gamma)}{\partial\theta}$ can be written as linear combinations of $S(\theta,\gamma)\Upsilon$, where $S(\theta,\gamma)$ is an element or the product of elements of the matrix set in (76) and $\Upsilon\in\left\{H,\mu^{*},u_{1},u_{2}\right\}$.
-Applying Corollary B.1 and Assumption 9, the gradient vector converges uniformly to its expected value.
-Appendix E GMM with Randomly Missing Variables
+Applying Corollary E.1 and Assumption 11, the gradient vector converges uniformly to its expected value.
+Appendix G GMM with Randomly Missing Variables
 If class-level characteristics like teacher’s race are missing and the variable is controlled in the model, then the whole class is dropped from the sample.
@@
 Below we discuss the case when individual level variables (control variable or outcome variable) are missing.
-Suppose class $c$ has $n_{c}$ students, and $n_{c}^{obs}\leqslant n_{c}$ of them are observed, which we define as having non-missing values for all variables.
+Suppose class $c$ has $n_{c}$ students, and $n_{c}^{obs}\leqslant n_{c}$ of them are observed.
+which We define the case where $n_{c}^{obs}=n_{c}$ as having non-missing values for all variables.
 Without loss of generality, we order the vectors for the class such that observed students are at the top and unobserved are at the bottom.
@@
 Note that $n_{c}^{unobs}=n_{c}-n_{c}^{obs}$.
-Similarly, $y_{c2}=(y_{c2}^{obs\prime},y_{c2}^{unobs\prime})^{\prime}$, $x_{c}=(x_{c}^{obs\prime},x_{c}^{unobs\prime})$, etc.
+Similarly, $y_{c2}=(y_{c2}^{obs\prime},y_{c2}^{unobs\prime})^{\prime}$, $x_{c}=(x_{c}^{obs\prime},x_{c}^{unobs\prime})^{\prime}$, etc.
 The model for class $c$ is
-$y_{1c}-f_{1}y_{2c}=\text{$\underline{X}$}_{c}\delta+(I_{c}+\rho M_{c})(u_{1c}-f_{1}u_{2c}).$ (82)
-[…]
+$y_{1c}-f_{1}y_{2c}=\text{\text@underline{X}}_{c}\delta+(I_{c}+\rho M_{c})(u_{1c}-f_{1}u_{2c}).$ (77)
+where $\underline{X}_{c}=\left(v_{c}^{c},w_{1,c}^{c},w_{2,c}^{c},v_{c}^{p},w_{1,c}^{p},w_{2,c}^{p},M_{c}v_{c}^{p},M_{c}w_{1,c}^{p},M_{c}w_{2,c}^{p}\right)$.
+Let $x_{c}^{c}=[v_{c}^{c},w_{1,c}^{c},w_{2,c}^{c}]$ be class level variables, $x_{c}^{p}=[v_{c}^{p},w_{1,c}^{p},w_{2,c}^{p}]$ be student personal level characteristics.
+The model for class $c$ then can be written more explicitly as
 $y_{1c}-f_{1}y_{2c}=x_{c}^{c}\delta_{c}+x_{c}^{p}\delta_{p}+M_{c}x_{c}^{p}\delta_{m}+(I_{c}+\rho M_{c})(u_{1c}-f_{1}u_{2c}).$ (78)
@@
 For example, some classes may have variable “$girl$” observed for all students, while race and test information may be missing for some.
-In this way we increase the efficiency of our approximation as having missing value for a variable is much less common than having missing values for all the variables.
+In this way we increase the efficiency of our approximation as observing a missing value for a particular variable is much less common than having missing values for all of the variables related to a particular student.
 Then $\bar{girl}_{c}^{p,obs}=\bar{girl}_{c}^{p}$ as we use all observed values of "$girl"$.
@@
 In all,
-$y_{1c}^{obs}-f_{1}y_{2c}^{obs}$ $=x_{c}^{c,obs}\delta_{c}+x_{c}^{p,obs}\delta_{p}+\tilde{x}_{c}^{m,obs}\delta_{m}+v_{c},$
-[…]
+$y_{1c}^{obs}-f_{1}y_{2c}^{obs}$ $=x_{c}^{c,obs}\delta_{c}+x_{c}^{p,obs}\delta_{p}+\tilde{x}_{c}^{m,obs}\delta_{m}+v_{c}=\text{\text@underline{x}}_{c}^{obs}\delta+v_{c},$
 […]
@@
 $E\left[\epsilon_{c}|x_{c}^{c,obs},x_{c}^{p,obs},\tilde{x}_{c}^{m,obs}\right]=0,$
-$\Omega^{obs}\equiv E\left[\epsilon_{c}\epsilon_{c}^{\prime}\right]=\sigma_{c}^{2}\left((\frac{n_{c}-1-\rho}{n_{c}-1})^{2}I_{n^{obs}}+\frac{1}{n_{c}}((1+\rho)^{2}-(\frac{n_{c}-1-\rho}{n_{c}-1})^{2})\iota_{n^{obs}}\iota_{n^{obs}}^{\prime}\right).$ (85)
-Note that $E\left[v_{c}|\text{$\underline{x}$}_{c}^{obs}\right]$ and $\textrm{var}(v_{c}|\text{$\underline{x}$}_{c}^{obs})$ requires additional assumptions on the distribution of $\xi_{c}$, in particular $E\left[\bar{x}_{c}^{p}-\bar{x}_{c}^{p,obs}|\bar{x}_{c}^{p,obs}\right]=0$.
+$\Omega_{c}^{obs}\equiv E\left[\epsilon_{c}\epsilon_{c}^{\prime}\right]=\sigma_{c}^{2}\left((\frac{n_{c}-1-\rho}{n_{c}-1})^{2}I_{n^{obs}}+\frac{1}{n_{c}}((1+\rho)^{2}-(\frac{n_{c}-1-\rho}{n_{c}-1})^{2})\iota_{n^{obs}}\iota_{n^{obs}}^{\prime}\right).$ (80)
+Note that $E\left[v_{c}|\text{\text@underline{x}}_{c}^{obs}\right]$ and $\textrm{var}(v_{c}|\text{\text@underline{x}}_{c}^{obs})$ requires additional assumptions on the distribution of $\xi_{c}$, in particular $E\left[\bar{x}_{c}^{p}-\bar{x}_{c}^{p,obs}|\bar{x}_{c}^{p,obs}\right]=0$.
 In practice, we ignore the term $\xi_{c}$.
@@
 The formula for the variance in (80) points to the importance for adjusting for original class size $n_{c}$ even if observations are missing randomly. We can then define
-$u_{c}^{+obs}=(\Omega^{obs})^{-1/2}\left(y_{1c}^{obs}-f_{1}y_{2c}^{obs}-\text{$\underline{x}$}_{c}^{obs}\delta\right)$
-and then continue with our GMM estimation strategy.
-Graham(2008) discusses missing values in the online appendix.
-For individual characteristics, “Twenty three kindergarten student records were missing information on free and reduced price school lunch eligibility, in these cases the missing values were replaced with either eligibility status for the same student in the closest of first, second or third grade (17 cases) or the median value among kindergarten students in their school (6 cases).
-In three cases missing student race values were replaced with school median values.” For test scores, he just adjust for the group size like what we do above “The analysis below assumes the pattern of missing test score data is indeed random and hence ignorable.
-Observed test scores are regarded as a random sample from different classrooms of known size.
-”
+$u_{c}^{+obs}=(\Omega_{c}^{obs})^{-1/2}\left(y_{1c}^{obs}-f_{1}y_{2c}^{obs}-\text{\text@underline{x}}_{c}^{obs}\delta\right)$
+and continue with our GMM estimation strategy.
+Appendix H Additional Empirical Results
+H.1 Additional Specification Tests
+Our analysis thus far has documented that the word and reading scores are highly correlated both contemporaneously as well as over the course of one school year.
+This evidence is consistent with restrictions imposed by Assumption 2 and the hypothesis that closely related scores depend on the same unobserved components.
+A caveat is that closely related scores may remove too much common variation when quasi-differences of these scores are used and thus render statistical evidence about peer effects imprecise.
+Panel A of Table 8 reports pseudo $R^{2}$ measures of model estimates for word and reading scores and model specifications (1) through (4) with column (1) only controlling for school fixed effects, and column (4) being the full specification with school and class type effects, student and teacher characteristics as well as peer characteristics.
+Results for other pairs of scores are presented in panels B-F as a comparison.
+The observed pseudo $R^{2}$ measures are all below 0.2 for the word-read pair.
+As expected the pseudo $R^{2}$ generally increases with larger models, or in other words by moving from Column (1) to (4).
+Compared to other pairs of scores, the word-read score pair often has the lowest $R^{2}$ which is an indication that the differential score in these specifications already controls for most of the variation explained by the covariates.
+We take this as further evidence that word-read best fits the model assumptions we impose.
+Table 9 presents standard deviations of $Q_{x}y$ for reading and word scores and $Q_{x}(y_{1}-f_{1}y_{2})$ for $y_{1}$ being word and $y_{2}$ being reading scores for Grade K to 3.
+The left and right panels are for small and regular classes respectively.
+Control variables for specifications (1)-(4) are the same as those in Table 8, as discussed above.
+We see that in Grade K to 3, the ratio of standard deviations of $Q_{x}(y_{1}-f_{1}y_{2})$ to that of $Q_{x}y$ is 1/5 to 1/4.This shows that quasi-differencing removes a significant part of the variation in the test scores, consistent with the term $\mu^{*}$ playing a major role in determining test performance.
+At the same time, the results in Table 9 indicate that the inclusion of controls does not eliminate much of the remaining variation not captured by fixed effects.
+The conclusion is that after differencing there is between 20 to 30 percent of test score variation remaining for the identification of $\rho$.
+H.2 Effects of Individual Characteristics
+We also provide the coefficients of peer averages of individual characteristics (age, black, free lunch, and girl) in Table 10.
+These four variables all pertain to test-invariant individual characteristics.
+The estimated coefficients can be expressed as $\rho(\beta_{v1}^{p}-f_{1}\beta_{v2}^{p})$, where $\rho\beta_{v1}^{p}$ and $\rho\beta_{v2}^{p}$ represent the exogenous peer effect of the variable on tests $t=1$ and $t=2$ respectively.
+Thus, these coefficients capture the differential exogenous peer effects across the two tests.
+We also scale the coefficients for dummies $girl$, $black$ and $free\,lunch$ by 1/100 so that the coefficients are the impact on the outcome of 1 percentage point changes in share of peers who are girl/black/free lunch receivers.
+Except for kindergarten where the presence of girls among peers positively impacts reading more than word study, none of the four variables show significant differential exogenous peer effects.
+Overall, the results indicate small and mostly insignificant peer effects on differential scores.
+This is consistent with a scenario where $y_{2}$ is essentially sufficient to control for both observed and unobserved test invariant determinants of test performance.
+In fact, when $f_{1}=1$ and $\beta_{v1}^{p}=\beta_{v2}^{p}$ these effects would be zero in our model and the overall evidence points to the conclusion that this is approximately the case in the sample as well.
 Table 4: Number of Missing Values for Each Variable
@@
   • 1. School ID, class ID, class size and class type have no missing values.
+Table 5: Summary Statistics
+Grade K Grade 1 Grade 2 Grade 3
+mean sd mean sd mean sd mean sd
+indiviudal characteristics
+free lunch 0.560 0.496 0.588 0.492 0.592 0.491 0.578 0.494
+black 0.326 0.469 0.327 0.469 0.347 0.476 0.333 0.471
+girl 0.486 0.500 0.480 0.500 0.484 0.500 0.480 0.500
+age 6.011 0.349 7.158 0.492 8.242 0.557 9.285 0.586
+teacher characteristics
+black teacher 0.165 0.371 0.174 0.379 0.204 0.403 0.209 0.406
+master degree 0.347 0.476 0.346 0.476 0.373 0.484 0.442 0.497
+years of experience 9.258 5.809 11.633 8.937 13.145 8.655 13.933 8.615
+SAT test scores
+math 485.377 47.698 530.528 43.109 580.613 44.574 617.970 39.841
+read 436.725 31.706 520.787 55.192 583.935 46.043 615.422 38.563
+list 537.475 33.140 567.487 33.674 595.476 34.908 624.119 32.265
+word 434.179 36.759 513.436 53.316 582.986 50.658 610.136 45.041
+Sample Size
+number of students 6325 6829 6840 6801
+number of classes 325 339 340 336
+# small 127 124 133 140
+# regular 99 115 100 89
+# regular/aide 99 100 107 107
+number of schools 79 76 75 75
+  • 1. Mean and standard deviations of student characteristics, teacher charcteristics, and SAT scores, as well as number of students and classes in each grade.
+Table 6: Correlation between other paris
+Rank Correlation Pseudo $R^{2}$ of 2SLS
+$y_{1},y_{2}$ $Q_{X}y_{1},Q_{X}y_{2}$ $y_{1},\hat{y}_{1}$ $y_{2},\hat{y}_{2}$
+$y_{1}=math$, $y_{2}=listening$
+Grade K 0.657 0.584 0.728 0.638
+Grade 1 0.721 0.654 0.766 0.716
+Grade 2 0.678 0.611 0.713 0.705
+Grade 3 0.649 0.610 0.689 0.673
+$y_{1}=reading$, $y_{2}=listening$
+Grade K 0.621 0.554 0.686 0.691
+Grade 1 0.624 0.534 0.704 0.601
+Grade 2 0.667 0.601 0.709 0.699
+Grade 3 0.652 0.623 0.690 0.683
+$y_{1}=reading$, $y_{2}=math$
+Grade K 0.741 0.672 0.697 0.763
+Grade 1 0.737 0.669 0.782 0.753
+Grade 2 0.724 0.673 0.756 0.758
+Grade 3 0.741 0.706 0.759 0.769
+$y_{1}=word$, $y_{2}=listening$
+Grade K 0.558 0.477 0.637 0.610
+Grade 1 0.570 0.483 0.645 0.562
+Grade 2 0.597 0.533 0.639 0.639
+Grade 3 0.524 0.478 0.581 0.552
+$y_{1}=word$, $y_{2}=math$
+Grade K 0.656 0.579 0.643 0.727
+Grade 1 0.686 0.619 0.726 0.715
+Grade 2 0.625 0.564 0.659 0.670
+Grade 3 0.683 0.643 0.712 0.708
+  • 1.
+The left panel is Spearman’s rank correlation between $y_{1}$, $y_{2}$ and between $Q_{X}y_{1}$, $Q_{X}y_{2}$, where $Q_{X}=I-\underline{X}(\underline{X}^{\prime}\underline{X})^{-1}\underline{X}^{\prime}$, $y_{1}$ and $y_{2}$ are raw SAT scores specified for each panel, $\underline{X}$ are the complete set of control variables, including school fixed effects, class type fixed effects, student characteristics, teacher characteristics and peer characteristics.
+  • 2.
+The right panel is pseudo $R^{2}$, i.e., Spearman’s rank correlation between $y_{1}$ and $\hat{y}_{1}$ in Column 3, and between $y_{2}$ and $\hat{y}_{2}$ in Column 4.
+Here $\hat{y}_{1}=\tilde{f}_{1}y_{2}+\text{$\underline{X}$}\tilde{\delta}$, with $\tilde{f_{1}}$ and $\tilde{\delta}$ defined in (14) and (15), $\underline{X}$ is the complete set of controls as described in note 1.
+That is, $\hat{y}_{1}$ is the predicted value of $y_{1}$ from 2SLS for $y_{1}=f_{1}y_{2}+\underline{X}\delta+\epsilon$ with $y_{2}$ instrumented by the constant term, $\hat{y}_{2}$ is defined in a similar manner by reversing the order of $y_{1},y_{2}$.
+Table 7: Estimates of $f_{1}$ and $\rho$ for alternate pairings
+$f_{1}$ $\rho$
+(1) (2) (3) (4) (1) (2) (3) (4)
+$y_{1}=math$, $y_{2}=listening$
+Grade K 0.915 0.912 0.833 0.779 0.721 0.691 0.694 0.688
+(0.009) (0.008) (0.019) (0.148) (0.079) (0.073) (0.075) (0.078)
+Grade 1 0.939 0.936 0.949 0.946 0.627 0.587 0.579 0.578
+(0.008) (0.007) (0.011) (0.063) (0.069) (0.073) (0.073) (0.074)
+Grade 2 0.946 0.944 0.993 1.109 0.637 0.621 0.554 0.525
+(0.005) (0.005) (0.011) (0.068) (0.073) (0.072) (0.071) (0.068)
+Grade 3 0.973 0.974 1.040 1.113 0.747 0.728 0.714 0.703
+(0.002) (0.002) (0.009) (0.068) (0.088) (0.087) (0.086) (0.090)
+$y_{1}=reading$, $y_{2}=listening$
+Grade K 0.812 0.812 0.822 0.989 0.434 0.415 0.407 0.377
+(0.005) (0.005) (0.014) (0.107) (0.079) (0.076) (0.078) (0.083)
+Grade 1 0.944 0.943 1.028 0.978 0.322 0.277 0.256 0.218
+(0.007) (0.007) (0.015) (0.074) (0.051) (0.049) (0.048) (0.046)
+Grade 2 0.970 0.970 1.108 1.187 0.568 0.553 0.571 0.561
+(0.005) (0.005) (0.012) (0.070) (0.073) (0.071) (0.074) (0.074)
+Grade 3 0.987 0.987 1.093 1.151 0.606 0.572 0.613 0.623
+(0.003) (0.003) (0.010) (0.060) (0.091) (0.088) (0.096) (0.101)
+$y_{1}=reading$, $y_{2}=math$
+Grade K 0.888 0.890 0.986 1.297 0.492 0.485 0.519 0.575
+(0.004) (0.004) (0.018) (0.236) (0.065) (0.065) (0.066) (0.076)
+Grade 1 1.005 1.007 1.092 1.032 0.441 0.434 0.434 0.360
+(0.005) (0.006) (0.014) (0.077) (0.063) (0.064) (0.061) (0.061)
+Grade 2 1.026 1.027 1.120 1.077 0.481 0.479 0.475 0.444
+(0.006) (0.006) (0.012) (0.055) (0.071) (0.071) (0.066) (0.066)
+Grade 3 1.015 1.015 1.052 1.036 0.462 0.460 0.452 0.440
+(0.003) (0.003) (0.009) (0.051) (0.078) (0.078) (0.082) (0.086)
+  • 1. To be continued on the next page.
+Table 7: Estimates of $f_{1}$ and $\rho$ for alternate pairings(Continued)
+$f_{1}$ $\rho$
+(1) (2) (3) (4) (1) (2) (3) (4)
+$y_{1}=word$, $y_{2}=listening$
+Grade K 0.817 0.816 0.833 0.954 0.398 0.378 0.373 0.359
+(0.005) (0.004) (0.016) (0.116) (0.066) (0.065) (0.065) (0.068)
+Grade 1 0.934 0.932 1.023 0.964 0.385 0.340 0.308 0.257
+(0.008) (0.008) (0.016) (0.081) (0.054) (0.052) (0.048) (0.047)
+Grade 2 0.978 0.977 1.145 1.284 0.537 0.521 0.512 0.497
+(0.006) (0.007) (0.014) (0.085) (0.070) (0.068) (0.070) (0.069)
+Grade 3 0.974 0.974 1.107 1.145 0.512 0.482 0.518 0.522
+(0.004) (0.004) (0.012) (0.069) (0.095) (0.091) (0.095) (0.097)
+$y_{1}=word$, $y_{2}=math$
+Grade K 0.893 0.895 1.001 1.261 0.517 0.513 0.538 0.578
+(0.006) (0.006) (0.021) (0.242) (0.067) (0.067) (0.070) (0.081)
+Grade 1 0.994 0.994 1.077 1.027 0.487 0.480 0.469 0.388
+(0.004) (0.005) (0.014) (0.089) (0.066) (0.065) (0.062) (0.064)
+Grade 2 1.034 1.034 1.159 1.160 0.493 0.492 0.478 0.459
+(0.007) (0.007) (0.014) (0.074) (0.071) (0.071) (0.066) (0.065)
+Grade 3 1.001 1.001 1.064 1.031 0.371 0.366 0.370 0.357
+(0.004) (0.004) (0.010) (0.054) (0.074) (0.072) (0.074) (0.078)
+school FE Y Y Y Y Y Y Y Y
+class type FE Y Y Y Y Y Y
+stu&tch char Y Y Y Y
+peer char Y Y
+  • 1.
+Estimates and standard errors (in the parentheses) for $f_{1}$ and $\rho$.
+Standard errors are clustered at the classroom level.
+All models allow for heteroscedasticity across small and regular classes.
+Estimates are adjusted for missing observations.
+  • 2.
+As indicated at the bottom of the table, the models may control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.
+Table 8: Pseudo $R^{2}$: $1-var(u_{1}-f_{1}u_{2})/var(y_{1}-f_{1}y_{2})$
+Small Classes Regualr Classes(w/wo Aide)
+(1) (2) (3) (4) (1) (2) (3) (4)
+Panel A: $\mathbf{y_{1}=word,y_{2}=reading}$
+Grade K 0.104 0.104 0.107 0.113 0.140 0.140 0.142 0.146
+Grade 1 0.088 0.088 0.092 0.090 0.093 0.093 0.098 0.093
+Grade 2 0.064 0.064 0.077 0.082 0.065 0.065 0.074 0.078
+Grade 3 0.093 0.093 0.092 0.096 0.033 0.034 0.035 0.038
+Panel B: $y_{1}=math$, $y_{2}=listening$
+Grade K 0.179 0.180 0.206 0.216 0.177 0.179 0.201 0.208
+Grade 1 0.138 0.140 0.143 0.144 0.113 0.115 0.123 0.124
+Grade 2 0.116 0.117 0.123 0.119 0.104 0.106 0.118 0.109
+Grade 3 0.135 0.136 0.147 0.144 0.085 0.087 0.092 0.089
+Panel C: $y_{1}=reading$, $y_{2}=listening$
+Grade K 0.143 0.144 0.153 0.150 0.145 0.146 0.150 0.147
+Grade 1 0.131 0.133 0.154 0.163 0.113 0.116 0.148 0.161
+Grade 2 0.087 0.087 0.144 0.141 0.083 0.084 0.125 0.125
+Grade 3 0.073 0.074 0.133 0.133 0.024 0.027 0.077 0.078
+Panel D: $y_{1}=reading$, $y_{2}=math$
+Grade K 0.148 0.148 0.177 0.219 0.126 0.127 0.155 0.192
+Grade 1 0.105 0.105 0.130 0.137 0.070 0.071 0.119 0.132
+Grade 2 0.092 0.092 0.134 0.137 0.086 0.087 0.110 0.109
+Grade 3 0.083 0.083 0.107 0.106 0.050 0.050 0.085 0.086
+Panel E: $y_{1}=word$, $y_{2}=listening$
+Grade K 0.124 0.125 0.132 0.129 0.132 0.133 0.136 0.134
+Grade 1 0.105 0.106 0.122 0.132 0.068 0.071 0.094 0.108
+Grade 2 0.084 0.084 0.131 0.132 0.070 0.072 0.107 0.109
+Grade 3 0.090 0.090 0.132 0.131 0.030 0.033 0.065 0.066
+Panel F: $y_{1}=word$, $y_{2}=math$
+Grade K 0.138 0.139 0.164 0.200 0.128 0.129 0.154 0.184
+Grade 1 0.092 0.092 0.114 0.118 0.058 0.059 0.101 0.109
+Grade 2 0.089 0.089 0.124 0.126 0.081 0.081 0.106 0.107
+Grade 3 0.085 0.085 0.107 0.106 0.039 0.040 0.071 0.070
+school FE Y Y Y Y Y Y Y Y
+class type FE Y Y Y Y Y Y
+stu&tch char Y Y Y Y
+peer char Y Y
+  • 1.
+The table reports one minus the ratio of the variance of $u_{1}-f_{1}u_{2}$ to that of $y_{1}-f_{1}y_{2}$ for small classes (in the left panel) and regular (with or without aide) classes (in the right panel).
+All models allow for heteroscedasticity across small and regular classes.
+Estimates are adjusted for missing observations.
+  • 2.
+As indicated at the bottom of the table, the models may control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.
+Table 9: Standard Deviations of $Q_{x}y$ and $Q_{x}(y_{1}-f_{1}y_{2})$
+Small Classes Regular Classes
+(1) (2) (3) (4) (1) (2) (3) (4)
+Grade K
+$Q_{x}y$, y=read 62.748 62.169 33.699 26.827 66.321 64.827 35.534 29.143
+$Q_{x}y$, y=word 65.196 64.629 38.017 31.883 68.723 67.294 39.693 33.916
+$Q_{x}(y_{1}-f_{1}y_{2})$ 14.131 14.131 14.103 14.302 14.705 14.707 14.651 14.831
+Grade 1
+$Q_{x}y$, y=read 83.720 82.984 57.172 46.261 88.372 86.584 58.298 49.448
+$Q_{x}y$, y=word 81.314 80.624 57.574 46.620 85.880 84.192 57.174 47.750
+$Q_{x}(y_{1}-f_{1}y_{2})$ 21.475 21.470 21.350 21.239 22.352 22.346 22.183 22.048
+Grade 2
+$Q_{x}y$, y=read 93.586 92.295 57.595 39.409 87.831 85.406 54.519 41.740
+$Q_{x}y$, y=word 96.864 95.555 62.274 44.613 90.149 87.782 59.070 46.967
+$Q_{x}(y_{1}-f_{1}y_{2})$ 24.071 24.070 24.070 24.312 24.496 24.497 24.539 24.796
+Grade 3
+$Q_{x}y$, y=read 102.223 100.223 52.813 33.948 85.025 82.325 51.772 38.114
+$Q_{x}y$, y=word 102.419 100.495 56.860 39.320 86.959 84.387 56.262 43.608
+$Q_{x}(y_{1}-f_{1}y_{2})$ 20.365 20.367 20.330 20.323 20.839 20.833 20.730 20.721
+  • 1.
+Standard deviations of $Q_{x}y$ and $Q_{x}(y_{1}-f_{1}y_{2})$ for small (on the left) and regular classes (on the right).
+For $Q_{x}y$, the test score $y$ is reading or word score For $Q_{x}(y_{1}-f_{1}y_{2})$, $y_{1}$ and $y_{2}$ are word and reading scores respectively.
+The control variables for models (1)-(4) are the same as those for the main models in Table 3.
+The estimates of $f_{1}$ are from our efficient GMM estimations for each specifications and presented in Table 3.
+Table 10: Estimates of Exogenous Peer Effects for the Word-Reading Pair
+           Age            Black            Poor            Girl
+Grade K            4.126            0.028            0.005            -0.067
+           (3.268)            (0.060)            (0.023)            (0.028)**
+Grade 1            0.701            0.032            -0.037            0.036
+           (3.307)            (0.063)            (0.042)            (0.048)
+Grade 2            -3.204            0.144            -0.027            -0.113
+           (2.997)            (0.093)            (0.053)            (0.058)*
+Grade 3            2.157            -0.029            -0.036            0.025
+           (2.486)            (0.061)            (0.039)            (0.044)
+  • 1.
+Estimates and standard errors (in the parentheses) for exogenous peer effects of peer’s age, race, free lunch status and gender.
+Coefficients for dummies black, free lunch and girl are divided by 100 so they can be interpreted as the impact of one percentage point change in these variables.
+Note that these four variables are all test-invariant personal characteristics.
+The reported coefficients are $\rho(\beta_{v1}^{p}-f_{1}\beta_{v2}^{p})$.
+  • 2.
+* 0.1, ** 0.05, *** 0.01.
+Standard errors are clustered at the classroom level.
+All models allow for heteroscedasticity across small and regular classes.
+Estimates are adjusted for missing observations.
+  • 3. The scores are raw SAT scores in word study skills(word) and reading (read) for $y_{1}$ and $y_{2}$ respectively.
+  • 4.
+All specifications control for (1) school fixed effects (excluding one); (2) class type fixed effects (excluding small); (3) a student’s characteristics (free lunch, black, girl, age) and teacher characteristics (black teacher, master, years of experience); (4) peer averages of student characteristics.